A Bibliography of Publications in ACM SIGSAM Bulletin and ACM Communications in Computer

Algebra

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 March 2024 Version 3.76

Title word cross-reference

#10 [Loo75]. #11 [WR79]. #2 [BH72a, Cam72, Hea72c, Tut73]. #3 [Fit73, Fit74c]. #4 [Bar73, Fit73, GJ74]. #5 [Cam73]. #6 [HJ74, McC73, Sta74]. #7 [CMR74, FMW74, GJY75a, JG74]. #8 [Fit74a, Hal75, Har77]. #9 [Kah75].

(2, 4, 4) [Sha15b]. (C^{m+1}) [Gae99]. 1/π [Wan13]. 1/π^{2} [Alm11]. 16 [DS15, DSS15]. 2 [GZ05b, Gol91]. 2W [Mur00]. 3 [GZ05b, Myl05, NO91, Sha15b]. 3/2 [Fri74]. 3 × 3 [HKS19]. 5 [vHK12]. _{2}F_{1} [vHK12, vHI14]. _{2n} [Cam72, Sun73]. α [STC22]. arccoth [CDJW00]. modp [Tro82]. C^{1} [JSS03]. C^{n×n} [Din09]. D [Rue05, Tab23, Tsa00]. δ [NS18a]. F [BHMM13]. F_{4} [iL15]. GF(q) [Vil98]. GL(nR) [Bro09]. i [LC19]. J_{1} [Gol91]. k [Bar15, POdCW19]. L [GR08, Zwi84]. λ [Fla88]. ↔ [Alv00]. P^{3} [HM99]. n [BC02, Fit74b, GHL15, KH81, PM19, POdCW19, WL08]. ω [FGHR12]. P [WGD82, Mor99, RZ20, Sla91, Tro82, Vol95, Yun74]. p^{n} [Vol95]. Q [Yas08, Zim97, BDM13, CX08, CDM15, Ghe10, Sho15]. Q(4; q), forq = 5, 7 [DHS04]. R_{n} [IP11]. S [Zim97]. sin^{2} x + cos^{2} x = 1 [Fat85b]. Sym^{n}P_{m} [Gae99]. T [HMMR22]. W [Din09, MS16a, MW18, SFG13, SFGZ14, LLG08]. X [DGG15, She10]. Y [She10]. y^{(3)} + ay′ + by = 0, a, b ∈ C[x] [Ber99a]. Y_{2n} [BH72a, Hea72c]. ye^{y} = x [Gos98]. ye^{y^{2}} = x [Gos98]. Z [HL10]. Z2 [GK16]. Z[x] [SO09, Zwi84]. Z_{4} [MMSY15]. Z_{4}[x]/⟨x^{2} − 2x⟩ [MMSY15].

-adic [Mor99, RZ20, Yun74, Vol95, WGD82]. -adically [Tro82]. -Algebraic [Tab23]. -ary [Mur00]. -calculus [Fla88]. -characteristic [MW18]. -D [WL08]. -deformed [CDM15]. -difference [Ghe10]. -equivariant [GK16]. -graded [HL10]. -Gröbner [LLG08]. -groups [NO91, Sla91]. -ideal [Tsa00]. -integral [Zim97]. -invariants [She10]. -LLL [NS18a]. -modular [Gol91]. -modules [Rue05]. -orbit [HMMR22]. -out-of- [POdCW19]. -polynomials [GR08]. -prefix [KH81]. -problem [Sun73]. -simple [Bar15]. -split [Sha15b]. -stuffle [BDM13]. -th [Vol95, Zwi84]. -theory [STC22]. -type [vHI14]. -Weierstrass [Sho15]. -WZ [CX08].

//www.orcca.on.ca/sca2002 [Rei02].

’05 [Kau05].

1 [CMR74, Ver75]. 1.0 [GPS97]. 10th [Ano08a, Mad92]. 130 [LM08]. 14 [Ano01a, MP10]. 14-15 [Sit06]. 17 [ECTB^{+}14, MP12, Pos13]. 1968 [Col69]. 1969 [Mos70b]. 1974 [Jen74a]. 1978 [Gak79, Wan78d]. 1979 [Ng79]. 1981 [Wan81]. 1992 [Mad92]. 1994 [Hon94]. 1998 [LR98]. 1999 [Lin98a, Lin98b, RL98, WRL99]. 1st [Gro95a].

2 [KP97, Lan86, Loo72b, MF83, Rod84]. 2-8849-113-9 [Qui97]. 2-8849-114-7 [Qui97]. 2-distribution [Edo08]. 2000 [Ano99c, Ano99e, Ano99d, Ano99b, Ano99g].

’2002 [Ano01b, Ano01c, BBS^{+}02, Buc03, Dav03, Gim03, Sch02, Sit02b]. 2005 [Ajw05]. 2006 [Tra06]. 2008 [Ano08a]. 2009 [Hea09]. 2010 [Ano10a, Ano10d, Ano10g, Ano10f]. 2011 [Fit12, JCMG11b, KZ11, May11]. 2012 [Ano12, KRR12, SS12b]. 2013 [Kal13, Roc13]. 2014 [Ano14a]. 2015 [Ano13a, IS16, Kal15, LW16]. 2017 [Ano17a, Lic17]. 2019 [Mon20, YFJS19]. 21st [DMP^{+}02, tW99]. 22nd [AB21]. 23rd [DMP^{+}02].

3 [GVR04]. 3-540-67733-X [GVR04]. 360 [Jen79a, Kal82a]. 370 [Whi78].

4 [DP15]. 40th [Eng15]. 4th [Ano99b].

5 [AB15]. 5.2.2 [AB17].

60th [RVMH11]. 6th [Ano99f, Mor89].

70th [Ano99a, Ano99i]. 73 [Bah77]. ’74 [Jen74a]. ’76 [Jen76]. 78 [Loo77c]. ’79 [Ng79]. 7th [Cha90].

80 [Sho82, Sto80]. 80th [KZ11]. ’81 [Wan81]. ’83 [Ano82a]. 85 [Zas84]. ’86 [Cha86]. ’88 [Gia88, Gia89]. ’89 [Gon89, ISS89]. 8th [Cha91, Kot02].

’90 [WN90]. ’91 [Wat91]. ’92

[Wan92, Sto98]. ’94

[ACM94, Hon94, ISS94, Nod96a]. ’95

[Lev95, Ano95b]. ’96 [Ano95d, CL96, Lak96, Ano95c, Ano95f, DL96, Emi96, Hon96a]. ’97 [Cha96a, Hon96c, Tec98, Sto98, Cha96b, Enc97a, Enc97b, Hon96b, Wes96]. 978 [DP15]. 978-981-270-942-4 [DP15]. ’98 [LR98, Wei97, CW98]. 981-270-942-8 [DP15]. ’99 [Ano98, HKS99, Jef98b, Kal98, RL98, Zim98a, Zim98b, Ber99b].

A&M [Sch02]. A-hypergeometric [CDS99]. A. [LeS70]. AAECC [Mor89, Ano01a, Mio88]. AAECC-14 [Ano01a]. AAECC-6 [Mor89]. Abel [CTKR99]. abelian [Lab15b, WSW13, GGSR99]. Abell [Cor96a]. abnormal [Joh75]. Abramov [CHL14]. Abramowitz [CDJW00]. Absolute [CEGW08, BCG09, Cap86a, Sto76, Sut18]. absolute-value [Sto76]. Abstract [Bar10a, BCD^{+}23, Bul10, CvHL10a, FvH10, Fon08, GRT10, Ley10, Man10, Mat10, MJF^{+}10, Rad10, Sau10, WB83, YS10, vH10, AsW08, Alm11, AP08, AM08, Ben08b, BM08, Bur08, BP20, CX08, DK98, DSJL08, Doe08, Ebe08, Edo08, Elk08, GR08, GZ08, GMGE08, HL08, Ibá08, Ili08, Iwa08, Jai08, Jon08, JM08, Kel08, LW08, LM08, Lew08, LLG08, MGR^{+}08, MS08, Nov08, Pet08, Pri08, Pud08, QH08, Ril08, RG08, RZ20, Sch08a, Sed08, SY22, SCWL08, Tha08, Thi08, Tsi08, VJ08, WL08, Wat08, Wis84, Wol08b, Yas08, YK08, ZT08, Zel08, Zho08]. abstraction [Pur80]. Abstracts [AB84, Abd84, Ano82b, Ano95a, Ano00a, Ano00b, Ano00c, Ano00d, Ano10a, Ano10b, Ano10c, Ano14a, Ano17b, Ber99b, CCA20, CCA13a, CCA13b, CCA14a, CCA14b, CCA14c, CCA14d, CCA15, Enc97b, Gie08a, Gim03, HKS99, HP18, IS16, Joh10, Joh11a, Joh11b, Joh12a, Joh12b, Joh12c, Joh12d, Joh13b, Kak98, KZ11, LR98, Loo77b, Loo77d, Men16, Mon01, Nod96b, RVMH11, Rob01, Roc13, Sau81a, Sit06, Ter11, Tho09, Tun17, Wan78a, Wan78b, Wan79a, Zie15, Zim99, Zim00, edi21, Ajw05, Cap03, GKL07, Hit07, KV04, KZ06, May11, SS12b, Sit02b, SVZ06, Wan06a, Wan06b, Wan80c, Wan80b, Loo77c]. ACA [RL98, Hon96a, Kot02, Wes96]. ACA’2000 [Ano99f]. Academia [Ano17c]. Academic [Aca20, sG06]. Accelerating [Zim18]. acceleration [Lew08]. accepted [Ano10d]. access [O’B91]. accessible [Mon21]. According [CDJW00, EZ23, Zim76]. accuracy [Sha05].

Accurate [JGB12, McN04, Nag04].

accurately [SK10]. achieved [SK00]. ACM [Gon89, Jen76, Wan81, Ano95f, Ano13b, Gie08b, Öre84, ISS89]. ACM-SIGSAM [Gon89, ISS89]. across [KPW06]. action [HK05, Lev15b, TK22]. actions [CC08, CG11, Gal15, HMMR22, Rue05]. activities [Mio88]. activity [Coo97]. actual [BCCR80]. Adapted [Met23]. Adaptive [Cel22, MBS15]. Adding [van86]. Additive [CG19, DGL19]. adic [Mor99, RZ20, WGD82, Yun74, Vol95]. adically [Tro82]. ADIL [Fat14]. Admissible [Kre88, Wei87]. Advanced [Kou09, dQ78]. Affine [BM08]. AG [Joy05a]. against [LM08]. aged [DP16]. AI [Gle92]. aided [Emi15, Lou84, Thi00a]. Aldes [Bun83]. Alexandros [Ano14a]. Alexandros-Stavros [Ano14a]. ALG [Kli74]. Algebra [ARS^{+}19, AB21, ABR^{+}23, Ajw05, Ano95a, Ano95c, Ano99a, Ano99f, Ano99h, Ano99i, Ano99j, Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01c, Ano10a, Ano14a, Asl96, Buc02, CCR00, CGG^{+}04, CCM^{+}14, Coo97, DMP^{+}02, Fèv98, GVR04, GPS97, HKS99, IS16, Kak98, Kal98, Kot02, KP97, sL17, Lic17, Lin98a, Lin98b, Lob97, May11, Mon01, Mor89, Ost99, Rec97, Rei02, RL98, Rob01, Roc13, Sau81b, Sen90, SS12b, SB97, Sit02a, SVV95, SJ05, SW97, Ste06, TO97, WRL99, dR90, Abb04, ABCR07, Abd84, ACS92, AS16, ARS^{+}21, Akr80a, Akr80b, Ano10b, Ano10c, Ano17b, Ayd15, BS15, Ban08, Bar02, Ber99b, BCLA82, BTL99, BDM13, CCA20, Cal84, CL87, Cas08, CCM^{+}16, CL09, CM02, Coo04, CJ88, Cor99a, D’A15, DS97, Ebe16, CCA13a, CCA13b, CCA14a]. algebra [CCA14b, CCA14c, CCA14d, CCA15, Far15, FP89, Fat96, Fat15, FCB^{+}22, FSH02, GL10, Gim03, GPS08, HSW89, Hoh88, HP18, JN04, Joh10, Joh11a, Joh11b, Joh12a, Joh12b, Joh12c, Joh12d, Joh13b, Joy08, Joy09b, JČMG11a, KS94, KPR16, Kli74, Kog15, Kre88, LW16, Las84, LS03, LW99, Lew00, Loo74, Lou84, Mac89, Men16, Mid13, Min15, Moe78, Mon21, MR84, NP88, Nod96b, jYO20, Ovc14, Pag07, Pri08, Ric09, Roa08, RSSRV89, Sau10, Sha03, Sit06, Sit14, Sta10, Sto80, Sto84, Sto11, Sto13a, SS07, Suz09, Ter11, Tif15, Too15, Tri84a, Tun17, Wil84, Win84, Xue15, Zen08, Zie15, Zip84, di 82, edi21, Gro95a, Kre01, Qui97, Sit02b, Cor01b]. Algebraic [ACM94, ACF^{+}15, Ano95a, Ano95d, Ano98, Ano99c, Ano99e, Ano99d, Ano99g, Ano01a, BF74, Bro93, BKR23, Cal01, Cap03, Cha86, Cha96a, DN90, Doo99, Enc97a, Eng15, GK85, Giu01, Glo98a, Gon89, GPS97, Gut04, HHKR20, HV98, HEW^{+}14, HN19, Jef98b, Jen76, Kau05, Küc97, LR98, Lak96, Laz01, Lev95, Lin98a, Lin98b, Loo72a, Mon93b, Mor89, Mor02, Mou01, Neu69, NMS97, OS23, POdCW19, PH81, PS98, Rei00, ISS89, ISS94, SST^{+}97, Sen03, Tab23, Tam15, Thi00a, Thi00b, Thi08, Tra06, Tra00, Tsi08, Wan78c, Wan81, Wan92, WN90, Wat91, Wat05, dRSE23, ÁS99, Arn81, Bah71, BS74, BK04, Bib84, Bos84, BDE^{+}21, Bro03, Bro04, BCLA82, Bul10, Can69, CH85, CE77, CP78, CGL05, CM76, Col74, CRDMBR19, Ebn86, Ele15, Emi15, FR18, Fat81, Fat84]. algebraic [Fat15, FCW13, FGVC15, Fre15, GMM87, Gia88, Gia89, GJY78, Grö09, GMGE08, HHS20, Hei13, Her11, HMS10, Hod11, Ili08, Iwa08, Jen79b, Jua15, KS97, KYA08, Ke00, KSZ11, Kra83, Küc82, KKM15, Laz80, Lee10, Llo84, MK96, Man15b, Man11, Maz15, McI85, Med15, Mig77, Mio81, Mio84, Mor87, Neu80, Ng74, Ng79, Ngô11, Pag22, Pon87, Pon88b, Pon88a, Pre15a, RSS15, Rum77, SSS05, Sán17, Sar15, SI08, Sch99, SV99, Ser99, Sha03, SH09, She07, SLW15, Sun73, Tri11, Wan75, Wee74, Wei15, Wib15, Xia07, YA07, ZL13, Zim84b, Zim84a, dCW09, dT84, PH83]. algebraically [HW75, Ull06]. algebras [Bre05, CDM15, DM12, KiL16, Lau15, LS03, Lev15a, Poi15, Shi99, Stu10, ZZ15]. ALGLIB [SW85]. Algorithm [Abr00, Hea72b, Nat92, Sas22, Sek23, SY14, UW95, AMORH15, AW80, AM10a, AM10b, Bar15, BGG15, BGM10, Ber83, Bla14, Bro09, But85, CT19, CL82, CsL08, Dav82b, DX07, EGP11, EL01, Fuc08, Gal15, GZ08, GK85, Göb93, HJS13, HW75, HM13, HZ15, HMSP18, HNRS18, IKRT89, Iwa05, JMV09, JM22, Jon08, Kel15, Kem81, KM11, Loo72a, McG84, Mos69, Mur00, Nag21, NS18a, NS17, PM19, Pér99, Pop00, Ren82, Rou09, RSS15, Sal12, San18, SL10, Sao10, SS92, Sed13, SS93, Sun73, SS07, Tri84b, Via13, Wan77, Wan80a, WB83, WCF12, Yun74, Zho82, ZvH19, Zim03, Fat14]. Algorithmic [Col69, Koe93, HHS12, KPR16, KB78, KV04, Rad10, Shi99]. algorithmically [SY22]. Algorithms [Ano01a, Ano08a, Ber98b, Cor00, DM21, EJP^{+}01, FT13, Hei70, Hel15, KW97, Knu75, Lin98a, Lin98b, Mor89, Mor99, NTW97, PST97, SV93, Sta10, TMCP22, ZGA^{+}23, Zim84b, ACF^{+}15, BCG09, BCCR80, But92, CK04, CL09, Dav12, Ede13, ELME22, FSS15, FT15, GTLN18, GK88, Hod11, Kli78, Li15, Lin07, Man77, MY74, Mio84, Mon99, Nag17, NS18b, POdCW19, PW03, Pon88b, Raj80, Sak19, Sch91a, SB89, SLW15, Win84, Xie07, YS10, Yun73, Zim18, dCW09, dQ78, vH75, vdHL20]. ALLTYPES [Sch07, Sch08b]. Almost [DMSP10, JPRZ23, PR15]. Almost-Commuting [JPRZ23]. alternative [Fat15]. ALTRAN [BH72b, HJ74, Hal74, Hal75, Fel75b]. Always [Ree91, Ree92]. ambiguous [Kor77]. American [Ano01b, Bra98, Abd84, Gon88, Sit06]. AMS [Akr80a, Akr80b, Joy09a]. Amsterdam [Loo77c]. ANALITIK [HS74, Kor76]. Analog [BSB^{+}23]. Analysis [Ano99a, Ano99i, Ano00c, Ano00d, Ban08, CC83, GK88, Rei02, Sch83, vH82, AP08, Arn81, Bel83, BZ12, Bun83, But92, DGK04, Ebe16, HJS13, HJS16, IOP21, IT12, Jeb18, Ke00, Lac92, Nag12, POdCW19, dCW09]. Analytic [Loo72b, Iwa05, Ksi15, Par15, Rag15, SSS^{+}11]. Analytical [SV93, Rod84, Sta10]. analytically [Bar89, Pea84]. Anatomy [Dav79]. Andrew [Dec14]. Andrews [Ano99c, Ano99d, Tra00]. Animation [BKR23]. Annihilating [Mid13]. annotated [Pon87]. Announcement [Ano98, Ano99a, Ano99c, Ano99e, Ano99d, Ano99b, Ano99h, Ano99g, Ano99i, Ano01c, Cal01, Cha96a, Giu01, Kre01, CAB22]. Announcements [Ano01a, Lin98a, Lin98b, RL98, Zim98a, Zim98b, Ano09]. Annual [Cha90, Mad92, Cha91]. anti [Ben08a, CJS18]. anti-derivatives [Ben08a, CJS18]. Antiquantization [SS17]. Antitranslator [Kry84]. ANTS [KV04]. any [HW75]. ApaTools [Zen08]. Apéry [BB96]. Apéry-like [BB96]. APL [Hoh88]. Apollonius [TE05]. appearing [Wol08a]. Application [BSB^{+}23, BC81, FG99, LPA15, Oku97, PH81, Sha05, Cas08, Fat15, Fel75a, FCW13, IOP21, Lem15, NP88, PDP00, Sao10, Wan91]. Applications [Ano99f, BMK98, DSJL08, Gro95a, KW97, Kot02, Mon01, Pon87, Rec97, RL98, Rob01, SI00, TO97, WRL99, tW99, BGLH^{+}10, Ber99b, BW98b, BRRT08, CL11, Dur15, Gim03, HI93, How70, HMMR22, Kou09, Lud69, Ple69, Poh81, Spe08, SVZ06, Tec11, Tsi08, Wan80c, Zho15, Zhu15]. Applied [Ano95f, Ano01a, Fra80, Gon88, WSW13, Lin98a, Lin98b, Mor89]. Applying [CL09]. approach [Bar10b, BKSS10, Dim71, Ebe83, HN19, Kou09, MCP15, MBS15, Por09, Rad10, SIK05]. approaches [Col69, EV05, Xie07, ZKK11]. Approximate [BA21, BD14, CKW03, CS19, Nag08, Nag21, SST^{+}97, SS97, SO09, SW06, Ste06, Fas06, KS97, Lee19, Nag12, Nag17, RSS15, Ter10, Ter13, Tri84b, UM12, Zen08]. Approximation [Pop97, Sak97, Sek23, CFRT20, EV05, Pop00, SS13]. April [HKS99, Sit06]. Arb [Joh13a]. arbitrary [Ben08a, Pur80]. Archimedes [Loo75, Loo77e]. architectures [Via13]. archive [Bee98]. arctangent [MM89]. argument [Row81]. Arising [Cor00, Bre05, Cam73, Raj80]. Arithmetic [CLZ23, KPS99, Mur14, Web90, Yan91, Joh13a, Joh74, JdM12, LTdD^{+}10, LMRS08, LIJ20, Mat10, MBS15, PM07, Pur80, SE22]. Arithmetics [Stu23]. arrangement [FGVC15]. art [RSSU79, Zel08]. article [MBKP81]. Articles [Jef98a, Lob96]. artificially [CPS11]. ary [Mur00]. ASCM [Ano03]. ASCM’2000 [Ano99b]. Asian [Ano99b, Ano03, tW99]. Asir [IS09, SS07]. AskConstants [Sto17]. aspect [Pre15a]. aspects [Kan76, MY74, Sta10]. assessing [HOPY19]. assessment [AsW08]. associated [APS15, GK16]. Association [Ano99f]. associative [BS15]. Asymptotic [SFG13, dQ77b, Sal91, dQ77a]. asymptotically [GK88]. Asymptotics [BFS23, FS97, Rai11]. Atkinson [But92]. attack [Duj08, Iwa08, LM08]. Attempt [Sas22]. Attempts [Sht77a]. attraction [Dem12]. August [Ano99c, Ano99d, BBS^{+}02, Glo98a, Jen74a, Jen76, LR98, Sen03, Tra00, Wan81, WN90]. Austria [BW98b, Hon94, Ano99h]. Austrian [AF84]. Author [Loo77e]. authoring [LW08, Meu03]. Automata [Ada96]. Automated [CCR00, Gie08b, KRV21, Nai22, Pei87a]. Automatic [Arn81, Cam75, Cor00, De 99, EJP^{+}01, MGR^{+}08, Sto76, ABK^{+}16, DS10a, DS10b, Gat85, HP67, LD11, Sal91, Ver74, vdR73]. automatically [Fat03b]. automating [CW72]. Automation [Zid23]. Automorphisms [Ksi15]. Available [Bah77, Sau80]. avoiding [Pud08, Ren82, Zho82]. award [Kal13, Kal15, Ano17a, Gie08b]. Axiom [CGG^{+}04, Dal02, Joy08, Pag07].

B [Bro03, Bro04]. B. [Abr03]. back [BDE^{+}21]. Background [Abr98, Abr03, Abr09]. Backward [Nag12]. Balanced [MX09]. ball [Joh13a]. Barcelona [Mon01]. Barker [BKM07]. Base [Stu23, Iwa05]. Based [FH96, MZ97, AsW08, Alv00, Ba12, BS15, CL87, CMPS10, CK03, Ede13, EF15, ELME22, FCB^{+}22, FIS16, GMRT^{+}15, HKN10, IMC^{+}22, Jen77, KS81, KD18, LW08, Li09, MCP15, MG15, Pea84, PW03, Sak19, San18, SLW15, Thi00b, Zho15, Zhu15, di 82]. Bases [Buc91, BW98a, GS95, GB97, GLGVTZ00, HNT14, Li92, Met23, Mon93b, Oku97, Sat97, Sch91b, Win97, Win98, BW98b, CC08, Cor96d, Dah12, DPFD15, FSS15, GK16, GHK18, GSGZ15, HRR19, HZ15, IS09, JS06, Lar19, Lev15a, LLG08, Mon99, MA08, MW18, Poh81, RZ11, Sal12, SK10, Sat00, Suz09, WW11, ZK11, Zha13, BB80, Buc76a, Ebe83, KB78]. Basic [CCM^{+}14, Kot01, Mon99, CCM^{+}16]. Basis [JPRZ23, McG04, NO23, Sas22, Buc76b, CLZ08, CS19, Ede13, EF15, ELME22, Fas06, FGVC15, GL13, IK20, Liu15, MC95, MSV08, Nag12, Ren15, Sak19, SI17, Sit89, Ste08, Ste00, SLW15, Tam15, Zho08, vdHL18]. Bates [Dec14]. Bath [Ano99e, Ano99g]. Baxter [GSGZ15]. BC [Doo99]. BCH [BSBC15]. be [CDJW00, Mos70a]. Beamforming [BSB^{+}23]. before [KV19]. Beijing [Ano03, Kau05]. Bell [Whe87]. benchmark [MH85]. benchmarking [HL15]. benchmarks [MP16]. Bendix [WB83]. bent [Joy15]. Bergman [JT91]. Berkeley [Sch02, Vol98]. Berkovich [Ksi15]. Bernd [Sch02]. Bernoulli [Har00]. Bernstein [CS19]. Bertini [Dec14, Bat13]. best [Moe78]. better [GMGE08, Nag17].

Betti [MSS18]. between [CK03, Dav82a, EMT08, MR84, NS09, Sas22]. beyond [Dav03]. Bézout [CT19, SC17]. Bezoutian [DD99]. Bibliography [Bee98, Fel77, Can91, Mio81, Pon87]. bidegree [CEGW08]. bids [Ano13a]. bifurcation [DGK04, GHK18]. bifurcations [Tot15]. big [AM08, KS81, Mat10]. big-float [KS81]. billion [vdHM20]. Binary [Gor99, BvzGG^{+}01, GMMCMMP22, GK88, SS93]. Bini [BD14, EG96]. Binomial [APS15, BFS23, Gao15, CC08, KM11, dCS99]. biological [BFR^{+}18]. biology [Ebe08]. Birkhoff [BTL99]. Birthday [Ano99a, Ano99i, KZ11, RVMH11]. Bit [ES19]. Bitslicing [YS10]. bivariate [Bou15, CEGW08, KN00, Kal82b, Lar19, MP11, PM19, SS10, WCF12, vdHL18]. Black [Ebe16, CM22]. blending [BCCL10]. Block [Liu19, Via13, Vil98, BFE20, Gol91, HJS16, HNRS18, Xam99]. Blocking [SIK14, DHS04]. Blow [Sak97]. Blow-up [Sak97]. blowups [ES11]. BLS [BP18]. Bohemian [CT16]. Bonn [Wat91]. Book [Ano79, Cet97, Dav03, Qui97, Lün84, Cet96]. Boolean [IS09, Sat00, ZK11, ZKK11]. Boor [Ste08]. Borwein [Ano16]. bound [ASP15, LW94a, LW94b, Sán17]. boundary [KRR10, KRR12, Mae87a, Mio74, Par15, Tec11]. boundary-value [Mae87a]. bounded [Gre15]. bounded-degree [Gre15]. Bounding [Fre15]. Bounds [Li92, Vas99, BSBC15, KM05, MS16b, Nag04, NS17]. box [Ebe16]. boxes [CM22]. brackets [Koh10]. Branch [ECTB^{+}14, Pat96, PBD10, RJ96, Sho15]. Branch-Cut [Pat96]. branches [JN04]. Branes} [Buc03]. Braselton [Cor96a]. Breach [Qui97]. Bregenz [Ano99h]. Bricard [Cou15]. Bridge [Sas22]. brief [Fel75b]. Brill [BCL23, Jen15]. Bring [AJ03]. bringing [LMRS08]. British [Ano98, Jef98b]. broader [JMV09].

Bronstein [Bro13]. Brouwer [Ben08b]. Brown [McG84]. Bruno [Gie08b]. Buchberger [Gie08b, Göb93, Sas22, Tri84b, Tri11]. bucket [CJ18]. Budan [Akr81]. bug [Bah73a]. build [Mon21]. Bulletin [Ano82b, Jef98a, Lob96]. Burnside [Bin87a]. burst [OVB15]. burst-error [OVB15]. Business [Coo96a, Coo96b, Lak98]. Bylaws [Kal96].

C [Coo04, Joh13a]. CA [Vol98]. Cache [CMX10, HHM10, Fat03b, Sal11, Sal12]. cache-efficient [Sal11]. cache-oblivious [Sal12]. CAD [NDSM19, WBD12]. CADO [Tho09]. CADs [Bro03]. Calculating [AM08, Ter13, Ter10]. Calculation [Koe93, Mur14, Coh74, Hei70, IKRT89, Kem81]. calculations [Bar89, Ebn86, Kra83, Rod84, SV10]. Calculator [Ano93]. Calculus [Gak79, Ged99, Pra90, Sch82, BGJ11, BCCR80, Fla88, Kog15, jYO20, Qi20, Ste81, Tif15, Vor80]. Calendar [Ano96]. Call [Ano82a, Ano99a, Ano99c, Ano99e, Ano99d, Ano99b, Ano99h, Ano99g, Ano99i, Ano99k, Ano01b, Ano01c, Ano03, Cha96a, Enc97a, Jef98b, Zim98a, Zim98b, Ano13a, Ano13e, CJ18, Mon21]. called [Mos70a]. Calmet [Sei20]. CAMAL [Bou72, Fit74c, Hie75]. Cambridge [Buc03, Dav03, HI93, BF74]. CAMP [Buc84]. can [vH74]. Canada [Doo99, Mou01, Rei02, Ano98, Ano00a, Ano00b, Ano00c, Ano00d, Jef98b, Lev95]. Cancellation [SS98]. Cannon [Sao10]. Canonical [Con90, Buc76b, Bul10, Cou15, DJ96, Man10, dO99]. canonicalization [Liu19]. can’t [Fat96]. Cantabria [Gut04]. Cantor [Stu23]. capacity [Cas08]. Cardinal [JMV09]. career [sG06]. Carpets [SFH00, FSH02]. CAS [BKG18, Sha05, Var15, Zid23]. CASC [Ano99j]. CASC-2000 [Ano99j]. Case [GS97, GLGVTZ00, WR79, Fat73b, Iwa05, KD18, Myl05, Sal93, Sla91, Vil98]. Castelnuovo [BGM99]. cattle [Loo75, Loo77e]. Cauchy [KN97, KN98]. Cauchy-type [KN98]. CAYLEY [Ano91a, Sch91a, CR84, Can91, Joy15, San91]. CBMS [Sch02]. celebrating [sL17]. Celebration [Ano99a, Ano99i]. celestial [GL10]. Celine [Zim03]. Cellular [Ada96, IdC21, Leg84]. Center [Pad02, Sit14]. centralizer [But85]. Centre [Ano00a, Ano00b, Ano00c, Ano00d]. Century [tW99]. Certain [Shi99, JM98, Kel15, LMS12]. Certicom [LM08]. certified [vdHLM^{+}11, Ley10]. Certifying [GLS97, HRR19, Lee19, ME18]. CFP [SW96]. CGAL [FT13, FT15]. CGB [SS07]. CGS [SS07]. CGSQE [FIS16]. CGSQE/SyNRAC [FIS16]. Chains [Bac97, AKM16, AM16, Gen71, KA99, MV11]. Chair [Kal93b, Kal95, Wat96, Ano14b, Ano14c, Bro17, Cha97, Cor99b, Cor01a, Cor02, Kot14a, Kot14b, Wat97b]. Chairman [Dav83]. Challenge [Tec98, Mon93a, Sto98, Tro97, YDH^{+}15, vzG92, LM08]. Challenges [FM97a, Ayd15, FM97b, FS97]. Change [Fra80, FGHR12]. chaos [HN19]. chaotic [LC19]. Character [DI91, Sla91]. Characteristic [FG99, DM21, GH08, HW75, Hel15, LM16, Lev15b, MW18, Pag22, PW03]. characteristics [Kor76]. characterization [BB80]. Charles [Dec14]. chasm [RZ20]. Cheater [KI11]. Chebyshev [Mur10]. checking [ÁAB^{+}16, AM10a, BGM10]. Checkpoints [AKR20]. Chemical [Abd84, FSH02]. chemistry [Bal84, Sch71]. Chen [SL10]. Chern [Hel15]. Chevalley [EL01]. Chiang [Ano99b]. Chicago [Pad02]. China [Kau05, Ano03, tW99]. Chionh [Ano10e]. Choose [HEW^{+}14]. chosen [Ste06]. Chow [DET23, Li15, Qi20]. Chuang [Dal01]. Church [Win84]. circle [Mig77, TE05]. Circles [Lub15]. Circuits [BCD^{+}23, Oku97, HYH15]. citation [Kal13, Kal15]. citations [Can91]. City [Sit14]. CL [LW95]. Clarifications [AE83]. class [BD15, Ben08a, DPFD15, IMSS08, JMV09, KKM15, KKM18, Mos69, Ng74, PS15, SS17]. Classes [Ber98b, CTKR99, Hel15, JR91]. classical [DP15, Pec07, Ren15, Vor80]. Classification [Oku97, CTKR99, WSW13]. classifications [GHK18]. Classifying [Bet13, Dou99]. Classroom [Oll84]. Clean [EZ23]. cleverly [Pud08]. Clifford [Buc03, Fèv98]. close [Sas04, Sto17]. Closed [CP12, Ull06, Col74, HW75, Mur09]. Closed-form [Ull06]. closet [Per15]. Closing [Ano95a]. closure [AKM16, Tsa00, Vas99]. cloud [Hor15]. CloudMath [Hor15]. Clustering [dRSE23]. CMS [Ano16]. Coast [Ajw05, Ano95c, Ano99a, Ano99i, Ano00a, Ano00b, Ano01c, Ano10a, Ano14a, Coo97, HKS99, IS16, Kal98, Lic17, May11, Roc13, SS12b, Sit02a, Sit02b]. Coates [Ber83]. CoCoA [Abb04, ABCR07, AB15, AB17]. CoCoA-5 [AB15]. CoCoA-5.2.2 [AB17]. CoCoALib [AB15, AB17]. Code [SvH81, APS15, Gat85, JP15, Lan86, MCP15]. code-based [MCP15]. Codes [DS15, DSS15, GTLN14, Lin98a, Lin98b, Mor89, ASP15, Ayd15, BGG15, BSBC15, Bro13, Bul10, DGG15, DMCMM15, Ele15, GMMCMMP22, GTLN18, Joy05a, Joy05b, MMSY15, MMNR15, Mio82, MG15, OVB15, PT99, Pre15b, Sar15, Xam99, Ano01a]. codimension [BGM99]. coding [JM08, Man15a, Por09, Xam99]. coefficient [HKS19, Iwa05, LW94a, LW94b]. Coefficients [AP23, HM77, SVV95, van80, CM83, CPS11, FG03, Glo98b, Grö09, Grö10, NS09, Qi22]. Cofactor [Cor96c]. Cohomology [HNT14]. collaboration [Hor15, KPW06]. Collaborative [Min15]. collection [Mae87b]. Collins [Ano17d]. Colloquium [Bai02]. collusion [GMRT^{+}15]. collusion-resistant [GMRT^{+}15]. colored [BM08]. Columbia [Ano98, Jef98b]. Column [Vol96, Vol97b, Vol97c, Dav83, Vol97a]. comb walks [JPBCL20]. Combinat [Thi08]. Combinatorial [Lab15c, Val20, Spe08, Tha08, ZT08]. Combinatorics [Thi08, KPR16, SSS^{+}11]. combinators [Kri84]. Comment [BH72a, Fat98, Sch88]. Commentary [FBOS88]. Comments [Fat85a, Gaw79, Mac89, Tri84a, Ebe83, Sit89]. Committee [Kal96, Wat98a]. common [BA21, BCLS11, Gen99, Glo98b, HM13, MY74, Sed08, FL91, Yan91]. Communicated [Ano14a, YFJS19]. communication [BKKS05]. Commutative [GGSR99, GVR04, GPS97, Ste06, Abb04, ABCR07, BLH10, Hei15, Lev10, MR84, Rod84, Stu10]. commutativity [DI91]. commutator [Gla91]. Commuting [JPRZ23, NOT18, Lab87]. Compact [Cam76, ES19]. compactly [Sto11]. comparative [Kli78]. Comparing [Fat03a, Lew08, FN73]. Comparison [HEW^{+}14, LW99, Lew00, CJ88, KA99, McN04, NM13, NW83, Sun73]. compartmental [BCCR80]. compatible [DMCMM15]. compatibly [DRR18]. Complementary [DR01]. complete [Akr80a, BGM10, Cap86a]. completely [Mad15a]. Completeness [Fèv98]. Completing [Zwi84]. completion [FR18, RSV02, Win84]. Complex [Asl96, Bel83, Emi15, HMS10, Ren82, Too15, Zho82]. complexes [ES10a, Rou09]. Complexity [BMK98, CMPS10, DET23, Mio84, BK12, CE77, CMX10, DSM^{+}05, ES19, FSS15, GK88, Ili08, KN87, LMS12, Man77, RZ20, vH10, vdHL20, GRT10]. componenets [AM16]. Component [Dir00]. components [BSS15]. composability [The10a]. composition [LLG08]. compositionality [SK00]. comprehensible [Sht77b]. Comprehensive [CJS18, Bul10, FIS16, Man10, Zho15, Zhu15]. compression [YS10]. Computation [ACM94, Ano98, Ano99c, Ano99d, Bro93, Cha86, CHJ^{+}97, CDM15, Enc97a, FG99, Ghe14, Glo98a, Gon89, Hon94, IMSS08, Jef98b, Jen76, KF99, KN98, LR98, Lak96, Lev95, LZ05, Mor02, Neu69, Pop98, Rei00, Rei02, ISS89, ISS94, SHS96, SST^{+}97, SS98, Sen03, Stu10, Sut18, Tra06, Wan81, WN90, Wat91, Wat98b, ÁAB^{+}16, AHCDT15, Ano91b, Bar03, Bar08, BvM06, Bib84, BGJ11, BCLS11, BCLA82, Buc84, CL96, CC72, CGG83, CKW03, Din09, ES19, ES10b, Fat84, Fat99, Fel75a, Fri74, FIS16, GMM87, Gia88, Gia89, GRT10, GL13, Gle86, GTLN18, Gra87, Gri76, GJY78, Gri79, HLM^{+}17, Har00, HJS13, HHS12, HMS10, How70, Hue84, KN00, KS97, KO13, Ke00, KS19, KRR12, Kri84, Lac92, LLT18, Leg84, Lic84, Llo84, Mio74, Mor10, Neu80, Ng79]. computation [Nor82, Poh81, RZ11, The10a, The10b, Sat00, SVZ06, Vas99, Vor80, Wal89, Wan76b, ZL13, Zho07, Zho08, dT84, vdHLM^{+}11, vdHL18, Ano95d, Cap03, Cha96a, Doo99, Eng15, Giu01, Gut04, Hon96c, KW97, Kau05, Küc97, Mou01, SB97, SW97, Tra00, Vol97c, Wan92]. Computational [Ano95a, Ano99k, BMK98, GVR04, Hoh88, Ili08, MPW97, MY74, Rei02, Sha03, van80, Arn87, FT15, Hei13, Mid11, SSS05, FT13]. Computations [BMK98, EG96, Sto07, Abb04, ABCR07, ACS92, AY16, Arn10, Bos84, Cam76, CM76, CRDMBR19, Ebe16, EF15, GPS08, KS16, LS03, LN76, MK96, Mat21, SS09, Sit89, Tsi08, UM12, Zho15, Zhu15]. Compute [dQ77b, Hel15, HM13, Mor99, Sak19, YFJS19, dQ77a]. Computer [ARS^{+}19, AB21, ABR^{+}23, Ajw05, Ano95c, Ano99a, Ano99b, Ano99f, Ano99h, Ano99i, Ano99j, Ano00a, Ano00b, Ano00c, Ano00d, Ano01c, Ano03, Ano10a, Ano14a, Asl96, Ayd15, Bar02, BU93a, BU93b, BCLA82, Buc02, Cal74, CCR00, CC72, CGG^{+}04, Coo97, Cor01b, DS97, DMP^{+}02, FCB^{+}22, GPS97, Gro95a, Ham14, HKS99, IS16, Kak98, Kal98, Kot02, Kot08, Kre01, sL17, Lic17, Lob97, May11, Mon01, Nor80, Qui97, Rec97, RL98, Rob01, RSSRV89, Roc13, Sen90, SS12b, SH09, Sit02a, Sit02b, SVV95, Ste81, Sto79, TO97, Tha08, WRL99, ZT08, dR90, Abd84, ACS92, AS16, ARS^{+}21, Akr80a, Akr80b, ÁS99, Ano10b, Ano10c, Ano17b, Ban08, Ber99b, CCA20, Cal84, CL87, CR84, CJ88, Cor99a, DP15, Ebn86, CCA13a, CCA13b, CCA14a]. computer [CCA14b, CCA14c, CCA14d, CCA15, Emi15, Far15, Fat81, FP89, Fat96, Fat15, FSH02, GL10, Gim03, GPS08, HK08, HP18, JN04, Joh10, Joh11a, Joh11b, Joh12a, Joh12b, Joh12c, Joh12d, Joh13b, Joy08, Joy09b, JČMG11a, KS94, Kre88, LW16, Las84, LS03, LW99, Lew00, Loo74, Lou84, Men16, Min15, Mon21, MR84, Nod96b, jYO20, Pag07, Pea84, Pec07, Ple69, Ric09, Roa08, Sta10, Sto80, Sto84, Sto11, Sto13a, SS07, Ter11, Thi00a, Thi00b, Tif15, Too15, Tri84a, Tun17, Wan80c, Wan80b, Win84, Xue15, Zie15, Zip84, di 82, edi21]. computer-aided [Emi15, Lou84]. computer-based [Pea84]. Computer-generated [Tha08, ZT08]. computerized [Jen79b]. Computers [Ano99f, Fel77, CP88, FB70, Ste81, di 82]. Computing [AMORH15, AKM16, Ano95f, Ano99j, Ano08a, Arr15, BG90, Ber99a, BCD^{+}23, BFS23, Bou15, BDD^{+}18, BKR^{+}91, CG11, CDL^{+}11, CLZ08, CM83, Dal01, Edo08, EMT08, FGVC15, GR08, GGMAVT21, GMMCMMP22, Gen99, HV98, HNT14, Hub00, IVV99, JPP22, JPRZ23, JM18, JS06, Kau03, LT18, Lar19, LM16, Le01, MQB98, Nie84, NMS97, Rec99, Ren15, Ren17, RS91, SK10, Suz09, US06, Ver75, Vol98, Web96, Wei15, YK08, vdHM20, AM10b, Bar15, BK12, Bel83, BGG15, Ber99c, Bla14, Bro74, Bro03, Bro04, Cav75, ELME22, Ele15, Gen71, Gre15, HI93, HZ15, JGB12, JM22, Jua15, Kel15, KKM18, LM12, LMS12, LMF12, Man11, Mon11, Nab07, Rai11, RRSS20, Sal12, Sed08, Ste00, Sut17, Wan89, Zen04]. concept [Pri08, WS83]. concise [Sto17]. Concurrent [Wil92]. Conditional [NO23]. conditioned [Mai99]. Conditioning [Mon93b]. conditions [Par15]. conduction [Sha05]. Conference [ABR^{+}23, Ano91a, Ano99f, Ano09, BW98b, GK20, Gro95a, Kot02, Mor89, Pad02, RL98, WRL99, tW99, AS16, ARS17, ARS^{+}19, ARS^{+}21, Loo77a, Wan79b, Gie08a]. Conferences [JK02, RVMH11]. configurations [Myl05]. conflict [TE05]. Conformal [JT91]. Congress [Ano84]. congruences [Ara15, Rad10]. Congruent [Zim84a]. conics [CFGVDT18, HM99]. Conjectural [Joy05a]. conjecture [BKG18]. conjectures [Com18, Tha08, ZT08]. conjugacy [BHMM13, JR91]. conjugated [Ber06]. conjugates [BvM06, Mig77]. connected [Kor77]. Connectivity [IP11, QH08]. conservative [NOT18]. considerations [Eng87, HKN10]. constant [FG03, Grö09, Grö10]. constants [CP78, Wan89]. Constraint [Sat97, BK04]. constraints [NDSM19, YA07]. Construct [Bar91]. constructible [CLM^{+}08, Ull06]. ConstructibleSetTools [CML^{+}08]. Constructing [FPS16, MGRR21, Roy91, Ayd15, GK85, Sha05]. Construction [BGG15, Gol06, SI00, CK03, De 99, DGG15, Ina05, KB78, Rag15, San18, SIK05, SI17, Ver74, Xam99]. Constructions [ZGA^{+}23, Ara15]. Constructive [Ser99, RRRA03]. containing [CHJ14, JRH15]. Content [Dav00a, Koh00, MSP05]. context [Fri70, Var15]. context-free [Fri70]. continuation [Ley16, RSV02, Ver10, VY15]. continued [Mur09]. continuing [Lic15]. continuity [CJS18]. continuous [MCJ15]. contraction [Man15b]. contributed [Ano10a]. Contributions [RRRA03]. control [KS03]. controversy [Akr81]. Convergence [SI08]. convergent [Bar89]. conversion [Mio82, SS92]. Converting [Lan86]. Convex [AV99, Ost99, AHCDT15]. convolution [BH15, Ste08]. convolutional [MG15]. cores [MX09]. Corner [Cor96d, CJ96]. correct [KS16, NS18a]. Correcting [Ano01a, Lin98a, Lin98b, Mor89, DGG15, Ele15, Joy05b, MCP15, Xam99]. Correction [BK14, OVB15]. corrective [MS08]. coset [Atk81, Dim71]. count [Pud08]. counter [JP11]. counter-examples [JP11]. Counting [EZ23, LH15, Mur10]. Course [BW98a, Fit75, Jen75, Rec97, Akr80a, Akr80b, Hei13, Pea84]. courses [Bea15, GGdR^{+}15]. cover [Mon11]. Covers [MBRD23]. CP [MF83]. CP/M [MF83]. CQF [Kop20]. CRACK [WB95, Wol05]. Crandall [Wol13]. Created [Ano82c]. Creative [CHL14, BCCL10, Zel08]. crimes [Sto12a]. Criterion [MSM23, Arr08, WB83]. critical [BFE20, ES19, Win84]. critical-pair [Win84]. critical-pair/completion [Win84]. critique [Can69]. crossings [Mad15a]. crunchers [Ste81]. cryptanalysis [Bul10, Tam15]. cryptographic [LMTV05]. cryptography [Dur15, KD18, MCP15, YDH^{+}15]. cryptosystem [BS15, Iwa08, MG15]. crystal [Fel75a, JM98]. crystallography [LN76]. CSL [Tri00]. CSL-hosted [Tri00]. CtCoq [Gro99]. CTP [HKN10]. CTP-based [HKN10]. cube [KM87]. cubic [DM15]. CUDA [HLM^{+}17]. CUMODP [HLM^{+}17]. current [Pon88b, Pon88a]. Curriculum [Ano84, Tif15, Zel08]. Curve [HV98, AHCDT15]. Curves [Elk08, Zim97, AHM15, Arn81, Ba12, Bes15, BSS15, BP18, Doe08, Ele15, FGVC15, GR08, Gut15, HMS10, Hod11, IP11, JSS03, KRTZ20, KLV03, Ksi15, LIJ20, Lub11, MSS18, Pag22, Qi20, Roq15, RSS15, SV99, Sha03, Sha15b, Sha15a, Sho15, Tra15, Wei15, YFJS19, Yas08]. Cut [Pat96]. cuts [ECTB^{+}14, PBD10, RJ96]. cycles [Qi20]. cyclic [Ayd15, OVB15]. cyclohexane [Cou15]. cyclotomic [AM08, AM10b]. Cylindrical [BKR23, HEW^{+}14, Lee10, dRSE23, BDE^{+}21, Bro04, Col74, CJ18].

D [Ano13e, Ano17a, Buc03, Dec14, Mon20, GZ05b, Myl05, WL08]. D-Branes} [Buc03]. D. [BY02]. D5 [DSM^{+}05]. DAE [HC97]. Daniel [Dec14]. Darboux [She10, Tsa98]. Data [MP14, PH83, Bun83, FP89, FN73, KV19, MP12, Pur80, Too15, YS10]. Database [BI91b, SPA^{+}99, BI91a]. Day [Ajw05, Ano99a, Ano99i, Ano00a, Ano00b, Ano00c, Ano00d, Ano01c, Ano10a, HKS99, Lic17, May11, SS12b, Coo97, Sit02b, Ano95c, Ano14a, Coo97, IS16, Kal98, Roc13, Sit02a]. DD [JP19]. DD-finite [JP19]. Death [DP16]. debugging [KR85b]. Dec [LW16]. decay [KA99]. December [Ano99b, Kot08, tW99, Col69]. Deciding [Wei00, PS15]. Decker [Hei13]. decoding [Bul10, CL11, Joy05a, Qui12, Xam99]. Decomposable [Sek23]. Decomposing [Pau15, Sch16]. Decomposition [BKR23, FG99, HEW^{+}14, Wan98, dRSE23, BGLH^{+}10, BDE^{+}21, Bro04, Col74, DGL19, GZ05a, JM98, Lee10, NY03, Wan77, Wat08]. decompositions [Bla14, GZ08, Li09, MA08]. Decoupling [BNN14]. Dedekind [DMK19]. Dedicated [BBS^{+}02, GL10]. Deduction [CCR00, SV93]. Deductive [BI91b, BI91a]. deep [LR22]. Defect [JP11]. defined [Gla91, Yas08]. defining [BMT99, Kat87, Liu15]. definite [AZ91, Bel83, CHJ14, Koe95, Lic11, Lic15, Zim03]. definition [Sha15a]. deformed [CDM15, SS17]. Degree [BH97, Kno97, Li92, BvzGG^{+}01, DMCMM15, Gre15, HHKR20, NS17]. degrees [Sla91]. DEMOCRACY [Bar10a]. Demonstration [BG97, Sit02b, Sut17]. demonstrations [KRR12]. demos [Ano10a]. Dempster [Dav03]. Denominator [MS16b, Mur14, Mur09]. dense [MX09]. depend [Glo98b]. dependence [LWZ07]. Derivation [EJP^{+}01]. Derivative [BC02]. Derivatives [Hv83, Ben08a, CJS18, Gol06, Kem81, Koe94, Pri08, Sas04]. Descartes [KM05]. descent [FvH10]. descente [Rec99]. Description [Cap86b, Fel75b, Whi78]. descriptions [Loo72a, O’B91]. Design [Hol92, MP14, Zid23, dR90, Emi15, Eng87, GMGE08, HKN10, Hel10, Jen77, Jen79a, CL96]. designs [KSZ11]. Detecting [Ghe14, RG08]. Detection [KU14, NO23, AHM15, CHS16]. determinantal [CFGVDT18, SS13]. determinants [US06]. determine [FSH02]. Determined [Cal01]. Determining [HM77, NS09, Gal15, dO99]. Deterministic [DPQ15]. Dethroning [Kel08]. Developer [CGG^{+}04]. Developing [CHH^{+}99]. Development [ABK^{+}16, Gaw77, Lem15, Thi08]. developments [AP08, AD10, Ter11]. DEWCAD [BDE^{+}21]. diagnostic [Fat99]. diagonal [VJ08]. diagram [GHK18]. Diagrams [Cal74]. Dickson [IK20]. Dictionaries [Dav00a]. Dictionary [Koh00]. dies [Ano16, MSV21, Sei20]. Difference [AB14, AP23, Ghe14, AKR20, BD15, CP12, CsL08, Gao15, Ghe10, GHL15, LOW08, Li15, Med15, OS16, PR15, Roq15, Sch08a, Wib15, ZZ15, Zha13, ZvH19]. difference-algebraic [Med15]. Differences [Ged99, KF99]. Different [Fat78, Kli74]. Differential [AB14, Ano95a, BBS^{+}02, Bai02, Bar10b, Ber98b, Cal01, Cor96a, JPRZ23, KS15, Kog15, Poi15, Pre15a, PS98, Rue15, SB97, Wil01, vHI14, AKR20, AvH19, AZ91, Arr15, Bar15, Ber06, BCLS11, BRRT08, Cha81, Cla69, DM15, FvH10, Fre15, Ghe10, GHL15, Goy21, Grö09, Grö10, Ili08, Jer15, KPR16, KRR10, LH15, Lem15, Lev15b, LOW08, LYG12, Li15, LMF12, Mad15b, Mai15, Mau16, Mid13, MBKP81, Ovc14, Pag22, Rag15, RSV02, RT88, Rue10, Sán17, SV10, Sch16, SW06, She07, Sit06, Sit14, Tsa98, Wol08a, ZZ15, Zha13, vHK12, vdR73]. Differential-Algebraic [PS98]. Differentiation [Fat14, Gol85, Cam73, CE77, EB70, Fat98, HHS12, Rag15, WS83, dQ78, vH75]. Digital [Hol92]. dimension [BGLGMM18, Dah12, FSH02, Lev15a, Lev15b, ZZ15]. dimensional [Ban08, CC08, FT12, GS95, GHK18]. Dimensions [Med15, SFH00, CsL18, JS06, Stu10, Zha13]. Dimentional [AP23]. Dimick [CGG^{+}04, JCMG11b]. diophantine [MT15]. Dirac [IKRT89]. Direct [DN90, Kal93a, Zim14, Bar15, MS18]. Director [Coo93]. discharge [CJ18]. discipline [Ste06]. DISCO [CL96]. DISCOVERER [Xia07]. Discovering [Mat21]. discovery [KRV21]. discrete [AMORH15, Bet13, ES11, Hol04, HL08, Wol08a, Yas08]. Discriminants [McC97]. discussion [Bul10, Man10, McG84]. diseases [Ebe08]. Disk [van80, Gle86]. dispatch [GMC^{+}21]. displacement [BJS06]. Display [FL91, KPW06]. dissertations [Ano10b, Ano10c, Ano17b, CCA20, CCA13a, CCA13b, CCA14a, CCA14b, CCA14c, CCA14d, CCA15, HP18, Joh10, Joh11a, Joh11b, Joh12a, Joh12b, Joh12c, Joh12d, Joh13b, Men16, Tun17, Zie15, edi21]. Distance [Sek23, GTLN18, Liu19]. Distinct [Kno97]. distinguished [Ano13b]. Distributed [Dir00, AsW08, Ber99c]. Distribution [Hol04, Edo08, LPA15].

Divergences [Cal74]. diversity [KPW06]. Divided [Ged99, KF99]. Divisibility [Wol08b, DX07, NS09]. Division [Abr00, KS92, FN75, MP10, Yun74]. divisor [BA21, Glo98b, HM13, LIJ20, MY74, Sed08]. divisors [Gen99]. Dixon [CK03, JM22, Lew08, LPA15]. Dixon-based [CK03]. DMAS [AsW08]. DNA [BGG15, MSP05]. do [Dav82c]. doctoral [Ano10b, Ano10c, Ano17b, CCA20, CCA13a, CCA13b, CCA14a, CCA14b, CCA14c, CCA14d, CCA15, HP18, Joh10, Joh11a, Joh11b, Joh12a, Joh12b, Joh12c, Joh12d, Joh13b, Men16, Tun17, Zie15, edi21]. documents [Meu03]. Dodunekov [Ano12]. does [CJ18]. Doing [BU93a, BU93b, Maz15]. Domain [Con90, Cap86a, SI08]. Domains [Sat97, DMK19, Tec11]. Dongming [Ano17c]. Doron [RVMH11]. Double [FN75, Tot15]. doubly [BDE^{+}21]. Douglas [WRW17]. Drexel [Sen03]. Drinfeld [ACLM23]. dsolve [CJ18]. Dual [DM12, MMSY15, Ste08]. Dubna [AB21]. Dynamic [KR85b, Sor97, CWZ12, IT12]. Dynamic-debugging [KR85b]. Dynamical [Lab15a, JPP22]. Dynamics [Qui97, Ban08, Bur08, Too15].

EACA [Mon01, Rob01, Ber99b, Gim03]. EACA-2000 [Mon01]. EACA-2001 [Rob01]. EACA-2002 [Gim03]. EACA-99 [Ber99b]. Early [Ghe14]. East [Ano99i, Ano14a, Coo97, IS16, Kal98, Roc13, Ajw05, Ano95c, Ano99a, Ano00a, Ano00b, Ano01c, Ano10a, HKS99, Lic17, May11, SS12b, Sit02a, Sit02b]. Eastern [Sit06]. easy [TK22]. EAT [LPR99]. ECC2K [LM08]. ECC2K-130 [LM08]. ECCAD [Ano95b, Ano95c, DL96, Hit07, Kal98, Sit02b]. Edge [Joy15]. Edge-weighted [Joy15]. edited [Dav03]. Editing [KO83, SCWL08, vdHGG^{+}13]. editor [Hie75, Cet96, Cor96d, CJ96, Cor96f, Cor96b, Cor96e, Cor97a, Cor97b, Cor98b, Cor98c, Cor98a, Cor99b, Dew00b, Gie99a, Gie99b, Gie99c, Gro95b, Jef03]. editorial [Fat73a]. education [Lic84, Mio84, Sto79, Sto84, Var15, Ano84]. EEZ [Wan80a]. EEZ-GCD [Wan80a]. efectiva [Ibá08]. Effect [NS18a]. Effective [Ano99e, Ano99g, Dou99, Jer15, MBRD23, KS97, Xam99]. Effectively [Tif15]. Effects [Ano84]. efficiency [Dav82a, SH09, Zas84, vdHLM^{+}11]. Efficient [Bac97, CHS16, Ric09, Row81, Sal93, Sch94, Zho08, dCW09, AM10a, CGKZ05a, CGKZ05b, HW75, PS15, Sak19, Sal11, Wan77, WCF12]. eigen [Pan05]. eigen-solving [Pan05]. eigenproblems [Cor96d, Ste96]. Eigenvalue [SN97, CT16]. eigenvalues [Mor99]. eighth [Gim03]. Eiichi [Ida05]. Eilenberg [RRSS20]. Einstein [Sch88, NP88]. elected [Ano17c]. Elections [Wat97a]. electric [HYH15]. electronic [O’B91]. elegant [Ste81]. Elementary [CJ98, Cor00, Jeb18, JGB12, Zim84a]. elements [LWZ07, Zel08]. Eleven [Fat85b]. eliminating [WB83]. Elimination [Grö98, KW97, Arn10, BFE20, Col74, FIS16, HW75, Rue15, ZKK11]. ellipse [Kah75]. Elliptic [SVV95, Zim97, Jai08, Roq15, Yas08]. embedded [DRR18]. embeddings [BDD^{+}18]. Emerging [Ano10d]. empirical [Mat10, Nag04]. encoding [GTLN18]. encryption [Küc87, Tam15]. End [Buc02]. Endomorphisms [Lau15]. Energy [Pop98]. Eng [Ano10e]. Engine [MZ97]. engineering [KS03, Sal12, Xie07, Zho07]. England [ACM94, ISS94]. Enhance [Sas22]. enhanced [Eng75]. Enhancing [SI17]. enjoying [HK08]. ensure [EGP11]. entering [SCWL08]. Entries [BK14, Fuc08]. enumerate [HM99]. Enumerating [VJ08]. Enumeration [EL01, Atk81, Dim71]. enumerator [JP15]. enveloping [DM12]. Environment [Ham14, HC97, CHH^{+}99, Gro99, Mos70a]. Environments [CHJ^{+}97, Arn87]. epistasis [IVV99]. equality [NDSM19]. Equation [Wil01, CFGVDT18, Cha81, GH08, HYH15, KYA08, LR22, RT88, Sha05, Wan16, vTG03]. Equations [BBS^{+}02, Ber98b, Buc91, Cor96a, GK85, Ghe14, Izu14, KN98, Laz01, Mon93b, NTW97, PS98, Sch02, SB97, vHI14, AvH19, AZ91, ÁS99, Arr15, Bar10a, BMT99, BPS12, CvHL10b, CP12, Cla69, FvH10, FBOS88, Fat96, Gae99, GK84, Ghe10, GCG83, Grö09, Grö10, Ili08, Kat87, KS15, KKM15, Lab87, LH15, Laz80, Ley10, LBL20, LMF12, Loo72b, Mai15, Mai99, Mat21, MS16b, MT15, MBKP81, Mor87, Nat92, Rag15, Rim84, RZ20, Roq15, Sán15, Sch16, SH09, SS17, Ste10, Tri11, Wol08a, YFJS19, Yun73, vHK12, vdR73, Bai02, Cal01]. equilibria [CG11, Tot15]. equivalents [Mio82]. equivariant [GK16]. Erick [sL17]. Erratum [Ree92]. Error [Ano01a, BK14, Lin98a, Lin98b, MCP15, Mor89, DGG15, Ele15, Hol04, Joy05b, Nag12, OVB15, Pop00, Xam99]. Error-Correcting [Lin98a, Lin98b, Mor89, MCP15, DGG15, Joy05b, Xam99]. Errors [SS98]. Escorial [RL98]. espectrales [Ibá08]. Essen [Ano91a]. essential [Nor82]. estimated [Hol04]. Estimating [Dem12]. Euclid [Sas22]. Euclidean [Abr00]. Eugeny [Ano08b]. Euler [Hel15, Wan16]. EUROCAL [Ano82a, Zas84]. Euroforum [RL98]. Europaea [Ano17c]. EUROSAM [Ng79, Jen74a]. EUROSAM’79 [Wan79b]. Evaluation [Har79, Pre15b, Bro09, CGKZ05a, CGKZ05b, Hir78, MW17, MM89, RJ96, Tob70]. even [Maz85, Sto80]. Events [JK02, Ano13c, Ano13d, Sta20]. EVERYTHING [Zei12]. Exact [KLV03, Ste10, Cas08, DDK^{+}11, Hei70, JPP22, KS16, MBS15, SS09, Sau10]. examination [Nei71]. Examples [Gon88, Koe93, Sal91, WL99, Ebn86, JP11, WBD12, Zim97]. excerpts [Fat76]. exchange [ADS96]. Exchanging [BG97]. Executing [KO83]. Execution [SHS96, Fat99, PH83]. Executive [Wat97a]. exercise [Pur80]. expanding [Myl05]. expansion [Iwa05, Mur10, Sto08]. Expansions [dQ77b, BG90, Sal91, SI08, Vor80, dQ77a]. Experience [Gen74, Sho82, Oll84]. Experiment [Web90, BK04]. experimental [HKN10]. Experimentation [SJ05]. Experiments [AF84, AW80]. Expert [dR90]. explanatory [Lic84]. explicit [Bro09, TK22, Wil84]. Exploiting [LBL20, NDSM19]. exploration [Thi00b]. Explorer [CHH^{+}99]. exponent [Duj08, FGHR12]. Exponential [BCD^{+}23, BDE^{+}21, CP78, Heb15, RC76, Wei00]. Exponents [CL15]. Expression [vH82, Zho07]. expressionals [MP16]. Expressions [BG97, BKR23, SV93, Web96, Bro74, Cam88, CPS11, EB70, Fat72b, Hal74, IMC^{+}22, KH81, Kor77]. ExQRGCD [NM13]. Extended [BCD^{+}23, KO83, PST97, SY22, BP20, Fla88, Ina05, IOS06, JP15, Mur00, RZ20, San18, SIK05, SI17, Wis84]. Extending [DGG15, JMV09, Ebe16, KSZ11]. Extensible [Wil01]. Extension [Iwa05, Cap86b, EVM21, WS83]. Extensions [Gla91, ÜHK82, DGL19, GGSR99, GR99, Lev15b]. Exterior [Sch82]. EZ [San18]. EZ-GCD [San18]. Ezcaray [Rob01].

F5 [Arr08, EGP11]. facial [WRW17]. Facilities [Hv83, vHvH87]. facility [BGJ70, Gat85, Gre72]. Factor [OS23]. Factored [Hal74, Bro74]. Factoring [BvzGG^{+}01, CM22, MJ14, Nov08, NvH08, Pag22, WR73, Wan75, Wan76a, CEGW08, GHL15, Jef10, Tro82, ZvH19, vH10]. factorisation [ABD85]. Factorization [Ber98b, Ber06, CJ97, HL10, Ina05, Len84, Mig80a, Mon93a, Tsa98, WL08, Zim76, BCG09, BDM13, DMK19, Goy21, Iwa05, Kal82b, Len81, MG89, NW83, SO09, Sch94, SW06, Tho09, vzG92]. Factorizations [Kno97, Sut18]. factorize [WCF12]. Factors [McC97, Gre15, HJS16, IOS06, LW94a, LW94b, Mig87, PM19]. Faculty [Ano08b]. Faithful [BKR23]. Fake [KU14]. Families [CMT14, DM12, Koe94, Lub11]. family [HKS19, SC17]. fan [DSJL08, DMCMM15, OS06]. fans [Hub00, Rue05]. Farewell [Zei12]. Fast [BCLS11, Dav82a, FGHR12, IKRT89, MW17, Sed13, Xie07, ZvH19, vdHL18, AM10b, BFE20, GK88, HMSP18, Kem81, LMRS08]. faster [CM02, Fat03b]. Fateman [Gaw79]. father [SSS^{+}11]. Faugère [EGP11]. feasible [CFRT20, LM08]. features [GMGE08, HS74]. February [BW98b]. Feedback [DR01, Thi08]. Fenchel [BvM06]. Fermat [Lew04, MP16]. few [Com18, JS06]. Feynman [BKSS10, Cal74]. FFCSA [ZGA^{+}23]. FFT [CMPS10]. FFT-based [CMPS10]. FFTs [MJF^{+}10]. FGLM [HNRS18]. Fibonacci [CC16, Kel08]. fibrations [MGRR21]. Field [Kno97, ZGA^{+}23, Far04, Lev15b, PW03, RS89, Sha15a]. Fields [BD14, MQB98, OS23, Rei02, AV99, Bar10a, BDD^{+}18, Col74, DM21, DPQ15, DRR18, Dem12, Ebe16, ES10b, GH08, Gut15, HW75, Huf15, JM10, NOT18, NY03, PR15, Ren15, RZ20, Sch08a, Ull06, Wan75]. Fifth [ABR^{+}23, Ber99b, Mos74]. Filters [Hol92]. Final [Fat76, Pfe21]. find [Mig77, Mur09, Per15, TK22]. Finding [Lan93, WRW17, vHK12, DPQ15, EV05, KKM15, MS18, Pan05, Uga12]. finds [AZ91]. Finite [FG99, Kno97, Mon93b, RS89, Rue05, ZGA^{+}23, Bar10a, BDD^{+}18, Dav84, DPQ15, DRR18, Erö10, Fre15, Fri74, GH08, GK85, Gut15, Huf15, JM10, JP19, Kas80, Lab15b, MBKP81, NY03, PW03, PR15, Raa11, Teo15]. Finitely [Abr98, Her98, MQB98, SFH00, BTL99, FSH02, Lev15a]. First [Ano99c, Ano99e, Ano99d, Ano99b, Cha96a, Bar15, Cla69, FG03, Hei13, MS18, MBKP81, Ngô11, Hon94]. first-order [Bar15, Ngô11]. Fix [Ste08]. fixed [Ben08b, EGK09, Hol04]. FJCC [Mos70b, Col69]. Flajolet [SSS^{+}11]. Fleshing [MS16a]. Flexible [Bac97, Cou15, Meu03]. FLINT [GR16, DRR18]. float [KS81]. Floating [SS98, KS97, MBS15, SK10]. Floating-point [SS98, KS97, MBS15, SK10]. floats [Sto17]. flow [Bun83, LBL20, MC95]. fluids [Rim84]. forgotten [RRRA03]. Form [Con90, CMT14, FG99, CP12, Li15, MBKP81, Ull06, WRW17]. FORMAC [Bah69, Bah73a, Bah73b, Bah74, Bah75b, Bah77, Ber70, vH74, EB70, Lud69]. Formal [Hv83, Izu14, Koe93, Lob96, Mad15a, Ost99, Pei87b, Pra90, Wee74, Loo74, Mad15b]. Formally [Jef98a]. FORMAT [GJY75b]. former [Ano16]. Forms [DET23, Gor99, Bar15, Bou15, Buc76b, DM21, DJ96, LM12, Mid11, Pre15a, TMV18, Wei87, dO99]. Formula [BD14, dQ77b, Che83, Don83, Nei71, Oll84, ST02, Sch71, dQ77a, vdR73]. formula-manipulation [vdR73]. Formulae [BB96]. Formulas [Bar91, Gak79, Ged99, Alm11, Bar89, Ben08a, DD99, Gae99, SCWL08, TK22, Ull06]. formulation [Jen79b]. FORTRAN [Bah69, Ber70, Fat78, Fri70]. forum [Öre84]. Fourier [AS14, KS14, Akr81]. fourth [ARS^{+}21]. Fourther [Coo97]. FP [Mur14]. FPGA [SL10]. fpLLL [EVM21]. FPS [TK22]. Fraction [MJ14, NTW97, Bar89, Mur09, Sto08]. Fraction-Free [MJ14, NTW97]. fractional [Sto13b]. fractions [Lem15, LOW08]. frame [FR18]. framework [Ber99c, KS13, KSZ11, LD11, Tec11]. France [Ng79, DMP^{+}02, Mor02]. Francisco [Col69]. Fraser [Ano98, Ber98a, Doo99, Jef98b]. Fredholm [Loo72b]. Free [FG99, MJ14, NTW97, Fri70, Gal15, GMMCMMP22, Hea09, Mau16, Mio74]. friendly [BP18, HHM10]. full [SW06, Tam15]. FUNARG [Mos70a, San71]. Function [BB96, Piq90, Pop97, RJ96, Yan91, Cap86a, CHJ14, DM21, Din09, ES10b, LPA15, MS16a, MM89, Mos70a, Pig90, Piq89, Pop00, Ren82, SFG13, SFGZ14, Smi84b, Sto76, Zho82, Mos70a]. Functional [Izu14, Kri84, BPS12, Wat08, WS83]. functionals [Ste08]. Functions [Asl96, BH72a, FL91, Gak79, GLGVTZ00, Hea72c, MSM23, Pop98, Tab23, AKM16, BH15, BCCL10, Cam72, Cam88, CDS99, CsL09, CsL10, Gol85, Gor98, Heb15, JN04, JGB12, JP19, Joy15, Jua15, Kou13, Lac92, LZ05, Man15b, Maz85, MCJ15, Mos69, Ng74, PS15, Raa11, Rai11, RC76, SI08, SPA^{+}99, Sta10, US06, VHC17, Wan13, Wan89, vdHL21]. functorial [Bar10b, Poi15]. functors [BRRT08]. funding [KO13]. funtf [FR18]. Further [AP08, Gaw77]. future [Bal84, Fat84, Hue84, Sun74, Zip84].

G [PT99]. G-sets [PT99]. G. [Hei13]. G2 [BP18]. G2V [Sal12]. Galerkin [Mio74]. Galois [Arr15, Ber99a, Roq15, Sán15, Sut17, Sut18]. game [Tha08, ZT08]. gamma [Heb15, IKRT89]. gamma-matrices [IKRT89]. GAP [AY16, DGS16, ES10a, Joy09b, Lin07, pro12]. Gathen [Cor01b, Mon93a]. Gauss [Fee99, Fuc08, GM94]. Gaussian [LPA15, Sto09]. Gaylord [Ada96]. GCD [BK12, CLZ08, CS19, Hea72b, KI11, Mur00, Nag08, Nag17, Nag21, PST97, San18, SS97, Sed13, Ter10, Ter13, UM12, Wan80a]. gcds [vdHL21, Sas22]. Gelfand [ZZ15, Zha13]. gene [DSJL08]. General [BRRT08, Kor76, Coh74, GK84, GCG83, Kra83, Mac89, MJ13, MBKP81, Ngô11, Ore99, Tec11]. general-size [MJ13]. Generalised [Stu23, BFE20]. generality [Dav82a]. generalization [Wat05]. Generalized [Lev15a, MSM23, NMS97, GZ05b, GMMCMMP22, JP15, Lau15, Lev15b, MS16a, Rag15, SFG13, SFGZ14, WW11]. Generated [MQB98, BTL99, Lev15a, MJF^{+}10, Tha08, ZT08]. Generating [Ged99, GTLN14, SVV95, Rai11, Wan13]. generation [GMRT^{+}15, Gat85, Llo84, MJ13, SS17, YDH^{+}15, dT84]. generators [Gra91, MG15, YK08]. generic [Lar19, LIJ20, vdHL18]. Genetic [PST97, GTLN18]. Genoa [Tra06]. GENOM3CK [HMS10]. GENTRAN [Gat85]. Genus [Doe08, Sha15b, Elk08, GR08, HMS10]. GeoGebra [ABK^{+}16, KRV21]. Geometric [Fèv98, RSV02, Bar10b, HSW89, Man11, Nie84, Sut18]. Geometry [Ano99e, Ano99g, BU93a, BU93b, FT13, GPS97, MA08, PBD10, Rei02, SJ05, BZ12, CWZ12, DP15, FT15, Hei13, KLPW17, LW08, Maz15, Pec07, SSS05, Tsi08, Wol08a, Ano95a, BU93a, BU93b]. GeometryEditor [LW08]. George [Ano17d]. GeoText [CWZ12]. Gerdt [AB21, MSV21]. Gerhard [Cor01b]. German [LR98]. Germany [CL96, Glo98a, Kot08, Vol97c, Wat91]. GF [Gra91]. Giesbrecht [Ano13b]. Gilbert [LC19]. Given [Ste99, Gra05, Mur09, RC05, vTG03]. Gjunter [RRRA03]. Global [Bun83, DM21]. GNU [tW99, vdH04, vdHGG^{+}13]. Goals [Eng87]. GOEDEL [Ebn86]. Goes [Bra98]. Golden [DR01]. gonal [Wei15]. Good [Moe99, Tra15]. Gordon [Qui97]. Goto [Ida05]. GPGCD [CT19, Ter10]. GPU [BGLM23, MP11]. GPUs [VY15]. Graded [PT99, CM15, HL10]. gradient [Ber70, Tob70]. Graduate [Sit14]. Graeffe [vdHM20]. GRAPES [Emi20]. graph [Nie70]. graphical [Dim71, LS85]. Graphing [CJ98]. graphs [Jen15, JM98, Joy15, KH81, Roy91, Thi00a, Thi00b]. Gravity [TO97]. gray [DS15]. greatest [BA21, Gen99, Glo98b, HM13, MY74, Sed08]. Greece [Kot02]. Greedy [CK04, BGG15]. Gregory [Mio82]. GRGEC [TO97]. Griesmer [Fit12]. Gröbner [BW98b, Abr09, BB80, Buc91, Buc76a, BW98a, CC08, Cor96d, Dah12, DSJL08, DMCMM15, DPFD15, Ebe83, Ede13, EF15, ELME22, FSS15, FIS16, GS95, GHK18, GLGVTZ00, GSGZ15, HRR19, HZ15, Hub00, IS09, JS06, KB78, LM12, Lar19, Lev15a, LLG08, Liu15, Man10, McG04, Mon93b, MC95, Mon99, Mon11, MA08, MW18, Nab07, Nag12, NO23, Oku97, RZ11, Sak19, Sal12, SK10, SI17, Sas22, Sat97, Sat00, Sch91b, Sit89, SLW15, Suz09, Tam15, WW11, Win97, Win98, ZK11, Zha13, vdHL18]. Gröbner-bases [BB80, Buc76a, Ebe83, KB78]. Groebner [Zho15, Zhu15]. Group [CC08, Fel77, LN76, ASP15, BHMM13, Ber99a, CR84, CG11, FB70, HK05, KM87, iL15, Lev15b, Man15b, O’B91, Pos13, Qi20, San91, Sao10]. Groups [BI91b, HNT14, KS14, RS91, Arr15, BI91a, But85, Gal15, GGSR99, Gla91, Gra91, JR91, Lab15b, Lin07, Med15, NO91, Öre84, Roq15, Sla91, Sut17, Sut18, Wib15]. Guangzhou [tW99]. GUAVA [Joy05b]. guessing [KV19]. Guest [Dew00b].

H [Dav03, Qui97]. Hadamard [Sch83]. Hagenberg [Hon94]. Hagenberg/Linz [Hon94]. Hakki [Mor10]. Hamiltonian [Man15b]. Hamming [GMMCMMP22, Sek23]. handled [Ber70]. Hankel [BK12, BA21, Sen90]. Hans [Ano99a, Ano99i]. Hansen [SVV95]. happen [vH74]. hardcover [Qui97]. Hardware [Zid23, Leg84]. Harmonic [DMH15]. Hashing [BP18, Pon87]. Haskell [VJ08]. Hauenstein [Dec14]. Haven [Jen75]. Having [AP23, Ghe10]. Hawaii [Küc97, Lin98a, Lin98b]. hazards [Lic15].

Heart [Ste96]. heat [Sha05]. Height [LW94a, LW94b]. Heights [Bes15, Jen76, Jai08]. Help [dR90]. Hensel [Ina05, IOS06, MY74, Mio82, OS06, PM19, Sal11, San18, SI00, SIK05, SI17]. Hensel-type [MY74]. Herb [KZ11]. Herbert [Zei12]. Hermite [BCCL10, Con90]. Hermitian [Edo08, EVM21]. Hessian [CHS16]. Heun [SS17]. heuristic [Bar10a]. Heuristics [HEW^{+}14, Sao10]. Hickory [Pad02]. High [GMC^{+}21, Web90, AF84, Dav82c, JS06, Li09, Lou84, PM07]. high-level [Dav82c]. High-performance [GMC^{+}21, Li09]. High-precision [Web90]. Higher [Gak79, MS16b]. Highly [BNN14]. Hilbert [CM15, GLGVTZ00, Gor98, LT18]. HIMMO [GMRT^{+}15]. HISC [Hec94]. Historical [Abr98, Abr03, Abr09, Tob70]. history [Fat73b]. Hölder [Wib15]. Holonomic [Kou13, Kou09, PS15]. Homogeneous [GLGVTZ00, Ghe10, Gor98, KM11, MG89]. Homogenized [FR18]. homographic [Myl05]. Homologa [Ibá08]. Homology [MBRD23, RS91, Dou99]. homomorphic [Tam15]. Homotopies [Kot01]. homotopy [Dem12, Ley16, Ver10, VY15]. Honolulu [Lin98a, Lin98b]. honor [KZ11, RVMH11]. Honours [Gie08b]. Hopf [BDM13]. Horner [CK04]. Horospherical [Man15b]. hosted [Tri00]. HTML [Soi97]. http [Rei02]. hull [AHCDT15]. human [KS94]. Hurwitz [DET23, KS15, Tre90]. Hybrid [KN97, KN98, KN00]. hydrodynamic [CJ88]. hyper [AM10a]. hyper-regularity [AM10a]. hyperbolic [Ben08a]. hyperelliptic [Bes15, GR08, Gut15, LIJ20, Sha15b]. hyperexponential [AZ91, LZ05, LWZ07]. Hypergeometric [CHL14, Le01, Wan13, CGKZ05a, CGKZ05b, MS18, OS16, TK22, Zei91, Zim18, CDS99]. hypersurfaces [ES19]. Hypothesis [MSM23].

idea [FN73, KS16]. Ideal [Abr98, Con90, GTLN14, Her98, APS15, BGM10, Gao15, GGSR99, LLG08, Liu15, MMNR15, Mon11, MA08, Per15, RRRA03, Tsa00]. Ideals [GS95, Li92, Wan98, BLH10, Buc76a, CC08, DMK19, FT12, GZ05a, GZ08, GGMAVT21, Hub00, IdC21, JS06, KM11, KB78, NY03, Rou09, Spe08, dCS99, vdHL18]. idempotent [Dur15]. Idempotents [GTLN14]. Identifiability [BCCR80, HOPY19, IOP21]. identification [Hir78, KI11]. Identities [BH97, CP88, HRR19, McI85, Zei91]. identity [GMRT^{+}15, Qi22]. identity-based [GMRT^{+}15]. II [Tha08, BH15, Lab15c, LW94b, Wan89, Zim00]. III [Dav82b]. Iliopoulos [Ano14a]. ill [Mai99, Ste06]. ill-chosen [Ste06]. ill-conditioned [Mai99]. Illinois [Pad02]. IMACS [Gro95a, RL98, Ano99f, Hon96a, WRL99]. IMACS-ACA [RL98, Hon96a]. IMACS-ACA’2000 [Ano99f]. image [FCB^{+}22]. image-based [FCB^{+}22]. images [MMSY15]. impact [Man77]. implement [Sto11]. Implementation [BHMM13, EJP^{+}01, FF82, IS09, NY03, Sch82, Smi84a, SS07, CL96, CMX10, Fri70, HS74, JP19, Koh10, Loo74, McG84, Raj80, SL10, Suz09, Whi78]. implementations [The10a]. Implementing [LM08, Thi08, Pur80, SLW15]. Implications [Wan76b]. implicit [CFGVDT18, FGVC15, YFJS19]. Implicitization [Ore99, EK12, JSS03, KLV03, Rue10]. importance [Grö10]. important [Ste06]. improve [HKS17]. Improved [CHL14, Hea72b, Jar12, KS14, LIJ20, Maz85, Iwa08, KB78]. Improvement [San18, But85]. improvements [Nam86]. Improving [Ede13, Tri84b]. Incidence [BU93a, BU93b]. including [Iwa05, Zel08]. inclusions [GK85]. Incoming [Gie99c]. incomplete [Heb15]. Inconsistency [GLS97]. Incremental [FCW13, Ede13, GZ08]. Indefinite [Zim14, Koe95, Lic11, Wan91, Zim18]. independence [CP78]. indeterminateness [Kem81]. Index [Ano82b, dCS99]. Indexed [Fel77]. indices [DMH15]. induced [HN19]. inductive [RZ11]. Industrial [BC81]. Inequalities [Mig87, Xue15]. inexact [NS09]. infectious [Ebe08]. Infinite [AP23, CX08]. Infinite-Dimentional [AP23]. INformatics [Ano01b]. Information [Coo93, Koh00, Pat96, ADS96, Arn10, HKS17, Dal01]. Infrastructure [Koh00]. Infrastructures [Fon08]. Initial [PS98, Sak97, Ili08]. Initiative [Bra98]. innermost [Sti83]. Innovator [Gie08b]. INRIA [Vol96, Vol97b, Vol97c]. INRIA-Rocquencourt [Vol97c]. inscribed [Mor10]. Institute [Jen74a, Rei02, Vol97c, Vol98, Sit06]. instruction [HLS84]. integer [JM18, Mig87, Mur00, Pfe21, SS92, Sed08, Tho09, YK08]. integer-to-rational [SS92]. Integers [Li92, Vol95, Cav75, Fit74b, Mig77, Nag08, Sed13, Sto09, WR73]. integrability [Tsa98]. integrable [CTKR99, Mad15a, WSW13]. Integral [KN97, KN98, LR22, Bar89, Dav79, Kal82b, Loo72b, SH09, Vas99, Zim97]. Integrals [Gak79, Piq90, AZ91, Bar03, BH15, Bel83, BKSS10, CHJ14, Grö09, Heb15, JRH15, KKM15, KKM18, MCJ15, Pig90, Raa15, Roa08]. Integration [Buc90, Dav84, Heb15, Kas80, KO83, Raa11, Sor97, Cam88, Col69, DDK^{+}11, Fat15, IMC^{+}22, JRH15, Jua15, Kan76, Koh10, Lab15c, Lem15, Lic11, Lic15, MBKP81, Mos69, Ng74, Piq89, Ren82, SS83, Zho82, Zim03]. integrator [Hel10]. integro [BRRT08, KRR10, Mid13]. integro-differential [BRRT08, KRR10, Mid13]. intelligent [CWZ12]. Intensive [BW98a]. interaction [KS94, Kli74]. Interactive [Dir00, KO83, KR85a, Wri00, BGJ70, Fat73b, Jen70, LaP72, LW08, LS85, Oll84]. interactively [Kal82a]. interdisciplinary [Far04]. interest [Öre84]. interesting [Bah73a]. Interface [ASW89, CRDMBR19, LS85, Sto13a, vH84]. interlaced [TK22]. intermediate [AB19, vTG03]. intermolecular [Bre05]. International [ACM94, Aca20, Ano84, Ano95d, Ano99c, Ano99e, Ano99d, Ano99f, Ano99g, Ano99j, Ano08a, Bai02, Bro93, CL96, Cha96a, Doo99, Enc97a, Eng15, Gia88, Gia89, Glo98a, Gon89, Gro95a, Gut04, Hon94, Hon96c, Kau05, Kot02, Küc97, LR98, Lak96, Lev95, Mor89, Mor02, Mou01, Ng79, Pad02, Rei00, ISS89, ISS94, Sen03, Tra06, Tra00, Wan92, WN90, Wat91, Ano98, Cap03, Giu01, Jef98b]. interplay [MR84]. Interpolation [Car03, KU14, vdHL14, AGR13, Cel22, CsL08, CsL09, CsL18, IK20, JM10, JM22, KN00, KS19, MT15, NS17, NS18b, vdHL21]. Intersection [MQB98, Ba12, BGM10]. INTERTWINED [Zei12]. Interval [BMK98, Mur14, OS23, Abb14, KS16, Mur09, NS18a]. Interval-Symbol [OS23, KS16, NS18a]. Intrinsic [CGL05, DMK19]. Introducing [Cal84]. Introduction [GM94, Smi74, Ebn86, Pea84, SS68, Zel08, Wan78c]. invariant [Göb93, HJS16, Kog15, Liu19]. Invariants [Gor99, SV10, Bar99, HK05, Kog15, Lab15a, Lab15b, Sar15, SW06, She10, Thi00a, Thi00b]. Inverse [NMS97, Ben08a, Cla89, Grö10, Pop00]. inverses [IMSS08]. Inversion [SY14]. invertible [Jef10]. Inverting [MV11]. Investigating [Too15, WB95]. Investigation [Pop97, CHH^{+}99, MBKP81, Win84]. Invited [Ano00a, Ano00c, Ano00d, HKS99, Ano10a, Wan78d]. Involutive [GS95, GB97, Ger98]. involving [Cam76, Cam88, Fri74]. Irreducible [BKM07, GZ05a, Gae99, GZ08]. ISAAC [Gut04]. ISBN [DP15, GVR04, Qui97]. ISBN-10 [DP15]. ISBN-13 [DP15]. Isn’t [CJ92]. Isogeny [KD18]. Isogeny-based [KD18]. Isolated [SS11]. Isolation [DK98, Akr78, CMX10, Rum77, Tsi16]. ISOLDE [BPS12]. isomorphism [Teo15]. isomorphisms [BDD^{+}18]. Isotopy [BCD^{+}23]. ISSAC [Gia89, Gon89, LR98, Wan92, ACM94, Ano95d, Ano98, Ano99c, Ano99d, Ano13a, Cha96a, Cha96b, Coo96a, Doo99, Enc97a, Enc97b, Gia88, GKL07, Glo98a, Jef98b, Kal96, Kau05, KRR12, Küc97, Lak96, Lev95, Tec98, Mor02, Mou01, Lak98, ISS89, ISS94, Sen03, Sto98, Tra00, Tro97, Wan06a, Wan06b, WN90, Wat91, Wei97, YFJS19, Zim98a, Zim98b, Dal01]. ISSAC’2001 [Rei00]. ISSAC’93 [Bro93]. ISSAC’98 [Tsa98]. ISSAC’99 [Zim99, Zim00]. Issue [Ano10g, Ano10f, Dew00b, KW97, Mon96, SB97, SW96, SW97]. Issues [Kat87, KS94, Lic11, Mio84]. Italy [Gia88, Gia89, Mor89, Tra06, Ano91b, GMM87]. Iterated [McC97, Koe94]. Iteration [Cor96c, SS13, WRW17]. Iterative [AHCDT15, Izu14, Sak97, Mau16, Pop00, Ter10]. itsol [Izu14].

J [Ano99a, Ano99i, Dec14]. Jacobi [CL82, SS93]. Jacobian [Pau15]. Jacobians [BSS15, Doe08, Sha15b]. Jacques [Sei20]. James [Fit12]. January [Hea09]. Japan [HI93, WN90]. Japanese [Kak98, Nod96b]. JAVA [CCR00, Ber99c]. Jenks [Ano17a, JCMG11b, Ano13e, CGG^{+}04, CAB22, Kal13, Kal15, Mon20]. Jerrard [AJ03]. Joachim [Cor01b]. Johnson [Buc03, ASP15]. Johnson-type [ASP15]. Joint [Ano99a, Ano99i]. JOME [Dir00]. Jonathan [Dec14, Ano16]. Jordan [Bre05, FG99, Wib15]. Journal [KW97, SB97, SW97]. Joux [AMORH15].

JSC [SW96]. Jügen [Cor01b]. Julia [KD18]. Julia/Nemo [KD18]. July [ACM94, Ano98, Bro93, Cha86, Doo99, Gia88, Gia89, Gon89, Gut04, Jef98b, Kau05, Küc97, Lak96, Lev95, Mor89, Mor02, Mou01, Rei02, ISS89, ISS94, Tra06, Wat91, LW16]. June [Ano99e, Ano99f, Ano99g, Kot02, Mad92, Ng79, Pad02, RL98, Ano10g].

K3 [Elk08, Kel15]. Kahan [Lau77]. Kalkbrener [Kal93a]. Kaltofen [sL17]. Kanellakis [Gie08b]. Kantorovich [LS16]. Karlsruhe [CL96, Kot08]. Kay [Fri70]. Keiper [MSM23]. Kenzo [CRDMBR19, RRSS20]. key [BS15, Dur15, GMRT^{+}15, Iwa08, Küc87]. Khachian [AW80, Raj80]. Khachian-type [AW80, Raj80]. Kiang [McC73, Sta74]. Kiev [Bro93]. kind [Loo72b]. Kingdom [Ano99e, Ano99g, ISS94, ACM94]. Kirillov [ZZ15, Zha13]. Kloosterman [Bro09]. knot [HMS10]. know [Sch75]. knowledge [CL87, Her11, Lee10]. knowledge-based [CL87]. Knuth [WB83]. Knuth-Bendix [WB83]. Kolchin [Ovc14, Sit14]. Koszul [Rou09]. Kovacic [UW95]. Krandick [Joh12e]. Kreuzer [GVR04]. Kutta [Cam73]. KWIC [Ano82b, Fel77]. Kyoto [Kak98, Nod96b].

L [Dal01]. label [APS15]. laboratory [Abb04, CW72, Eng87, Gra87, Mae87b]. lacunary [Gre15]. Lacunaryx [Gre15]. Lagrange [CLZ08, FGVC15]. Lambert [Din09, MS16a, MSM23, SFG13, SFGZ14]. Language [Fat90, Gle92, Kry84, Pra90, AD10, Dav82c, DS10a, DS10b, Ebn86, HS74, Jen74b, Jen79a, Kor76, LTdD^{+}10, Nei71]. languages [Can69, HKN10, HI93, Hue84]. Laplace [Cla89, Nor80]. Large [Pfe21, Ban08, BJS06, EB70, Lew04, Wol05, Wol08a, Zho07]. large-dimensional [Ban08]. larger [Wan76a]. laser [KSS15]. last [SK00]. LATIN [Ano01b, Ano01b]. lattice [APS15, Poh81, Rou09]. Lattices [DRR18, Len81, ASP15, EVM21]. Laurent [Koe93, LYG12, TMCP22, Wat08, WSW13]. Layer [CLZ23]. Lazard [NDSM19, dRSE23]. Lazutkin [BBS^{+}02]. lazy [MBS15]. LCM [CLZ08, Rou09]. leading [HJS16, Iwa05]. leaking [CJ18]. Learn [Buc90]. Learning [Ano93, Emi20, LR22]. least [EGK09]. Lecture [Sch02, Mon21]. lectures [Wan78d]. left [BCLS11]. leftmost [Sti83]. Legacy [Gao17]. Legendre [Fee99]. Leipzig [Las84]. LEMA [LTdD^{+}10]. length [FN75, NO91, Poh81]. Leningrad [RRRA03]. LeSchack [LeS70]. LET [Sht77b]. Letter [Coo93, Ebe06, Gaw79]. Letters [Bee98]. level [Dav82c]. Lexicographic [Ree91, Ree92, Dah12, MW18]. Li [MSM23]. Liapunov [Wan89]. libraries [AKM16]. Library [BKR^{+}91, FT13, MPW97, CML^{+}08, CDL^{+}11, EVM21, FT15, HLM^{+}17, HMS10, JR91, JM17, Joh13a, LMX05, LMRS08, Maz15, MJ13, Mon21, PM07, Ric09, Qui12]. LiDIA [MPW97]. Lie [Poi15, GJ74, Mig77, Bar73, Gal15, Shi99]. Lieb [KiL16]. Life [Gao17, Bar02]. lifting [CL11, PM19, Sal11, Ste10]. like [Alm11, BB96, Fat87, LS16, SS13]. likelihood [LPA15]. Lille [Mor02]. Limit [Har79, AM16, Kem81]. limited [FT12]. Limits [VHC17, AKM16, Kau03]. line [Sed08]. Linear [AB14, AP23, Bak93, BK14, Ger98, Ghe10, Ghe14, GLS97, GS97, Lob97, MMSY15, NTW97, Rue10, SIK14, AKR20, AW80, Arr15, BK04, BPS12, Bar15, Ber06, BJS06, BCLS11, Bou15, Bul10, Cas08, CvHL10a, CvHL10b, CP12, CL09, Cla69, DMSP10, DMCMM15, Ebe16, FvH10, Far15, GMMCMMP22, GTLN18, Goy21, GK88, Grö09, Grö10, JdM12, LWZ07, Mad15b, MS16b, MS18, Pet08, Raj80, SE22, Sau10, Sch94, SW06, She07, Ste10, SLW15, Suz09, Tsa98, Vil98, Wei87, Wei00, WSW13]. linear-exponential [Wei00]. Linearization [SC17]. linearizations [FPS16]. linearly [HMSP18]. linkages [Kou16]. linked [Moe78]. Links [BKR^{+}91]. Linz [BW98b, Hon94, Buc84, Vol96, Vol97b]. Liouville [Kas80]. Liouvillian [LMF12, Raa11]. Lisp [Fat14, Cam75, Fat78, FF82, Gre72, Hea69, Jen70, KS81, Mos70a, Nam86, Whi78, Fat90, FL91, FH96, Fat98, Fat03b, Kal82a, Lan86, MHGG80, MH85, Wil92, Yan91]. LISP-based [KS81, FH96]. Lisp-Language [Fat90]. Lisp/360 [Kal82a]. LISP/370 [Whi78]. list [Akr80a, CL11, Fat03b, LaP72, Moe78, NFIS69, Qui12, SS68, SVZ06]. live [HK08]. LLL [MSV08, NS18a]. LLL-reduced [MSV08]. load [MC95]. Local [DM15, SIK14, DS15, Sht77a]. localize [BLH10]. LocalizeRingForHomalg [BLH10]. locus [ELME22]. logarithm [Hol04, HL08, Yas08]. logarithmic [CP78, DGL19, Fat98, Jua15, Sán15]. logarithms [AMORH15]. Logic [Ano99k, Bak93, Bib84, DS97, Kan76]. London [Ano00a, Ano00b, Ano00c, Ano00d, Mou01, Rei02]. long [Ste81]. look [HL08]. Loop [RS91]. Lorenzen [Pra90]. Lorenzo [GVR04]. low [SS13]. LSSS [WSW13]. LU [Jef10, PW03]. lubrication [CJ88]. Ludeman [Ber70]. Lüroth [Jer15]. LWE [ACF^{+}15]. Lyon [DMP^{+}02]. Lyons [Sao10].

M [Ano16, RRRA03, Akr87, MF83]. M. [Dav03]. M4RI [SLW15]. MA [HI93]. Macaulay [DD99, Grö10, Lew08, Min04, Spe08]. Macaulay2 [HJN^{+}23, Lee19, Ley16]. Machine [Fat78, RSSRV89, SS83]. MacMahon [AP08]. Macpherson [Hel15]. MACSYMA [AW80, Cla89, Fat72a, FMW74, Fat87, Fit73, Loo77a, Mos74, Raj80, SB89, Wan75, WS83, Zip75, Gol85, Tec98, RS89].

MACSYMA-like [Fat87]. MAD [Meu03]. made [Sau81b]. Madrid [RL98]. Magma [Kop20, Sut17]. Mai [Ano99b]. main [Sar15]. man [SS83]. management [Arn10, Dav03, Her11]. Mandelbrot [CC16]. manifesto [Sto13a]. manipulating [Wan91]. Manipulation [Ng79, Ste82, dQ77b, BS74, BF74, Che83, Don83, Fla88, Hir78, Jen79b, Nei71, Oll84, Pon87, Pon88b, Pon88a, Sch71, SW85, Smi84a, Tut73, Wan80c, Wan80b, dQ77a, vdR73, Wan78c]. manipulations [Mio81, RS89]. manipulative [LW08]. Manual [SvH81]. Manuel [Bro13]. Many [Abr98, Her98, Elk08, Fri74, Sed13]. Maple [Cha90, Cha91, Cor96a, Mad92, KRTZ20, LMX05, LMRS08, AZ91, Ano95e, BPS12, CGG83, CML^{+}08, CDL^{+}11, CM99, ECTB^{+}14, GK20, Gon88, Hel10, IOP21, Izu14, JT91, JM17, JM98, KRR10, KRR12, ME18, Min04, MP10, MP12, MP14, MP16, PM07, Pos13, SE22, Sal93, YA07, Zei91]. MapleTech [Mon96]. mapping [JT91]. maps [DS15, HL08, MMSY15, Wei15]. Marc [DP16]. March [Ano99h, Ano10f]. Marichev [Gak79]. Mark [Ano13b]. Marko [Pet23]. Marseille [Ng79]. Martin [GVR04]. MAS [KP97]. massively [Ber99c]. Matching [KN87, Fat15, McI85, RSSU79]. MATCONT [DGK04]. Math [Pad02, Lic84, Sto79]. Mathemagix [Lar19, vdHLM^{+}11]. Mathematica [Ada96, Fat90, KA99, Kou13, Nag07, Nai22, PDP00, Tro97]. Mathematica-to-Lisp [Fat90]. Mathematical [Ano84, Ano95e, BG97, Bor98, Gon88, Ham14, Her11, MP14, Sit06, Vol98, Wat98b, ADS96, CC72, Eng87, GGdR^{+}15, Gri76, Gro99, Hor15, How70, Joy09a, LN76, Mae87b, Meu03, Mio84, Pea84, SPA^{+}99, SCWL08, Sun74, Wan80c, Wan80b, Zas84, Kot08]. mathematician [RRRA03]. Mathematics [Ano99b, Ano99f, Ano03, Ano08b, BMK98, Buc02, Dir00, FH96, Gie08b, Soi97, Ste82, Wri00, tW99, AsW08, Arn87, BKKS05, BGJ70, FC99, FB70, Fra80, HK08, HLS84, Hue84, Ril08, Sho82, Ste81, Var15, Wan76b, Zel08, Zho07]. MathLibre [Ham14]. MathML [Pad02, Alv00, Car00, Pad02, Soi97, WL99, Wri00]. Matlab [DGK04, Zen04, ZL13]. Matrices [CMT14, Sen90, AB19, AM10a, BK12, BA21, CC16, DJ96, Fri74, Fuc08, Gen99, HKS19, IKRT89, JMV09, Jef10, LMTV05, LM16, MV11, Mid11, Rue15, SC17, Uga12, WL08, WRW17, WW11, Wis84]. Matrix [BGLM23, BD14, Con90, CJ97, EG96, Ste96, SY14, Ba12, CHS16, CL09, CT19, Cor96d, DIS10, Ebe16, HHM10, LM12, Lew04, MK96, Mor99, Pan05, SH09, YS10, YK08, dO99, Cap86b, De 99]. matrix-vector [HHM10]. Matu [SK16]. Matu-tarow [SK16]. Maui [Küc97]. maxima [Joy06]. Maximal [Lan93, SS12a, BLH10, JR91]. maximization [LPA15]. maximum [WRW17, Xue15]. maximum/minimum [Xue15]. May [Ano00a, Ano00b, Ano00c, Ano00d, DMP^{+}02, Sch02]. McCrimmon [BH97]. McEliece [MG15]. MCR [MZ97]. MCR-Based [MZ97]. Means [SVV95, BCCR80, Jen79b, vdR73]. measure [MS08]. measurements [BS74]. mechanical [Ban08, Sau81b]. Mechanically [Bar91]. Mechanics [Ano08b, BC81, GL10]. Mechanization [Aca20]. Meeting [Coo96a, Coo96b, Mon01, Lak98, Rob01, Abd84, Ano99a, Ano99i, Ber99b, Coo97, Gim03, Nod96b, Rei02, Wan78d, Kak98, Sit06]. meets [DS97]. MEGA [Ano99e, Ano99g, AD10]. melding [Tif15]. Mellin [BFS23]. member [Ano17c]. Memorial [Ano17a, JCMG11b, Mon20]. Memoriam [Ano08b, CGG^{+}04, Joh12e, Ano10e, Ida05]. Memories [SK16]. Memory [AB21, Fat03b, Web90, BBS^{+}02, CM02]. merely [Ste81]. Mesh [Cor00]. Message [Ano14b, Ano14c, Bro17, Cet96, Cor96f, Cor96b, Cor96e, Cor97a, Cor97b, Cor98b, Cor98c, Cor98a, Cor99b, Cor01a, Cor02, Dew00b, Gie99a, Gie99b, Gie99c, Gro95b, Jef03, Kal93b, Kal95, Kot14a, Kot14b, Wat96, Wat97b, Cha97]. META [Jen70]. META/LISP [Jen70]. Method [BCL23, HM77, Lev15b, OS23, SN97, Akr78, Cam73, CX08, Fit74b, GH08, JT91, KS16, Koh10, KM05, LS16, Liu19, MS18, Mio74, Mur09, NS18a, Pei87a, Ren17, Rim84, SS09, Ter10, Tri84b, Tri11, Zim76, dO99, dRSE23, vdHM20, vTG03]. methodology [Var15]. Methods [AvH19, Ano99e, Ano99g, Ber98b, Kot08, NMS97, Bal84, BFR^{+}18, Dem12, FCB^{+}22, FSH02, KS03, KS19, Lic11, LMF12, Man11, McN04, She07, UM12, Xie07, vdR73]. metric [Jen15]. metrics [NP88, Sch88]. MICA [sL17, Gie08a]. Michael [Dal01]. micro [di 82]. micro-computers [di 82]. middle [AS14]. Mignotte [Akr87]. Milestones [sL17]. million [BvzGG^{+}01]. MIMD [RSSRV89]. minicomputers [CM76]. minima [Poh81]. Minimal [GB97, Jai08, Le01, Lub11, Sha15a, Bla14, DHS04, HJS13, Mur09, NFIS69, Poh81, YK08]. MiniMaple [KS13]. Minimax [Cor00]. Minimum [GTLN18, Xue15]. minisymposium [CW98]. Minsk [Gak79]. Minutes [Coo96a, Coo96b]. MIRA [Gaw77]. misdemeanors [Sto12b]. mixture [LPA15]. MKM [Far04]. MMICS [Kot08]. mmWave [BSB^{+}23]. Mobile [BKKS05]. Modal [BMK98]. Mode [PH81, AE83, CH85, Jen77, PH83]. mode-based [Jen77]. Model [DN90, Coo04]. modeler [HYH15]. Modelling [Ada96, LR22]. Models [Bur08, BCCR80, Cas08, IOP21, JPP22, Sha15a, Zas84]. Modern [Cor01b]. modes [KSS15]. Modified [Fuc08, AW80]. Modifying [EGP11]. MODLISP [DJ81]. modpn [LMRS08]. Modula [KP97].

Modula-2 [KP97]. Modular [BGLM23, BD14, CLZ23, Hea72b, SHS96, CL09, Ebe83, Gol91, HLM^{+}17, MJF^{+}10, Pre15a, Xie07]. modularity [vdHLM^{+}11]. module [RRSS20, Teo15]. Modules [ACLM23, CM99, Sor97, CML^{+}08, Lev15a, Mau16, Rue05, Zha13]. Moduli [HJN^{+}23, Qi20]. Modulo [Vol95, LOW08, MV11]. molecular [Bar89]. moment [WRW17]. monic [CM22]. monodromy [LBL20]. Monomial [Sch91b, BMT99, BGM99, BP20, D’A15, Fas06, GZ05a, GZ08, LLG08, Rou09]. monomialideal.lib [BGLG12]. monomials [KiL16, iL15, Mid13, Qi20]. monoterm [Liu19]. monotone [LLT18]. monotonicity [Joy15]. Monsters [EDS00]. Monthly [Gon88]. Montréal [Lev95]. Moore [RRSS20]. mordor [CJ18]. Moscow [Ano08b, AS16, ARS^{+}19, ARS^{+}21, ABR^{+}23]. Motion [Kou16]. Movers [Dav86]. moving [BGJ11, Ste81]. MP [BG97]. MPCR [Bac97]. MR [Min04]. MSSM [HJN^{+}23]. Multi [KSS15, Sch08a, DMH15, GHK18, MX09, MW17, POdCW19]. multi-cores [MX09]. multi-dimensional [GHK18]. multi-indices [DMH15]. Multi-parameter [KSS15]. multi-point [MW17]. multi-state [POdCW19]. Multi-summation [Sch08a]. multicore [CMX10, Via13]. multidimensional [Bos84]. multinomial [Qi22]. multinomials [Sto09]. Multiple [Asl96, Com18, DGL19, GRT10, GMC^{+}21, Mat21, OVB15, Ter10, Ter13]. Multiple-Valued [Asl96]. multiples [BCLS11]. Multiplication [BGLM23, BD14, CMPS10, CM02, DIS10, Fat03a, Gen71, HHM10, Kli78, MX09, MJ13, MP10]. multiplications [HKS19]. multiplicative [HMMR22, Rag15]. multiplicities [Zen04]. multiplying [HKS19]. Multiprocessors [Web90]. MultiRegeneration [CRWZ20]. multivalued [JN04]. Multivariable [MMNR15, jYO20]. Multivariate [KM19, KS92, SS98, Sto08, BSBC15, CK04, CsL08, CsL09, DD99, HM13, Ina05, Iwa05, Kal82b, KS19, Lev15b, MT15, NS17, NS18b, Pet08, Rai11, SI08, TMCP22, WR73, Wan75, Wan76a, Wat08, YDH^{+}15]. MuMath [AF84]. muMATH-80 [Sho82]. Mumford [BGM99]. MuPAD [HKM^{+}97, Sor97, Thi08]. muSIMP [Sho82]. muSIMP/muMATH [Sho82]. muSIMP/muMATH-80 [Sho82]. My [Fat84]. MZVs [CDM15].

N. [RRRA03]. NAClab [ZL13]. NAG [BKR^{+}91]. name [Ste06]. NaNs [Fat99]. NAOMI [Bra98]. national [Abd84, Las84, Wan78d]. natural [Jeb18]. natural-style [Jeb18]. Nature [Ada96]. Nauka [Gak79]. near [Mig77]. Nearest [Sek23, Ste99, Gra05, RC05]. nearring [AY16]. Need [Sto07]. needn’t [CDJW00]. negative [DMH15]. negativity [ME18]. Neighbors [EZ23]. Nemo [DRR18, KD18]. Nemo/Flint [DRR18]. Nested [BFS23, Raa15, Sht77a]. NETFORM [Smi74, SvH81]. Network [Man15a, DSJL08]. networks [Bin87b, BFR^{+}18, EV05, Wan91]. neural [EV05, Wan91]. newsletter [Loo77d]. Newton [Jal99, LS16, NOT18, Pér99, Ren17, SN97, SS13, Tri84b]. Newton-like [SS13]. NewtonSLRA [Nag21]. Nielsen [Mac89]. Nielson [Dal01]. Nikos [Ano14a]. NIL [Pur80]. nilpotency [dCS99]. Nishidate [Ada96]. No [Lau77, Ghe10, RS75]. Noda [SK16]. Noether [BCL23, BGLGMM18, Jen15]. Nominating [Wat98a]. nominations [Ano13e]. Non [Hea72b, Lev10, Mau16, Sat97, BS15, Bal84, CM15, CM22, CMPS10, CPS11, Fat72b, Fri74, Gae99, GR08, Hei15, Jef10, JdM12, KKM15, Lab87, ME18, Moe99, Raa11, Rod84, SE22, Sht77a, Sto17, Stu10, WSW13]. non-abelian [WSW13]. non-algebraic [KKM15]. non-associative [BS15]. Non-commutative [Lev10, Hei15, Rod84, Stu10]. non-commuting [Lab87]. non-floats [Sto17]. Non-free [Mau16]. non-hyperelliptic [GR08]. non-invertible [Jef10]. non-linear [JdM12, SE22]. non-Liouvillian [Raa11]. Non-Modular [Hea72b]. non-monic [CM22]. non-negativity [ME18]. non-numeric [CPS11]. Non-numerical [Sat97, Bal84, Fri74]. non-rational [Fat72b]. non-standard [CM15]. non-zeros [Moe99]. Nonassociative [CP88]. noncommutative [HRR19, LT18, LS03, Lev15a]. nondegenerate [ELME22]. nondimensional [Hir78]. Nonexistence [Ghe14, Li92]. Nonlinear [Ber98b, Oku97, Qui97, JPP22, KKM18, Nat92, Rag15, Tsa98]. nonnumerical [FB70]. Normal [tW99, DI91, ES11, LM12, Mad15a, Mid11, TMV18]. normalization [Sto09]. North [Bra98]. Note [Abr00, GK84, Nor80, UW95, Akr78, Akr87, CE77, CP78, Fat98, Gra05, IK20, Joh75, MM89, MG89, NP88, Sch88, Sti83, Tob70]. notes [Fit75]. notion [BSBC15]. Novel [BSB^{+}23, Cha81]. novelties [Bah75b]. November [Ano91a, Lin98a, Lin98b]. NSF [KO13]. NTOW [BS15]. NUCLEOL [NFIS69]. nucleotide [MSP05]. Nullspace [ES10b]. Nullstellensätze [KPS99]. Number [Abr00, EZ23, MPW97, Mon93b, OS23, Ano10g, Ano10f, CJ96, FCW13, Gle86, KS97, KV04, Mat10, Mur09, SS12a, Ste81, Tri84a, Wan75, Zim84b, Zim84a]. Numbers [Mur14, SS98, Sto07, Har00, Hir78, Hol04, MSS18, Nor79, Ren82, WGD82, Wee74, Zho82, Zim84a]. Numeric [Ano08a, BKR^{+}91, LMF12, Rei02, SV93, SW97, CPS11, CsL09, Goy21, Hod11, IMC^{+}22, NS17, RSV02, SVZ06, vH84, vdHLM^{+}11]. Numeric-symbolic [LMF12, RSV02]. Numerical [Ano99a, Ano99i, Ano00c, Ano00d, Bal84, BK14, Din09, EDS00, Ged99, MZ97, Mon93b, DGS16, DGK04, Fri74, Ili08, KS03, Ley10, MM89, Sat97, ZL13, vdHL20]. Numerically [Bat13, Dec14]. numericalsgps [DGS16]. Numerics [SFGZ14, GMC^{+}21]. Nyström [Cam73].

O [Gak79]. Objects [MP14, Bet13, Nab07, Nie84]. oblivious [Sal12, YS10]. obtain [SS09]. octahedra [Cou15]. October [Ano99j, Ano03, HI93, Vol98]. octonions [BS15]. odd [DM21, Maz85]. Odds [BY02]. ODE [BP20, CTKR99, Cor93, Ged99, HC97, IOP21, Sch94, Wil84]. ODE/DAE [HC97]. ODEs [DGK04, FG03, HOPY19, Ngô11, WSW13]. off [Dav82a]. Offsets [CFGVDT18, SSS05]. OGB [McG04]. OMDoc [Koh00]. OMEGA [Xam99]. One [Sun73, BM15, BFE20, BvzGG^{+}01, Coo97, CJ18, CsL18, vdHM20, Bra98]. one-algorithm [Sun73]. one-block [BFE20]. one-time [BM15]. Online [Bra98, McG04, SY14, SCWL08]. only [AsW08, Alm11, AP08, AM08, Bar10a, Ben08b, BM08, Bul10, Bur08, CvHL10a, CX08, DSJL08, Doe08, Ebe08, Edo08, Elk08, FvH10, Fon08, GR08, GZ08, GRT10, GMGE08, HL08, Ibá08, Ili08, Iwa08, Jai08, Jon08, JM08, Kel08, LW08, LM08, Lew08, Ley10, LLG08, Man10, Mat10, MGR^{+}08, MS08, MJF^{+}10, Nov08, Pet08, Pri08, Pud08, QH08, Rad10, Ril08, RG08, Sau10, Sch08a, Sed08, SCWL08, Tha08, Thi08, Tsi08, VJ08, WL08, Wat08, Wol08b, Yas08, YS10, YK08, ZT08, Zel08, Zho08, vH10]. Ontario [Ano99a, Ano99i, Ano00a, Ano00b, Ano00c, Ano00d, Mou01, Rei02, Cha86, Wat98b]. Open [Cor99a, Dal02, DMP^{+}02, Joy08, Joy09b, JČMG11a, Gaw79, Joy09a, Lic11, Mio82, Pag07]. OpenAxiom [LD11]. OpenMath [Bra98, ADS96, Alv00, Car00, Dav00b, Dav00a, Dew00a, Dew00b, Koh00, PDP00, SK00, Tri00]. Operations [Tab23, DMSP10, MBS15]. operator [HRR19, Rag15]. Operators [AB14, AP23, JPRZ23, AB19, Ber06, BCLS11, BRRT08, CH85, GHL15, Goy21, GSGZ15, KRR10, LLT18, Pag22, SW06, She07, Tsa98, ZvH19]. opportunities [KO13]. optics [KSS15]. Optimal [BP20, Gen71, Pop00, SV99]. Optimisation [PH81]. optimising [Emi20]. Optimization [Bin87b, DIS10, Met23, HHKR20, HHS12, HP67, TJ03]. Optimizer [SvH81]. optimizing [CK04]. Optimum [OVB15]. Option [Kal82a]. orbit [HMMR22]. Orbiter [Bet13]. orchestration [The10b]. Order [Ree91, Ree92, vHI14, Arr15, Bar15, Ben08a, Cla69, DS15, DSS15, FG03, GL13, MGR^{+}08, MS16b, MS18, MBKP81, Ngô11, Per15, Sán17, Sch16, Sha05, SW06, Zho08, Ree92]. ordered [AV99]. Ordering [HEW^{+}14, FGHR12, Sit89]. Orderings [Sch91b, Kre88]. orders [Wei87]. Ordinary [Ber98b, JPRZ23, PS98, AKR20, BCLS11, Cha81, KRR12]. ore [CL09, Jar12, CG19, Glo98b, Jar15, KM19, LM12, Mid11]. Oregon [ISS89, Gon89]. OREM [BRRT08]. oriented [LW99, Lew00]. Origami [Ril08]. orphan [vH74]. orthogonal [Gra91, KSZ11, Pre15a]. ORTOCARTAN [Kra83]. OSCAS [Joy06]. oscillations [Raj80]. other [SS07]. Outgoing [Cor99b]. outline [Jen75]. output [Fat87, Ste10]. outputs [Sht77b]. Over-Determined [Cal01]. Overcoming [CM02]. overdetermined [Wol05]. overlay [Gre72]. Overview [Dew00a, Sar15]. Oxford [ACM94, ISS94].

P [Ano14a, tW99]. P2P [GMGE08]. PA [Sen03]. Package [BFS23, FL91, HJN^{+}23, Izu14, Sch83, vH82, BPS12, Cap86a, Cap86b, DGS16, DGK04, FIS16, GJM09, Goy21, HSW89, JPBCL20, JM98, Joy05b, KRTZ20, Koe95, Kop20, KRR10, LW95, ME18, Min04, Nab07, Nag07, Nor75b, Pos13, San91, SW85, Smi84a, Sto80, Tut73, Wol05, YFJS19, Zen04]. packages [NW83]. PACLIB [HS92]. pads [BM15]. paging [SS68]. Painlevé [SS17]. pair/completion [Win84]. pairing [BP18]. pairing-friendly [BP18]. pairs [BD15, MCP15]. Pan [EG96]. panel [Col69]. Pankratiev [Ano08b]. paper [Abr09, Akr80a, Akr87, Mac89]. Papers [Ano82a, Ano84, Ano99c, Ano99e, Ano99d, Ano99b, Ano99h, Ano99g, Ano99k, Ano01b, Cha96a, Jef98b, Abd84, Abr03, Ano10d, HKS99]. paradigm [BKG18, CsL18]. Parallel [DK98, Hon94, Hon96c, Pon88a, SHS96, Sal11, Sat00, pro12, Coo04, Dav82b, HM13, Mat10, MJ13, MW17, The10b, Sed13, Wis84, Xie07, HI93, Vol98]. Parallelism [HKM^{+}97, Pon88b]. parallelization [LD11]. Parameter [GS97, SN97, Gle86, Glo98b, KSS15]. parameterization [Sch99]. Parameterized [Mai15, Arr15, Sar15]. Parameters [NO23, IMSS08, Sar15]. Parametric [BK14, CMT14, GS97, SIK14, BK04, Bul10, CM83, CJS18, FGVC15, GJM09, HOPY19, KRTZ20, Mon11, Nab07, Ore99, Man10]. ParametricSystemTools [CML^{+}08]. parametrizations [D’A15]. parametrize [RSS15]. Parametrized [BKR23, Sán15]. paraxial [KSS15]. parser [Fri70]. Parsing [FC99]. Part [BU93a, BU93b, BH15, Jua15, Lab15c, LW94a, LW94b]. partial [Bar89, Gol06, SW06, She07, Sto08, Tsa98, vdR73]. Participation [Ano01c, Ano03, Ano99a, Ano99i]. Partition [BZ12, AP08, Lac92]. Partitioning [Fat15]. partitions [BM08, Cam76]. Pascal [di 82]. PASCO [Hon94, Hon96b, Hon96c]. Passau [Vol97c]. Past [Wat97b]. patching [Mai15]. paths [MCJ15]. Pattern [McI85, CHS16, Pud08]. pattern-avoiding [Pud08]. Pavel [DP15].

PDEs [Ger98, Tsa98]. Pech [DP15]. Pedersen [Mac89]. pendulum [Tot15]. pentagon [Mor10]. perceived [Ste81]. perfect [Ara15, RG08]. performance [CMPS10, GMC^{+}21, Li09, Mac89, PM07]. permutation [But85, CM02, GTLN18, Joy05a]. permutations [JR91]. persistent [GHK18]. Personalizable [Ham14]. perspective [SSS05]. perturbation [Coh74, GL10, SS10, SS11, SS12a]. perturbations [Mae87a]. perturbed [Mad15b]. PetaFLOPS [Ree97]. Petersburg [Ano99f]. Petkovsek [CHL14, Pet23]. Pfaffian [Mad15a]. Pfister [Hei13]. PGB [Nab07]. Ph.D. [CC72, Win84]. PHCpack [Ver10]. PhD [Emi20]. Philadelphia [Sen03]. Philippe [SSS^{+}11]. Phrasebooks [CCR00]. physics [Fel75a, FB70, SPA^{+}99]. Piano [Dav86]. Picard [Hei15, Kel15]. PICOMATH [Sto80]. PICOMATH-80 [Sto80]. Piecewise [BH15]. Pioneer [CGG^{+}04]. Pitsianis [Ano14a]. PL [Bah71, Ber70, EB70, LaP72, Lud69]. PL/I [Bah71, Ber70, EB70, LaP72, Lud69]. Plaintext [BM15]. planar [JSS03, Kou16, NOT18]. Plane [HV98, AHCDT15, HMS10, Hod11, JPBCL20, Tra15]. planning [Cas08]. platform [vdHGG^{+}13]. Platforms [DN90, GZ05b]. PLDE [Par15]. PLMMS [Ano10d]. PLURAL [LS03]. Poincaré [BTL99]. Point [SI00, Ben08b, ES19, KS97, MBS15, MW17, Poi15, SS98, SK10]. Points [HV98, AM16, Edo08, Elk08, EGK09, Fas06, FT12, HHS20, Hol04, JM18, JS06, Kem81, Sho15, Zim97]. Poisson [SASG99]. Polar [Bar99]. poles [Nor82]. Polman [Tut73]. POLY [MP12, MP14]. polygon [Ren17]. Polyhedra [Ost99, DDK^{+}11, Jal99, JM17]. polyhedral [BZ12]. polyhedron [JM18]. polylogarithm [Mat21]. polylogarithms [DMH15]. polymake [KLPW17, JP11].

Polynomial [Abr98, AJ03, BNN14, BGLM23, BSB^{+}23, Buc91, CCM^{+}14, Dec14, EG96, FG03, Hea72b, Her98, KW97, Kal82b, Ley16, NTW97, Pan05, RT88, SS97, Sch02, Sek23, Ste96, Ste99, Uga12, Ver10, VY15, Wan98, vdHL14, Akr78, Bar10a, Bat13, BCG09, Bla14, BvzGG^{+}01, Buc76a, BRRT08, Bul10, CP88, CDL^{+}11, CCM^{+}16, CLZ08, CL09, CMPS10, CK03, CRWZ20, CsL08, Dem12, DPFD15, ELME22, Fat03a, FCB^{+}22, Gen99, Gen71, GJM09, Gle86, Glo98b, Gra05, GPS08, HLM^{+}17, HJS13, Hei70, HL10, IK20, Jar12, Joh74, Joh75, Kli74, Kli78, KB78, Kot01, KLV03, Lab87, Lee19, LS03, LW99, Lew00, Lew04, Ley10, LLG08, LMRS08, Li09, LW94a, LW94b, Liu15, Man10, MMNR15, MGR^{+}08, MX09, MP11, MJ13, MY74, MP10, MP12, Nag08, NOT18, NY03, PM07, PW03]. polynomial [RRRA03, RC05, Rim84, RG08, RZ20, Row81, ST02, Sas04, SS10, SS11, SS12a, SV99, Smi84b, Spe08, Ste08, Tec11, TMV18, Tut73, VHC17, WL08, Wol08a, Xie07, Yun73, Zen04, Zen08, vdHL18, vzG92, Bul10]. polynomial-oriented [LW99, Lew00]. Polynomials [Bak93, FL91, HNT14, Kno97, Li92, OS23, SW97, Tre90, AM08, AM10b, BK12, BA21, BKM07, Buc76b, BRRT08, CM15, CEGW08, CC16, CM22, CL15, CM83, DDK^{+}11, EGK09, GR08, Glo98b, Gre15, GR16, HL10, HM13, Ina05, Jar12, Jar15, KM19, KS19, Kou16, LM12, LM16, Len81, Len84, Lev15a, Lev15b, LYG12, Liu19, Mai99, Man77, Man15a, Mid11, Mig80a, Mig80b, Mig87, Moe99, MW17, MG89, Nag04, NS09, Nov08, NvH08, NS17, NS18b, Pfe21, Pre15a, Rum77, SO09, SH09, Sut17, Ter10, Ter13, Tro82, Ver75, WR73, Wan75, Wan76a, Wat08, Whe87, WCF12, Zim76, vH10]. polytopes [AV99, JP11]. Pommaret [IdC21]. population [Bur08, Ebe08]. PORT [Kli74]. PORT-ALG [Kli74]. portable [Kli74].

Portland [Gon89, ISS89]. Position

[Ber98a, Zas84]. Positions [Wat98b, Emi20]. positive [EGK09, Pag22]. positivity [PS15]. possibilities [GGdR^{+}15]. post [YDH^{+}15]. post-quantum [YDH^{+}15]. Poster [Enc97b, LR98, Tsa98, Zim99, Zim00, Ajw05, Cap03, DK98, GKL07, Hit07, KV04, Sit02b, Wan06a, Wan06b]. Posters [Ano00b, Enc97a, Zim98a, Zim98b, Ano10a]. Potential [Pop98]. Power [HM77, Koe93, KS92, AKM16, Ben08a, Gen71, Gen74, Gla91, LBL20, Mor87, Nor75b, Pei87b, Zwi84]. power-commutator [Gla91]. power-inverse [Ben08a]. Power-series [KS92]. powers [Gae99, RG08, Sto13b]. Practical [EB70, Sch71]. Practice [SST^{+}97, vdHL14, FS97]. pre [Hie75]. pre-editor [Hie75]. precision [FT12, Pur80, Web90]. preconditioner [Ebe16]. preconditioning [MG89]. predicate [Kan76]. prediction [EK12]. Preface [Gie08a]. prefix [KH81]. Preliminary [Ano98, Col74, RC76]. present [Bal84]. Presentation [Str74]. presentations [Gla91, JR91, YFJS19]. Presented [HKS99]. preserving [FPS16, MMSY15, SS12a]. President [Ano16]. Press [Buc03, Dav03, FBOS88]. Pretty [Hv83, vHvH87]. pretty-print [vHvH87]. previous [Mio88]. primal [WRW17]. primality [Nor79, Val20]. primary [MMR99]. Prime [Wan98, NY03]. primitive [RC76]. Principal [Con90, KN97, Gol91, HS74, MMNR15]. principle [DSM^{+}05]. principles [Kot01]. print [vHvH87]. Printing [Hv83]. Prize [JCMG11b, Kal13, Kal15, Ano13e, Mon20, CAB22]. Probabilistic [BCG09, HJS13, HJS16]. Problem [BH72a, Bar73, Cam72, Cam73, CHJ^{+}97, Cor00, Dav86, Fit74a, Har77, Hea72c, HC97, JG74, Kah75, Laz80, Loo75, McC73, Nor78, PST97, Rim84, SN97, WR79, van80, Bah75a, BHMM13, Ber70, BCCR80, CHH^{+}99, CMR74, CJ88, DX07, Fat73b, FMW74, Fat76, Fit74c, GCG83, GJ74, GJY75a, Hag77, HJ74, Hal75, Lau77, Loo77e, Mio82, MC95, Mos70a, Mur00, Nor75a, Par15, RS75, San71, Sch99, Sta74, SY22, Sun73, Teo15, Tut73, Yas08, Zip75]. problem-solving [Fat73b]. Problems [Buc91, FM97a, FM97b, Fit73, PS98, Sak97, SST^{+}97, Ste99, van80, ACF^{+}15, ÁS99, CC83, Cor99a, FB70, Gon88, Ili08, Jen79b, Kor77, KRR10, KRR12, LMTV05, Lew04, Mae87a, Mio74, Pea84, Rad10, RSSU79, SASG99, Shi99, Tec11, TJ03, WL08, Wei00, Xue15, Zho07, Tec98, Sto98]. procedure [PS15]. Procedures [Smi84b]. Proceedings [ACM94, Ano91a, Gia88, Glo98a, Kau05, Lak96, Lev95, Mor02, ISS89, ISS94, Sen03, Bro93, Doo99, Gut04, Jen76, Küc97, Mou01, Tra00, Wan79b, Wan81, WN90, Wat91, Cha86, Gon89, Jen74a, Mor89, Tra06, Wan92, CL96, Gia89, HI93]. process [AB19]. Processing [EJP^{+}01, Emi20, Fat03b, LaP72, Nie70]. processor [NFIS69, SS68]. processors [Pon88a, Wis84, SASG99]. Producing [van80, Sht77b]. product [AS14, Arn10]. Products [AB14, MP14, CP12, Fri74, JM98, OS16, Sto13b]. Professor [AB21, Ano99a, Ano99i, Ida05, SK16]. Program [BG90, Har79, RL98, RS91, WR79, AZ91, Bah71, Bro03, CC72, Kra83, LW16, Lab87, Sed08, Wan89, Zei91]. programmers [Sch75]. Programming [Gle92, HS92, AW80, AD10, BK04, Cam75, Can69, CW72, Fat81, HKN10, Kan76, Kri84, Lin07, Pfe21, Raj80, Ric09, Sal93, Zel08, pro12]. Programs [PH81, WB95, Fat98, Fat03a, Gre72, HJ74, KS13, KR85b, Loo72a, Lud69, PH83]. progress [PST97]. project [BFR^{+}18, BDE^{+}21, Buc84, CT16]. projecting [Ull06]. projection [UM12]. projects [Far15, Mae87b]. Proof [BCL23, Kal93a, Wan16, AD10, BB80, Gor98]. proofs [Fat85b, Jeb18, Nai22, Tha08, ZT08]. Properties [Pop97, Buc76a, Grö09]. property [Win84]. Proposal [BU93a, BU93b, HSW89, KS97, Soi97, ÜHK82, Sto79, Sto84]. proposed [Akr80a, Akr80b, Neu80, San71]. proposes [Sto17]. protocol [ADS96, The10a]. Prove [Xue15]. proves [AZ91]. Providing [O’B91]. Proving [Fèv98, MA08, Pec07, Pei87a, Win84, Zei91, DP15]. Provisional [FF82]. PRSs [Sas22]. Pseudo [SHS96, MG15]. Pseudo-Parallel [SHS96]. pseudo-random [MG15]. pseudovarieties [CKW03]. PSLQ [FCW13]. PTOPO [KRTZ20]. Public [Küc87, BS15, Dur15, Iwa08]. public-key [Iwa08]. Publication [Bor98]. publicly [Mon21]. Publishers [Qui97]. Puiseux [BG90, Pér99]. Puisieux [Koe93]. punctured [MG15]. pure [GR99, Ver74]. purpose [vH75]. pushing [BDE^{+}21]. Puzzles [FM97a, FM97b]. PVM [LW95].

QD [Sch83]. QEPCAD

[Arn10, Bro03, Bro04]. QRGCD [NM13]. Quadratic [Abb14, DM21, SV10, YDH^{+}15]. quadratization [BP20]. quadrature [Fee99]. quadrics [CFGVDT18]. quadtrees [Wis84]. qualitative [Lee10]. Quantifier [Col74, Arn10, BFE20, FIS16, HW75, ZKK11]. quantization [Por09]. Quantum [Dal01, Ele15, DGG15, Lau15, Roa08, YDH^{+}15, KS14]. quarter [JPBCL20]. Quasi [IdC21, AM16, CM15, CDM15]. quasi-componenets [AM16]. quasi-polynomials [CM15]. quasi-shuffle [CDM15]. Quasi-stable [IdC21]. quasilinear [Sch16]. quaternionic [Liu15]. queries [IP11]. query [DX07]. quest [vdHLM^{+}11]. Question [Abr98, Her98]. questions [Bea15]. Quickly [BCD^{+}23]. Quite [CJ92]. Quotient [KS92, CsL08, Fas06, MMNR15]. quotient-difference [CsL08]. quotients [VHC17].

R [LeS70, tW99, De 99]. Rachford [WRW17]. Radical [BH97, Web96, Sto79, Sto84]. Radicals [Jar15, MMR99, Wan98, Sht77a]. Radio [BSB^{+}23]. radioactive [KA99]. Radius [Mor10]. Ramanujan [Alm11, Rad10]. Ramanujan-like [Alm11]. Ramified [SFH00, FSH02]. Rand [Qui97]. Random [Raj80, SFH00, Fit74a, MG15]. Rank [Fon08, BJS06, JP15, SS13, WRW17]. rankings [Gol06]. ranks [Kel15]. Rapid [Har00]. Rational [BCCL10, FL91, Ghe14, HK05, KU14, Lab15b, Mur14, Ngô11, Yan91, AKR20, AHCDT15, AHM15, AKM16, Bro74, CDS99, CsL09, DJ96, Elk08, EMT08, ES10b, Fat72b, Ghe10, Hal74, KN00, Mio82, MCJ15, MP16, Mur09, Ren82, SS92, Smi84b, US06, WGD82, YFJS19, Zho82, vdHL21]. Rational-functions [BCCL10]. Rationally [Fat72b]. rationals [Nov08, NvH08, WCF12, vH10]. re [JRH15]. re-writing [JRH15]. Real [AM16, BKR23, DK98, GK85, Mai99, Rum77, Abb14, Akr78, ÁS99, Arn81, Cap86a, CMX10, CDL^{+}11, CGL05, CM83, Col74, ES19, FIS16, Gle86, MBS15, SE22, SS11, Tsi16, Ver75, YA07]. Real-Time [BKR23]. RealCertify [ME18]. Realization [KO83]. Really [KO83, AM08]. realm [Ste81]. rearranging [Fat03b]. Reasoning [Bar08, ABK^{+}16, Gro99, KRV21, Lee10]. Recognition [Cal74, Tre90, Smi84b]. recombination [Bre05]. Reconstructing [CsL10]. reconstruction [DSJL08, FCW13, WGD82]. recordings [Mon21]. Recovering [Gut15]. recovery [BM15]. Rectangular [CJ97]. Recurrence [SVV95, Bar03, CC83, CvHL10a, CvHL10b, MS16b, MS18, RT88, Ver74]. Recurrences [Bac97, AZ91, Pet08]. recurrent [HMSP18]. Recursive [AGR13, Kem81, AM10b, Gae99]. REDLOG [DS97]. REDUCE [Loo77d, Cap86b, AE83, ASW89, BC81, CH85, Dav82a, Fit73, Gat85, Har77, Har79, Hea72a, Hea72c, Hea09, Hol92, Hv83, Kal82a, Koe95, KO83, KR85b, KR85a, MF83, MH85, Maz85, Nor78, PH81, PH83, Sch82, Sch83, ÜHK82, WR79, vH82, van86, vHvH87, Alv00, LS85, Loo72b, NW83, Rod84, Tri00]. Reduce-2 [Rod84]. Reduce-based [Alv00]. reduced [MSV08, Poh81]. reducible [Ber06]. reducing [KS16, SS09]. Reduction [CHL14, AB19, BCCL10, Buc76b, BRRT08, Kal82b, Küc82, Mad15b, Sak19, Sti83, Tra15, WRW17, vdHL18]. reductions [IdC21, JPP22, WB83]. Rees [D’A15]. references [Akr80a]. refinement [Abb14]. Refiner [HC97]. reflections [Jen79a]. region [Dem12]. regression [IMC^{+}22]. regular [AKM16, AM16, MV11]. RegularChains [CML^{+}08, CDL^{+}11, LMX05, Maz15]. regularity [AM10a, BGM99]. regulators [Jai08]. Related [Fel77, WB95, Har00, Nab07, Ple69, Rad10, Teo15]. Relation [Pra90, RT88]. Relations [SVV95, CC83, CvHL10a, Gra91, LOW08, Ver74]. relationship [CK03]. relative [HZ15]. relatives [Sto76]. Relativistic [How70]. relativity [Coh74, Kra83, Mac89]. Relaxed [SY14, Nag21]. Relaxing [GL13]. Release [Bah77]. relevant [Mio84]. reliability [POdCW19, dCW09]. reliable [LTdD^{+}10]. Remainder [SS97, Jar12, Joh75, Yun74]. remark [ABD85, Akr80a, Akr80b]. remarks [Neu80]. Remembering [Wol13]. remove [KV19]. removing [vTG03]. Ren [YFJS19]. Renschuch [Abr03]. reparameterization [SV99]. repeller [HN19]. replacement [HK05]. Reply [Fat73a]. Report [Arn87, Cha90, Cha91, CW98, Eng15, Hec94, HLS84, Lab87, Las84, Mad92, MHGG80, Mio88, Sch82, Vol96, Vol97b, Vol97c, Wat98a, ADS96, Col74, RC76, Vol97a]. repository [WBD12]. Representation [OS16, Pat96, Sch91a, CFGVDT18, CLM^{+}08, CGL05, Gol91, Kat87]. Representations [BFS23, Ba12, CDM15, Emi15, Gae99, JM98, Kli74, Wee74]. represented [CM22]. Representing [MP14, Soi97, Wis84]. reprint [Hea69]. Research [Ano00a, Ano00b, Ano00c, Ano00d, Emi15, Ham14, Liu15, Vol98, tW99, Ano91b, Far15, Far04, GMM87, Joy09a, sL17, Ter11, Vol97c]. residue [DPFD15, IMSS08, KKM15, KKM18]. residues [Nor82]. Resistance [SFH00]. resistant [GMRT^{+}15]. resisting [Tam15]. resolution [IdC21, Pei87a]. resolutions [BGLGMM18, Gal15, GMMCMMP22]. Resolvent [DJ96]. Response [LeS70]. restrictions [JM98]. Result [KS14]. Resultant [SS98, CK03, DD99, LPA15, LYG12, Min04]. Resultants [McC97, JM22, LMS12, Lew08, RZ11, Rue10]. results [AB19, CMPS10, Huf15, SS09, Teo15, Tri84b]. Retreat [Mad92]. Reversed [HM77]. Review [Ber96, BY02, Buc03, Cet96, Cor01b, Dal01, DP15, Dav03, Dec14, Gak79, GVR04, Hei13, Lob96, Qui97, Ano79]. Reviewed [Jef98a]. Reviews [Cet97]. revised [Arr08]. revision [Bah74]. Revisited [MJ14, RC05]. Revisiting [CJ18, LC19, NM13]. revolutionary [Fra80]. Rewrite [Fèv98, Raa15]. Rewriting [OS23, GSGZ15, KS16, Küc82, NS18a]. REX [McG84]. Rhine [Ano99h, Kre01]. RICA [BFS23]. Riccati [KYA08]. Richard [Ano17a, Mon20, Qui97, Ano13e, CGG^{+}04, JCMG11b, Wol13]. Ridge [Pad02]. Riemann [BB96, CJ98, Liu19, MSM23, MCJ15]. right [Glo98b]. RIMS [Nod96a, Ter11]. ring [Liu15, Pér99, San91]. rings [BLH10, CM15, DS15, DSS15, DPFD15, HL10, IMSS08, MMNR15, OS16, Teo15]. Rioja [Rob01]. Risa [IS09, SS07].

Risa/Asir [IS09, SS07]. RISC

[Vol96, Vol97b]. RISC-Linz [Vol96, Vol97b]. Risch [Dav82b, Mos69, SB89]. risk [Dav03]. Ritz [Mio74]. RLISP [Kry84]. Robbiano [GVR04]. Robust [KS19, Man11, NS18b, Fas06, Ke00]. Rocquencourt [Vol97c]. Romania [Ano08a]. Rome [Gia88, Gia89, Mor89]. Root [CL11, DK98, EV05, Akr78, CMX10, DPQ15, Gra05, Hei70, Mur10, Pan05, Rum77, Tsi16, Uga12]. root-finding [Pan05, Uga12]. Roots [Vol95, Web96, Abb14, Cav75, Fit74b, JN04, KYA08, ST02, Sas04, Zen04, vdHM20]. Rosser [Win84]. Rostock [Glo98a, LR98, Wei97]. Rota [GSGZ15]. Rotating [van80, Rim84]. rotation [Fri74, KM87]. Rounding [CPS11]. routines [CL09, HP67, SLW15]. Row [AB19]. Royal [Jen74a]. RSA [Duj08, Jon08]. RUBI [JRH15]. Rubik [KM87]. Rule [DR01]. Rules [Buc90, Raa15]. Runge [Cam73]. Running [Kal82a]. Rupert [SY22]. Russia [BBS^{+}02, Ano99f, Ano08b]. RWCA’00 [Ano99h]. Rybowicz [DP16].

SAC

[Ano95f, CMR74, Hor72, Lan86, Ver75]. SAC-1 [CMR74, Ver75, Hor72]. SAC-2 [Lan86]. SACLIB [Ric09]. saga [Lic15]. SAGE [JM08, Joy09a, SJ05, Goy21, JP19, JPBCL20, Erö10]. SageMath [ACLM23, CRDMBR19, KM19]. Saint [Ano99f]. Samarkand [Ano99j]. Same [Bea15, Ano82c]. SAMPLE [Gle92]. Sampling [CFRT20]. San [Col69]. Saneblidze [VJ08]. Sanpou [Mor10]. Sanpou-Hakki [Mor10]. Santander [Gut04]. SAT [BKG18, ZK11]. Satisfiability [ÁAB^{+}16]. satisfied [AZ91]. Sato [CG19, SS07]. saturated [CC08]. saturation [KM11]. Saul [Zei12]. Saving [Atk81]. sca2002 [Rei02]. scalar [AB19]. scale [CsL18]. Scaling [SFH00, Lab15a]. Scheme [KU14, GMRT^{+}15, KI11]. Schemes [Sch83, Bar10b, CK04, HKS19, Sha05]. schools [AF84, Lou84]. SCHOONSCHIP [RS75, Str74]. Schubert [HM99]. Schützenberger [BDM13]. Schwartz [Hel15]. Science [Kot08, CC72]. Sciences [Ano95e, Vol98, Bar02]. Scientific [Ano99j, Ano08a, CHJ^{+}97, DP15, Qui97, RL98, Fat99, vdHGG^{+}13]. scientist [Ano13b]. Scotland [Ano99c, Ano99d, Tra00]. Scratch [BCL23]. SCRATCHPAD [GJY75b, GJY75a, Jen72, Jen74b, Jen79a, Nor75b, Wan89]. SCRATCHPAD/360 [Jen79a]. SDeval [HL15]. SDPs [CFRT20]. Search [ZGA^{+}23, AD10, CDS99, CTKR99, LC19]. Searching [BB96, AKR20]. Second [Ano99h, Ano99g, Hon96c, vHI14, ARS17, Arr15, Loo72b, Sch16]. second-order [Arr15, Sch16]. Secret [KU14, Duj08, KI11]. Section [Sit06, Rim84]. seeing [JN04]. Seeking [Nag17]. Segrè [Hel15]. Seiler [IdC21]. Selected [Pec07, DP15]. Selection [Cor00, MGR^{+}08]. Self [Lic84, Ber06]. self-conjugated [Ber06]. Self-explanatory [Lic84]. Semi [BKR23, Bro03, Bro04, HHS20, Vor80, Xia07]. Semi- [BKR23]. semi-algebraic [Bro03, Bro04, HHS20, Xia07]. semi-classical [Vor80]. semialgebraic [QH08]. semigroups [DGS16, GGMAVT21, Ple69, Pre15b]. Seminar [Wan78c, Ovc14, Sit14]. Seminumerical [Knu75]. semirings [Dur15]. semisimple [Gal15]. SENAC [SB89]. sensitive [Ste10]. separating [Bou15, ST02, Sas04, Ste00]. separation [Nag04]. September [Ano08a, CL96, Hon94, Mon01, Rob01]. Sequence [AP23, Kel08, SS97, Ara15, MSP05, RRSS20]. sequences [BKM07, HMSP18, Jar12, Joh75, Kau03, TK22, YK08]. Series [HM77, Koe93, SASG99, Sch02, Sch83, Sto12a, Sto12b, AKM16, Bar89, CGKZ05a, CGKZ05b, GL10, Gen74, KS15, KS92, LT18, Mor87, Nor75b, Pei87b, SI08, SFG13, TMCP22, Wan13, Wan16, Wil84, vdR73]. Server [Bac97]. service [FCB^{+}22]. servoing [FCB^{+}22]. Session [LR98, Zim99, Zim00, Col69, Joy09a, Mos70b, Sit06]. Sessions [Ano95a]. Set [Sat97, BD15, Fat76, GH08]. sets [Bro03, Bro04, CFRT20, CLM^{+}08, DMSP10, DHS04, GK85, HHS20, KH81, LPR99, Lev15b, MW18, PR15, QH08, Ull06, PT99]. Seventh [Ano99h, Rob01]. several [BA21]. shapes [Emi15, Emi20]. Shared [Web90]. Shares [KU14]. Sharing [KU14, Kal82a, KI11]. shell [Jen79b]. shift [CsL18]. Shirshov [GSGZ15]. short [Akr78, Akr80a, Akr80b, Fat98, LS16, RSSU79, Sht77b, Whi78]. shortened [OVB15]. Should [Buc90, KV19, Mos70a, Sch75]. shuffle [CDM15]. SIAM [CW98, Wan78d]. SIAN [HOPY19]. Sierpinski [FSH02, SFH00]. SIGART [Mos70b]. Signal [EJP^{+}01]. signature [Ede13, EF15, ELME22, Sak19, SLW15]. signature-based [Ede13, EF15, ELME22, Sak19, SLW15]. Significance [Ost99]. SIGSAM [Gon89, Mos70b, Ano82b, Ano14b, Ano14c, Bro17, Cha97, Col69, Coo96b, Cor99b, Cor01a, Cor02, FS97, HJ74, Kot14a, Kot14b, Lau77, Lak98, Neu69, Nor78, RS75, ISS89, Wat97a, Wat98a]. SIGSAM/ISSAC [Lak98]. Similar [Ste99, Loo72b]. Simon [Ano98, Doo99, Jef98b, Ber98a]. simpcomp [ES11, ES10a]. Simple [Bar91, BI91b, CJ92, FL91, Bar15, BI91a, Fat96, Fit74b, Hag77, JR91, NP88, SW85, vHvH87]. Simplicial [ES11, LPR99, BGM10, ES10a, MGRR21, Rou09]. Simplification [Yan91, Bah71, Cap86a, Col69, Maz85, McG84, Nat92, Sht77a, Sto76]. simplified [BB80]. Simplifying [LOW08, Sto13b, AvH19, Fat72b]. simply [CJ18]. Simulation [Ano99f]. Simulations [HC97]. sine [Mai99]. sine-polynomials [Mai99]. single [FN75]. Singular [KN98, SI00, Iwa05, Vil98, GPS08, LS03, GPS97, Lev10]. singularities [GK16, Nor82, vHK12]. Singularity [GPS97, SIK05]. singularly [Mad15b]. singularly-perturbed [Mad15b]. Sister [Zim03]. six [Sun73]. Sixth [Ano99e, Ano99g, Ano03, Mon01]. size [BD14, DHS04, Fre15, MJ13]. sized [Nor79]. skeletons [Ksi15, Per15]. skew [Ayd15, CL15, HL10, LOW08, Man15a]. slice [Rou09]. SLIP [HP67, SS68]. slow [KS81]. slowly [Bar89]. SLV [Tsi16]. Smale [STC22]. Small [BI91b, Dav00b, Dav12, KO83, BI91a, DIS10, Duj08, Ebe16, NY03, ST02]. smaller [Bah69, Sto80]. Smallest [Mur14]. Smith [Con90]. Smooth [HHS20, ES19]. SMP [Fat85a]. SMT [SE22]. SNAP [CW98, Emi96, Nag07]. snapback [HN19]. Snowbird [Wan81]. Society [Abd84, Sit06]. softcover [Qui97]. Software [DDK^{+}11, Dir00, Mon11, Nie70, Sor97, Ano10a, HSW89, Hea09, HKS17, Joy09a, KRR12, jYO20, Rai11, The10a, Tsi16, Xie07, YFJS19, Zen08]. solitary [Rag15]. soliton [CHH^{+}99, CHH^{+}99]. Solution [AB14, AP23, Bah75a, Cor00, FMW74, GK85, GCG83, Har77, Hea72c, KYA08, RS75, Sto98, Tut73, BHMM13, Cla69, CMR74, CJ88, Cor93, Fit74c, GK84, GJ74, GJY75a, Hal75, Jen79b, KA99, Lau77, Loo77e, Mor87, Nor78, Rim84, RT88, San71, Sta74, SCWL08, Vil98, Wan16, Zip75, vdR73]. Solutions [Ghe14, Izu14, Tec98, Mon93b, Nor75a, vHI14, AKR20, CP12, CDL^{+}11, FG03, Ghe10, Gos98, Ili08, LH15, Lee19, LMF12, Mad15a, MS18, Ngô11, Par15, Rag15, SS12a, Ste10, Tro97, Wil84, vHK12].

Solvability [SS10]. solve

[BCCR80, Fat96, Lew04, Mio82, MT15].

Solved [WR79]. Solver

[Sat97, Wil01, Cha81, SE22, WSW13]. Solvers [Ber96, BSB^{+}23, Li09]. Solving [BK04, BGLM23, BJS06, BK14, Buc91, CvHL10b, CHJ^{+}97, Dec14, HC97, JdM12, Laz01, Mae87a, MK96, MP11, MC95, Pet08, Roa08, SASG99, SIK14, Sch02, Ste96, Bat13, Bul10, CHH^{+}99, CRWZ20, Fat73b, FBOS88, FCB^{+}22, GH08, GJM09, GK88, HMSP18, Kot01, Lab87, Ley10, Mai99, MGR^{+}08, Mio74, Pan05, Pea84, RZ20, Sch16, TMV18, Tri11, Wol05, Wol08a, Xia07, Xie07, Xue15, YA07, Yun73, ZK11, Zho07, CvHL10a, FvH10, SY22]. Some [Bar91, BKR^{+}91, Buc91, Buc76a, CHJ14, Dur15, Ebe83, HKS17, Huf15, Hv83, KS94, Kan76, Kor77, Küc82, Nam86, Neu80, PR15, Sit89, Thi08, vHvH87, ÁS99, Bar89, Bel83, Ber98b, D’A15, GGdR^{+}15, Gra91, Joy05a, Lev15a, NO91, Sar15]. sometimes [Fat96]. Sommese [Dec14]. Son [Knu75]. Sonata [AY16]. Source [Dal02, DMP^{+}02, Joy08, Joy09a, Joy09b, JČMG11a, Pag07]. Southern [Ano99a, Ano99i, Ano00c, Ano00d]. Soviet [Mio81]. Space [CGKZ05a, CGKZ05b, HJN^{+}23, AHM15, Atk81, Dou99, KH81, Nai22, Qi20]. Space-efficient [CGKZ05a, CGKZ05b]. Spaces [AB14, AP23, RS91, HMMR22]. Spain [Gut04, Mon01, RL98, Rob01]. Spanish [Mon01, Rob01, Ber99b, Gim03]. Sparse [Bak93, CsL08, EK12, FSS15, GLS97, GR16, HNT14, HNRS18, Joh74, LYG12, MP10, SIK14, vdHL14, AGR13, CsL09, CsL10, CsL18, Ebe16, Fat03a, HHM10, HM13, IK20, IMC^{+}22, JM10, Kli78, KS19, Man77, MT15, MW17, NS17, NS18b, PM07, RZ20, Row81, Vil98, WSW13, WCF12, vdHL21]. sparsest [GRT10]. sparsity [CHS16]. spatial [Lee10]. Special [Dew00b, Gra91, KW97, Mon96, Piq89, Piq90, SB97, SW96, SW97, vH75, Joy09a, Mos69, Öre84, Pig90, Rei02, Sit06, SPA^{+}99, Sta10, Win84, Ano95a]. Specialization [GLGVTZ00, Mon99, Rue15]. specific [SC17]. spectral [MGRR21, RRSS20]. Speed [FP89, Bah73b, Fat03a]. speed-up [Bah73b]. Speeding [FH96]. sphere [MCJ15]. spin [Fri74]. Spinning [Tot15]. Spinor [Lac92]. SPIRAL [MJF^{+}10]. SPIRAL-generated [MJF^{+}10]. spline [JSS03]. split [Doe08, Sha15b]. sponsored [Col69]. sporadic [Sao10]. Spreadsheet [ACS92]. Spring [Sit06]. Springer [GVR04]. Springer-Verlag [GVR04]. Square [FG99]. Square-Free [FG99]. squarefree [Wan77]. St [Ano99c, Ano99d, Tra00]. St.Petersburg [BBS^{+}02]. Stability [Myl05, DM15, Ke00]. stable [IdC21, Qi20]. Stages [Ghe14]. Standalone [MZ97]. Standard [FF82, KiL16, iL15, CM15, GK16, Kat87, Nam86, Ren15, Hea69, MHGG80]. starting [Pop00]. State [Ano99j, Ano08b, KH81, Gri79, POdCW19, RSSU79]. State-space [KH81]. statement [GJY75b, Zas84]. statements [Sht77b, van86]. statistical [HL08, Mig80a]. Statistics [Tro82]. Status [Sch82]. Stavros [Ano14a]. Steering [Kal96]. Stefan [Ano12]. Stegun [CDJW00]. Steps [Abr00, Abr98, Her98, HKS17]. Stetter [Ano99a, Ano99i]. Stevens [Sit06]. Stewart [DN90, GZ05b]. Stickelberger [Ara15]. stimulate [Coo97]. Stockholm [Jen74a]. storage [Cam76, Moe78]. storing [AB19]. straight [Sed08]. Strang [LC19]. strategies [Küc82]. strategy [Sak19]. String [Alm11]. strong [FPS16]. strongly [Wil84]. structural [HOPY19, IOP21]. structure [CDM15, Dah12, FPS16, Moe78, MP12, RC76]. structure-preserving [FPS16]. Structured [BNN14, HHS12, BJS06, JMV09, LM16, SS13, Uga12]. Structures [MP14, FP89, FN73, PH83, Ple69, Neu69]. Student [Far15]. Students [Buc90, Cal84]. Study [WR79, CW72, KD18, Kli78, Mat10, Mig80a, Sal93, Sla91, Thi00a, Thi00b]. stuffle [BDM13]. Sturm [Mur10]. Sturmfels [Sch02]. style [Jeb18, Rue15]. Subcommittee [Neu69]. subexpressions [Fat15]. Subfields [GR99, Lan93]. subgroups [DI91, JR91]. submodules [GK16, MMR99]. Subprograms [CCM^{+}14, CCM^{+}16]. subroutine [Tut73]. Subscripts [Hv83]. substitutions [Row81]. subsystems [Lev10]. successive [Poh81]. sucesiones [Ibá08]. Sum [MP14, Bar89, KYA08]. Sum-of-Products [MP14]. Summary [Mos70b, Lak98, WRL99, Col69]. Summation [Erö10, Zim14, BKSS10, Car03, ES10b, Koe95, McN04, Sch08a, Zim18, Zim03]. summations [Wan91]. Summer [Sit14]. Sums [BCD^{+}23, BFS23, Le01, Bro09, CX08, DMH15, Fri74, Smi84a]. sumset [GGMAVT21]. Sun [Ano14a]. superelliptic [Bes15, BSS15, Sha15a, Sho15]. superhigh [Sha05]. support [EK12, LaP72, Tri00, pro12]. supports [CK03]. surface [CHH^{+}99, Iwa08]. Surfaces [CJ98, BGJ11, CGL05, Edo08, ES11, Elk08, Jai08, Kel15, KLV03, Lub11, Lub15, Sch99, YFJS19]. survey [EF15, Gri76, LS16, Ple69, RSSU79, Sau80]. Suzuki [SS07]. Sweden [Jen74a]. Sweedler [Kal93a]. Switzerland [Lak96]. syllabus [Akr80a, Akr80b]. Sylow [JR91]. Sylvester [Rue15, WW11]. SYMBAL [Eng75, Nei71, Sun72]. SYMBIONT [BFR^{+}18]. SymbMath [Ano93]. Symbol [Che83, Don83, OS23, Ste82, Vor80, CL82, KS16, NS18a, SS93, SW85]. symbol-manipulation [SW85]. SYMBOLANG [Ber70]. Symbolic [ACM94, Ano91b, Ano93, Ano95d, Ano98, Ano99c, Ano99d, Ano08a, Bar03, BvM06, Ber96, BFS23, Bos84, Bro93, BKR^{+}91, Cam88, Cap03, CC72, Cha86, Cha96a, CHJ^{+}97, Cla69, CsL09, Doo99, Enc97a, Eng15, Fat81, FH96, Fat99, Fla88, Ged99, GMM87, Gia88, Giu01, Gle86, Glo98a, Gon89, Gri76, Gut04, Har79, Hod11, Hon94, Hon96c, IMC^{+}22, IT12, Jef98b, Jen76, KF99, KW97, Kau05, Ke00, Küc97, LR98, Lak96, LLT18, Lev95, Lic11, Mor02, Mou01, Ng74, Ng79, Nor82, Pop98, Rei00, Rei02, The10a, ISS89, ISS94, SN97, Sen03, She07, SB97, SW97, Sun74, SVZ06, Tra06, Tra00, Vol97c, Vol98, Wan78c, Wan81, Wan92, WN90, Wat91, Wat98b, Zho07, AE83, ÁAB^{+}16, Bar08, Bel83, Ben08a, Ber99c, BGJ70, BKSS10, BGJ11, BCCR80, BFR^{+}18, BCLA82]. symbolic [Buc84, CL96, Car03, Cha81, CGG83, Col69, CM02, Coo04, Din09, Ebn86, EB70, ES10b, Fat84, FBOS88, Fat15, Fel75a, FS97, Fuc08, Gen71, GMC^{+}21, Goy21, Gra87, Gri79, HI93, HYH15, HLS84, How70, Hue84, Jen77, KO13, Kan76, KPR16, KA99, Koh10, Kri84, Lac92, Leg84, Lic15, Lic84, Llo84, LMF12, Mat21, Mio74, NP88, NS17, RSV02, The10b, SV10, Sho82, SS17, Sto79, Tec11, Vor80, Wan76b, Wan80c, Wan80b, dT84, vH84, vdHLM^{+}11, Ano84, BU93a, BU93b, Gia89, KRR12, Ng79, Rei02]. Symbolic-algebraic [She07]. Symbolic-Numeric [SW97, CsL09, Hod11, IMC^{+}22, SVZ06, Goy21, vH84]. symbolic-numerics [GMC^{+}21]. Symbolically [BB96, SASG99, Xie07]. SymbolicData [Grä18]. SymGrid [The10b]. SymGrid-Par [The10b]. Symmetric [KS14, CP12, FPS16, Gae99, GMRT^{+}15, JGB12, SS68]. Symmetry [AHM15, Met23, LBL20]. symplectic [Gra91]. Symposium [ACM94, Aca20, Ano95d, Ano95f, Ano98, Ano99c, Ano99e, Ano99d, Ano99b, Ano99g, Ano03, Ano08a, Bro93, CL96, Cap03, Cha86, Cha96a, Doo99, Enc97a, Eng15, Gia88, Gia89, Giu01, Glo98a, Gon89, Gut04, Hon94, Hon96c, Jef98b, Jen76, Kau05, Küc97, LR98, Lak96, Lev95, Mor02, Mou01, Ng79, Rei00, ISS89, ISS94, Sen03, Tra06, Tra00, Wan81, Wan92, WN90, Wat91, KV04]. SymPy [JČMG11a]. Symsac [Cha86, Wan81, Jen76]. synergy [Öre84]. SyNRAC [FIS16, YA07]. Synthesizing [CC83]. SYPAC [di 82]. System [BGLM23, BK14, Dav00b, Fèv98, GK85, GPS97, KO83, Mon93b, Pop97, SVV95, SJ05, Ste96, TO97, Yan91, dR90, AsW08, ÁS99, Arn10, Bah69, BS74, BGJ11, Bro04, Bul10, CL87, CGG83, Cla69, CRWZ20, ELME22, Eng75, Far15, FCB^{+}22, GL10, GK84, GCG83, GK88, GPS08, Grö10, HK08, HLS84, JT91, Jen70, Jen77, KS81, Kli74, Kot01, KR85b, LaP72, LW08, LPR99, LS03, Li09, Lou84, Man10, MGR^{+}08, Meu03, Moe78, Mor87, NP88, Pag07, Rim84, Roa08, Rod84, SS11, SS12a, SH09, SW06, Sho82, Str74, Tro97, Wil84, Xue15, Zho15, Zhu15, dCW09, di 82, KP97]. Systems [Ano84, BNN14, Buc91, Cal01, Dec14, Ger98, GLS97, GS97, Laz01, SIK14, Sch02, Stu23, dQ77b, dR90, AKR20, BK04, Ban08, Bar10a, BPS12, Bar15, Bat13, Bos84, BJS06, Bou15, BP20, CL96, CDL^{+}11, CK03, CJ88, DM15, Dav12, EGK09, Fat81, Fat87, FP89, Fat96, Fat15, FIS16, Gae99, GJM09, GSGZ15, HI93, IT12, Joy08, Joy09b, JČMG11a, KS03, Kou09, Kre88, KKM15, KKM18, Lab15a, Laz80, Lee19, LW99, Lew00, Ley10, Llo84, LMF12, Mac89, Mad15b, Mad15a, MK96, Man10, MP11, MS16b, MS18, MGRR21, Min15, NOT18, Nat92, POdCW19, Pon88a, Por09, RSV02, The10b, Sau80, SV10, SS13, SS10, Ste10, Sun73, Sun74, SS07, TMV18, Tif15, Tri11, Vil98, Wol05, Wol08a, WSW13, Xia07, Xie07, Yun73, dQ77a, dT84]. syzygies [Ter13, Wal89].

T [CHJ14]. Table [JK02, Pig90, Piq90].

tableaux [CP88]. Tables

[Gak79, Nor80, DI91]. taking [Fit74b].

Talks [Ano00a, Ano00c, Ano00d, HKS99, Ano10a]. tangent [JRH15, vdHM20]. tangents [Dav82b]. tarow [SK16]. taxonomy [GJY78, Neu80]. Taylor [Kal93a, Wil84, van80, vdR73]. Teaching [jYO20, tW99, Buc84, CR84, Wan76b]. Technical [Vol96, Vol97b, Vol97c, Vol97a, Mos70b, Whi78]. technique [SS68]. Techniques [Jeb18, PH81, CL11, DIS10, Lew08, Ren15, Ric09, Zho07]. Technologies [Pad02]. Technology [Gle92, Jen74a, tW99, GGdR^{+}15, Sit06]. Tekhnika [Gak79]. Telescoper [Le01]. telescopers [BCCL10]. Telescoping [CHL14, BCCL10]. Temperley [KiL16]. tensor [JM98, Liu19, jYO20]. term [JRH15, Kre88, Küc82, Sit89, Sti83]. term-ordering [Sit89]. termination [EGP11, HZ15]. Terms [CHL14, Le01, CP12, CPS11, Dav84, Erö10, Heb15, Kas80, LMTV05, Raa11, vTG03]. test [Val20]. Testers [Bsh12]. Testing [LWZ07, Nor79, Par15]. Tests [Mig80b, Wol08b, dQ78]. TEX-REDUCE [ASW89]. Texas [Sch02]. TEXmacs [vdHGG^{+}13, vdH04]. textbook [CWZ12]. th [BC02, Fit74b, Vol95, BRRT08, Zwi84]. Thailand [Ano99b]. Their [Ano84, EZ23, CFGVDT18, DS15, Kli74, Ksi15, Mio82, Sán15, Sha15a]. Thematic [LW16]. Theorem [Kal93a, Sch83, BB80, Ben08b, Bin87a, CG19, Gor98, Jer15, MA08, Pei87a, RC76, Sas04, Spe08, Wib15, Win84]. Theorema [Tec11]. Theorems [Fèv98, LS16]. Theoretical [Ano01b, Buc91, Buc76b, FB70, LN76, Sch71]. Theories [Li15]. Theory [Abr98, Alm11, Fel77, Gak79, GPS97, Grö98, Her98, MPW97, Ost99, Pra90, TO97, Bos84, CR84, Göb93, Grö10, Hei15, HMS10, Jen15, JM08, Kas80, KV04, Man77, Pos13, RRRA03, Sán15, SV10, Sch91a, Spe08, STC22, Tha08, Tri84a, Ull06, ZT08, Zim84b, Zim84a]. thesis [Win84]. Thiele [Cel22]. Third [Ano99j, Ano91a]. Thirteenth [Bai02]. Thomas [BGLH^{+}10]. Thompson [BHMM13]. Thought [Aca20]. Three [HEW^{+}14, Zim97, CJ88, Loo72b, SW06, VHC17]. Thursday [DMP^{+}02]. TI [Sto98]. TI-92 [Sto98]. Time [BKR23, Sak97, BM15, Bla14, CJ18, DMSP10, Kal82a]. times [PH83]. Timisoara [Ano08a]. Titles [Wan80c, Wan80b]. Tlisp [Wil92]. TOCL [Ano99k]. today [vH74]. Tokyo [WN90]. tool [AY16, CDM15, HOPY19, Pea84, SS17, Xia07]. toolbox [ES10a, YA07, Zen08, ZL13]. toolkit [HL15, Wan91]. Tools [CHJ^{+}97, Gie08b, ABK^{+}16, Ban08, Fat99, KRV21, Sch91a, Smi84a]. TOP [Coo04]. TOP-C [Coo04]. Topics [Fel77, Qui97, DP15, Mio84, Pec07]. topological [CGL05]. topology [Bar99, CRDMBR19, FGVC15, Her11, KRTZ20, Ser99]. tori [Rue05]. Toric [KLPW17, BGM10, Gao15, Hub00]. torsion [DPFD15]. traces [IKRT89]. tracking [Fat99]. trade [Dav82a]. trade-off [Dav82a]. Transactions [Ano99k]. transalpyne [DS10a, DS10b]. transcendence [Ste00]. Transcendental [Gak79, GR99, KKM18]. Transform [Nor80, Bar73, Cla89, GJ74, AS14, KS14]. transformation [AB19, Ste08]. Transformations [LMTV05, AJ03, She10]. transforms [IVV99]. transitive [Roy91]. Translation [WR79]. Translator [Alv00, Fat90, Jen70]. transposition [DS10a, DS10b]. Transputers [Wil92]. Treasurer [Ebe06]. treating [FB70]. treatment [Loo72b]. Tree [CHJ14]. trends [Ano10d]. triangular [DMSP10, Li09]. Tricoloring [MS08]. TRIGMAN [Jef72]. Trigonometric [Met23, CsL10]. trinomial [Wan16]. TRIP [GL10]. tropical [Ksi15, Ren15, Ren17]. true [Zei12]. Truncated [TMV18, Gen74, AS14]. Tschirnhaus [AJ03]. Tsun [Aca20, sG06, Gao17]. Tuesday [DMP^{+}02]. Tuning [BS74]. Turing [CJ97]. twice [BM15]. Twisted [HNT14]. Two [CLZ23, Fon08, GS97, SN97, BK12, BGM99, LMTV05]. Two-Layer [CLZ23]. Two-Parameter [GS97, SN97]. Type [BCD^{+}23, Dav00b, ASP15, BTL99, GSGZ15, KN98, MY74, Mur09, Zim03, vHK12, vHI14, AW80, Raj80]. typed [Wil84].

ubiquity [Jal99]. UC [Sch02]. UIUC [Gra87]. Ukraine [Bro93]. Ultimate [vdHL20]. Ultrametric [STC22]. Umble [VJ08]. Unboundedness [MSS18]. uncouth [CDJW00]. Undecidability [Par15]. Under- [Cal01]. underflowing [CPS11]. undergraduate [Far15, HLS84, Pea84]. unification [KN87, RSSU79]. Unified [Ben08a, dO99]. union [Ore99]. uniqueness [Par15]. Unit [Fon08, Sto09, Gle86, Mig77]. United [ACM94, Ano99e, Ano99g, ISS94]. Unity [Web96]. univariate [CMX10, CMPS10, IMC^{+}22, LW94a, LW94b, Mig80a, MY74, Nov08, NvH08, Ter10, Ter13, vH10]. Universal [MBRD23]. Universität [Ano91a]. Université [Mor02]. University [Ano98, Ano99c, Ano99e, Ano99d, Ano99g, Ano99j, Ano00a, Ano00b, Ano00c, Ano00d, Ano08b, Buc03, Dav03, Doo99, Glo98a, Gut04, Jef98b, Mou01, Rei02, Sen03, Sit14, Tra00, tW99, Ber98a, Buc84, Jen75, Sch02, Vol97c, Wat98b]. unknown [Gol85, NS17]. Unmixed [Wan98]. unnecessary [WB83]. unspecified [Cam88]. unusual [Rim84]. Unwinding [MCJ15, CJ96]. upcoming [Ano13c, Ano13d, Sta20]. upper [Sán17]. US/Japan [HI93]. USA [HI93, Küc97, Lin98a, Lin98b, Sen03, Vol98]. Use [Bar91, Fel77, FB70, Nor80, dQ77b, CH85, Dav82b, HLS84, Lou84, Mio74, Sht77b, Zid23, dQ77a]. used [BM15, Kre88, LaP72]. Useful [Sto07]. user [Loo77a, Nei71, Sto13a, Sun73]. Users [Ano91a, Cal84]. Using [Bar89, BD14, Buc91, CM76, FSH02, GGdR^{+}15, Göb93, HM99, Hol92, KN97, LS03, Lew04, MT15, Mon93b, OS23, RZ11, Xam99, Ada96, AMORH15, BG97, Bro03, EB70, EK12, Erö10, FvH10, Far15, FMW74, Fèv98, Fit74c, HMS10, HNRS18, JT91, JM98, Joy09a, KI11, KYA08, KA99, LBL20, MK96, MBS15, MC95, MA08, Mur10, Mur14, Nai22, Pri08, RSV02, Ren15, Ren17, RS75, SASG99, SI17, Ste10, SLW15, SS07, Ter13, Wri00, Xue15, vdR73, vdHM20]. Utah [Wan81, Nam86]. utility [Lud69]. Utilizing [Bah75b]. Uzbekistan [Ano99j].

V [Buc03, Cor96a]. Vacuum [HJN^{+}23]. Validated [KS03]. valuation [Ren15]. Value [KN97, PS98, Sak97, Cap86a, Dav03, Ili08, Mae87a, Sto76]. Valued [Asl96]. Values [BB96, BFE20, Com18]. Vancouver [Doo99, Ano98, Jef98b, Zim98a, Zim98b]. Variable [HEW^{+}14, UM12, MGR^{+}08]. variables [GHL15, Lab87, Liu15, Rod84, VHC17]. variant [Duj08]. variational [Hel10, Kog15]. Varieties [DN90, BMT99, BGM99, FR18, Ore99, Pau15, Ren15, Ren17]. variety [Fre15, Gao15, Rec99, Sán17]. VarInt [Hel10]. Various [Li92]. VAX [FF82]. Vaxima [NW83]. vector [Dem12, HHM10, Lou84, NOT18, Nai22, Por09]. vector-quantization [Por09]. vectors [Poh81]. Veltman [Gaw79]. verification [KS13]. Verlag [GVR04]. Version [KP97]. versus [Hue84]. vertex [Roy91]. vertex-transitive [Roy91]. vertices [Roy91]. Vessiot [Hei15]. VI [KV04]. via [BK12, BZ12, Bro04, BKR23, CsL08, Elk08, Fat03b, GMMCMMP22, Ged99, GTLN18, GMC^{+}21, HRR19, Kan76, Man15a, Mat21, PR15, Too15, Uga12, Wri00, Zha13]. view [Fat84, Mid11, Poi15]. vision [ÁS99]. VISIONARY [Zei12]. Visual [KU14, FCB^{+}22]. Visualization [IOS06, MSP05]. Vladimir [AB21, BBS^{+}02, MSV21]. Volos [Kot02]. volume [CFRT20]. vs [DP15, Pec07].

W [Dec14, Zei12]. W. [Hei13]. Walk [SFH00, Fit74a, LM12]. walks [JPBCL20]. wall [BDE^{+}21, CM02]. Walsh [IVV99]. Wampler [Dec14]. Wang [Ano17c]. want [Dav82c]. was [Zei12]. Waterloo [Cha86]. Way [BFS23, TK22]. Ways [Bar91, Sto11]. Web [Pad02, Sch07, AsW08, Bor98, LW08, Sch08b, Wri00]. Web-based [AsW08, LW08]. wedges [Pet08]. Wee [Ano10e]. Weierstrass [HV98, Pre15b, Sho15]. Weighing [BY02]. weight [JP15]. weighted [Joy15]. weights [GMMCMMP22, Sho15]. Weil [Rec99]. Well [CJ92]. Wen [Aca20, sG06, Gao17]. Wen-Tsun [Aca20, sG06, Gao17]. Werner [Joh12e]. Western [Ano00a, Ano00b, Ano00c, Ano00d, Mou01, Rei02, Wat98b]. Weyl [Lau15, iL15, Mid13, Tsa00]. whether [BGM10]. which [ÁS99, Glo98b]. whom [Zei12]. Whose [Mur14, Mig77, Mur09, NS17]. Wiedemann [Ebe16, HJS13, HJS16, HNRS18, Via13]. Wielandt [NO91]. Wiener [Duj08]. Wilf [Zei12, KZ11]. Williams [BY02]. Williamson [BKG18]. within [Hor15, Suz09]. Witt [BTL99]. Word [BD14, Nor79]. Word-size [BD14]. word-sized [Nor79]. Words [EZ23, Pud08]. Work [Rod84, PST97, Pon88b, Pon88a]. workbench [Arn10]. works [Mio81, RRRA03]. Workshop [AB21, Ano99h, Ano99j, KPR16, Kre01, Ree97, Vol98, Arn87, Emi15, HI93, Hec94, Las84, sL17, Ter11, Tho09, BBS^{+}02, Cal01, Cha90, Cha91, DMP^{+}02]. World [DP15]. Worst [Ree91, Ree92]. Write [Lün84]. Writing [Dav00a, JRH15, Jen70, Zel08]. Wu [Aca20, sG06, Gao17]. WWCA [KZ06, KZ11]. www.orcca.on.ca [Rei02]. WZ [CX08].

X [GVR04]. Xiaobai [Ano14a]. XSL

[Car00]. XXIII [RRRA03].

Yale [Jen75]. Year [Ano95a, Mos74]. Years [Grä18, Win97, Win98, BW98b]. yield [ÁS99]. York [Jen76, Sit14]. Yorktown [Jen76]. Young [CP88]. Yue [YFJS19].

Zariski  [AKM16].  Zassen-

haus [Zim76]. Zeilberger [RVMH11]. Zero [FT12, GS95, OS23, Ste99, CC08, Dah12, GK85, KS16, NS18a, Qi20, RC05, SS11]. Zero-dimensional [FT12, GS95, CC08]. zeroes [Gut15]. zeros [CM83, Gle86, KV19, Moe99, Ver75]. Zeta [BB96, Com18]. Zigzag [CMT14]. zur [Mon93a]. Zurich [Lak96].

References

Abraham:2016:SCS

[ÁAB^{+}16] E. Ábrahám, J. Abbott, B. Becker, A. M. Bigatti, M. Brain, B. Buchberger, A. Cimatti, J. H. Davenport, M. England, P. Fontaine, S. Forrest, A. Griggio, D. Kroening, W. M. Seiler, and T. Sturm. Satisfiability checking and symbolic computation. ACM Communications in Computer Algebra, 50(4):145–147, December 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Abdali:1984:A

[AB84] S. Kamal Abdali and Bruno Buchberger. Abstracts. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(1):2–7, February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Abramov:2014:SSP

[AB14] Sergei A. Abramov and Moulay A. Barkatou. On solution spaces of products of linear differential or difference operators. ACM Communications in Computer Algebra, 48(3/4):155– 165, September 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Abbott:2015:WNC

[AB15] John Abbott and Anna Maria Bigatti. What is new in CoCoALib and CoCoA-5? ACM Communications in Computer Algebra, 49(4): 113–116, December 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Abbott:2017:CC

[AB17] John Abbott and Anna M.

Bigatti. CoCoA-5.2.2 and CoCoALib. ACM Communications in Computer Algebra, 51(3):95–97, September 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Abramov:2019:RRP

[AB19] Sergei A. Abramov and Moulay A. Barkatou. Row reduction process for matrices of scalar operators: storing the intermediate results of row transformation. ACM Communications in Computer Algebra, 53(1): 23–30, March 2019. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Abramov:2021:WCA

[AB21] Sergei A. Abramov and Alla Bogolubskaya. 22nd Workshop on Computer Algebra in Memory of Professor Vladimir Gerdt in Dubna. ACM Communications in Computer Algebra, 55(2): 21–29, June 2021. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3493492.3493493.

Abbott:2004:CLC

[Abb04]  John Abbott.  CoCoA:

a laboratory for computations in commutative algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38 (1):18–19, March 2004. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Abbott:2014:QIR

[Abb14] John Abbott. Quadratic interval refinement for real roots. ACM Communications in Computer Algebra, 48(1):3–12, March 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Abbott:2007:CCC

[ABCR07] John Abbott, Anna Bigatti, Massimo Caboara, and Lorenzo Robbiano. Co-CoA: computations in commutative algebra. ACM Communications in Computer Algebra, 41(3):111– 112, September 2007. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Abdali:1984:ACA

[Abd84] S. Kamal Abdali. Abstracts of computer algebra papers from 1984 national meeting of American Chemical Society. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18 (3):26–30, August 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Abbott:1985:RF

[ABD85] J. A. Abbott, R. J. Bradford, and J. H. Davenport. A remark on factorisation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19 (2):31–33, 37, May 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Abanades:2016:DAR

[ABK^{+}16] Miguel Abánades, Francisco Botana, Zoltán Kovács, Tomás Recio, and Csilla Sólyom-Gecse. Development of automatic reasoning tools in GeoGebra. ACM Communications in Computer Algebra, 50(3):85–88, September 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Abramson:1998:HBQ

[Abr98] Michael Abramson. Historical background: “The Question of Finitely Many Steps in Polynomial Ideal Theory”. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (3):8, September 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Abramov:2000:NND

[Abr00] S. A. Abramov. A note on the number of division steps in the Euclidean algorithm. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34 (4):1–2, December 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Abramson:2003:HBB

[Abr03] Michael Abramson. Historical background for B. Renschuch’s papers. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(2):33–34, June 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Abramson:2009:HBG

[Abr09] Michael P. Abramson.

Historical background to Gröbner’s paper. ACM Communications in Computer Algebra, 43(2):22–23, June 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Abramov:2023:FCC

[ABR^{+}23] Sergei A. Abramov, Alexander B. Batkhin, Anna Ryabenko, Leonid A. Sevastianov, Yvette Zonn, and Min Wu. The fifth conference “Computer Algebra” in Moscow. ACM Communications in Computer Algebra, 57(2):72– 84, June 2023. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3614408.3614418.

AMSS:2020:ISW

[Aca20] Academy of Mathematics and Systems Science. International symposium on Wen-Tsun Wu’s academic thought and mathematics mechanization. ACM Communications in Computer Algebra, 53(4):155– 181, March 2020. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3388974.3388975.

Albrecht:2015:AAL

[ACF^{+}15] Martin R. Albrecht, Carlos Cid, Jean-Charles Faugère, Robert Fitzpatrick, and Ludovic Perret. Algebraic algorithms for LWE problems. ACM Communications in Computer Algebra, 49(2):62, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ayotte:2023:DMS

[ACLM23] David Ayotte, Xavier Caruso, Antoine Leudière, and Joseph Musleh. Drinfeld modules in SageMath. ACM Communications in Computer Algebra, 57(2): 65–71, June 2023. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3614408.3614417.

ACM:1994:IPI

[ACM94] ACM, editor. ISSAC ’94:

Proceedings of the 1994 International Symposium on Symbolic and Algebraic Computation: July 20– 22, 1994, Oxford, England, United Kingdom. ACM Press, New York, NY, USA, 1994. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994.

Abdali:1992:SCC

[ACS92] S. Kamal Abdali, Guy W.

Cherry, and Neil Soiffer. Spreadsheet computations in computer algebra. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26(2):10–18, April 1992. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Autexier:2010:RDM

[AD10] Serge Autexier and Dominik Dietrich. Recent developments in Mega’s proof search programming language. ACM Communications in Computer Algebra, 44(2):52–59, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Adami:1996:GNM

[Ada96] Chris Adami. Gaylord and Nishidate’s modelling nature with cellular automata using Mathematica. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(3):24–25, September 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Abbott:1996:ROP

[ADS96] John Abbott, Angel Díaz, and Robert S. Sutor. A report on OpenMath. A protocol for the exchange of mathematical information. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(1):21–24, March 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Abi-Ezzi:1983:CSM

[AE83] Salim S. Abi-Ezzi. Clarifications to the symbolic mode in REDUCE. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(3–4):43–47, August/November 1983. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Aspetsberger:1984:EMA

[AF84] K. Aspetsberger and G. Funk. Experiments with MuMath in Austrian high schools. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18–19(4–1): 4–7, November/February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Arnold:2013:RSI

[AGR13] Andrew Arnold, Mark Giesbrecht, and Dan Roche. Recursive sparse interpolation. ACM Communications in Computer Algebra, 47(3–4):104–105, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Alcazar:2015:ICC

[AHCDT15] Juan G. Alcázar, Carlos Hermoso, Jorge Caravantes, and Gema M. Díaz-Toca. Iterative computation of the convex hull of a rational plane curve. ACM Communications in Computer Algebra, 49(2):52, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Alcazar:2015:SDR

[AHM15] Juan G. Alcázar, Carlos Hermoso, and Georg Muntingh. Symmetry detection of rational space curves. ACM Communications in Computer Algebra, 49(2): 51, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Adamchik:2003:PTT

[AJ03] Victor S. Adamchik and David J. Jeffrey. Polynomial transformations of Tschirnhaus, Bring and Jerrard. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37 (3):90–94, September 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ajwa:2005:ECC

[Ajw05] Iyad Ajwa. East Coast Computer Algebra Day 2005: poster abstracts. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(1):15–25, March 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Alvandi:2016:CLR

[AKM16] Parisa Alvandi, Mahsa Kazemi, and Marc Moreno Maza. Computing limits with the regular chains and power series libraries: from rational functions to Zariski closure. ACM Communications in Computer Algebra, 50(3):93–96, September 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Akritas:1978:SNN

[Akr78] Alkiviadis G. Akritas. A short note on a new method for polynomial real root isolation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 12 (4):12–13, August 1978. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Akritas:1980:CLR

[Akr80a] Alkiviadis G. Akritas. A complete list of references for the paper: “A remark on the proposed syllabus for an AMS short course on computer algebra”. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14(3):32, August 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Akritas:1980:RPS

[Akr80b] Alkiviadis G. Akritas. A remark on the proposed syllabus for an AMS short course on computer algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14 (2):24–25, May 1980. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Akritas:1981:BFC

[Akr81] Alkiviadis G. Akritas. On the Budan–Fourier controversy. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15 (1):8–10, February 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Akritas:1987:NPM

[Akr87] Alkiviadis G. Akritas. A note on a paper by M. Mignotte. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21 (4):23, November 1987. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Abramov:2020:CSR

[AKR20] Sergei A. Abramov, Denis E. Khmelnov, and Anna A. Ryabenko. Checkpoints in searching for rational solutions of linear ordinary difference and differential systems. ACM Communications in Computer Algebra, 54(2):18–29, September 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3427218.3427219.

Almkvist:2011:RLF

[Alm11] Gert Almkvist. Ramanujanlike formulas for 1/π^{2} and string theory [abstract only]. ACM Communications in Computer Algebra, 45(2):92, June 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). To appear in Proceedings of WWCA 2011.

AlvarezSobreviela:2000:RBO

[Alv00] Luis Alvarez Sobreviela. A Reduce-based OpenMath ↔ MathML translator. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(2):31– 32, June 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Arnold:2008:CRB

[AM08] Andrew Arnold and Michael Monagan. Calculating really big cyclotomic polynomials (abstract only). ACM Communications in Computer Algebra, 42(1–2):48– 49, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Antritter:2010:EAC

[AM10a] Felix Antritter and Johannes Middeke. An efficient algorithm for checking hyper-regularity of matrices. ACM Communications in Computer Algebra, 44(3):84–86, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Arnold:2010:FRA

[AM10b] Andrew Arnold and Michael Monagan. A fast recursive algorithm for computing cyclotomic polynomials. ACM Communications in Computer Algebra, 44(3):89–90, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Alvandi:2016:RLP

[AM16] Parisa Alvandi and Marc Moreno Maza. Real limit points of quasi-componenets of regular chains. ACM Communications in Computer Algebra, 50(4):148–150, December 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Adj:2015:CDL

[AMORH15] Gora Adj, Alfred Menezes, Thomaz Oliveira, and Francisco Rodríguez-Henríquez. Computing discrete logarithms using Joux’s algorithm. ACM Communications in Computer Algebra, 49(2):60, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:1979:BR

[Ano79] Anonymous. Book review.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 13(3):??, August 1979. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1982:ECP

[Ano82a] Anonymous. EUROCAL ’83 call for papers. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16(2):37, May 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1982:KIS

[Ano82b] Anonymous. KWIC index of SIGSAM Bulletin abstracts (1978–1981). SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16(2):15–36, May 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1982:SC

[Ano82c] Anonymous. SAME created. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16 (2):38, May 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1984:PSM

[Ano84] Anonymous. Papers from Symbolic Mathematical Systems and Their Effects on the Curriculum — International Congress on Mathematical Education. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(4):1–62, November/February 1984. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Issue labeled vol. 18, no. 4, November 1984 (issue #72) and vol. 19, no. 1, February 1985 (issue #73).

Anonymous:1991:PTC

[Ano91a] Anonymous. Proceedings of the Third CAYLEY Users Conference, November 17– 19, 1988, Universität Essen. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(1):25–81, January 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1991:SCR

[Ano91b] Anonymous. Symbolic computation research in Italy. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(3):38– 48, July 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1993:SSC

[Ano93] Anonymous. SymbMath 2.2: a symbolic calculator with learning. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27(3):25–??, September 1993. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1995:ACS

[Ano95a] Anonymous. Abstracts of closing sessions of the Special Year in Computational Differential Algebra and Algebraic Geometry. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(3–4):13–18, December 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1995:E

[Ano95b] Anonymous. ECCAD ’95.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(2S (special issue)):21–28, June 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). 2S (special issue).

Anonymous:1995:EEC

[Ano95c] Anonymous. ECCAD ’96: East Coast Computer Algebra Day. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29 (3–4):21, December 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1995:IIS

[Ano95d] Anonymous. ISSAC ’96: International Symposium on Symbolic and Algebraic Computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(3–4):19, December 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1995:MMS

[Ano95e]  Anonymous.  Maple in

the mathematical sciences. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(3–4):20, December 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1995:SAS

[Ano95f]  Anonymous.  SAC ’96:

ACM symposium on applied computing. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(2S (special issue)): 19–20, June 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). 2S (special issue).

Anonymous:1996:CIB

[Ano96]  Anonymous.  Calendar.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(3):??, September 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Inside back cover.

Anonymous:1998:PAI

[Ano98] Anonymous. Preliminary announcement: ISSAC ’99 Vancouver: International Symposium on Symbolic and Algebraic Computation, July 1999, Simon Fraser University, British Columbia, Canada. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(1):64, March 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (elec-

tronic). URL http://www.
cecm.sfu.ca/ISSAC99/.

Anonymous:1999:ECC

[Ano99a] Anonymous. East Coast Computer Algebra Day 2000: Joint meeting with Southern Ontario Numerical Analysis Day 2000 and the 70th Birthday Celebration for Professor Hans J. Stetter; Announcement and Call for Participation. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(2):59, June 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1999:FACd

[Ano99b]  Anonymous.  First an-

nouncement and call for papers: ASCM’2000: The 4th Asian Symposium on Computer Mathematics, Chiang Mai, Thailand, December 17–21, 2000. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (2):62, June 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). URL http://ww.

mmrc.iss.ac.cn/~ascm.

Anonymous:1999:FACa

[Ano99c]  Anonymous.  First an-

nouncement and call for papers: ISSAC 2000: International Symposium on Symbolic and Algebraic Computation, University of St. Andrews, Scotland, August 7–9, 2000. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(1):28, March 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1999:FACc

[Ano99d]  Anonymous.  First an-

nouncement and call for papers: ISSAC 2000: International Symposium on Symbolic and Algebraic Computation, University of St. Andrews, Scotland, August 7–9, 2000. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (2):60, June 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1999:FACb

[Ano99e]  Anonymous.  First an-

nouncement and call for papers: MEGA 2000: The Sixth International Symposium on Effective Methods in Algebraic Geometry, Bath University (United Kingdom), June 20–24, 2000. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (1):30, March 1999. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1999:IAI

[Ano99f]  Anonymous.  IMACS-

ACA’2000: International Association for Mathematics and Computers in Simulation: 6th International Conference on Applications of Computer Algebra, June 25–28, 2000, Saint Petersburg, Russia. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (2):58, June 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1999:SACb

[Ano99g] Anonymous. Second announcement and call for papers: MEGA 2000: The Sixth International Symposium on Effective Methods in Algebraic Geometry, Bath University (United Kingdom), June 20–24, 2000. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (2):61, June 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1999:SACa

[Ano99h] Anonymous. Second announcement and call for papers: RWCA’00: Seventh Rhine Workshop on Computer Algebra, March 22 – March 24, 2000, Bregenz, Austria. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(1):29, March 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1999:SON

[Ano99i] Anonymous. Southern Ontario Numerical Analysis Day 2000: Joint meeting with East Coast Computer Algebra Day 2000 and the 70th Birthday Celebration for Professor Hans J. Stetter; Announcement and Call for Participation. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(2):63, June 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1999:TIW

[Ano99j] Anonymous. The Third International Workshop on Computer Algebra in Scientific Computing, CASC-2000, October 5–9, 2000, Samarkand State University, Samarkand, Uzbekistan. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (1):31, March 1999. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1999:TCL

[Ano99k]  Anonymous.  Transac-

tions on Computational Logic (TOCL): Call for papers. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (2):64, June 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (elec-

                                        tronic). URL http://www.
                                        acm.org/tocl.

Anonymous:2000:ECCa

[Ano00a] Anonymous. East Coast Computer Algebra Day: May 13, 2000, Ontario Research Centre for Computer Algebra, University of Western Ontario, London, Ontario, Canada: Abstracts of invited talks. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(3):21, September 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:2000:ECCb

[Ano00b] Anonymous. East Coast Computer Algebra Day: May 13, 2000, Ontario Research Centre for Computer Algebra, University of Western Ontario, London, Ontario, Canada: Abstracts of posters. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(3):22–27, September 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:2000:SONa

[Ano00c] Anonymous. Southern Ontario Numerical Analysis Day: May 12, 2000, Ontario Research Centre for Computer Algebra, University of Western Ontario, London, Ontario, Canada: Abstracts of invited talks. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(3):28, September 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:2000:SONb

[Ano00d] Anonymous. Southern Ontario Numerical Analysis Day: May 12, 2000, Ontario Research Centre for Computer Algebra, University of Western Ontario, London, Ontario, Canada: Abstracts of invited talks. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(3):29–31, September 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:2001:AAA

[Ano01a] Anonymous. Announcements: AAECC-14 Applied Algebra, Algebraic Algorithms and Error Correcting Codes. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35 (2):35, June 2001. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:2001:CPL

[Ano01b]  Anonymous.  Call for

papers: Latin American Theoretical INformatics — LATIN ’2002. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35(2):36, June 2001. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:2001:ECC

[Ano01c] Anonymous. East Coast Computer Algebra Day 2002: Announcement and call for participation. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35(4):40, December 2001. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:2003:CPA

[Ano03] Anonymous. Call for participation: ASCM 2003: The Sixth Asian Symposium on Computer Mathematics, October 23–25, 2003 Beijing, China. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37 (2):64, June 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:2008:ISS

[Ano08a] Anonymous. 10th International Symposium on Symbolic and Numeric Algorithms for Scientific Computing, Timisoara, Romania, September 26–29, 2008. ACM Communications in Computer Algebra, 42(1–2): 104, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2008:MEP

[Ano08b] Anonymous. In memoriam Eugeny Pankratiev: Faculty of Mechanics and Mathematics, Moscow State University, Moscow, Russia. ACM Communications in Computer Algebra, 42(1–2):23–26, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2009:CA

[Ano09] Anonymous. Conference announcements. ACM Communications in Computer Algebra, 43(1):17–20, March 2009. CODEN ????

ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2010:AIT

[Ano10a] Anonymous. Abstracts of invited talks, contributed posters, and software demos at the East Coast Computer Algebra Day 2010. ACM Communications in Computer Algebra, 44(2): ??, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2010:ARDa

[Ano10b] Anonymous. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 44(2): ??, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2010:ARDb

[Ano10c] Anonymous. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 44(2): ??, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2010:ETP

[Ano10d]  Anonymous.  Emerging

trends papers accepted for PLMMS 2010. ACM Communications in Computer Algebra, 44(2):??, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2010:MCE

[Ano10e] Anonymous. In memoriam: Chionh Eng Wee. ACM Communications in Computer Algebra, 44(2): 18–19, June 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2010:INM

[Ano10f]  Anonymous.  Issue 171

number 1, March 2010. ACM Communications in Computer Algebra, 44(2): ??, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2010:INJ

[Ano10g] Anonymous. Issue 172 number 2, June 2010. ACM Communications in Computer Algebra, 44(2):??, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2012:SD

[Ano12] Anonymous. Stefan Dodunekov 1945–2012. ACM Communications in Computer Algebra, 46(3–4):70– 71, September 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2013:ICB

[Ano13a] Anonymous. ISSAC 2015: call for bids. ACM Communications in Computer Algebra, 47(1–2):17, March 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2013:MGA

[Ano13b] Anonymous. Mark Giesbrecht ACM distinguished scientist. ACM Communications in Computer Algebra, 47(3–4):188, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2013:RUEa

[Ano13c] Anonymous. Recent and upcoming events. ACM Communications in Computer Algebra, 47(1–2):24– 25, March 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2013:RUEb

[Ano13d] Anonymous. Recent and upcoming events. ACM Communications in Computer Algebra, 47(1–2):71– 72, March 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2013:RDJ

[Ano13e] Anonymous. Richard D. Jenks prize 2013: call for nominations. ACM Communications in Computer Algebra, 47(1–2):18, March 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2014:AEC

[Ano14a] Anonymous. Abstracts of the 2014 East Coast Computer Algebra Day: Communicated by Xiaobai Sun, Nikos P. Pitsianis, and Alexandros-Stavros Iliopoulos. ACM Communications in Computer Algebra, 48(2): 59–63, March 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2014:MSCa

[Ano14b] Anonymous. Message from the SIGSAM chair. ACM Communications in Computer Algebra, 48(1):1–2, March 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2014:MSCb

[Ano14c] Anonymous. Message from the SIGSAM chair. ACM Communications in Computer Algebra, 48(3/4):77, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2016:JMB

[Ano16] Anonymous. Jonathan M.

Borwein, former CMS President, dies at 65. ACM Communications in Computer Algebra, 50(3):121– 122, September 2016. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2017:RDJ

[Ano17a]  Anonymous.  The 2017

Richard D. Jenks Memorial Award. ACM Communications in Computer Algebra, 51(4):111, December 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2017:ARD

[Ano17b] Anonymous. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 51(2): 70–72, June 2017. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2017:DWE

[Ano17c]  Anonymous.  Dongming

Wang elected member of the Academia Europaea. ACM Communications in Computer Algebra, 51(4): 112, December 2017. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anonymous:2017:GC

[Ano17d] Anonymous. George E. Collins (1928–2017). ACM Communications in Computer Algebra, 51(4):113, December 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Andrews:2008:FDM

[AP08] George E. Andrews and Peter Paule. Further developments in MacMahon’s partition analysis (abstract only). ACM Communications in Computer Algebra, 42(1–2):18, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Abramov:2023:LDO

[AP23] Sergei Abramov and Gleb Pogudin. Linear difference operators with sequence coefficients having infinitedimentional solution spaces. ACM Communications in Computer Algebra, 57(1): 1–4, March 2023. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3610377.3610378.

Aliasgari:2015:BIA

[APS15] Malihe Aliasgari, Daniel Panario, and Mohammad-Reza Sadeghi. Binomial ideal associated to a lattice and its label code. ACM Communications in Computer Algebra, 49(1): 16–17, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Arasu:2015:SCP

[Ara15]  K. T. Arasu.  Stickel-

berger’s congruences and perfect sequence constructions. ACM Communications in Computer Algebra, 49(1):17, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Arnon:1981:AAR

[Arn81] Dennis S. Arnon. Automatic analysis of real algebraic curves. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15 (4):3–9, November 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Arnon:1987:RWE

[Arn87] Dennis S. Arnon. Report of the workshop on environments for computational mathematics. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21(4):42–48, November 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Arnon:2010:PIM

[Arn10] Dennis S. Arnon. A product information management system as a workbench for quantifier elimination computations with QEPCAD. ACM Communications in Computer Algebra, 44(3):87–88, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Arri:2008:FCR

[Arr08]  Alberto Arri.  The F5

criterion revised. ACM Communications in Computer Algebra, 42(3):169– 170, September 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Arreche:2015:CDG

[Arr15] Carlos E. Arreche. Computing differential Galois groups of parameterized second-order linear differential equations. ACM Communications in Computer Algebra, 49(1):23, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Abramov:2017:SC

[ARS17] Sergei A. Abramov, Anna A.

Ryabenko, and Timur M. Sadykov. The second conference. ACM Communications in Computer Algebra, 51(4):103–110, December 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Abramov:2019:CCA

[ARS^{+}19] Sergei A. Abramov, Anna Ryabenko, Leonid A. Sevastianov, Min Wu, and Yvette Zonn. The conference “Computer Algebra” in Moscow. ACM Communications in Computer Algebra, 53(2):65–77, June 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Abramov:2021:FCC

[ARS^{+}21] Sergei A. Abramov, Anna Ryabenko, Leonid A. Sevastianov, Yvette Zonn, and Min Wu. The fourth conference “computer algebra” in Moscow. ACM Communications in Computer Algebra, 55(2):30–38, June 2021. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3493492.3493494.

Alvarez:1999:SRP

[ÁS99] L. Álvarez and J. Sánchez. On some real problems in computer vision which yield to algebraic system of equations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):13–14, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Arnold:2014:TFT

[AS14] Andrew Arnold and Éric Schost. A Truncated Fourier Transform middle product. ACM Communications in Computer Algebra, 48(3/4):98–99, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Abramov:2016:CCA

[AS16] Sergei A. Abramov and Leonid A. Sevastianov. The conference “computer algebra” in Moscow. ACM Communications in Computer Algebra, 50(2):61–68, June 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Aslaksen:1996:MVC

[Asl96] Helmer Aslaksen. Multiplevalued complex functions and computer algebra. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(2):12–20, June 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Aliasgari:2015:JTB

[ASP15] Malihe Aliasgari, Mohammad-Reza Sadeghi, and Daniel Panario. A Johnson-type bound for group codes and lattices. ACM Communications in Computer Algebra, 49(1):16, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Antweiler:1989:TRI

[ASW89] Werner Antweiler, Andreas Strotmann, and Volker Winkelmann. A TEX-REDUCE interface. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23(2):26–33, April 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Al-shomrani:2008:DWB

[AsW08] Saleh Al-shomrani and Paul Wang. DMAS: a Web-based distributed mathematics assessment system (abstract only). ACM Communications in Computer Algebra, 42(1–2):67–68, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Atkinson:1981:SSC

[Atk81]  M. D. Atkinson.  Sav-

ing space in coset enumeration. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15 (4):12–14, November 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Andradas:1999:CPO

[AV99] C. Andradas and M. P.

Vélez. Convex polytopes over ordered fields. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):14, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Aldossari:2019:MSD

[AvH19] Shayea Aldossari and Mark van Hoeij. Methods for simplifying differential equations. ACM Communications in Computer Algebra, 53(3):93–95, September 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Anderson:1980:MEM

[AW80] Ned Anderson and Paul S.

Wang. MACSYMA experiments with a modified Khachian-type algorithm in linear programming. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14(1):8–13, February 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Aichinger:2016:SGT

[AY16] Erhard Aichinger and Rika Yatchak. Sonata: a GAP tool for nearring computations. ACM Communications in Computer Algebra, 50(3):89–92, September 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Aydin:2015:CAC

[Ayd15] Nuh Aydin. Computer algebra challenges for constructing skew cyclic codes. ACM Communications in Computer Algebra, 49(1): 13, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Almkvist:1991:MPF

[AZ91] Gert Almkvist and Doron Zeilberger. A Maple program that finds, and proves, recurrences and differential equations satisfied by hyperexponential definite integrals. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25 (3):14–17, July 1991. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ba:2012:MBR

[Ba12] Thang Luu Ba. The matrix based representations of the intersection curves. ACM Communications in Computer Algebra, 46(3–4):114– 115, September 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Belhaj:2021:AGC

[BA21] Skander Belhaj and Abdulrahman Alsulami. Approximate greatest common divisor of several polynomials from Hankel matrices. ACM Communications in Computer Algebra, 55(3):77–81, September 2021. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3511528.3511532.

Bachmann:1997:MEF

[Bac97] Olaf Bachmann. MPCR:

An efficient and flexible chains of recurrences server. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(1):15–21, March 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bahr:1969:SFF

[Bah69]  K. Bahr.  A smaller

FORTRAN–FORMAC system. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (12):59–60, July 1969. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bahr:1971:ASP

[Bah71] Knut Bahr. An algebraic simplification program in PL/I. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (17):43–50, January 1971. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bahr:1973:IBF

[Bah73a] Knut Bahr. An interesting bug in FORMAC. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(25):23–24, March 1973. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bahr:1973:SF

[Bah73b] Knut A. Bahr. A speedup of FORMAC. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(27):26–29, September 1973. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bahr:1974:TRF

[Bah74] Knut A. Bahr. Toward a revision of FORMAC. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(1):10–16, February 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bahr:1975:SP

[Bah75a] Knut Bahr. Solution to problem 8. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9 (2):6, May 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bahr:1975:UFN

[Bah75b] Knut Bahr. Utilizing the FORMAC novelties. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(1):21–24, February 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bahr:1977:NRF

[Bah77] K. A. Bahr. New release of FORMAC 73 available. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11–12(4–1):6, November and February 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Issue labeled vol. 11, no. 4, November 1977 (issue #44) and vol. 12, no. 1, February 1978 (issue #45).

Bainov:2002:TIC

[Bai02]  Drumi Bainov.  Thir-

teenth International Colloquium on Differential Equations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36 (1):32, March 2002. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Baker:1993:SPL

[Bak93] Henry G. Baker. Sparse polynomials and linear logic. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27 (4):10–14, December 1993. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Balaban:1984:NNN

[Bal84] Alexandru T. Balaban. Numerical and non-numerical methods in chemistry: present and future. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(2):29–30, May 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Banshchikov:2008:ADL

[Ban08] Andrey V. Banshchikov. Analysis of dynamics of large-dimensional mechanical systems by the tools of computer algebra. ACM Communications in Computer Algebra, 42(3):166, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Barton:1973:PLT

[Bar73] David Barton. Problem #4:

the Lie transform. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(25):12–13, March 1973. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Barnett:1989:UPF

[Bar89]  M. P. Barnett.  Using

partial fraction formulas to sum some slowly convergent series analytically for molecular integral calculations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23 (3):13–18, July 1989. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Barnett:1991:SSW

[Bar91] Michael P. Barnett. Some simple ways to construct and to use formulas mechanically. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25 (2):28–36, April 1991. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Barroso:1999:PIT

[Bar99] E. García Barroso. Polar invariants and topology. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):16, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Barnett:2002:CAL

[Bar02] Michael P. Barnett. Computer algebra in the life sciences. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36 (4):5–32, December 2002. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Barnett:2003:SCI

[Bar03] Michael P. Barnett. Symbolic computation of integrals by recurrence. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(2):49–63, June 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Barnett:2008:RSC

[Bar08] Michael P. Barnett. Reasoning in symbolic computation. ACM Communications in Computer Algebra, 42(1–2):1–17, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bard:2010:AOD

[Bar10a] Gregory V. Bard. Abstract only: DEMOCRACY: a heuristic for polynomial systems of equations over finite fields. ACM Communications in Computer Algebra, 44(2):25, June 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bardavid:2010:DSG

[Bar10b] Colas Bardavid. Differential schemes: geometric approach and functorial approach. ACM Communications in Computer Algebra, 44(3):157–159, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Barkatou:2015:DAC

[Bar15] Moulay A. Barkatou. A direct algorithm for computing k-simple forms of firstorder linear differential systems. ACM Communications in Computer Algebra, 49(1):23, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bates:2013:NSP

[Bat13] Daniel J. (Daniel James)

Bates. Numerically solving polynomial systems with Bertini. Software, environments, and tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2013. ISBN 1-61197-269-8. xx + 352 pp. LCCN QA161.P59 B38 2013. URL http:

//www.loc.gov/catdir/
enhancements/fy1403/2013025584-
b.html;   http://www.
loc.gov/catdir/enhancements/
fy1403/2013025584-d.html;
 http://www.loc.gov/catdir/
enhancements/fy1403/2013025584-
t.html.

Bachmair:1980:SPC

[BB80] L. Bachmair and B. Buchberger. A simplified proof of the characterization theorem for Gröbner-bases. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14(4):29–34, November 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Borwein:1996:SSA

[BB96] Jonathan Borwein and David Bradley. Searching symbolically for Apérylike formulae for values of the Riemann zeta function. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(2):2– 7, June 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (elec-

tronic). URL http://doi.
acm.org/10.1145/235699.
235700.

Babich:2002:WDE

[BBS^{+}02] V. M. Babich, Yu. N.

Bibikov, S. Yu. Slavyanov, N. V. Svanidze, T. F. Pankratova, and V. Novikova. Workshop on Differential Equations: Dedicated to the memory of Vladimir Lazutkin, St.Petersburg, Russia, August 18–20, 2002. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36(1):18, March 2002. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bordoni:1981:ARI

[BC81] Luciana Bordoni and Attilio Colagrossi. An application of REDUCE to industrial mechanics. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15(2):8–12, May 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Benghorbal:2002:D

[BC02] Mhenni Benghorbal and Robert M. Corless. The n th derivative. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36(1):10–14, March 2002. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bostan:2010:RFT

[BCCL10] Alin Bostan, Shaoshi Chen, Frédéric Chyzak, and Ziming Li. Rational-functions telescopers: blending creative telescoping with Hermite reduction. ACM Communications in Computer Algebra, 44(1):9–10, March 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bossi:1980:ICM

[BCCR80] A. Bossi, L. Colussi, C. Cobelli, and G. Romanin Jacur. Identifiability of compartmental models: algorithms to solve an actual problem by means of symbolic calculus. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14(4):35–40, November 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bihan:2023:QCI

[BCD^{+}23] Frédéric Bihan, Erika Croy, Weixun Deng, Kaitlyn Phillipson, Robert J. Rennie, and J. Maurice Rojas. Quickly computing isotopy type for exponential sums over circuits (extended abstract). ACM Communications in Computer Algebra, 57(3):152– 155, September 2023. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3637529.3637538.

Bertone:2009:PAP

[BCG09] C. Bertone, G. Chéze, and A. Galligo. Probabilistic algorithms for polynomial absolute factorization. ACM Communications in Computer Algebra, 43(3):77–78, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Berardini:2023:PBN

[BCL23] Elena Berardini, Alain Couvreur, and Grégoire Lecerf. A proof of the Brill–Noether method from scratch. ACM Communications in Computer Algebra, 57(4):200– 229, December 2023. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3653002.3653004.

Buchberger:1982:CAS

[BCLA82] B. Buchberger, G. E.

Collins, R. Loos, and R. Albrecht. Computer algebra symbolic and algebraic computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16 (4):5, November 1982. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bostan:2011:FCC

[BCLS11] Alin Bostan, Frédéric Chyzak, Ziming Li, and Bruno Salvy. Fast computation of common left multiples of linear ordinary differential operators. ACM Communications in Computer Algebra, 45(2):111–112, June 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Boyer:2014:MMW

[BD14] Brice Boyer and Jean-Guillaume Dumas. Matrix multiplication over wordsize modular fields using Bini’s approximate formula. ACM Communications in Computer Algebra, 48(3/4): 100–102, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bakshi:2015:CDS

[BD15] Ankita Bakshi and Deeksha. On a class of difference set pairs. ACM Communications in Computer Algebra, 49(1):17, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Brieulle:2018:CIE

[BDD^{+}18] Ludovic Brieulle, Luca De Feo, Javad Doliskani, Jean Pierre Flori, and Éric Schost. Computing isomorphisms and embeddings of finite fields. ACM Communications in Computer Algebra, 52(4):117–119, December 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bradford:2021:DPP

[BDE^{+}21] R. Bradford, J. H. Davenport, M. England, A. Sadeghimanesh, and A. Uncu. The DEWCAD project: pushing back the doubly exponential wall of cylindrical algebraic decomposition. ACM Communications in Computer Algebra, 55(3):107– 111, September 2021. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3511528.3511538.

Bui:2013:SFQ

[BDM13] C. Bui, G. H. E. Duchamp, and Hoang Ngoc Minh. Schützenberger’s factorization on q-stuffle Hopf algebra. ACM Communications in Computer Algebra, 47(3–4):90–91, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Beaudin:2015:SCN

[Bea15]  Michel Beaudin.  Same

courses, new questions. ACM Communications in Computer Algebra, 49(2): 48, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Beebe:1998:LBA

[Bee98] Nelson H. F. Beebe. Letters: Bibliography archive. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(1):1–3, March 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Belovari:1983:CAS

[Bel83] G. Belovari. Complex analysis in symbolic computing of some definite integrals. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17 (2):6–11, May 1983. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Benghorbal:2008:UFA

[Ben08a] Mhenni M. Benghorbal. Unified formulas for arbitrary order symbolic derivatives and anti-derivatives of the power-inverse hyperbolic class 1. ACM Communications in Computer Algebra, 42(3):156, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Benkiran:2008:BFP

[Ben08b] Amine Benkiran. Brouwer’s fixed point theorem (abstract only). ACM Communications in Computer Algebra, 42(1–2):68–69, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bernstein:1970:LPF

[Ber70] Herbert J. Bernstein. Ludeman’s PL/I FORMAC gradient problem as handled by FORTRAN SYMBOL-ANG. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (14):19–34, January 1970. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Berry:1983:CA

[Ber83] T. G. Berry. On Coates algorithm. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17 (2):12–17, May 1983. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bernardin:1996:RSS

[Ber96] Laurent Bernardin. A review of symbolic solvers.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(1):9–20, March 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Berggren:1998:PSF

[Ber98a] L. Berggren. Position at Simon Fraser University. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(2):2, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Berkovich:1998:FSC

[Ber98b] L. M. Berkovich. Factorization of some classes of nonlinear ordinary differential equations: Methods and algorithms. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (2):52, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Abstract only.

Berman:1999:CGG

[Ber99a]  Peter Berman.  Com-

puting the Galois group of y^{(3)} + ay′ + by = 0, a, b ∈ C[x]. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):20, September 1999.

CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bermejo:1999:AFS

[Ber99b] Isabel Bermejo. Abstracts of the fifth Spanish meeting on computer algebra and applications EACA-99. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):12–19, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bernardin:1999:JFM

[Ber99c]  Laurent Bernardin.  A

Java framework for massively distributed symbolic computing. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):20–21, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Berkovich:2006:FSC

[Ber06] L. M. Berkovich. Factorization of self-conjugated and reducible linear differential operators. ACM Communications in Computer Algebra, 40(2):35–37, June 2006. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Beshaj:2015:HHS

[Bes15] L. Beshaj. Heights on hyperelliptic and superelliptic curves. ACM Communications in Computer Algebra, 49(2):58, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Betten:2013:CDO

[Bet13] Anton Betten. Classifying discrete objects with Orbiter. ACM Communications in Computer Algebra, 47(3–4):183–186, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Barton:1974:AMC

[BF74] D. Barton and J. P. Fitch. Algebraic manipulation in Cambridge. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(1): 8–9, February 1974. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Berthomieu:2020:TFO

[BFE20] Jérémy Berthomieu, Andrew Ferguson, and Mohab Safey El Din. Towards fast one-block quantifier elimination through generalised critical values. ACM Communications in Computer Algebra, 54(3):109– 113, September 2020. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3457341.3457348.

Boulier:2018:SPS

[BFR^{+}18] François Boulier, François Fages, Ovidiu Radulescu, Satya Swarup Samal, Andreas Schuppert, Werner Seiler, Thomas Sturm, Sebastian Walcher, and Andreas Weber. The SYM-BIONT project: symbolic methods for biological networks. ACM Communications in Computer Algebra, 52(3):67–70, September 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Blumlein:2023:CMR

[BFS23] Johannes Blümlein, Nikolai Fadeev, and Carsten Schneider. Computing Mellin representations and asymptotics of nested binomial sums in a symbolic way: The RICA package. ACM Communications in Computer Algebra, 57(2): 31–34, June 2023. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3614408.3614410.

Baladi:1990:PCP

[BG90] V. Baladi and J. P. Guillement. A program for computing Puiseux expansions. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 24(4):33–41, October 1990. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bachmann:1997:DEM

[BG97] Olaf Bachmann and Simon Gray. A demonstration of exchanging mathematical expressions using MP. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):50, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Bennenni:2015:CCD

[BGG15] Nabil Bennenni, Kenza Guenda, and Aaron Gulliver. Construction of codes for DNA computing by the greedy algorithm. ACM Communications in Computer Algebra, 49(1): 14, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Blair:1970:IFS

[BGJ70] F. W. Blair, J. H. Griesmer, and R. D. Jenks. An interactive facility for symbolic mathematics. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(14):17–18, January 1970. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Boady:2011:SCS

[BGJ11] Mark Boady, Pavel Grinfeld, and Jeremy Johnson. A symbolic computation system for the calculus of moving surfaces. ACM Communications in Computer Algebra, 45(2): 109–110, June 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bermejo:2012:ML

[BGLG12] I. Bermejo, E. García-Llorente, and Ph. Gimenez.

monomialideal.lib. ACM

Communications in Computer Algebra, 46(3–4):160– 161, September 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bermejo:2018:NRD

[BGLGMM18] I. Bermejo, E. García-Llorente, I. García-Marco, and M. Morales. Noether resolutions in dimension 2. ACM Communications in Computer Algebra, 52(4): 114–116, December 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bachler:2010:TDA

[BGLH^{+}10] Thomas Bächler, Vladimir Gerdt, Markus Lange-Hegermann, Wilhelm Plesken, and Daniel Robertz. Thomas decomposition and its applications. ACM Communications in Computer Algebra, 44(3):91–92, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Berthomieu:2023:MMM

[BGLM23] Jérémy Berthomieu, Stef Graillat, Dimitri Lesnoff, and Theo Mary. Modular matrix multiplication on GPU for polynomial system solving. ACM Communications in Computer Algebra, 57(2):35–38, June 2023. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3614408.3614411.

Bermejo:1999:CMR

[BGM99] I. Bermejo, Ph. Gimenez, and M. Morales. On Castelnuovo–Mumford regularity of codimension two monomial varieties. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):14, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bermejo:2010:ACW

[BGM10] Isabel Bermejo and Ignacio García-Marco. An algorithm for checking whether a simplicial toric ideal is a complete intersection. ACM Communications in Computer Algebra, 44(3):93–94, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Barton:1972:CPF

[BH72a] David Barton and Anthony C. Hearn. Comment on problem #2 — the Y_{2n} functions. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15 (??):??, ???? 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Brown:1972:A

[BH72b] W. S. Brown and A. D.

Hall. ALTRAN. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(24):4–7, October 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bondari:1997:MRI

[BH97] Siamack Bondari and Irvin Roy Hentzel. The McCrimmon radical for identities of degree 3. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (1):2–5, March 1997. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Beaudin:2015:PFC

[BH15] Michel Beaudin and Frédéric Henri. Piecewise functions and convolution integrals (part I, part II). ACM Communications in Computer Algebra, 49(1):34, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Belk:2013:ISC

[BHMM13] James Belk, Nabil Hossain, Francesco Matucci, and Robert McGrail. Implementation of a solution to the conjugacy problem in Thompson’s group F. ACM Communications in Computer Algebra, 47(3–4):120– 121, September 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bulter:1991:TDD

[BI91a] Greg Bulter and Sridhar S. Iyer. Towards a deductive database for small simple groups. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25 (4):7–17, October 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Butler:1991:TDD

[BI91b] G. Butler and S. S. Iyer. Towards a deductive database for small simple groups. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(4):7–17, October 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bibel:1984:LAC

[Bib84]  Wolfgang Bibel.  Logic

and algebraic computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(2):26, May 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Binger:1987:BT

[Bin87a]  D. Binger.  Burnside’s

theorem. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21 (3):19–23, August 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Binger:1987:ON

[Bin87b] D. Binger. Optimization on networks. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21 (3):54–60, August 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bostan:2006:SSL

[BJS06] Alin Bostan, Claude-Pierre Jeannerod, and Éric Schost. Solving structured linear systems of large displacement rank. ACM Communications in Computer Algebra, 40(2):42–44, June 2006. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ballarin:2004:SPL

[BK04] Clemens Ballarin and Manuel Kauers. Solving parametric linear systems: an experiment with constraint algebraic programming. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(2):33–46, June 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Belhaj:2012:CCG

[BK12] Skander Belhaj and Haïthem Ben Kahla. On the complexity of computing the GCD of two polynomials via Hankel matrices. ACM Communications in Computer Algebra, 46(3–4):74–75, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Boyer:2014:NLS

[BK14] Brice Boyer and Erich L.

Kaltofen. Numerical linear system solving with parametric entries by error correction. ACM Communications in Computer Algebra, 48(3/4):103–105, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bright:2018:SPW

[BKG18] Curtis Bright, Ilias Kotsireas, and Vijay Ganesh. The SAT+CAS paradigm and the Williamson conjecture. ACM Communications in Computer Algebra, 52(3):82–84, September 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Belov:2005:MMC

[BKKS05] Nadya Belov, Colin Koeck, Werner Krandick, and Joshua Shaffer. Mobile mathematics communication. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39 (3):99, 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). ISSAC 2005 poster abstract.

Borwein:2007:IPB

[BKM07] Peter Borwein, Erich Kaltofen, and Michael J. Mossinghoff. Irreducible polynomials and Barker sequences. ACM Communications in Computer Algebra, 41(4): 118–121, December 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Broughan:1991:SSC

[BKR^{+}91] K. A. Broughan, G. Keady, T. D. Robb, M. G. Richardson, and M. C. Dewar. Some symbolic computing links to the NAG numeric library. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25 (3):28–37, July 1991. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Brown:2023:FRT

[BKR23] Christopher Brown, Zoltán Kovács, and Tomás Recio. Faithful real-time animation of parametrized (semi- ) algebraic expressions via cylindrical algebraic decomposition. ACM Communications in Computer Algebra, 57(2):43–46, June 2023. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3614408.3614413.

Blumlein:2010:SSA

[BKSS10] Johannes Blümlein, Sebastian Klein, Carsten Schneider, and Flavia Stan. A symbolic summation approach to Feynman integrals. ACM Communications in Computer Algebra, 44(3):95–96, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Blankertz:2014:PTA

[Bla14] Raoul Blankertz. A polynomial time algorithm for computing all minimal decompositions of a polynomial. ACM Communications in Computer Algebra, 48(1):13–23, March 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Barakat:2010:LLC

[BLH10] Mohamed Barakat and Markus Lange-Hegermann. LocalizeRingForHomalg: localize commutative rings at maximal ideals. ACM Communications in Computer Algebra, 44(4):201– 204, December 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Billey:2008:ACP

[BM08] Sara C. Billey and Stephen A.

Mitchell. Affine colored partitions (abstract only). ACM Communications in Computer Algebra, 42(1–2): 19, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bard:2015:PRO

[BM15] Gregory V. Bard and Theodore McDonnough. Plaintext recovery for onetime pads used twice. ACM Communications in Computer Algebra, 49(1):17–18, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bouchon-Meunier:1998:ICM

[BMK98] Bernadette Bouchon-Meunier and Vladik Kreinovich. From interval computations to modal mathematics: Applications and computational complexity. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(2):7–11, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Barile:1999:EDM

[BMT99] M. Barile, M. Morales, and A. Thoma. On equations defining monomial varieties. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):14, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bates:2014:DHS

[BNN14] Daniel J. Bates, Andrew J.

Newell, and Matthew E. Niemerg. Decoupling highly structured polynomial systems. ACM Communications in Computer Algebra, 48(3/4):133, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Borwein:1998:MPW

[Bor98] Jonathan Borwein. Mathematical publication on the Web. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (1):4–7, March 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bose:1984:SAC

[Bos84] N. K. Bose. Symbolic and algebraic computations in multidimensional systems theory. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18 (2):31–32, May 1984. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bourne:1972:C

[Bou72] S. R. Bourne. CAMAL.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(24):8–11, October 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bouzidi:2015:CSL

[Bou15] Yacine Bouzidi. Computing separating linear forms for bivariate systems. ACM Communications in Computer Algebra, 49(2):51–52, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Budroni:2018:HGB

[BP18] Alessandro Budroni and Federico Pintore. Hashing to G2 on BLS pairingfriendly curves. ACM Communications in Computer Algebra, 52(3):63–66, September 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bychkov:2020:OMQ

[BP20] Andrey Bychkov and Gleb Pogudin. Optimal monomial quadratization for ODE systems: extended abstract. ACM Communications in Computer Algebra, 54(3):119–123, September 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3457341.3457350.

Barkatou:2012:IMP

[BPS12] Moulay A. Barkatou, Eckhard Pflügel, and Flavia Stan. ISOLDE: a Maple package for systems of linear functional equations. ACM Communications in Computer Algebra, 46(3–4): 157–159, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Braham:1998:NON

[Bra98] Stephen P. Braham. NAOMI One: North American OpenMath Initiative goes online. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (2):23–29, June 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bremner:2005:JAA

[Bre05] Murray R. Bremner. Jordan algebras arising from intermolecular recombination. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(4):106–117, December 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Brown:1974:CFR

[Bro74] W. S. Brown. On computing with factored rational expressions. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8 (3):26–34, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bronstein:1993:IPI

[Bro93] Manuel Bronstein, editor. ISSAC’93: proceedings of the 1993 International Symposium on Symbolic and Algebraic Computation, July 6–8, 1993, Kiev, Ukraine. ACM Press, New York, NY, USA, 1993. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. ACM order number: 505930.

Brown:2003:QBP

[Bro03] Christopher W. Brown. QEPCAD B: a program for computing with semialgebraic sets using CADs. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(4):97–108, December 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Brown:2004:QBS

[Bro04] Christopher W. Brown. QEPCAD B: a system for computing with semialgebraic sets via cylindrical algebraic decomposition. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38 (1):23–24, March 2004. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Broughan:2009:AEE

[Bro09] Kevin A. Broughan. An algorithm for the explicit evaluation of GL(nR) Kloosterman sums. ACM Communications in Computer Algebra, 43(1):1–10, March 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bronstein:2013:MBC

[Bro13] Manuel Bronstein. Manuel Bronstein’s codes. ACM Communications in Computer Algebra, 47(1–2):16, March 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Brown:2017:MSC

[Bro17] Christopher W. Brown.

Message from the SIGSAM chair. ACM Communications in Computer Algebra, 51(4):101–102, December 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Buchberger:2008:GPR

[BRRT08] Bruno Buchberger, Georg Regensburger, Markus Rosenkranz, and Loredana Tec. General polynomial reduction with TH 9 OREM 8 functors: applications to integro-differential operators and polynomials. ACM Communications in Computer Algebra, 42(3):135– 137, September 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bahr:1974:TAM

[BS74] Knut Bahr and Jaap Smit. Tuning an algebraic manipulation system through measurements. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(3):17–23, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bagheri:2015:NNA

[BS15] Kadijeh Bagheri and Mohammad-Reza Sadeghi. A new non-associative cryptosystem based on NTOW public key cryptosystem and octonions algebra. ACM Communications in Computer Algebra, 49(1):13, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bhayani:2023:NAP

[BSB^{+}23] Snehal Bhayani, Praneeth Susarla, S. S. Krishna Chaitanya Bulusu, Olli Silven, Markku Juntti, and Janne Heikkila. A novel application of polynomial solvers in mmWave analog radio beamforming. ACM Communications in Computer Algebra, 57(3):148– 151, September 2023. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3637529.3637537.

Bernal:2015:NMB

[BSBC15] José Joaquín Bernal, Juan Jacobo Simón, and Diana H. Bueno-Carreño. A notion of multivariate BCH bounds and codes. ACM Communications in Computer Algebra, 49(1):18, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bshouty:2012:T

[Bsh12] Nader H. Bshouty. Testers.

ACM Communications in Computer Algebra, 46(3– 4):78–79, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Beshaj:2015:JCS

[BSS15] L. Beshaj, T. Shaska, and C. Shor. On Jacobians of curves with superelliptic components. ACM Communications in Computer Algebra, 49(2):52, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bueso:1999:WFG

[BTL99] J. L. Bueso, J. Gómez Torrecillas, and F. J. Lobillo. When is a finitely generated algebra of Poincaré– Birkhoff–Witt type? SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):14–15, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Beutelspacher:1993:SIGa

[BU93a] Albrecht Beutelspacher and Johannes Ueberberg. Symbolic Incidence Geometry: Proposal for doing geometry with a computer (part 1 of 2). SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27 (2):19–29, April 1993. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Beutelspacher:1993:SIGb

[BU93b] Albrecht Beutelspacher and Johannes Ueberberg. Symbolic Incidence Geometry: Proposal for doing geometry with a computer (part 2 of 2). SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27 (3):9–24, September 1993. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Buchberger:1976:SPG

[Buc76a] B. Buchberger. Some properties of Gröbner-bases for polynomial ideals. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 10(4):19–24, November 1976. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Buchberger:1976:TBR

[Buc76b] B. Buchberger. A theoretical basis for the reduction of polynomials to canonical forms. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 10 (3):19–29, August 1976. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Buchberger:1984:CTP

[Buc84] B. Buchberger. CAMP: a teaching project in symbolic computation at the University of Linz. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18–19(4–1):8–9, November/February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Buchberger:1990:SSL

[Buc90] Bruno Buchberger. Should students learn integration rules? SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 24 (1):10–17, January 1990. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Buchanan:1991:STP

[Buc91] S. Alasdair Buchanan. Some theoretical problems when solving systems of polynomial equations using Gröbner bases. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(2):24–27, April 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Buchberger:2002:CAE

[Buc02] Bruno Buchberger. Computer algebra: The end of mathematics? SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36(1):3–9, March 2002. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Buchel:2003:RDB

[Buc03] Alex Buchel. Review of D-Branes by Clifford V. Johnson. Cambridge University Press 2002. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(4):125–127, December 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bulygin:2010:AOP

[Bul10] Stanislav Bulygin. Abstract only: Polynomial system solving for decoding linear codes and algebraic cryptanalysis parametric polynomial system discussion: canonical comprehensive. ACM Communications in Computer Algebra, 44(2): 72, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bundschuh:1983:GDF

[Bun83] Florian Bundschuh. Global data flow analysis in Aldes. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(1):6–11, February 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Burkhart:2008:MPD

[Bur08] Brandon Burkhart. Models of population dynamics (abstract only). ACM Communications in Computer Algebra, 42(1–2):69, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Butler:1985:ICA

[But85] Gregory Butler. An improvement to the centralizer algorithm for permutation groups. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19 (2):14–18, May 1985. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Butler:1992:AAA

[But92] Greg Butler. An analysis of Atkinson’s algorithms. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26(2):1–9, April 1992. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bauschke:2006:SCF

[BvM06] Heinz H. Bauschke and Martin v. Mohrenschildt. Symbolic computation of Fenchel conjugates. ACM Communications in Computer Algebra, 40(1):18–28, March 2006. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Bonorden:2001:FBP

[BvzGG^{+}01] Olaf Bonorden, Joachim von zur Gathen, Jürgen Gerhard, Olaf Müller, and Michael Nöcker. Factoring a binary polynomial of degree over one million. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35(1):16–18, March 2001. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Buchberger:1998:ICG

[BW98a] B. Buchberger and F. Winkler. Intensive course on Gröbner bases. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(2):38, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Buchberger:1998:YGB

[BW98b] Bruno Buchberger and Franz Winkler, editors. 33 years of Gröbner bases: Gröbner bases and applications: Conference — February 1998, Linz, Austria, number 251 in London Mathematical Society Lecture Note Series. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-63298-6. LCCN QA251.3.G76 1998.

Braun:2002:RWO

[BY02] W. John Braun and Hao Yu. Review: Weighing the Odds, by D. Williams. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36(1):15–17, March 2002. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Breuer:2012:PAP

[BZ12] Felix Breuer and Zafeirakis Zafeirakopoulos. Partition analysis via polyhedral geometry. ACM Communications in Computer Algebra, 46(3–4):76–77, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Cohen:2022:JPA

[CAB22] Henri Cohen, Bill Allombert, and Karim Belabas. Jenks Prize announcement. ACM Communications in Computer Algebra, 56(2):92–94, June 2022. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3572867.3572884.

Calmet:1974:CRD

[Cal74] Jacques Calmet. Computer recognition of divergences in Feynman diagrams. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(3):74–75, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Calmet:1984:ICA

[Cal84] Jacques Calmet. Introducing computer algebra to users and to students. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18–19(4–1):10–14, November/February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Calmet:2001:AWU

[Cal01] J. Calmet. Announcement:

Workshop on Under- and Over-Determined Systems of Algebraic or Differential Equations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35 (3):39, September 2001. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Campbell:1972:PF

[Cam72] J. A. Campbell. Problem #2 — the Y_{2n} functions. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(22):8– 9, March 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Campbell:1973:PPD

[Cam73] J. A. Campbell. Problem #5: a problem in differentiation arising from the Runge–Kutta–Nyström method. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (25):13–14, March 1973. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Campbell:1975:APL

[Cam75] J. A. Campbell. Automatic programming in LISP. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(2):11–13, May 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Campbell:1976:CSC

[Cam76] J. A. Campbell. Compact storage for computations involving partitions. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 10(4):46–47, November 1976. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Campbell:1988:SIE

[Cam88] Graham H. Campbell. Symbolic integration of expressions involving unspecified functions. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22 (1):25–27, January 1988.

CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Cannon:1969:CAP

[Can69] John J. Cannon. A critique of algebraic programming languages. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (12):18–27, July 1969. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Cannon:1991:BCC

[Can91] John Cannon. A bibliography of Cayley citations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(1):75–81, January 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Caprasse:1986:CSP

[Cap86a] H. Caprasse. A complete simplification package for the absolute value function in the real domain. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 20(1–2):18–21, February/May 1986. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Implementation for REDUCE 3.2 of the function ABS.

Caprasse:1986:DEM

[Cap86b] H. Caprasse. Description of an extension of the Matrix package of “Reduce”. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 20(4):7–10, December 1986. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Caprotti:2003:ISS

[Cap03]  Olga Caprotti.  Inter-

national Symposium on Symbolic and Algebraic Computation: poster abstracts 2003. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37 (3):65, September 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Carlisle:2000:OMX

[Car00] David Carlisle. OpenMath, MathML, and XSL. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(2):6–11, June 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Caruso:2003:ISS

[Car03] Fabrizio Caruso. Interpolation in symbolic summation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37 (3):66, September 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Casale:2008:AEL

[Cas08] Giuliano Casale. An application of exact linear algebra to capacity planning models. ACM Communications in Computer Algebra, 42(4):202–205, December 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Caviness:1975:MCR

[Cav75]  B. F. Caviness.  More

on computing roots of integers. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9 (3):18–20, August 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Caviness:1972:SMC

[CC72] B. F. Caviness and G. E.

Collins. Symbolic mathematical computation in a Ph.D. Computer Science program. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (23):25–28, July 1972. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Calmet:1983:SRR

[CC83] Jacques Calmet and Ian Cohen. Synthesizing recurrence relations. I. analysis of the problems. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(2):18–25, May 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Caboara:2008:GAG

[CC08] Massimo Caboara and Fabrizio Caruso. Group actions on Gröbner bases of saturated zero-dimensional binomial ideals. ACM Communications in Computer Algebra, 42(3):138– 139, September 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Chan:2016:FMP

[CC16] Eunice Y. S. Chan and Robert M. Corless. Fibonacci– Mandelbrot polynomials and matrices. ACM Communications in Computer Algebra, 50(4):155–157, December 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Editors:2013:ARDa

[CCA13a] CCA Editors. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 47 (3–4):124–125, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Editors:2013:ARDb

[CCA13b] CCA Editors. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 47 (3–4):189–190, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Editors:2014:ARDa

[CCA14a] CCA Editors. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 48 (1):28, March 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Editors:2014:ARDb

[CCA14b] CCA Editors. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 48(2): 64–66, March 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Editors:2014:ARDc

[CCA14c] CCA Editors. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 48 (3/4):148–150, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Editors:2014:ARDd

[CCA14d] CCA Editors. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 48(3/4): 202, September 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Editors:2015:ARD

[CCA15] CCA Editors. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 49 (1):36, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

CCA:2020:ARD

[CCA20] CCA editors. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 53(4): 183–188, March 2020. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/abs/10.
1145/3388974.3388977.

Chen:2014:BPA

[CCM^{+}14] Changbo Chen, Svyatoslav Covanov, Farnam Mansouri, Marc Moreno Maza, Ning Xie, and Yuzhen Xie. Basic polynomial algebra subprograms. ACM Communications in Computer Algebra, 48(3/4):197–201, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Chen:2016:BPA

[CCM^{+}16] Changbo Chen, Svyatoslav Covanov, Farnam Mansouri, Robert H. C. Moir, Marc Moreno Maza, Ning Xie, and Yuzhen Xie. The basic polynomial algebra subprograms. ACM Communications in Computer Algebra, 50(3):97– 100, September 2016. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Caprotti:2000:JPC

[CCR00] Olga Caprotti, Arjeh M.

Cohen, and Manfred Riem. JAVA phrasebooks for computer algebra and automated deduction. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(2):33–37, June 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:2000:AAS

[CDJW00] Robert M. Corless, James H. Davenport, David J. Jeffrey, and Stephen M. Watt. “according to Abramowitz and Stegun” or arccoth needn’t be uncouth. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34 (2):58–65, June 2000. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Chen:2011:CRS

[CDL^{+}11] Changbo Chen, James H. Davenport, François Lemaire, Marc Moreno Maza, Bican Xia, Rong Xiao, and Yuzhen Xie. Computing the real solutions of polynomial systems with the RegularChains library in Maple. ACM Communications in Computer Algebra, 45(3–4):166–168, September 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Chien:2015:CTD

[CDM15] Bui Van Chien, G. H. E. Duchamp, and Hoang Ngoc Minh. Computation tool for the q-deformed quasishuffle algebras and representations of structure of MZVs. ACM Communications in Computer Algebra, 49(4):117–120, December 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Cattani:1999:SRH

[CDS99] E. Cattani, A. Dickenstein, and B. Sturmfels. The search for rational Ahypergeometric functions. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):15, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Caviness:1977:NCA

[CE77] B. F. Caviness and H. I. Epstein. A note on the complexity of algebraic differentiation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11 (3):4–6, August 1977. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

ChAze:2008:AFB

[CEGW08] G. Chèze, M. Elkadi, A. Galligo, and M. Weimann. Absolute factoring of bidegree bivariate polynomials. ACM Communications in Computer Algebra, 42(3):151–153, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Celis:2022:ATI

[Cel22] Oliver Salazar Celis. Adaptive Thiele interpolation. ACM Communications in Computer Algebra, 56(3): 125–132, September 2022. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3594252.3594254.

Cetinkaya:1996:MBR

[Cet96]  Cetin Cetinkaya.  Mes-

sage from the Book Review Editor. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (3):1, September 1996. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Cetinkaya:1997:BR

[Cet97] Cetin Cetinkaya. Book reviews. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (1):26, March 1997. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Caravantes:2018:OCQ

[CFGVDT18] Jorge Caravantes, Mario Fioravanti, Laureano Gonzalez-Vega, and Gema Diaz-Toca. Offsets to conics and quadrics: a new determinantal representation for their implicit equation. ACM Communications in Computer Algebra, 52(3):85–88, September 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Chalkis:2020:SFS

[CFRT20] Apostolos Chalkis, Vissarion Fisikopoulos, Panagiotis Repouskos, and Elias Tsigaridas. Sampling the feasible sets of SDPs and volume approximation. ACM Communications in Computer Algebra, 54(3):114– 118, September 2020. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3457341.3457349.

Chatterji:2011:CEG

[CG11] Samaresh Chatterji and Ratnik Gandhi. Computing equilibria with group actions. ACM Communications in Computer Algebra, 45(2):115–116, June 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Chen:2019:AOS

[CG19] Shaoshi Chen and Jing Guo.

Additive Ore–Sato theorem. ACM Communications in Computer Algebra, 53(3):96–98, September 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Char:1983:MSC

[CGG83] Bruce Char, Keith Geddes, and Gaston Gonnet. The Maple symbolic computation system. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(3–4):31–42, August/November 1983. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Caviness:2004:MRD

[CGG^{+}04] Bob Caviness, Barbara Gatje, James H. Griesmer, Tony Hearn, Manual Bronstein, and Erich Kaltofen. In memoriam: Richard Dimick Jenks: Axiom developer and computer algebra pioneer. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(1):30, March 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). URL

                                        http://savannah.nongnu.
                                        org/projects/axiom/.

Cheng:2005:SEEa

[CGKZ05a] Howard Cheng, Barry Gergel, Ethan Kim, and Eugene Zima. Space-efficient evaluation of hypergeometric series. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39 (2):41–52, June 2005. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Cheng:2005:SEEb

[CGKZ05b] Howard Cheng, Barry Gergel, Ethan Kim, and Eugene Zima. Space-efficient evaluation of hypergeometric series. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39 (3):81–83, 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). ISSAC 2005 poster abstract.

Cheng:2005:ITR

[CGL05] Jin-San Cheng, Xiao-Shan Gao, and Ming Li. Intrinsic topological representation of real algebraic surfaces. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39 (3):80–81, September 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). ISSAC 2005 poster abstract.

Caprasse:1985:NUO

[CH85] H. Caprasse and M. Hans.

A new use of operators in the algebraic mode of REDUCE. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19 (3):46–52, August 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Chan:1981:NSO

[Cha81] Woon Cheung Chan. A novel symbolic ordinary differential equation solver. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15(3):9–14, August 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Char:1986:PSS

[Cha86] Bruce W. Char, editor. Proceedings of the 1986 Symposium on Symbolic and Algebraic Computation: Symsac ’86, July 21–23, 1986, Waterloo, Ontario. ACM Press, New York, NY, USA, 1986. ISBN 0-89791-199-7 (paperback). LCCN QA155.7.E4 A281 1986. ACM order no. 505860.

Char:1990:RAM

[Cha90] Bruce W. Char. Report on the 7th Annual Maple Workshop. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 24 (1):18, January 1990. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Char:1991:RAM

[Cha91] Bruce W. Char. Report on the 8th Annual Maple Workshop. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25 (1):24, January 1991. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Char:1996:FAC

[Cha96a] Bruce Char. First announcement and call for papers: ISSAC ’97: International Symposium on Symbolic and Algebraic Computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (2):44–46, June 1996. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Char:1996:I

[Cha96b] Bruce Char. ISSAC ’97.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(3):30, September 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Char:1997:SCM

[Cha97]  Bruce Char.  SIGSAM

Chair’s message. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):2–3, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Chen:1983:SMF

[Che83] Dongyue Chen. Symbol manipulation and formula manipulation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17 (2):26–27, May 1983. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Char:1999:DSE

[CHH^{+}99] Bruce W. Char, Anthony Harrison, Thomas Hewett, Ron Perline, and Muksim Rakhimov. Developing the Soliton Explorer: a problem solving environment for soliton surface investigation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):22–23, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Char:1997:SCT

[CHJ^{+}97] Bruce Char, Tom Hewett, Jeremy Johnson, Lakshman Yagati, Ron Perline, Raji Venkatesan, Hoi Man Chang, Michael Miller, and Clint Hepner. Symbolic computation tools in scientific problem solving environments. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):58, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Corless:2014:SDI

[CHJ14] Robert M. Corless, Junrui Hu, and D. J. Jeffrey. Some definite integrals containing the Tree T function. ACM Communications in Computer Algebra, 48(2): 33–41, March 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Chen:2014:IAP

[CHL14] Shaoshi Chen, Hui Huang, and Ziming Li. Improved Abramov–Petkovsek’s reduction and creative telescoping for hypergeometric terms. ACM Communications in Computer Algebra, 48(3/4):106–108, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Carter:2016:EDH

[CHS16] Richard G. Carter, Shahadat Hossain, and Marzia Sultana. Efficient detection of Hessian matrix sparsity pattern. ACM Communications in Computer Algebra, 50(4):151–154, December 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Corless:1988:CTC

[CJ88] R. M. Corless and D. J. Jeffrey. A comparison of three computer algebra systems for the solution of a problem in hydrodynamic lubrication. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22 (2):50–62, April 1988. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1992:WLI

[CJ92] Robert M. Corless and David J. Jeffrey. Well … it isn’t quite that simple. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26(3):2–6, August 1992. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1996:ECU

[CJ96] Robert M. Corless and David J. Jeffrey. Editor’s corner: The unwinding number. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (2):28–35, June 1996. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1997:TFR

[CJ97] R. M. Corless and D. J.

Jeffrey. The Turing factorization of a rectangular matrix. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):20–28, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1998:GER

[CJ98] Robert M. Corless and David J. Jeffrey. Graphing elementary Riemann surfaces. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (1):11–17, March 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:2018:RDT

[CJ18] Robert M. Corless and Julia E. Jankowski. Revisiting the discharge time of a cylindrical leaking bucket: or, “one does not simply

call dsolve into mordor.”.

ACM Communications in Computer Algebra, 52(1): 1–10, March 2018. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Corless:2018:CAD

[CJS18] Robert M. Corless, David J. Jeffrey, and David R. Stoutemyer. Comprehensive anti-derivatives and parametric continuity. ACM Communications in Computer Algebra, 52(2):32–33, June 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Chtcherba:2003:RBD

[CK03] Arthur D. Chtcherba and Deepak Kapur. On the relationship between the Dixonbased resultant construction and the supports of polynomial systems. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(3):66–67, September 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ceberio:2004:GAO

[CK04] Martine Ceberio and Vladik Kreinovich. Greedy algorithms for optimizing multivariate Horner schemes. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(1):8–15, March 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:2003:ACP

[CKW03] Robert M. Corless, Hiroshi Kai, and Stephen M. Watt. Approximate computation of pseudovarieties. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(3):67–71, September 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Collins:1982:JSA

[CL82] George E. Collins and Rüdiger G. K. Loos. The Jacobi symbol algorithm. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16(1):12–16, February 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Calmet:1987:KBS

[CL87] Jacques Calmet and Denis Lugiez. A knowledgebased system for computer algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21 (1):7–13, February 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Calmet:1996:DIS

[CL96] Jacques Calmet and Carla Limongelli, editors. Design and implementation of symbolic computation systems: International Symposium, DISCO ’96, Karlsruhe, Germany, September 18–20, 1996: proceedings, volume 1128 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61697-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88I576 1996.

Cheng:2009:ALA

[CL09] Howard Cheng and George Labahn. Applying linear algebra routines to modular ore polynomial matrix algorithms. ACM Communications in Computer Algebra, 43(3):78–79, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Chowdhury:2011:RLT

[CL11] Muhammad F. I. Chowdhury and Romain Lebreton. Root lifting techniques and applications to list decoding. ACM Communications in Computer Algebra, 45(2): 113–114, June 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Cherchem:2015:ESP

[CL15] Ahmed Cherchem and André Leroy. Exponents of skew polynomials. ACM Communications in Computer Algebra, 49(1):14, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Clark:1969:SSS

[Cla69] S. James Clark. Symbolic solution of a system of first order linear differential equations. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(11):13–18, April 1969. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Clarkson:1989:MIL

[Cla89] M. Clarkson. MACSYMA’s inverse Laplace transform. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23(1):33–38, January 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Chen:2008:RCS

[CLM^{+}08] Changbo Chen, Liyun Li, Marc Moreno Maza, Wei Pan, and Yuzhen Xie. On the representation of constructible sets. ACM Communications in Computer Algebra, 42(3):162– 163, September 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Cheng:2008:CPL

[CLZ08] Howard Cheng, George Labahn, and Wei Zhou. Computing polynomial LCM and GCD in Lagrange basis. ACM Communications in Computer Algebra, 42(3):129–130, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Chen:2023:TLM

[CLZ23] Benjamin Chen, Yu Li, and Eugene Zima. On a twolayer modular arithmetic. ACM Communications in Computer Algebra, 57(3): 133–136, September 2023. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3637529.3637534.

Cioni:1976:UMA

[CM76] Gianna Cioni and Alfonso Miola. Using minicomputers for algebraic computations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 10 (4):50–52, November 1976. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Colagrossi:1983:CRZ

[CM83] Attilio Colagrossi and Alfonso M. Miola. Computing real zeros of polynomials with parametric coefficients. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17 (1):12–15, February 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Clark:1999:MM

[CM99] David Clark and James Mc-Carron. Modules for Maple. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):24–25, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Cooperman:2002:OMW

[CM02] Gene Cooperman and Xiaoqin Ma. Overcoming the memory wall in symbolic algebra: a faster permutation multiplication. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36(4):1–4, December 2002. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Caboara:2015:HQP

[CM15] Massimo Caboara and Carla Mascia. On the Hilbert quasi-polynomials for non-standard graded rings. ACM Communications in Computer Algebra, 49(3):101–104, September 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Chen:2022:FNM

[CM22] Tian Chen and Michael Monagan. Factoring nonmonic polynomials represented by black boxes. ACM Communications in Computer Algebra, 56(2):80– 83, June 2022. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572867.3572881.

Chen:2008:CPM

[CML^{+}08] Changbo Chen, Marc Moreno Maza, François Lemaire, Wei Pan, Liyun Li, and Yuzhen Xie. The ConstructibleSetTools and ParametricSystemTools modules of the RegularChains library in Maple. ACM Communications in Computer Algebra, 42(3):182– 184, September 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Chowdhury:2010:CPR

[CMPS10] Muhammad F. I. Chowdhury, Marc Moreno Maza, Wei Pan, and Éric Schost. Complexity and performance results for non FFTbased univariate polynomial multiplication. ACM Communications in Computer Algebra, 44(3):99– 100, September 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Collins:1974:SSP

[CMR74] G. E. Collins, D. R.

Musser, and M. Rothstein. SAC-1 solution of problem #7. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(2): 17–19, May 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:2014:ZFF

[CMT14] Robert M. Corless, Marc Moreno Maza, and Steven E. Thornton. Zigzag form over families of parametric matrices. ACM Communications in Computer Algebra, 48 (3/4):109–112, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Chen:2010:CCM

[CMX10] Changno Chen, Marc Moreno Maza, and Yuzhen Xie. Cache complexity and multicore implementation for univariate real root isolation. ACM Communications in Computer Algebra, 44(3):97–98, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Cohen:1974:PCG

[Coh74] H. Ian Cohen. A perturbation calculation in general relativity. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8 (3):78–79, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Collins:1969:AAS

[Col69] G. E. Collins. Algorithmic approaches to symbolic integration and simplification: summary of the panel session sponsored by SIGSAM at the FJCC in San Francisco, December 10, 1968. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (12):5–16, July 1969. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Collins:1974:QER

[Col74] George E. Collins. Quantifier elimination for real closed fields by cylindrical algebraic decomposition– preliminary report. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(3):80–90, August 1974. CODEN SIGSBZ.

ISSN 0163-5824 (print), 1557-9492 (electronic).

Combariza:2018:FCA

[Com18] Germán A. G. Combariza.

A few conjectures about the multiple zeta values. ACM Communications in Computer Algebra, 52(1): 11–20, March 2018. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Conti:1990:HCF

[Con90] Pasqualina Conti. Hermite canonical form and Smith canonical form of a matrix over a principal ideal domain. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 24 (3):8–16, July 1990. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Cooperman:1993:LID

[Coo93] Gene Cooperman. Letter from the information director. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27 (3):3–??, September 1993. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Cooperman:1996:IBM

[Coo96a] Gene Cooperman. ISSAC business meeting minutes.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(4):2–3, December 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Cooperman:1996:SBM

[Coo96b] Gene Cooperman. SIGSAM business meeting minutes. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(4):1–2, December 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Cooperman:1997:FEC

[Coo97] G. Cooperman. The Fourther East Coast Computer Algebra Day: a one day meeting to stimulate activity in computer algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (1):29, March 1997. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Cooperman:2004:TCP

[Coo04] Gene Cooperman. The TOP-C parallel model and symbolic algebra. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(1):16–17, March 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1993:WSO

[Cor93] Robert M. Corless. What is a solution of an ODE? SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27(4):15–19, December 1993. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1996:ABD

[Cor96a]  R. M. Corless.  Abell

and Braselton’s Differential Equations with Maple V. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(1):57–60, March 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1996:MEa

[Cor96b] R. M. Corless. Message from the Editor. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(1):2–3, March 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1996:CI

[Cor96c] Robert M. Corless. Cofactor iteration. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (1):34–38, March 1996. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1996:ECG

[Cor96d] Robert M. Corless. Editor’s corner: Gröbner bases and matrix eigenproblems. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(4):26–32, December 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1996:MEb

[Cor96e] Robert M. Corless. Message from the Editor. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(2):1, June 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1996:ME

[Cor96f] Robert M. Corless. Message from the Editor. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(3):1, September 1996. CODEN SIGSBZ.

ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1997:MEa

[Cor97a] R. M. Corless. Message from the Editor. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(1):1, March 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1997:MEb

[Cor97b] Robert M. Corless. Message from the Editor. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):1, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1998:MEb

[Cor98a] Rob Corless. Message from the Editor. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (4):1, December 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1998:EM

[Cor98b] Robert M. Corless. Editor’s message. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(3):1, September 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1998:MEa

[Cor98c] Robert M. Corless. Message from the Editor. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(2):1–2, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1999:OPC

[Cor99a]  R. M. Corless.  Open

problems in computer algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):15, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1999:MOE

[Cor99b] Rob Corless. Message from the outgoing Editor and SIGSAM Chair. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(2):1–2, June 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:2000:ESM

[Cor00] Robert M. Corless. An elementary solution of a minimax problem arising in algorithms for automatic mesh selection. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(4):7–15, December 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:2001:MSC

[Cor01a] Robert M. Corless. Message from the SIGSAM Chair. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35 (1):1–2, March 2001. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:2001:RMC

[Cor01b] Robert M. Corless. Review of Modern Computer Algebra, by Joachim von zu Gathen and Jügen Gerhard. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35 (1):8–13, March 2001. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:2002:MSC

[Cor02] Robert M. Corless. Message from the SIGSAM Chair. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36 (1):1–2, March 2002. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Coutsias:2015:BFO

[Cou15] Evangelos A. Coutsias. Bricard flexible octahedra and the canonical cyclohexane. ACM Communications in Computer Algebra, 49(2):56, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Caviness:1978:NAI

[CP78] B. F. Caviness and M. J. Prelle. A note on algebraic independence of logarithmic and exponential constants. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 12 (2):18–20, May 1978. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Cattaneo:1988:NPI

[CP88] G. M. Piacentini Cattaneo and M. Pittaluga. Nonassociative polynomial identities, Young tableaux and computers. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22 (2):17–26, April 1988. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Cha:2012:CFS

[CP12] Yongjae Cha and Veronika Pillwein. Closed form solutions of linear difference equations in terms of symmetric products. ACM Communications in Computer Algebra, 46(3–4):80– 81, September 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Corless:2011:RCA

[CPS11] Robert Corless, Erik Postma, and David R. Stoutemyer. Rounding coefficients and artificially underflowing terms in non-numeric expressions. ACM Communications in Computer Algebra, 45(1):17–48, March 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Cannon:1984:CTG

[CR84] John J. Cannon and Jim Richardson. Cayley: teaching group theory by computer. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18– 19(4–1):15–18, November/ February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Cuevas-Rozo:2019:KIA

[CRDMBR19] Julián Cuevas-Rozo, Jose Divasón, Miguel Marco-Buzunáriz, and Ana Romero. A Kenzo interface for algebraic topology computations in SageMath. ACM Communications in Computer Algebra, 53(2):61–64, June 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Crowley:2020:MPS

[CRWZ20] Colin Crowley, Jose Israel Rodriguez, Jacob Weiker, and Jacob Zoromski. MultiRegeneration for polynomial system solving. ACM Communications in Computer Algebra, 54(2):39–43, September 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3427218.3427221.

Corless:2019:AGB

[CS19] Robert M. Corless and Leili Rafiee Sevyeri. Approximate GCD in Bernstein basis. ACM Communications in Computer Algebra, 53(3):103–106, September 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Cuyt:2008:SMP

[CsL08] Annie Cuyt and Wen shin Lee. Sparse multivariate polynomial interpolation via the quotient-difference algorithm. ACM Communications in Computer Algebra, 42(3):154–155, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Cuyt:2009:SNS

[CsL09] Annie Cuyt and Wen shin Lee. Symbolic-numeric sparse interpolation of multivariate rational functions. ACM Communications in Computer Algebra, 43(3): 79–80, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Cuyt:2010:RST

[CsL10] Annie Cuyt and Wen shin Lee. Reconstructing sparse trigonometric functions. ACM Communications in Computer Algebra, 44(3):101–102, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Cuyt:2018:SSP

[CsL18] Annie Cuyt and Wen shin Lee. A scale and shift paradigm for sparse interpolation in one and more dimensions. ACM Communications in Computer Algebra, 52(3):75–77, September 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Corless:2016:BEP

[CT16] Robert M. Corless and Steven E. Thornton. The Bohemian eigenvalue project. ACM Communications in Computer Algebra, 50(4): 158–160, December 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Chi:2019:GAB

[CT19] Boming Chi and Akira Terui. The GPGCD algorithm with the Bézout matrix. ACM Communications in Computer Algebra, 53(3):99–102, September 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Cheb-Terrab:1999:SCI

[CTKR99] Edgardo S. Cheb-Terrab, Theodore Kolokolnikov, and Austin D. Roche. The search for and classification of integrable Abel ODE classes. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):23–24, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Cha:2010:AOS

[CvHL10a] Yongjae Cha, Mark van Hoeij, and Giles Levy. Abstract only: Solving linear recurrence relations.

ACM Communications in Computer Algebra, 44(2): 26, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Cha:2010:SLR

[CvHL10b] Yongjae Cha, Mark van Hoeij, and Giles Levy. Solving linear recurrence equations. ACM Communications in Computer Algebra, 44(4):183–185, December 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Cheatham:1972:LSA

[CW72] T. E. Cheatham, Jr. and Ben Wegbreit. A laboratory for the study of automating programming. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(21):8–26, January 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Corless:1998:RSM

[CW98] Robert M. Corless and Stephen M. Watt. Report on the SNAP minisymposium at SIAM ’98. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(2):35–37, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Chen:2012:GID

[CWZ12] Xiaoyu Chen, Dongming Wang, and Ting Zhao. Geo-Text: an intelligent dynamic geometry textbook. ACM Communications in Computer Algebra, 46(3–4): 171–175, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Chen:2008:WMI

[CX08] William Y. C. Chen and Ernest X. W. Xia. The q-WZ method for infinite sums (abstract only). ACM Communications in Computer Algebra, 42(1–2):19, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

DAndrea:2015:RAS

[D’A15] Carlos D’Andrea. The Rees algebra of some monomial parametrizations. ACM Communications in Computer Algebra, 49(2):52–53, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Dahan:2012:SLG

[Dah12] Xavier Dahan. On the structure of lexicographic Gröbner bases in dimension zero. ACM Communications in Computer Algebra, 46(3–4):82–83, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Daley:2001:RQC

[Dal01] Mark Daley. Review of Quantum Computing and Quantum Information, by Michael A. Nielson and Issac L. Chuang. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35(1):14–15, March 2001. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Daly:2002:AOS

[Dal02] Tim Daly. Axiom as open source. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36 (1):28, March 2002. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Davenport:1979:AI

[Dav79] J. H. Davenport. Anatomy of an integral. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 13(4):16–18, November 1979. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Davenport:1982:FRT

[Dav82a] James H. Davenport. Fast REDUCE: The trade-off between efficiency and generality. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16 (1):8–11, February 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Davenport:1982:PRA

[Dav82b] James H. Davenport. On the parallel Risch algorithm (III): use of tangents. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16(3):3–6, August 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Davenport:1982:WDW

[Dav82c] James H. Davenport. What do we want from a highlevel language? SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16(4):6–9, 23, November 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Davenport:1983:CC

[Dav83] James H. Davenport. Chairman’s column. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(2):1–5, May 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Davenport:1984:IFT

[Dav84] James Davenport. Integration in finite terms. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(2):20–21, May 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Davenport:1986:PMP

[Dav86] J. H. Davenport. A “piano movers” problem. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 20(1–2):15–17, February/May 1986. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Davenport:2000:WOC

[Dav00a] James H. Davenport. On writing OpenMath content dictionaries. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34 (2):12–15, June 2000. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Davenport:2000:SOT

[Dav00b] James H. Davenport. A small OpenMath type system. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34 (2):16–21, June 2000. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Davison:2003:BRR

[Dav03] Matt Davison. Book review: Review of risk management: value at risk and beyond, edited by M. A. H. Dempster. Cambridge University Press 2002. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(1):28–29, March 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Davenport:2012:SAS

[Dav12] James H. Davenport. Small algorithms for small systems. ACM Communications in Computer Algebra, 46(1–2):1–9, March/June 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

deCastilla:1999:INB

[dCS99] I. Ojeda Martínez de Castilla and R. Piedra Sánchez. Index of nilpotency of binomial ideals. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):18, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

deCabezon:2009:EAA

[dCW09] Eduardo Sáenz de Cabezón and Henry P. Wynn. Efficient algorithms for the algebraic analysis of system reliability. ACM Communications in Computer Algebra, 43(3):98–99, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

DAndrea:1999:BFM

[DD99] C. D’Andrea and A. Dickenstein. Bezoutian formulas à la Macaulay for the multivariate resultant. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):15, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

DeLoera:2011:SEI

[DDK^{+}11] J. A. De Loera, B. Dutra, M. Köppe, S. Moreinis, G. Pinto, and J. Wu. Software for exact integration of polynomials over polyhedra. ACM Communications in Computer Algebra, 45 (3–4):169–172, September 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

DeWit:1999:ACR

[De 99]  David De Wit.  Auto-

matic construction of an R Matrix. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):25, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Decker:2014:RNS

[Dec14] Wolfram Decker. Review of “Numerically Solving Polynomial Systems with Bertini” by Daniel J. Bates, Jonathan D. Hauenstein, Andrew J. Sommese, and Charles W. Wampler. ACM Communications in Computer Algebra, 48(2):57–58, March 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). See reviewed book [Bat13].

Demenkov:2012:ERA

[Dem12] Max Demenkov. Estimating region of attraction for polynomial vector fields by homotopy methods. ACM Communications in Computer Algebra, 46(3–4):84– 85, September 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Dogan:2023:CCH

[DET23] M. Levent Dogan, Alperen A. Ergür, and Elias Tsigaridas.

On the complexity of Chow and Hurwitz forms. ACM Communications in Computer Algebra, 57(4):167– 199, December 2023. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3653002.3653003.

Dewar:2000:OO

[Dew00a] Mike Dewar. OpenMath:

An overview. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(2):2–5, June 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Dewar:2000:SIO

[Dew00b] Mike Dewar. Special issue on OpenMath: Message from the Guest Editor. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(2):1, June 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Degwekar:2015:ECX

[DGG15] Akshay Degwekar, Kenza Guenda, and T. Aaron Gulliver. Extending construction X for quantum errorcorrecting codes. ACM Communications in Computer Algebra, 49(1):18, March 2015. CODEN ????

ISSN 1932-2232 (print), 1932-2240 (electronic).

Dhooge:2004:MMP

[DGK04] A. Dhooge, W. Govaerts, and Yu. A. Kuznetsov. MATCONT: a Matlab package for numerical bifurcation analysis of ODEs. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(1):21–22, March 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Du:2019:ADM

[DGL19] Hao Du, Jing Guo, and Ziming Li. An additive decomposition in multiple logarithmic extensions. ACM Communications in Computer Algebra, 53(3):110– 113, September 2019. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Delgado:2016:NGP

[DGS16] M. Delgado and P. A. García-Sánchez. numericalsgps, a GAP package for numerical semigroups. ACM Communications in Computer Algebra, 50(1):12–24, March 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

DeBeule:2004:SMB

[DHS04] J. De Beule, A. Hoogewijs, and L. Storme. On the size of minimal blocking sets of Q(4; q), forq = 5, 7. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(3):67–84, September 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

diScala:1982:SPB

[di 82] R.-M. di Scala. SYPAC:

a Pascal based computer algebra system for microcomputers. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16 (3):7–10, August 1982. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Dockx:1991:CTC

[DI91] Kris Dockx and Paul Igodt.

Character tables and commutativity of normal subgroups. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25 (1):28–31, January 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Dimino:1971:GAC

[Dim71] Lucien A. Dimino. A graphical approach to coset enumeration. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (19):8–43, July 1971. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ding:2009:NSC

[Din09] Hui Ding. Numerical and symbolic computation of the Lambert W function in C^{n×n}. ACM Communications in Computer Algebra, 43(4):121, December 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Dirat:2000:JSC

[Dir00] Laurent Dirat. JOME, a software component for interactive and distributed mathematics. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(2):38–42, June 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Drevet:2010:OTS

[DIS10] Charles Éric Drevet, Md. Nazrul Islam, and Éric Schost. Optimization techniques for small matrix multiplication. ACM Communications in Computer Algebra, 44(3):107–108, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Davenport:1981:M

[DJ81] James H. Davenport and Richard D. Jenks. MOD-LISP. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15 (1):11–20, February 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Dora:1996:RRC

[DJ96] J. Della Dora and F. Jung.

Resolvent and rational canonical forms of matrices. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (3):4–10, September 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Decker:1998:PRR

[DK98] Thomas Decker and Werner Krandick. Parallel real root isolation: poster abstract. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (2):53–54, June 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Abstract only.

Dooley:1996:E

[DL96] S. Dooley and A. Lobo. EC-CAD ’96. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (1):39–49, March 1996. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Duchamp:2012:DFE

[DM12] Gàérard H. E. Duchamp and Vincel Hoang Ngoc Minh. Dual families in enveloping algebras. ACM Communications in Computer Algebra, 46(3–4):86– 87, September 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Dali:2015:LSC

[DM15] Dahira Dali and Jugurta Mahrez. Local stability of cubic differential systems. ACM Communications in Computer Algebra, 49(1): 29, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Darkey-Mensah:2021:AQF

[DM21] Mawunyo Kofi Darkey-Mensah. Algorithms for quadratic forms over global function fields of odd characteristic. ACM Communications in Computer Algebra, 55(3):68–72, September 2021. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL

https://dl.acm.org/doi/
10.1145/3511528.3511530.

Duck:2015:DCG

[DMCMM15] Natalia Dück, Irene Márquez-Corbella, and Edgar Martínez-Moro. The degree compatible Gröbner fan for linear codes. ACM Communications in Computer Algebra, 49(1):18–19, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Duchamp:2015:HSP

[DMH15] Gérard H. E. Duchamp, Hoang Ngoc Minh, and Ngo Quoc Hoan. Harmonic sums and polylogarithms at negative multiindices. ACM Communications in Computer Algebra, 49(3):70–73, September 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Darkey-Mensah:2019:IFI

[DMK19] Mawunyo Kofi Darkey-Mensah and Przemysław Koprowski. Intrinsic factorization of ideals in Dedekind domains. ACM Communications in Computer Algebra, 53(3):107–109, September 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Duparc:2002:WOS

[DMP^{+}02] Daniel Duparc, Bernard Mourrain, Bernard Parisse, Fabrice Rouillier, Marie-Françoise Roy, Nicolas Thiéry, and Paul Zimmermann. Workshop on Open Source Computer Algebra, Tuesday 21st – Thursday 23rd of May 2002, Lyon, France. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36 (1):24–27, March 2002. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Dahan:2010:ALT

[DMSP10] Xavier Dahan, Marc Moreno Maza, Éric Schost, and Adrien Poteaux. Almost linear time operations with triangular sets. ACM Communications in Computer Algebra, 44(3):103– 104, September 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Dedieu:1990:SVD

[DN90] J. P. Dedieu and G. Norton. Stewart varieties: a direct algebraic model for Stewart platforms. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 24 (4):42–59, October 1990. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

deOlazabal:1999:UMD

[dO99] J. M. de Olazábal. Unified method for determining canonical forms of a matrix. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (1):6–20, March 1999. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Doerksen:2008:GCS

[Doe08] Kevin Doerksen. Genus 2 curves with split Jacobians (abstract only). ACM Communications in Computer Algebra, 42(1–2):49– 50, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Dongyue:1983:SMF

[Don83] Chen Dongyue. Symbol manipulation and formula manipulation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17 (2):26–27, May 1983. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Dooley:1999:IJS

[Doo99] Sam Dooley, editor. IS-SAC 99: July 29–31, 1999, Simon Fraser University, Vancouver, BC, Canada: proceedings of the 1999 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY, USA, 1999. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999.

Dousson:1999:EHC

[Dou99] Xavier Dousson. Effective homology of a classifying space. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):2–9, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Dana-Picard:2015:RPP

[DP15] Thierry N. Dana-Picard. Review of “Pavel Pech: Selected topics in geometry with classical vs. computer proving”, World Scientific 2007 ISBN-10: 981-270-942-8, ISBN-13: 978-981-270-942-4. ACM Communications in Computer Algebra, 49(2):65–66, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Duval:2016:DMR

[DP16] Dominique Duval and Adrien Poteaux. Death of Marc Rybowicz, aged 52. ACM Communications in Computer Algebra, 50(4):191, December 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Dukkipati:2015:GBR

[DPFD15] Ambedkar Dukkipati, Nithish Pai, Maria Francis, and Abhishek Dubey. On Gröbner bases over rings and residue class polynomial rings with torsion. ACM Communications in Computer Algebra, 49(2):60–61, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

DeFeo:2015:DRF

[DPQ15] Luca De Feo, Christophe Petit, and Michaël Quisquater. Deterministic root finding in finite fields. ACM Communications in Computer Algebra, 49(3):87–89, September 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

deQueiroz:1977:UFM

[dQ77a] F. Teixeira de Queiroz. The use of formula manipulation systems to compute asymptotic expansions. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11–12(4–1):11– 15, November and February 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

deQueroz:1977:UFM

[dQ77b] F. T. de Queroz. The use of formula manipulation systems to compute asymptotic expansions. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11–12(4–1):11– 16, November and February 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

deQueiroz:1978:ATA

[dQ78] F. Teixeira de Queiroz. Advanced tests for algorithms of differentiation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 12(4):19–22, August 1978. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

dosSantos:1990:DEH

[dR90] R. P. dos Santos and W. L. Roque. On the design of an expert help system for computer algebra systems. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 24(4):22–25, October 1990. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Dambacher:2001:GRC

[DR01] Jeffrey M. Dambacher and Philippe A. Rossignol. The golden rule of complementary feedback. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35(4):1–9, December 2001. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

DeFeo:2018:LCE

[DRR18] Luca De Feo, Hugues Randriambololona, and Édouard Rousseau. Lattices of compatibly embedded finite fields in Nemo/Flint. ACM Communications in Computer Algebra, 52(2): 38–41, June 2018. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

delRio:2023:CLM

[dRSE23] Tereso del Río, AmirHosein Sadeghimanesh, and Matthew England. Clustering in the Lazard method for cylindrical algebraic decomposition. ACM Communications in Computer Algebra, 57(3):126– 132, September 2023. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3637529.3637533.

Dolzmann:1997:RCA

[DS97] Andreas Dolzmann and Thomas Sturm. REDLOG: Computer algebra meets computer logic. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(2):2–9, June 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

DeFeo:2010:TLAa

[DS10a] Luca De Feo and Eric

                                        Schost.  transalpyne:

a language for automatic transposition. ACM Communications in Computer Algebra, 44(2):59–71, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

DeFeo:2010:TLAb

[DS10b] Luca De Feo and Éric

                                        Schost.  transalpyne:

a language for automatic transposition. ACM Communications in Computer Algebra, 44(3):105–106, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Dougherty:2015:CLR

[DS15] Steven T. Dougherty and Esengül Salturk. Codes over local rings of order 16 and their gray maps. ACM Communications in Computer Algebra, 49(1): 14, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Dimitrova:2008:AGF

[DSJL08] Elena S. Dimitrova, Brandilyn Stigler, Abdul Salam Jarrah, and Reinhard Laubenbacher. Applications of the Gröbner fan to gene network reconstruction (abstract only). ACM Communications in Computer Algebra, 42(1–2):69, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Dahan:2005:CDP

[DSM^{+}05] X. Dahan, É. Schost, M. Moreno Maza, W. Wu, and Y. Xie. On the complexity of the D5 principle. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(3):97–98, 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). IS-SAC 2005 poster abstract.

Dougherty:2015:CRO

[DSS15] Steven Dougherty, Esengül Saltürk, and Steve Szabo. Codes over rings of order 16. ACM Communications in Computer Algebra, 49 (1):15, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

deTrias:1984:NGS

[dT84] Regina Llopis de Trias. A new generation of symbolic and algebraic computation systems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18 (2):12–14, May 1984. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Dujella:2008:VWA

[Duj08] Andrej Dujella. A variant of Wiener’s attack on RSA with small secret exponent. ACM Communications in Computer Algebra, 42(1–2):50–51, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Durcheva:2015:SAI

[Dur15] Mariana Durcheva. Some applications of idempotent semirings in public key cryptography. ACM Communications in Computer Algebra, 49(1):19, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Dumitrescu:2007:QAD

[DX07] Adrian Dumitrescu and Guangwu Xu. On a query algorithm for a divisibility problem. ACM Communications in Computer Algebra, 41(4):122–124, December 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Eisenpress:1970:PSD

[EB70] Harry Eisenpress and Abel Bomberault. Practical symbolic differentiation of large expressions using PL/ I Formac. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (14):36–61, January 1970. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ebert:1983:SCM

[Ebe83] G. L. Ebert. Some comments on the modular approach to Gröbner-bases. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(2):28– 32, May 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Eberly:2006:LT

[Ebe06] Wayne Eberly. Letter from the Treasurer. ACM Communications in Computer Algebra, 40(1):1, March 2006. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ebersole:2008:PBI

[Ebe08] Caroline Ebersole. The population biology of infectious diseases (abstract only). ACM Communications in Computer Algebra, 42(1–2):

70, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Eberly:2016:BBL

[Ebe16] Wayne Eberly. Black box linear algebra: extending Wiedemann’s analysis of a sparse matrix preconditioner for computations over small fields. ACM Communications in Computer Algebra, 50(4):164– 166, December 2016. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ebner:1986:GCL

[Ebn86] Dieter W. Ebner. GOEDEL: a computer language for symbolic algebraic calculations — an introduction by examples. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 20 (3):16–19, September 1986. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

England:2014:BCM

[ECTB^{+}14] M. England, E. Cheb-Terrab, R. Bradford, J. H. Davenport, and D. Wilson. Branch cuts in Maple 17. ACM Communications in Computer Algebra, 48(1): 24–27, March 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Eder:2013:IIS

[Ede13] Christian Eder. Improving incremental signaturebased Gröbner basis algorithms. ACM Communications in Computer Algebra, 47(1–2):1–13, March 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

editors:2021:ARD

[edi21] CCA editors. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 55(3): 117–124, September 2021. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3511528.3511540.

Edoukou:2008:CDP

[Edo08] Frédéric A. B. Edoukou.

Computing the 2-distribution of points on Hermitian surfaces (abstract only). ACM Communications in Computer Algebra, 42(1–2):52– 53, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Essex:2000:NM

[EDS00] Christopher Essex, Matt Davison, and Christian Schulzky. Numerical monsters. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34 (4):16–32, December 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Eder:2015:SSB

[EF15] Christian Eder and Jean-Charles Faugère. A survey on signature-based Gröbner basis computations. ACM Communications in Computer Algebra, 49(2):61, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Emiris:1996:BPP

[EG96] I. Z. Emiris and A. Galligo. Bini and Pan’s polynomial and matrix computations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (3):21–23, September 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Esparza:2009:LFP

[EGK09] Javier Esparza, Andreas Gaiser, and Stefan Kiefer. On least fixed points of systems of positive polynomials. ACM Communications in Computer Algebra, 43(3):81–83, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Eder:2011:MFF

[EGP11] Christian Eder, Justin Gash, and John Perry. Modifying Faugère’s F5 algorithm to ensure termination. ACM Communications in Computer Algebra, 45(2):70–89, June 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Egner:2001:ADI

[EJP^{+}01] S. Egner, J. Johnson, D. Padua, M. Püschel, and J. Xiong. Automatic derivation and implementation of signal processing algorithms. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35 (2):1–19, June 2001. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Emiris:2012:SIU

[EK12] Ioannis Z. Emiris and Tatjana Kalinka. Sparse implicitization using support prediction. ACM Communications in Computer Algebra, 46(3–4):88–89, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Egorychev:2001:ECA

[EL01] G. P. Egorychev and V. M. Levchuk. Enumeration in the Chevalley algorithm. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35(2):20– 34, June 2001. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Elezi:2015:QCE

[Ele15] A. Elezi. Quantum computing, error correcting quantum codes and algebraic curves. ACM Communications in Computer Algebra, 49(2):58, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Elkies:2008:CGM

[Elk08] Noam D. Elkies. Curves of genus 2 with many rational points via K3 surfaces (abstract only). ACM Communications in Computer Algebra, 42(1–2):53– 55, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Eder:2022:TSB

[ELME22] Christian Eder, Pierre Lairez, Rafael Mohr, and Mohab Safey El Din. Towards signature-based Gröbner basis algorithms for computing the nondegenerate locus of a polynomial system. ACM Communications in Computer Algebra, 56(2):41– 45, June 2022. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572867.3572872.

Emiris:1996:S

[Emi96] I. Emiris. SNAP ’96. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(1):62, March 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Emiris:2015:RWA

[Emi15] Ioannis Emiris. Research workshop on algebraic representations in computeraided design for complex shapes. ACM Communications in Computer Algebra, 49(4):134, December 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Emiris:2020:PPG

[Emi20] Ioannis Emiris. 15 PhD positions in GRAPES: learning, processing & optimising shapes. ACM Communications in Computer Algebra, 53(4):189, March 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3388974.3388978.

Emiris:2008:CRB

[EMT08] Ioannis Emiris, Bernard Mourrain, and Elias Tsigaridas. Computing a rational in between. ACM Communications in Computer Algebra, 42(3):160– 161, September 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Encarnacion:1997:CPI

[Enc97a] M. J. Encarnacion. Call for posters: ISSAC ’97: International symposium on symbolic and algebraic computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (1):33, March 1997. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Encarnacion:1997:IPA

[Enc97b] M. J. Encarnación. IS-SAC ’97 poster abstracts. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):29–61, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Engeli:1975:ESS

[Eng75] Max Engeli. An enhanced SYMBAL system. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(4):21–29, November 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Engeler:1987:GDC

[Eng87]  Erwin Engeler.  Goals

and design considerations for a mathematical laboratory. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21 (3):7–8, August 1987. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

England:2015:RIS

[Eng15] Matthew England. Report on the 40th International Symposium on Symbolic and Algebraic Computation. ACM Communications in Computer Algebra, 49(3):69, September 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Erocal:2010:SFT

[Erö10] Burçin Eröcal. Summation in finite terms using sage. ACM Communications in Computer Algebra, 44(4):190–193, December 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Effenberger:2010:SGT

[ES10a] Felix Effenberger and Jonathan Spreer. simpcomp: a GAP toolbox for simplicial complexes. ACM Communications in Computer Algebra, 44(4):186–189, December 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Erocal:2010:NCR

[ES10b] Burçin Eröcal and Arne Storjohann. Nullspace computation over rational function fields for symbolic summation. ACM Communications in Computer Algebra, 44(3):109–110, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Effenberger:2011:SBD

[ES11] Felix Effenberger and Jonathan Spreer. Simplicial blowups and discrete normal surfaces in simpcomp. ACM Communications in Computer Algebra, 45(3–4):173– 176, September 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Elliott:2019:BCC

[ES19] Jesse Elliott and Éric Schost. Bit complexity for critical point computation in smooth and compact real hypersurfaces. ACM Communications in Computer Algebra, 53(3):114– 117, September 2019. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Epitropakis:2005:RFA

[EV05] M. G. Epitropakis and M. N. Vrahatis. Root finding and approximation approaches through neural networks. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(4):118–121, December 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Elbaz-Vincent:2021:EFL

[EVM21] Philippe Elbaz-Vincent and Etienne Marcatel. An extension of the fpLLL library to Hermitian lattices. ACM Communications in Computer Algebra, 55(2): 54–58, June 2021. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3493492.3493498.

Ekhad:2023:CCW

[EZ23] Shalosh B. Ekhad and Doron Zeilberger. Counting clean words according to the number of their clean neighbors. ACM Communications in Computer Algebra, 57(1):5–9, March 2023. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3610377.3610379.

Farmer:2004:MNI

[Far04] William M. Farmer. MKM: a new interdisciplinary field of research. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38 (2):47–52, June 2004. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Farley:2015:SUR

[Far15] Rosemary Carroll Farley. Student undergraduate research projects in linear algebra using a computer algebra system. ACM Communications in Computer Algebra, 49(2):50, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Fassino:2006:RMQ

[Fas06] Claudia Fassino. A robust monomial quotient basis for approximate points. ACM Communications in Computer Algebra, 40(2): 50–51, June 2006. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Fateman:1972:M

[Fat72a] R. J. Fateman. MAC-SYMA. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (24):12–13, October 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1972:RSN

[Fat72b] Richard J. Fateman. Rationally simplifying nonrational expressions. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(23):8–9, July 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1973:RE

[Fat73a] R. J. Fateman. Reply to an editorial. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (25):9–11, March 1973. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1973:CHI

[Fat73b] Richard J. Fateman. A case history in interactive problem-solving. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(28):30–32, December 1973. CODEN SIGSBZ.

ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1976:FPS

[Fat76] Richard J. Fateman. Final problem set excerpts. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 10(3):14, August 1976. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1978:LMD

[Fat78] Richard J. Fateman. Is a LISP machine different from a FORTRAN machine? SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 12 (3):8–11, August 1978. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1981:SAC

[Fat81] Richard J. Fateman. Symbolic and algebraic computer programming systems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15 (1):21–32, February 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1984:MVF

[Fat84] Richard J. Fateman. My view of the future of symbolic and algebraic computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18 (2):10–11, May 1984. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1985:CS

[Fat85a] Richard J. Fateman. Comments on SMP. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19(3):5–7, August 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1985:EPS

[Fat85b] Richard J. Fateman. Eleven proofs of sin^{2} x + cos^{2} x = 1. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19(2):25– 28, May 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1987:TOM

[Fat87] Richard J. Fateman. TEX output from MACSYMAlike systems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21 (4):1–5, November 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1990:LLM

[Fat90] Richard J. Fateman. A Lisp-language Mathematicato-Lisp translator. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 24(2):19–21, April 1990. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1996:WCA

[Fat96] Richard J. Fateman. Why computer algebra systems sometimes can’t solve simple equations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(2):8–11, June 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1998:SNS

[Fat98] Richard Fateman. A short note on short differentiation programs in Lisp, and a comment on logarithmic differentiation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(3):2–7, September 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1999:SEN

[Fat99] Richard J. Fateman. Symbolic execution and NaNs: diagnostic tools for tracking scientific computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):25–26, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:2003:CSP

[Fat03a] Richard Fateman. Comparing the speed of programs for sparse polynomial multiplication. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(1):4–15, March 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:2003:MCL

[Fat03b] Richard Fateman. Memory cache and Lisp: faster list processing via automatically rearranging memory. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(4):109–116, December 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:2014:ADL

[Fat14]  Richard Fateman.  Al-

gorithm Differentiation in Lisp: ADIL. ACM Communications in Computer Algebra, 48(3/4):78– 89, September 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Fateman:2015:PAS

[Fat15] Richard Fateman. Partitioning of algebraic subexpressions in computer algebra systems: an alternative to matching with an application to symbolic integration. ACM Communications in Computer Algebra, 49(2):38–47, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Flodmark:1970:UCT

[FB70] Stig Flodmark and Esko Blokker. Use of computers in treating nonnumerical mathematics and group theoretical problems in physics. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (15):28–63, July 1970. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1988:CSS

[FBOS88] Richard Fateman, Alan Bundy, Richard O’Keefe, and Leon Sterling. Commentary on: solving symbolic equations with PRESS. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22(2):27– 40, April 1988. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1999:PTM

[FC99] Richard J. Fateman and Eylon Caspi. Parsing TEX into mathematics. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):26, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fontan:2022:CAM

[FCB^{+}22] Jorge García Fontán, Alessandro Colotti, Sébastien Briot, Alexandre Goldsztejn, and Mohab Safey El Din. Computer algebra methods for polynomial system solving at the service of image-based visual servoing. ACM Communications in Computer Algebra, 56 (2):36–40, June 2022. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572867.3572871.

Feng:2013:IPA

[FCW13] Yong Feng, Jingwei Chen, and Wenyuan Wu. Incremental PSLQ with application to algebraic number reconstruction. ACM Communications in Computer Algebra, 47(3–4):112– 113, September 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Fee:1999:GLQ

[Fee99] Greg Fee. Gauss–Legendre quadrature. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):26, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Feldman:1975:ASC

[Fel75a] Stuart I. Feldman. An application of symbolic computation to crystal physics. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(2): 19–24, May 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Feldman:1975:BDA

[Fel75b] Stuart I. Feldman. A brief description of Altran. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(4):12–20, November 1975. CODEN SIGSBZ.

ISSN 0163-5824 (print), 1557-9492 (electronic).

Felsch:1977:KIB

[Fel77] V. Felsch. A KWIC indexed bibliography on the use of computers in group theory and related topics. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11–12(4–1):23– 86, November and February 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fevre:1998:CRS

[Fèv98] Stéphane Fèvre. Completeness of a rewrite system for proving geometric theorems using Clifford algebra. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(2):30–34, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fogelholm:1982:SLV

[FF82] Rabbe Fogelholm and Inge B.

Frick. Standard LISP for the VAX: a provisional implementation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16(4):10–12, November 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fortuna:1999:SFD

[FG99] Elisabetta Fortuna and Patrizia Gianni. Squarefree decomposition in finite characteristic: an application to Jordan form computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (4):14–32, December 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Feng:2003:PSF

[FG03] Ruyong Feng and Xiao-Shan Gao. Polynomial solutions for first order ODEs with constant coefficients. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(3):71–72, September 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Faugare:2012:FCO

[FGHR12] Jean-Charles Faugàre, Pierrick Gaudry, Louise Huot, and Guénaël Renault. Fast change of ordering with exponent ω. ACM Communications in Computer Algebra, 46(3–4):92–93, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Fioravanti:2015:CTA

[FGVC15] Mario Fioravanti, Laureano Gonzalez-Vega, and Jorge Caravantes. Computing the topology of an arrangement of parametric or implicit algebraic curves in the Lagrange basis. ACM Communications in Computer Algebra, 49 (2):53, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Fateman:1996:SLB

[FH96] Richard J. Fateman and Mark Hayden. Speeding up Lisp-based symbolic mathematics. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (1):25–30, March 1996. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fukasaku:2016:CSR

[FIS16] Ryoya Fukasaku, Hidenao Iwane, and Yosuke Sato. CGSQE/SyNRAC: a real quantifier elimination package based on the computation of comprehensive Gröbner systems. ACM Communications in Computer Algebra, 50(3):101– 104, September 2016. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Fitch:1973:PRM

[Fit73] John Fitch. Problems #3 and #4 in REDUCE and MACSYMA. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (28):10–11, December 1973. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fitch:1974:PRW

[Fit74a] John Fitch. Problem #8: random walk. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(4):11, November 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fitch:1974:SMT

[Fit74b]  John Fitch.  A simple

method of taking n th roots of integers. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(4): 26, November 1974. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fitch:1974:SPU

[Fit74c]  John Fitch.  A solu-

tion of problem #3 using CAMAL. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8 (4):14–25, November 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fitch:1975:CN

[Fit75] John Fitch. Course notes.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(3):4–8, August 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fitch:2012:JG

[Fit12] John Fitch. James Griesmer 1929–2011. ACM Communications in Computer Algebra, 46(1–2):10– 11, March/June 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Fateman:1991:SDP

[FL91] Richard J. Fateman and Derek T. Lai. A simple display package for polynomials and rational functions in Common Lisp. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(4):1–3, October 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Flaherty:1988:SME

[Fla88] Terry Flaherty. Symbolic manipulation in an extended λ-calculus. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22(1):21–24, January 1988. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fee:1997:PPCa

[FM97a] G. J. Fee and M. B. Monagan. Problems, puzzles, challenges. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (1):22–23, March 1997. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fee:1997:PPCb

[FM97b] G. J. Fee and M. B. Monagan. Problems, puzzles, challenges. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (4):48–49, December 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1974:SPU

[FMW74] R. Fateman, J. Moses, and P. Wang. Solution to problem #7 using MAC-SYMA. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(2): 14–16, May 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fitch:1973:ICD

[FN73] John Fitch and Arthur Norman. An idea for comparing data structures. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(27):17–18, September 1973. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fitch:1975:DSL

[FN75] John Fitch and Arthur Norman. Double by single length division. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(2):7, May 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fontein:2008:AIU

[Fon08] Felix Fontein. Abstract infrastructures of unit rank two (abstract only). ACM Communications in Computer Algebra, 42(1–2):55– 56, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Fateman:1989:SDS

[FP89] R. J. Fateman and C. G.

Ponder. Speed and data structures in computer algebra systems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23 (2):8–11, April 1989. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fassbender:2016:CSS

[FPS16] Heike Fassbender, Javier Pérez, and Nikta Shayanfar. Constructing symmetric structure-preserving strong linearizations. ACM Communications in Computer Algebra, 50(4):167– 169, December 2016. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Farnsworth:2018:HFV

[FR18] Cameron Farnsworth and Jose Israel Rodriguez. Homogenized funtf varieties and algebraic frame completion. ACM Communications in Computer Algebra, 52(3):108–111, September 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Frauenthal:1980:CAM

[Fra80] James C. Frauenthal. Change in applied mathematics is revolutionary. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14(2):23, May 1980.

CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Freitag:2015:BSF

[Fre15] James Freitag. Bounding the size of a finite differential algebraic variety. ACM Communications in Computer Algebra, 49(1): 23–24, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Friedman:1970:FIK

[Fri70] Joyce Friedman. Fortran implementation of the Kay context-free parser. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(16):19–45, October 1970. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Frick:1974:NNC

[Fri74] I. Frick. A non-numerical computation involving sums of many products of finite rotation matrices for spin 3/2. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8 (3):76–77, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Flajolet:1997:SCS

[FS97] Philippe Flajolet and Bruno Salvy. The SIGSAM challenges: symbolic asymptotics in practice. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(4):36–47, December 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Franz:2002:UCA

[FSH02] Astrid Franz, Christian Schulzky, and Karl Heinz Hoffmann. Using computer algebra methods to determine the chemical dimension of finitely ramified Sierpinski carpets. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36(2):18–30, June 2002. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Faugere:2015:SGB

[FSS15] Jean-Charles Faugère, Pierre-Jean Spaenlehauer, and Jules Svartz. Sparse Gröbner bases: algorithms and complexity. ACM Communications in Computer Algebra, 49(2):61–62, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Fassino:2012:ZDI

[FT12] C. Fassino and M. Torrente. Zero-dimensional ideals of limited precision points. ACM Communications in Computer Algebra, 46(3–4):90–91, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Fogel:2013:CGA

[FT13] Efi Fogel and Monique Teillaud. The Computational Geometry Algorithms Library CGAL. ACM Communications in Computer Algebra, 47(3–4):85– 87, September 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Fogel:2015:CGA

[FT15] Efi Fogel and Monique Teillaud. The computational geometry algorithms library CGAL. ACM Communications in Computer Algebra, 49(1):10–12, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Fuchssteiner:2008:MGA

[Fuc08] Benno Fuchssteiner. Modified Gauss algorithm for matrices with symbolic entries. ACM Communications in Computer Algebra, 42(3):108–121, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Fang:2010:AOS

[FvH10] Tingting Fang and Mark van Hoeij. Abstract only: Solving linear differential equations by using descent. ACM Communications in Computer Algebra, 44(2): 26, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gaeta:1999:NNR

[Gae99] F. Gaeta. New non recursive formulas for irreducible representations of GL(C^{m+1}) and systems of equations for the symmetric powers Sym^{n}P_{m}. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):15, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gakhov:1979:RCI

[Gak79] F. D. Gakhov. Review of “Calculus of Integrals of Higher Transcendental Functions (The Theory and Tables of Formulas), by O. I. Marichev”, Minsk, Nauka i Tekhnika, 1978. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 13(3):7, August 1979. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Galetto:2015:ADA

[Gal15] Federico Galetto. An algorithm for determining actions of semisimple Lie groups on free resolutions. ACM Communications in Computer Algebra, 49(2): 59, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gao:2015:BDI

[Gao15] Xiao-Shan Gao. Binomial difference ideal and toric difference variety. ACM Communications in Computer Algebra, 49(1):24, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gao:2017:WTW

[Gao17] Xiao-Shan Gao. Wen-Tsun Wu: His life and legacy. ACM Communications in Computer Algebra, 51(2): 73–79, June 2017. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gates:1985:GAC

[Gat85] Barbara L. Gates. GEN-TRAN: an automatic code generation facility for RE-DUCE. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19 (3):24–42, August 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gawlik:1977:FDM

[Gaw77] H. J. Gawlik. The further development of MIRA. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11(2):22– 28, May 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gawlik:1979:COL

[Gaw79] H. James Gawlik. Comments on the open letter from Fateman to Veltman. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 13(1):14–15, February 1979. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gerdt:1997:MIB

[GB97] Vladimir P. Gerdt and Yuri A. Blinkov. Minimal involutive bases. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):44, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Gonnet:1983:SGS

[GCG83] Gaston H. Gonnet, Bruce W.

Char, and Keith O. Geddes. Solution of a general system of equations (problem). SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(3–4):48–49, August/November 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Geddes:1999:GNO

[Ged99] K. O. Geddes. Generating numerical ODE formulas via a symbolic calculus of divided differences. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(2):29–42, June 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gentleman:1971:OMC

[Gen71] W. Morven Gentleman. Optimal multiplication chains for computing a power of a symbolic polynomial. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(18):23–30, April 1971. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gentleman:1974:ETP

[Gen74] W. Morven Gentleman. Experience with truncated power series. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8 (3):61–62, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gentle:1999:CGC

[Gen99] Cassidy Gentle. Computing greatest common divisors of polynomial matrices. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):26–27, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gerdt:1998:ISL

[Ger98] Vladimir P. Gerdt. Involutive systems of linear PDEs. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (2):55–56, June 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Abstract only.

Garcia:2015:UTM

[GGdR^{+}15] Alfonsa García, Francisco García, Ángel Martín del Rey, Gerardo Rodríguez, and Agustín de la Villa. Using technology in mathematical courses: some possibilities. ACM Communications in Computer Algebra, 49(2):49, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Garcia-Garcia:2021:CIS

[GGMAVT21] J. I. García-García, D. Marín-Aragón, and A. Vigneron-Tenorio. Computing the ideals of sumset semigroups. ACM Communications in Computer Algebra, 55(3): 73–76, September 2021. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3511528.3511531.

Garcia-Garcia:1999:CIE

[GGSR99] J. I. García-García, P. A. García Sánchez, and J. C. Rosales. Commutative ideal extensions of Abelian groups. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):16, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gao:2008:CSM

[GH08] Xiao-Shan Gao and Zhenyu Huang. A characteristic set method for equation solving over finite fields. ACM Communications in Computer Algebra, 42(3):149– 150, September 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gheffar:2010:LDD

[Ghe10] A. Gheffar. Linear differential, difference and q-difference homogeneous equations having no rational solutions. ACM Communications in Computer Algebra, 44(3):78–83, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gheffar:2014:DNR

[Ghe14] Amal Gheffar. Detecting nonexistence of rational solutions of linear difference equations in early stages of computation. ACM Communications in Computer Algebra, 48(3/4):90– 97, September 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gazor:2018:GBM

[GHK18] Majid Gazor, Amir Hashemi, and Mahsa Kazemi. Gröbner bases and multi-dimensional persistent bifurcation diagram classifications. ACM Communications in Computer Algebra, 52(4):120– 122, December 2018. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Giesbrecht:2015:FDD

[GHL15] Mark Giesbrecht, Albert Heinle, and Viktor Levandovskyy. On factoring differential and difference operators in n variables. ACM Communications in Computer Algebra, 49(1): 33, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gianni:1988:SAC

[Gia88] P. Gianni, editor. Symbolic and algebraic computation: International Symposium ISSAC ’88, Rome, Italy, July 4–8, 1988: Proceedings, volume 358 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-51084-2. LCCN QA76.95 .I571 1988.

Gianni:1989:SAC

[Gia89] P. (Patrizia) Gianni, editor. Symbolic and algebraic computation: International Symposium ISSAC ’88, Rome, Italy, July 4– 8, 1988: proceedings, volume 358 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51084-2. LCCN QA76.95 .I571 1988. Conference held jointly with AAECC-6.

Giesbrecht:1999:MEa

[Gie99a] Mark Giesbrecht. Message from the Editor. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):1, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Giesbrecht:1999:MEb

[Gie99b] Mark Giesbrecht. Message from the Editor. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(4):1, December 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Giesbrecht:1999:MIE

[Gie99c] Mark Giesbrecht. Message from the incoming Editor. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(2):2, June 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Giesbrecht:2008:MCP

[Gie08a]  Mack Giesbrecht.  The

MICA 2008 Conference, preface and abstracts. ACM Communications in Computer Algebra, 42(1–2):92– 102, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Giesbrecht:2008:AKA

[Gie08b] Mark Giesbrecht. ACM 2007 Kanellakis Award honours Bruno Buchberger, innovator of automated tools for mathematics. ACM Communications in Computer Algebra, 42(1–2):37– 38, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gimenez:2003:AES

[Gim03] Philippe Gimenez. Abstracts of the eighth Spanish meeting on computer algebra and applications: EACA-2002. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37 (1):16–27, March 2003. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Giusti:2001:AIS

[Giu01]  M. Giusti.  Announce-

ment: International Symposium on Symbolic and Algebraic Computation. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35(3):40, September 2001. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Griesmer:1974:SPL

[GJ74] James H. Griesmer and Richard D. Jenks. A solution to problem #4: the lie transform. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8 (4):12–13, November 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gerhard:2009:PSP

[GJM09] Jürgen Gerhard, D. J. Jeffrey, and Guillaume Moroz. A package for solving parametric polynomial systems. ACM Communications in Computer Algebra, 43(3):61–72, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Griesmer:1975:SSP

[GJY75a] J. H. Griesmer, R. D. Jenks, and D. Y. Y. Yun. A SCRATCHPAD solution to problem #7. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9 (3):13–17, August 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Griesmer:1975:FSS

[GJY75b] J. H. Griesmer, R. D. Jenks, and David Y. Y. Yun. A FORMAT statement in SCRATCHPAD. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(3):24–25, August 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Griesmer:1978:TAC

[GJY78] J. H. Griesmer, R. D. Jenks, and D. Y. Y. Yun. A taxonomy for algebraic computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 12 (3):25–28, August 1978. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gebauer:1984:NSG

[GK84] R. Gebauer and H. Kredel. Note on “solution of a general system of equations”. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18 (3):5–6, August 1984. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gebauer:1985:RSS

[GK85] R. Gebauer and H. Kredel. Real solution system for algebraic equations. an algorithm system for constructing finite inclusions of zero sets of algebraic equations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19 (2):2, May 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gregory:1988:ABC

[GK88] Brent Gregory and Erich Kaltofen. Analysis of the binary complexity of asymptotically fast algorithms for linear system solving. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22(2):41–49, April 1988. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gazor:2016:ZES

[GK16] Majid Gazor and Mahsa Kazemi. Z2-equivariant standard bases for submodules associated with Z2-equivariant singularities. ACM Communications in Computer Algebra, 50(4):170–172, December 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gerhard:2020:MC

[GK20] Jürgen Gerhard and Ilias S. Kotsireas. Maple conference 2019. ACM Communications in Computer Algebra, 53(4):182, March 2020. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.
                                        1145/3388974.3388976.

Giesbrecht:2007:IPA

[GKL07] Mark W. Giesbrecht, Ilias Kotsireas, and Austin Lobo. ISSAC 2007 poster abstracts. ACM Communications in Computer Algebra, 41(1–2):38–72, March– June 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gastineau:2010:TCA

[GL10] Mickaël Gastineau and Jacques Laskar. TRIP: a computer algebra system dedicated to celestial mechanics and perturbation series. ACM Communications in Computer Algebra, 44(4):194–197, December 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Giorgi:2013:ROB

[GL13] Pascal Giorgi and Romain Lebreton. Relaxing order basis computation. ACM Communications in Computer Algebra, 47(3–4):100– 101, September 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Glasby:1991:EGD

[Gla91] S. P. Glasby. Extensions of groups defined by powercommutator presentations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(1):32–35, January 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gleyse:1986:SCN

[Gle86] B. Gleyse. Symbolic computation and number of zeros of a real parameter polynomial in the unit disk. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 20(1–2):22–28, February/May 1986. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gleibman:1992:SNP

[Gle92] Andrew H. Gleibman. SAMPLE: New programming technology and AI language. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26 (3):21–29, August 1992. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gonzalez-Lopez:2000:GBS

[GLGVTZ00] M.-J. Gonzalez-Lopez, L. Gonzalez-Vega, C. Traverso, and A. Zanoni. Gröbner bases specialization through Hilbert functions: The homogeneous case. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):1–8, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gloor:1998:IPI

[Glo98a] Oliver Gloor, editor. IS-SAC 98: Proceedings of the 1998 International Symposium on Symbolic and Algebraic Computation, August 13–15, 1998, University of Rostock, Germany. ACM Press, New York, NY, USA, 1998. ISBN 1-58113-002-3. LCCN QA155.7.E4 E88 1998.

Glotov:1998:GCR

[Glo98b] Peter E. Glotov. On the greatest common right divisor of Ore polynomials with polynomial coefficients which depend on a parameter. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (2):57, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Abstract only.

Giesbrecht:1997:CIS

[GLS97] Mark Giesbrecht, Austin Lobo, and David Saunders. Certifying inconsistency of sparse linear systems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):40, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Gruntz:1994:IG

[GM94] D. Gruntz and M. Monagan. Introduction to Gauss. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 28 (2):3–19, August 1994. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gowda:2021:HPS

[GMC^{+}21] Shashi Gowda, Yingbo Ma, Alessandro Cheli, Maja Gwó’zzd’z, Viral B. Shah, Alan Edelman, and Christopher Rackauckas. Highperformance symbolic-numerics via multiple dispatch. ACM Communications in Computer Algebra, 55(3):92–96, September 2021. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3511528.3511535.

Guzide:2008:NAD

[GMGE08] Osman Guzide, Kenneth May, James Getz, and Adam Edgeson. New algebraic design for better P2P features (abstract only). ACM Communications in Computer Algebra, 42(1–2): 70–71, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gianni:1987:SAC

[GMM87] P. Gianni, A. Miola, and T. Mora. Symbolic and algebraic computation research in Italy. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21(2):8–15, February 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Garcia-Marco:2022:CGH

[GMMCMMP22] Ignacio García-Marco, Irene Márquez-Corbella, Edgar Martínez-Moro, and Yuriko Pitones. Computing generalized Hamming weights of binary linear codes via free resolutions. ACM Communications in Computer Algebra, 56(2):19– 24, June 2022. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572867.3572868.

Garcia-Morchon:2015:HCR

[GMRT^{+}15] Oscar García-Morchón, Ronald Rietman, Ludo Tolhuizen, Domingo Gómez, and Jaime Gutiérrez. HIMMO: a collusion-resistant identitybased scheme for symmetric key generation. ACM Communications in Computer Algebra, 49(1):19, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gobel:1993:UBA

[Göb93]  Manfred Göbel.  Using

Buchberger’s algorithm in invariant theory. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27(4):3–9, December 1993. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Golden:1985:DUF

[Gol85] Jeffrey P. Golden. Differentiation of unknown functions in Macsyma. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19(2):19–24, May 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gollan:1991:MRP

[Gol91] Holger W. Gollan. The 2modular representation of J_{1} in the principal block. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(1):36–39, January 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Golubitsky:2006:CRP

[Gol06] Oleg Golubitsky. Construction of rankings on partial derivatives. ACM Communications in Computer Algebra, 40(2):38–42, June 2006. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gonnet:1988:EMA

[Gon88] Gaston H. Gonnet. Examples of Maple applied to problems from the American Mathematical Monthly. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22(2):8–16, April 1988. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gonnet:1989:PAI

[Gon89] Gaston H. Gonnet, editor.

Proceedings of the ACM-SIGSAM 1989 International Symposium on Symbolic and Algebraic Computation: ISSAC ’89 / July 17–19, 1989, Portland, Oregon. ACM Press, New York, NY, USA, 1989. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

Gordan:1998:NPH

[Gor98]  Paul Gordan.  A new

proof of Hilbert’s theorem on homogeneous functions. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(2):47–48, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Translation by Michael Abramson of the original German article Neure Beweis des Hilbertschen Satzes über homogene Funktionen. Nachr. der Königsl. Ges. der Wiss. zu Göttingen 3 (1899), 240– 242.

Gordan:1999:IBF

[Gor99] P. Gordan. Invariants of binary forms. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):28–32, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Translation by Michael Abramson of the article Les invariants des formes binaires, originally published in J. de Mathématiques Pures et Appliqués, 6, (1900), 141-156.

Gosper:1998:S

[Gos98] R. William Gosper, Jr. The solutions of ye^{y^{2}} = x and ye^{y} = x. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (1):8–10, March 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Goyer:2021:SPS

[Goy21] Alexandre Goyer. A Sage package for the symbolicnumeric factorization of linear differential operators. ACM Communications in Computer Algebra, 55(2): 44–48, June 2021. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3493492.3493496.

Greuel:1997:SCA

[GPS97] G.-M. Greuel, G. Pfister, and H. Schönemann. Singular 1.0: a computer algebra system for singularity theory, algebraic geometry and commutative algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):48–50, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Greuel:2008:SCA

[GPS08] G. M. Greuel, G. Pfister, and H. Schönemann. SIN-GULAR: a computer algebra system for polynomial computations. ACM Communications in Computer Algebra, 42(3):180– 181, September 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gutierrez:1999:SPT

[GR99] J. Gutierrez and R. Rubio. Subfields in pure transcendental extensions. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):16, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Galbraith:2008:CPN

[GR08] Steven Galbraith and Raminder S. Ruprai. Computing L-polynomials of nonhyperelliptic genus 4 and 5 curves (abstract only). ACM Communications in Computer Algebra, 42(1–2): 56–57, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Groves:2016:SPF

[GR16] A. Whitman Groves and Daniel S. Roche. Sparse polynomials in FLINT. ACM Communications in Computer Algebra, 50(3): 105–108, September 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gray:1987:SCL

[Gra87] John W. Gray. The symbolic computation laboratory at UIUC. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21(3):9–11, August 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Grams:1991:SGR

[Gra91] Gerhard Grams. Special generators and relations for some orthogonal and symplectic groups over GF(2). SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(1):39–45, January 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Graillat:2005:NNP

[Gra05] Stef Graillat. A note on a nearest polynomial with a given root. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39 (2):53–60, June 2005. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Grabe:2018:YS

[Grä18]  Hans-Gert Gräbe.  20

years SymbolicData. ACM Communications in Computer Algebra, 52(3):45–54, September 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Greenawalt:1972:OFL

[Gre72] E. M. Greenawalt. An overlay facility for LISP programs. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (21):27–31, January 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Grenet:2015:LCB

[Gre15] Bruno Grenet. Lacunaryx: computing bounded-degree factors of lacunary polynomials. ACM Communications in Computer Algebra, 49(4):121–124, December 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Griesmer:1976:SMC

[Gri76] James H. Griesmer. Symbolic mathematical computation — a survey. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 10(2):30–32, ???? 1976. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Griesmer:1979:SSC

[Gri79] James H. Griesmer. The state of symbolic computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 13 (3):25–28, August 1979. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Grossman:1995:IIC

[Gro95a] Robert Grossman. 1st International IMACS Conference on Applications of Computer Algebra. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(3–4):9–12, December 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Grossman:1995:ME

[Gro95b] Robert Grossman. Message from the Editor. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(3–4):1, December 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Grobner:1998:ET

[Grö98] W. Gröbner. On elimination theory. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (2):40–46, June 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Translation by Michael Abramson of the original German article Über die Eliminationstheorie, Monatshefte für Mathematik 5 (1950), 71–78.

Group:1999:CEM

[Gro99] Lemme Group. CtCoq: an environment for mathematical reasoning. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):21–22, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Grobner:2009:API

[Grö09] Wolfgang Gröbner. On the algebraic properties of integrals of linear differential equations with constant coefficients. ACM Communications in Computer Algebra, 43(2):24–46, June 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Grobner:2010:MIS

[Grö10] Wolfgang Gröbner. On the Macaulay inverse system and its importance for the theory of linear differential equations with constant coefficients. ACM Communications in Computer Algebra, 44(2):20–23, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Giesbrecht:2010:AOC

[GRT10] Mark Giesbrecht, Daniel S.

Roche, and Hrushikesh Tilak. Abstract only: Complexity of sparsest multiple computation. ACM Communications in Computer Algebra, 44(2):25, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Garcia-Sanchez:1995:GIB

[GS95] Pedro A. García-Sánchez.

Gröbner and involutive bases for zero-dimensional ideals. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29 (2):12–15, April 1995. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Giesbrecht:1997:PLS

[GS97] Mark Giesbrecht and David Saunders. Parametric linear systems: The twoparameter case. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):40, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Guo:2015:RBT

[GSGZ15] Li Guo, William Sit, Xing Gao, and Shanghua Zheng.

Rota–Baxter type operators, rewriting systems and Gröbner–Shirshov bases. ACM Communications in Computer Algebra, 49(1): 33, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gomez-Torrecillas:2014:GII

[GTLN14]  José  Gómez-Torrecillas,

F. J. Lobillo, and Gabriel Navarro. Generating idempotents in ideal codes. ACM Communications in Computer Algebra, 48(3/4):113– 115, September 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gomez-Torrecillas:2018:MDC

[GTLN18]  José  Gómez-Torrecillas,

F. J. Lobillo, and Gabriel Navarro. Minimum distance computation of linear codes via genetic algorithms with permutation encoding. ACM Communications in Computer Algebra, 52(3):71–74, September 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gutierrez:2004:IJU

[Gut04] Jaime Gutierrez, editor. ISAAC 2004: July 4– 7, 2004, University of Cantabria, Santander, Spain: proceedings of the 2004 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY, USA, 2004. ISBN 1-58113-827-X.

Gutierrez:2015:RZH

[Gut15]  J. Gutierrez.  Recover-

ing zeroes of hyperelliptic curves over finite fields. ACM Communications in Computer Algebra, 49(2): 58, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gonzalez-Vega:2004:RCC

[GVR04] Laureano González-Vega and Tomás Recio. Review of “Computational Commutative Algebra I” by Martin Kreuzer and Lorenzo Robbiano, Springer-Verlag, 2000, ISBN 3-540-67733-X. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(2):53– 55, June 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gao:2005:IDM

[GZ05a] Shuhong Gao and Mingfu Zhu. Irreducible decomposition of monomial ideals. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(3):99, 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). ISSAC 2005 poster abstract.

Gao:2005:GSP

[GZ05b] Xiao-Shan Gao and Gui-Fang Zhang. 2D and 3D generalized Stewart Platforms. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39 (3):92–93, 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). ISSAC 2005 poster abstract.

Gao:2008:IAI

[GZ08] Shuhong Gao and Mingfu Zhu. An incremental algorithm for irreducible decompositions of monomial ideals (abstract only). ACM Communications in Computer Algebra, 42(1–2):70, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hagen:1977:SP

[Hag77] G. Tim Hagen. A simple problem. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11 (2):20–21, May 1977. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hall:1974:FRE

[Hal74] Andrew D. Hall, Jr. Factored rational expressions in ALTRAN. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8 (3):35–45, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hall:1975:ASP

[Hal75] A. D. Hall. ALTRAN solution to problem #8. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(4):39–40, November 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hamada:2014:MPC

[Ham14] Tatsuyoshi Hamada. Math-Libre: Personalizable computer environment for mathematical research. ACM Communications in Computer Algebra, 48(3/4):116– 117, September 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Harrington:1977:RSP

[Har77] S. J. Harrington. RE-DUCE solution to problem #8. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11– 12(4–1):7–8, November and February 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Harrington:1979:SLE

[Har79] Steven J. Harrington. A symbolic limit evaluation program in REDUCE. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 13(1):27–31, February 1979. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hare:2000:RCB

[Har00] K. Hare. Rapid computation of Bernoulli and related numbers. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):18–19, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hunt:1997:RPS

[HC97] Kenny Hunt and James Cremer. Refiner: a problem solving environment for ODE/DAE simulations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):42–43, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Hearn:1969:SLR

[Hea69] Anthony C. Hearn. Standard LISP (reprint). SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(13):28–49, December 1969. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hearn:1972:R

[Hea72a] A. C. Hearn. REDUCE.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(24):14–15, October 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hearn:1972:INM

[Hea72b] Anthony C. Hearn. An improved non-modular polynomial GCD algorithm. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(23):10– 15, July 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hearn:1972:RSP

[Hea72c] Anthony C. Hearn. A RE-DUCE solution of problem #2 — the Y_{2n} functions. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14(??):14–??, ???? 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hearn:2009:RFS

[Hea09] Anthony C. Hearn. RE-DUCE is free software as of January 2009. ACM Communications in Computer Algebra, 43(1):15–16, March 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hebisch:2015:ITE

[Heb15] Waldemar Hebisch. Integration in terms of exponential integrals and incomplete gamma functions. ACM Communications in Computer Algebra, 49(3):98– 100, September 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Heck:1994:RHW

[Hec94] André Heck. Report on the HISC workshop. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 28(1):2–8, March 1994. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Heindel:1970:AEP

[Hei70] Lee E. Heindel. Algorithms for exact polynomial root calculation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (16):9, October 1970. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Heinle:2013:RFC

[Hei13] Albert Heinle. Review of “A first course in computational algebraic geometry” by W. Decker and G. Pfister. ACM Communications in Computer Algebra, 47(1– 2):62–63, March 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Heiderich:2015:TNC

[Hei15] Florian Heiderich. Towards a non-commutative Picard– Vessiot theory. ACM Communications in Computer Algebra, 49(1):24, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hellstrom:2010:VVI

[Hel10] Christian Hellström. VarInt:

variational integrator design with Maple. ACM Communications in Computer Algebra, 44(3):115, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Helmer:2015:ACC

[Hel15]  Martin Helmer.  Algo-

rithms to compute Chern– Schwartz–Macpherson and Segrè classes and the Euler characteristic. ACM Communications in Computer Algebra, 49(2):56, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hermann:1998:QFM

[Her98]  Grete Hermann.  The

question of finitely many steps in polynomial ideal theory. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (3):8–30, September 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Translation by Michael Abramson of the article Die Frage der endlich vielen Schritte in der Theorie der Polynomideale, originally published in Math. Ann. 95 (1926), 736–788 (Received 29 May 1925).

Heras:2011:MKM

[Her11] Jónathan Heras. Mathematical knowledge management in algebraic topology. ACM Communications in Computer Algebra, 45 (3–4):236–237, September 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Huang:2014:CTH

[HEW^{+}14] Zongyan Huang, Matthew England, David Wilson, James H. Davenport, and Lawrence C. Paulson. A comparison of three heuristics to choose the variable ordering for cylindrical algebraic decomposition. ACM Communications in Computer Algebra, 48(3/4):121– 123, September 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Harkonen:2020:AOD

[HHKR20] Marc Härkönen, Benjamin Hollering, Fatemeh Tarashi Kashani, and Jose Israel Rodriguez. Algebraic optimization degree. ACM Communications in Computer Algebra, 54(2):44–48, September 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3427218.3427222.

Haque:2010:CFS

[HHM10] Sardar Anisul Haque, Shahadat Hossain, and M. Moreno Maza. Cache friendly sparse matrix-vector multiplication. ACM Communications in Computer Algebra, 44(3):111–112, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hascoet:2012:SCO

[HHS12] Laurent Hascoët, Shahadat Hossain, and Trond Steihaug. Structured computation in optimization and algorithmic differentiation. ACM Communications in Computer Algebra, 46(3–4):94–95, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Harris:2020:SPS

[HHS20] Katherine Harris, Jonathan D. Hauenstein, and Agnes Szanto. Smooth points on semi-algebraic sets. ACM Communications in Computer Algebra, 54(3):105– 108, September 2020. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3457341.3457347.

Halstead:1993:PSC

[HI93] Robert H. Halstead and Takayasu Ito, editors. Parallel symbolic computing: languages, systems, and applications: US/Japan workshop, Cambridge, MA, USA, October 14–17, 1992: proceedings, number 748 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-57396-8, 3-540-57396-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.P3785 1993.

Hiemstra:1975:PEC

[Hie75] Bauke Hiemstra. A preeditor for CAMAL. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(3):30–34, August 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hirschberg:1978:EMI

[Hir78] Morton A. Hirschberg. The evaluation, manipulation and identification of nondimensional numbers. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 12(2):21–26, May 1978. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hitz:2007:EPA

[Hit07] Markus A. Hitz. ECCAD 2007 poster abstracts. ACM Communications in Computer Algebra, 41(1–2):12– 24, March–June 2007. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hall:1974:APS

[HJ74] A. D. Hall and S. C. Johnson. ALTRAN programs for SIGSAM problem #6. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(2):12–36, May 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

He:2023:MMP

[HJN^{+}23] Yang-Hui He, Vishnu Jejjala, Brent D. Nelson, Hal Schenck, and Michael Stillman. MSSM: a Macaulay2 package for the vacuum moduli space. ACM Communications in Computer Algebra, 57(2):39– 42, June 2023. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3614408.3614412.

Harrison:2013:PAW

[HJS13] Gavin Harrison, Jeremy Johnson, and B. David Saunders. Probabilistic analysis of Wiedemann’s algorithm for minimal polynomial computation. ACM Communications in Computer Algebra, 47(3–4):118– 119, September 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Harrison:2016:PAB

[HJS16] Gavin Harrison, Jeremy Johnson, and B. David Saunders. Probabilistic analysis of block Wiedemann for leading invariant factors. ACM Communications in Computer Algebra, 50(4):173–175, December 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hubert:2005:RRI

[HK05] Evelyne Hubert and Irina Kogan. Rational and replacement invariants of a group action. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39 (3):85–87, 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). ISSAC 2005 poster abstract.

Hamada:2008:KML

[HK08] Tatsuyoshi Hamada and KNOPPIX/Math committers. KNOPPIX/Math: a live system for enjoying mathematics with computer. ACM Communications in Computer Algebra, 42(3):175–176, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Heckler:1997:PM

[HKM^{+}97] Christian Heckler, Oliver Kluge, Torsten Metzner, Andreas Sorgatz, and Paul Zimmermann. Parallelism in MuPAD. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):48, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Haftmann:2010:CBP

[HKN10] Florian Haftmann, Cezary Kaliszyk, and Walther Neuper. CTP-based programming languages?: considerations about an experimental design. ACM Communications in Computer Algebra, 44(2):27–41, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hong:1999:ECC

[HKS99] Hoon Hong, Erich Kaltofen, and Michael Singer. East Coast Computer Algebra Day ’99 (April 24, 1999): Abstracts of Invited Talks and Presented Papers. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(2):43–52, June 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Heinle:2017:SSI

[HKS17] Albert Heinle, Wolfram Koepf, and Wolfram Sperber. Some steps to improve software information. ACM Communications in Computer Algebra, 51(1): 1–11, March 2017. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Heule:2019:FSM

[HKS19] Marijn Heule, Manuel Kauers, and Martina Seidl. A family of schemes for multiplying 3 × 3 matrices with 23 coefficient multiplications. ACM Communications in Computer Algebra, 53(3):118–121, September 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Holden:2008:SLM

[HL08] Joshua Holden and Nathan Lindle. A statistical look at maps of the discrete logarithm (abstract only). ACM Communications in Computer Algebra, 42(1–2):57– 59, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Heinle:2010:FPG

[HL10] Albert Heinle and Viktor Levandovskyy. Factorization of polynomials in Zgraded skew polynomial rings. ACM Communications in Computer Algebra, 44(3):113–114, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Heinle:2015:SBT

[HL15] Albert Heinle and Viktor Levandovskyy. The SDeval benchmarking toolkit. ACM Communications in Computer Algebra, 49(1): 1–9, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Haque:2017:CCL

[HLM^{+}17] S. Anisul Haque, X. Li, F. Mansouri, M. Moreno Maza, D. Mohajerani, and W. Pan. CUMODP: a CUDA library for modular polynomial computation. ACM Communications in Computer Algebra, 51(3):89–91, September 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hosack:1984:RUS

[HLS84] John Hosack, Kenneth Lane, and Donald Small. Report on the use of a symbolic mathematics system in undergraduate instruction. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18– 19(4–1):19–22, November/ February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hall:1977:MDC

[HM77] David L. Hall and B. Craig Meyers. A method for determining the coefficients of a reversed power series.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11–12(4–1):16– 19, November and February 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hernandez:1999:USE

[HM99] X. Hernández and J. M.

Miret. Using Schubert to enumerate conics in P^{3}. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):16, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hu:2013:PAC

[HM13] Jiaxiong Hu and Michael Monagan. A parallel algorithm to compute the greatest common divisor of sparse multivariate polynomials. ACM Communications in Computer Algebra, 47(3–4):108–109, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hubert:2022:OSM

[HMMR22] Evelyne Hubert, Tobias Metzlaff, Philippe Moustrou, and Cordian Riener. T -orbit spaces of multiplicative actions and applications. ACM Communications in Computer Algebra, 56(2):72–75, June 2022. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3572867.3572879.

Hodorog:2010:GLG

[HMS10] Mădălina Hodorog, Bernard Mourrain, and Josef Schicho. GENOM3CK: a library for genus computation of plane complex algebraic curves using knot theory. ACM Communications in Computer Algebra, 44(4):198–200, December 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hyun:2018:FAS

[HMSP18] Seung Gyu Hyun, Stephen Melczer, and Catherine St-Pierre. A fast algorithm for solving linearly recurrent sequences. ACM Communications in Computer Algebra, 52(3):100–103, September 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Huang:2019:AAC

[HN19] Bo Huang and Wei Niu. Algebraic approach to chaos induced by snapback repeller. ACM Communications in Computer Algebra, 53(3):122–125, September 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hyun:2018:SFU

[HNRS18] Seung Gyu Hyun, Vincent Neiger, Hamid Rahkooy, and Éric Schost. Sparse FGLM using the block Wiedemann algorithm. ACM Communications in Computer Algebra, 52(4):123– 125, December 2018. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hibi:2014:CBT

[HNT14] Takayuki Hibi, Kenta Nishiyama, and Nobuki Takayama. Computing bases of twisted cohomology groups for sparse polynomials. ACM Communications in Computer Algebra, 48(3/4):118– 120, September 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hodorog:2011:SNA

[Hod11] Madalina Hodorog. Symbolicnumeric algorithms for plane algebraic curves. ACM Communications in Computer Algebra, 45(3–4): 239–240, September 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hohti:1988:CAA

[Hoh88] Aarno Hohti. Computational algebra with APL. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22(1):12–20, January 1988. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hollenhorst:1992:DDF

[Hol92] Manfred Hollenhorst. The design of digital filters using REDUCE. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26 (1):10–12, January 1992. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Holden:2004:DEE

[Hol04] Joshua Holden. Distribution of the error in estimated numbers of fixed points of the discrete logarithm. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(4):111–118, December 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hong:1994:FIS

[Hon94] Hoon Hong, editor. First International Symposium on Parallel Symbolic Computation, PASCO ’94, Hagenberg/Linz, Austria, September 26–28, 1994, volume 5 of Lecture notes series in computing. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1994. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Hong:1996:IA

[Hon96a] H. Hong. IMACS-ACA ’96.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(1):61, March 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hong:1996:P

[Hon96b] Hoon Hong. PASCO ’97.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(3):31, September 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hong:1996:PSI

[Hon96c] Hoon Hong. PASCO ’97:

Second International Symposium on Parallel Symbolic Computation. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(2):47–48, June 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hong:2019:STA

[HOPY19] Hoon Hong, Alexey Ovchinnikov, Gleb Pogudin, and Chee Yap. SIAN: a tool for assessing structural identifiability of parametric ODEs. ACM Communications in Computer Algebra, 53(2): 37–40, June 2019. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Horowitz:1972:S

[Hor72] E. Horowitz. SAC-1. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(24):22–24, October 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hornecker:2015:CMC

[Hor15] Achim Hornecker. Cloud-Math: mathematical collaboration within the cloud. ACM Communications in Computer Algebra, 49(4): 125–126, December 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Howard:1970:RAS

[How70] James C. Howard. Relativistic applications of symbolic mathematical computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??

(15):64–83, July 1970. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Heindel:1967:AOS

[HP67] Lee E. Heindel and Paul W.

Purdom, Jr. The automatic optimization of SLIP routines. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (8):21–30, December 1967. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hu:2018:ARD

[HP18] Jiaxiong (Lucas) Hu and Anton Panferov. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 52(4):154–155, December 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hofstadler:2019:COI

[HRR19] Clemens Hofstadler, Clemens G.

Raab, and Georg Regensburger. Certifying operator identities via noncommutative Gröbner bases. ACM Communications in Computer Algebra, 53(2): 49–52, June 2019. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Husberg:1974:APF

[HS74] Nisse Husberg and Jouko Seppänen. ANALITIK: principal features of the language and its implementation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8 (3):24–25, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hong:1992:PP

[HS92] Hoon Hong and Wolfgang Schreiner. Programming in PACLIB. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26 (4):1–6, November 1992. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Havel:1989:PGA

[HSW89] T. F. Havel, B. Sturmfels, and N. White. Proposal for a geometric algebra software package. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23(1):13–15, January 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Huber:2000:CGF

[Hub00] Birkett Huber. Computing Gröbner fans of toric ideals.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):19–20, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Huet:1984:FSC

[Hue84] Gérard Huet. The future of symbolic computation: mathematics versus languages. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18 (2):2–5, May 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hufford:2015:SRF

[Huf15] James Hufford. Some results on finite fields. ACM Communications in Computer Algebra, 49(1):20, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hulshof:1983:SRF

[Hv83] B. J. A. Hulshof and J. A.

van Hulzen. Some RE-DUCE facilities for pretty printing subscripts and formal derivatives. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(1):16–20, February 1983. CODEN SIGSBZ.

ISSN 0163-5824 (print), 1557-9492 (electronic).

Heiligman:1998:CWP

[HV98] Mark Heiligman and Emil Volcheck. On computing the Weierstrass points of a plane algebraic curve. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(2):58, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Abstract only.

Heintz:1975:EQE

[HW75] J. Heintz and R. Wüthrich. An efficient quantifier elimination algorithm for algebraically closed fields of any characteristic. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(4):11, November 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hirata:2015:SEM

[HYH15] Takuya Hirata, Kazuya Yamaguchi, and Ichijo Hodaka. A symbolic equation modeler for electric circuits. ACM Communications in Computer Algebra, 49(3):80–82, September 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Huang:2015:TAC

[HZ15] Guanli Huang and Meng Zhou. On the termination of algorithm for computing relative Gröbner bases. ACM Communications in Computer Algebra, 49(1): 28, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ibanez:2008:HES

[Ibá08] Ana Romero Ibáñez Ibáñez.

Homologa efectiva y sucesiones espectrales (abstract only). ACM Communications in Computer Algebra, 42(1–2):21, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ida:2005:MPE

[Ida05] Tetsuo Ida. In memoriam:

Professor Eiichi Goto. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(3):71–72, September 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Iglesias:2021:CRP

[IdC21] Rodrigo Iglesias and Eduardo Sáenz de Cabezón. Cellular reductions of the Pommaret–Seiler resolution for quasi-stable ideals. ACM Communications in Computer Algebra, 55(3): 102–106, September 2021. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3511528.3511537.

Imamoglu:2020:NSP

[IK20] Erdal Imamoglu and Erich L. Kaltofen. A note on sparse polynomial interpolation in Dickson polynomial basis. ACM Communications in Computer Algebra, 54(4): 125–128, December 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3465002.3465003.

Ilyin:1989:FAC

[IKRT89] V. A. Ilyin, A. P. Kryukov, A. Ya. Rodioniov, and A. Yu. Taranov. Fast algorithm for calculation of Dirac’s gamma-matrices traces. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23 (4):15–24, October 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lee:2015:SMW

[iL15] Dong il Lee. Standard monomials for the Weyl group F_{4}. ACM Communications in Computer Algebra, 49(3):74–76, September 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ilie:2008:CCN

[Ili08] Silvana Ilie. Computational complexity of numerical solutions of initial value problems for differential algebraic equations (abstract only). ACM Communications in Computer Algebra, 42(1–2):21–22, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Iravanian:2022:SNI

[IMC^{+}22] Shahriar Iravanian, Carl Julius Martensen, Alessandro Cheli, Shashi Gowda, Anand Jain, Yingbo Ma, and Chris Rackauckas. Symbolicnumeric integration of univariate expressions based on sparse regression. ACM Communications in Computer Algebra, 56(2):84– 87, June 2022. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572867.3572882.

Inoue:2008:CIR

[IMSS08] Shutaro Inoue, Kohei Mukaida, Akira Suzuki, and Yosuke Sato. Computation of inverses in residue class rings with parameters. ACM Communications in Computer Algebra, 42(3):140– 141, September 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Inaba:2005:FMP

[Ina05] Daiju Inaba. Factorization of multivariate polynomials by extended Hensel construction. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39 (1):2–14, March 2005. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ilmer:2021:MAS

[IOP21] Ilia Ilmer, Alexey Ovchinnikov, and Gleb Pogudin. Maple application for structural identifiability analysis of ODE models. ACM Communications in Computer Algebra, 55(2):49– 53, June 2021. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3493492.3493497.

Inaba:2006:VEH

[IOS06] Daiju Inaba, Takeshi Osoekawa, and Tateaki Sasaki. Visualization of extended Hensel factors. ACM Communications in Computer Algebra, 40(2):44–45, June 2006. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Islam:2011:CQC

[IP11] Md. Nazrul Islam and A. Poteaux. Connectivity queries on curves in R_{n}. ACM Communications in Computer Algebra, 45(2): 117–118, June 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Inoue:2009:IBG

[IS09] Shutaro Inoue and Yosuke Sato. Implementation of Boolean Gröbner bases in Risa/Asir. ACM Communications in Computer Algebra, 43(3):83–84, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ilie:2016:AEC

[IS16] Silvana Ilie and Arne Storjohann. Abstracts of the 2015 East Coast Computer Algebra Day. ACM Communications in Computer Algebra, 50(1):35–39, March 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

SIGSAM:1989:PAI

[ISS89] Proceedings of the ACM-SIGSAM 1989 International Symposium on Symbolic and Algebraic Computation: ISSAC ’89, July 17– 19, 1989, Portland, Oregon. ACM Press, New York, NY, USA, 1989. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$22.00. English and French. Includes bibliographical references and index.

SIGSAM:1994:IPI

[ISS94] ISSAC ’94: Proceedings of the 1994 International Symposium on Symbolic and Algebraic Computation: July 20–22, 1994, Oxford, England, United Kingdom. ACM Press, New York, NY, USA, 1994. ISBN 0-89791-638-7. LCCN QA76.95 .I59 1994.

Irtegov:2012:SAD

[IT12] V. D. Irtegov and T. N. Titorenko. Symbolic analysis of dynamic systems. ACM Communications in Computer Algebra, 46(3– 4):98–99, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Iglesias:1999:CET

[IVV99] M. T. Iglesias, C. Vidal, and A. Verschoren. Computing epistasis through Walsh transforms. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):16–17, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Iwami:2005:EEB

[Iwa05] Maki Iwami. Extension of expansion base algorithm for multivariate analytic factorization including the case of singular leading coefficient. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(4):122–126, December 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Iwami:2008:AIA

[Iwa08] Maki Iwami. An attack on improved algebraic surface public-key cryptosystem (abstract only). ACM Communications in Computer Algebra, 42(1–2):71– 74, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Izumi:2014:MPI

[Izu14] Hideaki Izumi. Maple package itsol for formal solutions of iterative functional equations. ACM Communications in Computer Algebra, 48(3/4):192– 196, September 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jain:2008:MHR

[Jai08] Sonal Jain. Minimal heights and regulators for elliptic surfaces (abstract only). ACM Communications in Computer Algebra, 42(1–2):

59–61, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jalabert:1999:UNP

[Jal99] M. Lejeune Jalabert. On the ubiquity of Newton polyhedra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):12, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jaroschek:2012:IPR

[Jar12] Maximilian Jaroschek. Improved polynomial remainder sequences for ore polynomials. ACM Communications in Computer Algebra, 46(3–4):100–101, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jaroschek:2015:ROP

[Jar15] Maximilian Jaroschek. Radicals of Ore polynomials. ACM Communications in Computer Algebra, 49(1): 24, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Joyner:2011:OSC

[JČMG11a] David Joyner, Ondřej Čertík, Aaron Meurer, and Brian E. Granger. Open source computer algebra systems: SymPy. ACM Communications in Computer Algebra, 45(3–4):225– 234, September 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Joyner:2011:RDJ

[JCMG11b] David Joyner, Ondrej Certík, Aaron Meurer, and Brian E. Granger. Richard Dimick Jenks Memorial Prize 2011. ACM Communications in Computer Algebra, 45(3–4):235, September 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jovanovic:2012:SNL

[JdM12] Dejan Jovanović and Leonardo de Moura. Solving nonlinear arithmetic. ACM Communications in Computer Algebra, 46(3–4):104– 105, September 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jebelean:2018:TNS

[Jeb18] Tudor Jebelean. Techniques for natural-style proofs in elementary analysis. ACM Communications in Computer Algebra, 52(3):92–95, September 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jeffreys:1972:T

[Jef72] W. H. Jeffreys. TRIGMAN. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(24):20–21, October 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jeffrey:1998:FRA

[Jef98a] D. J. Jeffrey. Formally reviewed articles in the Bulletin. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (2):5–6, June 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jeffrey:1998:IVC

[Jef98b] D. J. Jeffrey. ISSAC ’99 Vancouver call for papers: International Symposium on Symbolic and Algebraic Computation, 28–31 July 1999, Simon Fraser University, British Columbia, Canada. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (2):63–64, June 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jeffrey:2003:ME

[Jef03] David J. Jeffrey. Message from the Editor. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(2):63, June 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jeffrey:2010:LFN

[Jef10] D. J. Jeffrey. LU factoring of non-invertible matrices. ACM Communications in Computer Algebra, 44(1):1–8, March 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jenks:1970:MLI

[Jen70]  R. D. Jenks.  META/

LISP: an interactive translator writing system. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(16):10–18, October 1970. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jenks:1972:S

[Jen72] R. D. Jenks. SCRATCH-PAD. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (24):16–17, October 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jenks:1974:PER

[Jen74a] R. D. Jenks, editor. Proceedings of Eurosam ’74, Royal Institute of Technology, Stockholm, Sweden, August 1–2, 1974. ACM Press, New York, NY, USA, 1974. Published in ACM SIGSAM Bulletin, volume 8, number 3.

Jenks:1974:SL

[Jen74b] R. D. Jenks. The SCRATCH-PAD language. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(2):20–30, May 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jenks:1975:COY

[Jen75] Richard D. Jenks. Course outline: Yale University, New Haven. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9 (3):9–10, August 1975. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jenks:1976:SPA

[Jen76] Richard D. Jenks, editor. Symsac ’76: proceedings of the 1976 ACM Symposium on Symbolic and Algebraic Computation, August 10–12, 1976, Yorktown Heights, New York. ACM Press, New York, NY, USA, 1976. LCCN QA155.7.E4 A15 1976. US$20.00.

Jenks:1977:DMB

[Jen77] Richard D. Jenks. On the design of a mode-based symbolic system. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11(1):16–19, February 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jenks:1979:SRL

[Jen79a] R. D. Jenks. SCRATCH-PAD/360: reflections on a language design. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 13(1):16–26, February 1979. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jensen:1979:FSS

[Jen79b] Jarl Jensen. On the formulation and solution of shell problems by means of computerized algebraic manipulation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 13 (2):13, May 1979. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jensen:2015:BNT

[Jen15]  David Jensen.  Brill–

Noether theory for metric graphs. ACM Communications in Computer Algebra, 49(2):60, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jeronimo:2015:EDL

[Jer15] Gabriela Jeronimo. Effective differential Lüroth theorem. ACM Communications in Computer Algebra, 49(1):24–25, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Johnson:1974:P

[JG74] S. C. Johnson and R. L. Graham. Problem #7. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(1):4, February 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jiang:2012:ACE

[JGB12] Hao Jiang, Stef Graillat, and Roberto Barrio. Accurate computing elementary symmetric functions. ACM Communications in Computer Algebra, 46(3–4):102– 103, September 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jeffrey:2002:TCE

[JK02] D. J. Jeffrey and Ilias S.

Kotsireas. Table of conferences and events. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36(1):17, March 2002. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Joyner:1998:MPD

[JM98] David Joyner and Roland Martin. A Maple package for the decomposition of certain tensor products and restrictions of representations using crystal graphs. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (2):19–22, June 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Joyner:2008:SCT

[JM08] David Joyner and Robert Miller. SAGE and coding theory (abstract only). ACM Communications in Computer Algebra, 42(1–2): 74–78, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Javadi:2010:SIF

[JM10] Seyed Mohammad Mahdi Javadi and Michael Monagan. On sparse interpolation over finite fields. ACM Communications in Computer Algebra, 44(3):116– 117, September 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jing:2017:PLM

[JM17] Rui-Juan Jing and Marc Moreno Maza. The polyhedra library in Maple. ACM Communications in Computer Algebra, 51(3):86–88, September 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jing:2018:CIP

[JM18] Rui-Juan Jing and Marc Moreno Maza. Computing the integer points of a polyhedron. ACM Communications in Computer Algebra, 52(4):126–129, December 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jinadu:2022:NIA

[JM22] Ayoola Jinadu and Michael Monagan. A new interpolation algorithm for computing Dixon resultants. ACM Communications in Computer Algebra, 56(2): 88–91, June 2022. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572867.3572883.

Jeannerod:2009:ECA

[JMV09] Claude-Pierre Jeannerod, Christophe Mouilleron, and Gilles Villard. Extending Cardinal’s algorithm to a broader class of structured matrices. ACM Communications in Computer Algebra, 43(3):84–85, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jeffrey:2004:SRB

[JN04] D. J. Jeffrey and A. C. Norman. Not seeing the roots for the branches: multivalued functions in computer algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38 (3):57–66, September 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Johnson:1974:SPA

[Joh74] Stephen C. Johnson. Sparse polynomial arithmetic. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(3):63–71, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Johnson:1975:NAP

[Joh75] S. C. Johnson. A note on abnormal polynomial remainder sequences. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(3):28–29, August 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Johnson:2010:ARD

[Joh10] Jeremy Johnson. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 44(3): 155, September 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Johnson:2011:ARDa

[Joh11a] Jeremy Johnson. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 45 (1):62, March 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Johnson:2011:ARDb

[Joh11b] Jeremy Johnson. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 45(2): 143–145, June 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Johnson:2012:ARDa

[Joh12a] Jeremy Johnson. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 46(1–2):14–16, March/June 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Johnson:2012:ARDb

[Joh12b] Jeremy Johnson. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 46(1–2):60–63, March/June 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Johnson:2012:ARDc

[Joh12c] Jeremy Johnson. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 46 (3–4):127–129, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Johnson:2012:ARDd

[Joh12d] Jeremy Johnson. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 46 (3–4):176–179, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Johnson:2012:MWK

[Joh12e] Jeremy Johnson. In memoriam: Werner Krandick. ACM Communications in Computer Algebra, 46(1–2): 52–54, March/June 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Johansson:2013:ACL

[Joh13a] Fredrik Johansson. Arb: a C library for ball arithmetic. ACM Communications in Computer Algebra, 47(3–4):166–169, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Johnson:2013:ARD

[Joh13b] Jeremy Johnson. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 47(1–2): 19–23, March 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jones:2008:RAA

[Jon08] Josh Jones. The RSA algorithm (abstract only). ACM Communications in Computer Algebra, 42(1–2):74, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Joyner:2005:CPD

[Joy05a]  David Joyner.  Conjec-

tural permutation decoding of some AG codes. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(1):26–32, March 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Joyner:2005:GEC

[Joy05b] David Joyner. GUAVA:

an error-correcting codes package. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39 (2):65–68, June 2005. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Joyner:2006:OM

[Joy06] David Joyner. OSCAS:

maxima. ACM Communications in Computer Algebra, 40(3–4):108–111, September–December 2006. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Joyner:2008:OSC

[Joy08] David Joyner. Open source computer algebra systems: Axiom. ACM Communications in Computer Algebra, 42(1–2):39–47, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Joyner:2009:ASS

[Joy09a] David Joyner. AMS special session on SAGE and mathematical research using open source software. ACM Communications in Computer Algebra, 43(2): 49–54, June 2009. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Joyner:2009:OSC

[Joy09b] David Joyner. Open source computer algebra systems: GAP. ACM Communications in Computer Algebra, 43(4):110–118, December 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Joyner:2015:EWC

[Joy15]  David Joyner.  Edge-

weighted Cayley graphs, monotonicity and bent functions. ACM Communications in Computer Algebra, 49(1):20–21, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Joswig:2011:DPC

[JP11] Michael Joswig and Andreas Paffenholz. Defect polytopes and counterexamples with polymake. ACM Communications in Computer Algebra, 45(3–4): 177–179, September 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jurrius:2015:EGR

[JP15] Relinde Jurrius and Ruud Pellikaan. The extended and generalized rank weight enumerator of a code. ACM Communications in Computer Algebra, 49(1):21, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jimenez-Pastor:2019:SID

[JP19] Antonio Jiménez-Pastor. A Sage implementation for DD-finite functions. ACM Communications in Computer Algebra, 53(2):53–56, June 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jimenez-Pastor:2020:SPC

[JPBCL20] Antonio Jiménez-Pastor, Alin Bostan, Frédéric Chyzak, and Pierre Lairez. The Sage package comb walks for walks in the quarter plane. ACM Communications in Computer Algebra, 54(2): 30–38, September 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3427218.3427220.

Jimenez-Pastor:2022:CEN

[JPP22] Antonio Jiménez-Pastor and Gleb Pogudin. Computing exact nonlinear reductions of dynamical models. ACM Communications in Computer Algebra, 56 (2):25–31, June 2022. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572867.3572869.

Jimenez-Pastor:2023:CAC

[JPRZ23] Antonio Jiménez-Pastor, Sonia L. Rueda, and Maria-Angeles Zurro. Computing almost-commuting basis of ordinary differential operators. ACM Communications in Computer Algebra, 57(3):111– 118, September 2023. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3637529.3637531.

Jamali:1991:LSG

[JR91] A. R. Jamali and E. F. Robertson. A library of simple groups: permutations, presentations, conjugacy classes, maximal subgroups and Sylow subgroups. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25 (1):46–49, January 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jeffrey:2015:RIT

[JRH15] David Jeffrey, Albert Rich, and Junrui Hu. RUBI and integration as term rewriting: integrals containing tangent. ACM Communications in Computer Algebra, 49(1):34, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Just:2006:CGB

[JS06] Winfried Just and Brandilyn Stigler. Computing Gröbner bases of ideals of few points in high dimensions. ACM Communications in Computer Algebra, 40(3–4):67–78, September– December 2006. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Juttler:2003:SIP

[JSS03] B. Jüttler, J. Schicho, and M. Shalaby. C^{1} spline implicitization of planar curves. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37 (3):65, September 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jank:1991:CMU

[JT91] Gerhard Jank and Lutz H.

Tack. Conformal mapping using Bergman’s method and the Maple system. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(2):18–23, April 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Juan:2015:IAF

[Jua15] Lourdes Juan. On the integration of algebraic functions: computing the logarithmic part. ACM Communications in Computer Algebra, 49(1):25, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ou:2020:TMC

[jYO20] Miao jung Yvonne Ou. Teaching multivariable calculus and tensor calculus with computer algebra software. ACM Communications in Computer Algebra, 54(4):134–135, December 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3465002.3465005.

Khan:1999:CSS

[KA99] Rao F. H. Khan and N. Ahmad. A comparison of symbolic solution of radioactive decay chains using Mathematica. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):20, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kahan:1975:PEP

[Kah75] W. Kahan. Problem #9:

an ellipse problem. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(3):11–12, August 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kako:1998:AJC

[Kak98] Fujio Kako. Abstracts of Japanese Computer Algebra Meeting in Kyoto. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(1):56–63, March 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kalina:1982:RRL

[Kal82a] Zdenêk Kalina. Running REDUCE and Lisp/360 interactively under a Time Sharing Option. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16(2):3, May 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kaltofen:1982:PRM

[Kal82b] Erich Kaltofen. Polynomial reduction from multivariate to bivariate integral polynomial factorization. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16(4):4, November 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kaltofen:1993:DPT

[Kal93a] Erich Kaltofen. Direct proof of a theorem by Kalkbrener, Sweedler, and Taylor. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27(4):2, December 1993. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kaltofen:1993:MC

[Kal93b] Erich Kaltofen. Message from the chair. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27(3):1–??, September 1993. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kaltofen:1995:MC

[Kal95] Erich Kaltofen. Message from the chair. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(2):1, April 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kaltofen:1996:ISC

[Kal96] E. Kaltofen. ISSAC steering committee bylaws. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(1):31–33, March 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kaltofen:1998:EEC

[Kal98] E. Kaltofen. ECCAD ’99:

East Coast Computer Algebra Day. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (2):64, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kaltofen:2013:JPA

[Kal13] Erich Kaltofen. Jenks Prize 2013 award citation. ACM Communications in Computer Algebra, 47(3–4):187, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kaltofen:2015:JPA

[Kal15] Erich Kaltofen. Jenks Prize 2015 award citation. ACM Communications in Computer Algebra, 49(4):112, December 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kanoui:1976:SAS

[Kan76] Henry Kanoui. Some aspects of symbolic integration via predicate logic programming. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 10 (4):29–42, November 1976. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kasper:1980:IFT

[Kas80] Toni Kasper. Integration in finite terms: the Liouville theory. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14 (4):2–8, November 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Katz:1987:IDE

[Kat87] Alan Katz. Issues in defining an equations representation standard. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21(2):19–24, February 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kauers:2003:CLS

[Kau03] Manuel Kauers. Computing limits of sequences. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(3):74–77, September 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kauers:2005:IJB

[Kau05] Manuel Kauers, editor. IS-SAC ’05: July 24–27, 2005, Beijing, China: Proceedings of the 2005 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY, USA, 2005. ISBN 1-59593-095-7. LCCN ???? ACM Order Number 505050.

Kollreider:1978:IAC

[KB78] C. Kollreider and B. Buchberger. An improved algorithmic construction of Gröbner-bases for polynomial ideals. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 12 (2):27–36, May 1978. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kieffer:2018:IBC

[KD18] Jean Kieffer and Luca De Feo. Isogeny-based cryptography in Julia/Nemo: a case study. ACM Communications in Computer Algebra, 52(4):130–132, December 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ke:2000:SAC

[Ke00]  Nainn-Ping Ke.  Sym-

bolic and algebraic computation in robust stability analysis. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34 (1):21, March 2000. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kelsey:2008:DFS

[Kel08] Michael A. Kelsey. Dethroning Fibonacci sequence (abstract only). ACM Communications in Computer Algebra, 42(1–2): 78, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kelly:2015:ACP

[Kel15]  Tyler Kelly.  An algo-

rithm for computing Picard ranks of certain K3 surfaces. ACM Communications in Computer Algebra, 49(2):59, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kempelmann:1981:RAF

[Kem81] Helmut Kempelmann. Recursive algorithm for the fast calculation of the limit of derivatives at points of indeterminateness. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15(4):10–11, November 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kahan:1999:SCD

[KF99] W. Kahan and Richard J.

Fateman. Symbolic computation of divided differences. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (2):7–28, June 1999. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kofalusi:1981:SSS

[KH81] Victor Köfalusi and Edit Halmay. State-space sets, state-space graphs and nprefix expressions. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15(1):33–36, February 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kai:2011:CIS

[KI11] Hiroshi Kai and Shigenobu Inoue. Cheater identification on a secret sharing scheme using GCD. ACM Communications in Computer Algebra, 45(2):119– 120, June 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kim:2016:SMT

[KiL16] SungSoon Kim and Dong il Lee. Standard monomials for Temperley–Lieb algebras. ACM Communications in Computer Algebra, 50(4):179–181, December 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kytmanov:2015:FRI

[KKM15] A. A. Kytmanov, A. M. Kytmanov, and E. K. Myshkina. On finding residue integrals for a class of systems of non-algebraic equations. ACM Communications in Computer Algebra, 49(3):77–79, September 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kytmanov:2018:CRI

[KKM18] A. A. Kytmanov, A. M. Kytmanov, and E. K. Myshkina. On computing residue integrals for a class of nonlinear transcendental systems. ACM Communications in Computer Algebra, 52(4):133–135, December 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Klip:1974:DPR

[Kli74] Dorothea A. Klip. Different polynomial representations and their interaction in the portable algebra system PORT-ALG. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(3):72–73, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Klip:1978:CSA

[Kli78] Dorothea A. Klip. A comparative study of algorithms for sparse polynomial multiplication. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 12 (3):12–24, August 1978. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kastner:2017:TGP

[KLPW17] Lars Kastner, Benjamin Lorenz, Andreas Paffenholz, and Anna-Lena Winz. Toric geometry in polymake. ACM Communications in Computer Algebra, 51(3):92–94, September 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kotsireas:2003:EIP

[KLV03] Ilias Kotsireas, Edmond Lau, and Richard Voino. Exact implicitization of polynomial curves and surfaces. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37 (3):78, September 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Knill:1987:RGR

[KM87] O. Knill and R. E. Mäder. The rotation group of Rubik’s cube. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21 (3):33–43, August 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Krandick:2005:NBD

[KM05] Werner Krandick and Kurt Mehlhorn. New bounds for the Descartes method. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(3):94, 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). ISSAC 2005 poster abstract.

Kesh:2011:SAH

[KM11] Deepanjan Kesh and Shashank K. Mehta. A saturation algorithm for homogeneous binomial ideals. ACM Communications in Computer Algebra, 45(2):121– 122, June 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kauers:2019:MOP

[KM19] Manuel Kauers and Marc Mezzarobba. Multivariate Ore polynomials in Sage-Math. ACM Communications in Computer Algebra, 53(2):57–60, June 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kapur:1987:MUC

[KN87] Deepak Kapur and Paliath Narendran. Matching, unification and complexity. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21(4):6–9, November 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kai:1997:CPV

[KN97] Hiroshi Kai and Matu-Tarow Noda. Cauchy principal value integral using hybrid integral. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):37–38, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Kai:1998:HCC

[KN98] Hiroshi Kai and Matu-Tarow Noda. Hybrid computation of Cauchy-type singular integral equations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(2):59– 60, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Abstract only.

Kai:2000:HCB

[KN00] Hiroshi Kai and Matu-Tarow Noda. Hybrid computation of bivariate rational interpolation. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):20–21, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Knopfmacher:1997:DDF

[Kno97] Arnold Knopfmacher. Distinct degree factorizations for polynomials over a finite field. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):38–39, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Knuth:1975:SSA

[Knu75] Donald E. Knuth. Son of seminumerical algorithms. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(4):10–11, November 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kruse:1983:ARE

[KO83] Hans-Guenther Kruse and Karin Ohlsen. About the realization of an extended, but really interactive RE-DUCE by integration of a small editing and executing system. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17 (1):21–25, February 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kaltofen:2013:NFO

[KO13] Erich Kaltofen and Alexey Ovchinnikov. NSF funding opportunities for symbolic computation. ACM Communications in Computer Algebra, 47(3–4):84, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Koepf:1993:EAC

[Koe93] Wolfram Koepf. Examples for the algorithmic calculation of formal Puisieux, Laurent and power series. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27(1):20–32, January 1993. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Koepf:1994:FID

[Koe94] Wolfram Koepf. On families of iterated derivatives. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 28(3–4):10–26, December 1994. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Koepf:1995:RPI

[Koe95] Wolfram Koepf. REDUCE package for the indefinite and definite summation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(1):14–30, January 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kogan:2015:DAI

[Kog15] Irina A. Kogan. Differential algebra of invariants and invariant variational calculus. ACM Communications in Computer Algebra, 49 (1):25, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kohlhase:2000:OIO

[Koh00] Michael Kohlhase. OM-Doc: An infrastructure for OpenMath content dictionary information. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(2):43–48, June 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kohl:2010:IMB

[Koh10] Karen T. Kohl. An implementation of the method of brackets for symbolic integration. ACM Communications in Computer Algebra, 44(3):118–119, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Koprowski:2020:CMP

[Kop20]  Przemysław  Koprowski.

CQF Magma package. ACM Communications in Computer Algebra, 54(2): 53–56, September 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3427218.3427224.

Korpela:1976:GCA

[Kor76]  Jukka Korpela.  Gen-

eral characteristics of the ANALITIK language. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 10(3):30–48, August 1976. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Korpela:1977:SPC

[Kor77] Jukka Korpela. Some problems connected with ambiguous expressions. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11(3):7–9, August 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kotsireas:2001:HPS

[Kot01] Ilias S. Kotsireas. Homotopies and polynomial system solving I: Basic principles. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35 (1):19–32, March 2001. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kotsireas:2002:AIC

[Kot02] Ilias S. Kotsireas. ACA 2002 8th International Conference on Applications of Computer Algebra June 25– 28, 2002, Volos, Greece. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36(1):22, March 2002. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kotsireas:2008:MMC

[Kot08] Ilias S. Kotsireas. Mathematical Methods in Computer Science 2008: MMICS 2008, December 17–19, 2008, Karlsruhe, Germany. ACM Communications in Computer Algebra, 42(1–2): 103, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kotsireas:2014:MSCa

[Kot14a] Ilias S. Kotsireas. Message from the SIGSAM Chair. ACM Communications in Computer Algebra, 48(2): 32, March 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kotsireas:2014:MSCb

[Kot14b] Ilias S. Kotsireas. Message from the SIGSAM Chair. ACM Communications in Computer Algebra, 48(3/4): 154, September 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Koutschan:2009:AAH

[Kou09] Christoph Koutschan. Advanced applications of the holonomic systems approach. ACM Communications in Computer Algebra, 43(4):119, December 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Koutschan:2013:HFM

[Kou13] Christoph Koutschan. Holonomic functions in Mathematica. ACM Communications in Computer Algebra, 47(3–4):179–182, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Koutschan:2016:MPP

[Kou16] Christoph Koutschan. Motion polynomials and planar linkages. ACM Communications in Computer Algebra, 50(3):109–112, September 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kredel:1997:MMA

[KP97] Heinz Kredel and Michael Pesch. MAS, the Modula-2 Algebra System, version 1.00. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):52–54, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Kauers:2016:WSC

[KPR16] Manuel Kauers, Peter Paule, and Greg Reid. Workshop on symbolic combinatorics and algorithmic differential algebra. ACM Communications in Computer Algebra, 50(1):27–34, March 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Krick:1999:AN

[KPS99] T. Krick, L. M. Pardo, and M. Sombra. Arithmetic Nullstellensätze. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):17, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Krandick:2006:TCA

[KPW06] Werner Krandick, Shawn S.

Pierce, and Zhendong Wan. Towards collaboration across display diversity. ACM Communications in Computer Algebra, 40 (2):49–50, June 2006. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kryukov:1985:IR

[KR85a] A. P. Kryukov and A. Ya.

Rodinov. Interactive RE-DUCE. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19 (3):43–45, August 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kryukov:1985:DDS

[KR85b] A. P. Kryukov and A. Ya. Rodionov. Dynamicdebugging system for the REDUCE programs. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19(2):34–37, May 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Krasinski:1983:OPA

[Kra83] Andrzej Krasiński. OR-TOCARTAN: a program for algebraic calculations in general relativity. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(3–4):12–18, August/November 1983. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kredel:1988:ATO

[Kre88] Heinz Kredel. Admissible term orderings used in computer algebra systems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22(1):28–31, January 1988. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kredel:2001:ARW

[Kre01] H. Kredel. Announcement: Rhine Workshop on Computer Algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35(3):38, September 2001. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Krishnamurthy:1984:FPC

[Kri84] E. V. Krishnamurthy. Functional programming with combinators for symbolic computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18 (2):24, May 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Korporal:2010:MPI

[KRR10] Anja Korporal, Georg Regensburger, and Markus Rosenkranz. A Maple package for integro-differential operators and boundary problems. ACM Communications in Computer Algebra, 44(3):120–122, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Korporal:2012:ISD

[KRR12] Anja Korporal, Georg Regensburger, and Markus Rosenkranz. ISSAC 2012 software demonstrations: Symbolic computation for ordinary boundary problems in Maple. ACM Communications in Computer Algebra, 46(3–4):154–156, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Katsamaki:2020:PMP

[KRTZ20] Christina Katsamaki, Fabrice Rouillier, Elias Tsigaridas, and Zafeirakis Zafeirakopoulos. PTOPO: a MAPLE package for the topology of parametric curves. ACM Communications in Computer Algebra, 54(2):49–52, September 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3427218.3427223.

Kovacs:2021:ART

[KRV21] Zoltán Kovács, Tomás Recio, and M. Pilar Vélez. Automated reasoning tools in GeoGebra discovery. ACM Communications in Computer Algebra, 55(2):39– 43, June 2021. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3493492.3493495.

Kryukov:1984:ARL

[Kry84] A. P. Kryukov. An antitranslator of the RLISP language. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18 (3):12–15, August 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kanada:1981:LBB

[KS81] Yasumasa Kanada and Tateaki Sasaki. LISPbased “big-float” system is not slow. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15 (2):13–19, May 1981. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kolar:1992:MQP

[KS92] M. Kolár and T. Sasaki.

Multivariate quotient by power-series division. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26(3):17–20, August 1992. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kajler:1994:SHI

[KS94] Norbert Kajler and Neil Soiffer. Some human interaction issues in computer algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 28 (1):18–28, March 1994. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kako:1997:PEF

[KS97] Fujio Kako and Tateaki Sasaki. Proposal of “effective floating-point number” for approximate algebraic computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):31, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only. Discusses fuzzy and interval floatingpoint arithmetic.

Kanno:2003:VNM

[KS03] Masaaki Kanno and Malcolm C. Smith. Validated numerical methods for systems and control engineering. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37 (3):72–73, September 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Khan:2013:VFM

[KS13] Muhammad Taimoor Khan and Wolfgang Schreiner. A verification framework for MiniMaple programs. ACM Communications in Computer Algebra, 47(3–4):98– 99, September 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kawano:2014:QFT

[KS14] Yasuhito Kawano and Hiroshi Sekigawa. Quantum Fourier Transform over symmetric groups: Improved result. ACM Communications in Computer Algebra, 48(3/4):127–129, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Keigher:2015:DEH

[KS15] William F. Keigher and V. Ravi Srinivasan. Differential equations and Hurwitz series. ACM Communications in Computer Algebra, 49(1):31, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Katayama:2016:NII

[KS16] Akiyuki Katayama and Kiyoshi Shirayanagi. A new idea on the interval-symbol method with correct zero rewriting for reducing exact computations. ACM Communications in Computer Algebra, 50(4):176– 178, December 2016. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kondo:2019:RCM

[KS19] Kazuki Kondo and Hiroshi Sekigawa. Robust computation methods for sparse interpolation of multivariate polynomials. ACM Communications in Computer Algebra, 53(3):130– 133, September 2019. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ksir:2015:ATC

[Ksi15] Amy Ksir. Automorphisms of tropical curves and Berkovich analytic curves and their skeletons. ACM Communications in Computer Algebra, 49(2):60, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Koutschan:2015:MPL

[KSS15] Christoph Koutschan, Erwin Suazo, and Sergei K. Suslov. Multi-parameter laser modes in paraxial optics. ACM Communications in Computer Algebra, 49 (1):33, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Koukouvinos:2011:AFE

[KSZ11] Christos Koukouvinos, Dimitris E. Simos, and Zafeirakis Zafeirakopoulos. An algebraic framework for extending orthogonal designs.

ACM Communications in Computer Algebra, 45(2): 123–124, June 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kai:2014:FSD

[KU14] Hiroshi Kai and Keita Ueda. Fake shares detection on a visual secret sharing scheme by rational interpolation. ACM Communications in Computer Algebra, 48(3/4):124–126, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kuchlin:1982:SRS

[Küc82] Wolfgang Küchlin. Some reduction strategies for algebraic term rewriting. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16(4):13–23, November 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kuchlin:1987:PKE

[Küc87] W. Küchlin. Public key encryption. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21 (3):69–73, August 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kuchlin:1997:PPS

[Küc97] Wolfgang W. Küchlin, editor. ISSAC 97: July 21– 23, 1997, Maui, Hawaii, USA: proceedings of the 1997 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY, USA, 1997. ISBN 0-89791-875-4. LCCN QA76.95.

                                        URL  http://www.acm.
                                        org/pubs/contents/proceedings/
                                        issac/258726/.

Kotsireas:2004:AVA

[KV04] Ilias Kotsireas and Emil Volcheck. ANTS VI: algorithmic number theory symposium poster abstracts. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(3):93–107, September 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kauers:2019:WYS

[KV19] Manuel Kauers and Thibaut Verron. Why you should remove zeros from data before guessing. ACM Communications in Computer Algebra, 53(3):126–129, September 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kalkbrener:1997:JSC

[KW97] M. Kalkbrener and D. Wang. Journal of Symbolic Computation special issue on polynomial elimination: Algorithms and applications. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(1):31, March 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Kanno:2008:SAR

[KYA08] Masaaki Kanno, Kazuhiro Yokoyama, and Hirokazu Anai. Solution of the algebraic Riccati equation using the sum of roots. ACM Communications in Computer Algebra, 42(3):144– 145, September 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kotsireas:2006:WA

[KZ06] Ilias S. Kotsireas and Eugene A. Zima. WWCA 2006 abstracts. ACM Communications in Computer Algebra, 40(2):52–59, June 2006. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kotsireas:2011:AWH

[KZ11] Ilias Kotsireas and Eugene Zima. Abstracts of WWCA 2011 in honor of Herb Wilf’s 80th birthday. ACM Communications in Computer Algebra, 45(2):92–99, June 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Labonte:1987:RPS

[Lab87] Gilles Labonté. Report on a program for solving polynomial equations in noncommuting variables. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21(2):4–7, February 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Labahn:2015:DSS

[Lab15a] George Labahn. Dynamical systems and scaling invariants. ACM Communications in Computer Algebra, 49(1):25, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Labahn:2015:RIF

[Lab15b] George Labahn. Rational invariants of finite abelian groups. ACM Communications in Computer Algebra, 49(1):31–32, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Labelle:2015:CIP

[Lab15c] Gilbert Labelle. Combinatorial integration (part I, part II). ACM Communications in Computer Algebra, 49(1):35, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lacolle:1992:SAS

[Lac92] B. Lacolle. Spinor analysis and symbolic computation of partition functions. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26(4):7–14, November 1992. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

LakshmanYN:1996:IPI

[Lak96] Lakshman Y.N., editor.

ISSAC ’96: Proceedings of the 1996 International Symposium on Symbolic and Algebraic Computation, July 24–26, 1996, Zurich, Switzerland. ACM Press, New York, NY, USA, 1996. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996.

N:1998:SIB

[Lak98] Lakshman Y.N. SIGSAM/ ISSAC business meeting summary. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (2):3–4, June 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Langemyr:1986:CSC

[Lan86] Lars Langemyr. Converting SAC-2 code to Lisp. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 20(4):11–13, December 1986. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Landau:1993:FMS

[Lan93] Susan Landau. Finding maximal subfields. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27(3):4–8, September 1993. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

LaPlace:1972:PLP

[LaP72] André M. L. LaPlace. PL/I list processing used as interactive system support. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(22):10–24, March 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Larrieu:2019:CGB

[Lar19] Robin Larrieu. Computing generic bivariate Gröbner bases with Mathemagix. ACM Communications in Computer Algebra, 53(2): 41–44, June 2019. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lassner:1984:RNC

[Las84] W. Lassner. Report on a national computer algebra workshop in Leipzig. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(3):22–23, August 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lauer:1977:SKP

[Lau77] Markus Lauer. A solution to Kahan’s problem (SIG-SAM problem no. 9). SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11(2):16–20, May 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Launois:2015:EQG

[Lau15] Stephane Launois. Endomorphisms of quantum generalized Weyl algebras. ACM Communications in Computer Algebra, 49(1): 31, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lazard:1980:PSA

[Laz80] D. Lazard. Problem 7 and systems of algebraic equations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14 (2):26–29, May 1980. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lazard:2001:SSA

[Laz01] D. Lazard. Solving systems of algebraic equations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35(3):11–37, September 2001. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Originally published as “Résolution des Systémes d’Équations Algébraiques”, Theoretical Computer Science 15 (1981), 77–110. Translated by Michael Abramson.

Lindberg:2020:ESP

[LBL20] Julia Lindberg, Nigel Boston, and Bernard C. Lesieutre. Exploiting symmetry in the power flow equations using monodromy. ACM Communications in Computer Algebra, 54(3):100– 104, September 2020. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3457341.3457346.

Li:2019:RGS

[LC19] Ao Li and Robert M. Corless. Revisiting Gilbert Strang’s “a chaotic search for i”. ACM Communications in Computer Algebra, 53(1):1–22, March 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Li:2011:APF

[LD11] Yue Li and Gabriel Dos Reis. An automatic parallelization framework for OpenAxiom. ACM Communications in Computer Algebra, 45(2):125– 126, June 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Le:2001:CMT

[Le01]  H. Q. Le.  Comput-

ing the minimal telescoper for sums of hypergeometric terms. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35 (3):2–10, September 2001. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lee:2010:CAD

[Lee10] Jae Hee Lee. Cylindrical algebraic decomposition for reasoning about qualitative spatial knowledge. ACM Communications in Computer Algebra, 44(3):123– 124, September 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lee:2019:CAS

[Lee19]  Kisun Lee.  Certifying

approximate solutions to polynomial systems on Macaulay2. ACM Communications in Computer Algebra, 53(2):45–48, June 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Legendi:1984:CHS

[Leg84] T. Legendi. Cellular hardware and symbolic computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18 (2):33, May 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lemaire:2015:NDA

[Lem15] François Lemaire. New development and application of integration of differential fractions. ACM Communications in Computer Algebra, 49(1):26, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lenstra:1981:LFP

[Len81] A. K. Lenstra. Lattices and factorization of polynomials. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15 (3):15–16, August 1981.

CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lenstra:1984:FP

[Len84] Arjen K. Lenstra. Factorization of polynomials. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(2):16–18, May 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

LeSchack:1970:RRL

[LeS70] A. R. LeSchack. Response from A. R. LeSchack. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(14):13–16, January 1970. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Levelt:1995:IPI

[Lev95] A. H. M. Levelt, editor. ISSAC ’95: Proceedings of the 1995 International Symposium on Symbolic and Algebraic Computation: July 10–12, 1995, Montréal, Canada, ISSAC -PROCEEDINGS- 1995. ACM Press, New York, NY, USA, 1995. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. ACM order number: 505950.

Levandovskyy:2010:NCS

[Lev10] Viktor Levandovskyy. Noncommutative subsystems of Singular. ACM Communications in Computer Algebra, 44(3):125– 126, September 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Levin:2015:GGB

[Lev15a] Alexander Levin. Generalized Gröbner bases and dimension polynomials of modules over some finitely generated noncommutative algebras. ACM Communications in Computer Algebra, 49(1):26, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Levin:2015:MGC

[Lev15b] Alexander Levin. Method of generalized characteristic sets and multivariate dimension polynomials of differential field extensions with a group action. ACM Communications in Computer Algebra, 49(1):34, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lewis:2000:CPO

[Lew00] Robert H. Lewis. Comparison of polynomial-oriented computer algebra systems.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):24, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lewis:2004:UFS

[Lew04] Robert H. Lewis. Using Fermat to solve large polynomial and matrix problems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(1):27–28, March 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lewis:2008:CAT

[Lew08] Robert Lewis. Comparing acceleration techniques for the Dixon and Macaulay resultants (abstract only). ACM Communications in Computer Algebra, 42(1–2): 79–81, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Leykin:2010:AOC

[Ley10] Anton Leykin. Abstract only: Certified numerical solving of systems of polynomial equations. ACM Communications in Computer Algebra, 44(2):24, June 2010. CODEN ????

ISSN 1932-2232 (print), 1932-2240 (electronic).

Leykin:2016:PHC

[Ley16] Anton Leykin. Polynomial homotopy continuation in Macaulay2. ACM Communications in Computer Algebra, 50(3):113– 116, September 2016. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lange-Hegermann:2015:CSD

[LH15] Markus Lange-Hegermann. Counting solutions of differential equations. ACM Communications in Computer Algebra, 49(1):25–26, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Li:1992:NDB

[Li92] Wei Li. Nonexistence of degree bounds of various bases for ideals of polynomials over the integers. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26(1):4–9, January 1992. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Li:2009:THP

[Li09]  Xin Li.  Toward high-

performance polynomial system solvers based on triangular decompositions. ACM Communications in Computer Algebra, 43(4): 120, December 2009. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Li:2015:TAD

[Li15]  Yinghong Li.  Theories

and algorithms for difference and differential Chow form. ACM Communications in Computer Algebra, 49(3):108–109, September 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lichtenberger:1984:SES

[Lic84]  F. Lichtenberger.  Self-

explanatory symbolic computation for math education. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18– 19(4–1):23–24, November/ February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lichtblau:2011:SDI

[Lic11] Daniel Lichtblau. Symbolic definite (and indefinite) integration: methods and open issues. ACM Communications in Computer Algebra, 45(1):1–16, March 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lichtblau:2015:HSD

[Lic15] Daniel Lichtblau. The hazards of symbolic definite integration (a continuing saga). ACM Communications in Computer Algebra, 49(1):35, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lichtblau:2017:ECC

[Lic17]  Daniel Lichtblau.  East

Coast Computer Algebra Day 2017. ACM Communications in Computer Algebra, 51(2):66–69, June 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lindner:2020:IDA

[LIJ20] Sebastian Lindner, Laurent Imbert, and Michael J. Jacobson. Improved divisor arithmetic on generic hyperelliptic curves. ACM Communications in Computer Algebra, 54(3):95–99, September 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3457341.3457345.

Lin:1998:AAAa

[Lin98a]  Shu Lin.  Announce-

ments: Applied Algebra Algebraic Algorithms Error-Correcting Codes, Honolulu, Hawaii (USA) — 14– 19 November 1999. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(3):31, September 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lin:1998:AAAb

[Lin98b]  Shu Lin.  Announce-

ments: Applied Algebra Algebraic Algorithms Error-Correcting Codes, Honolulu, Hawaii (USA) — 14– 19 November 1999. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(4):31, December 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Linton:2007:GGA

[Lin07] Steve Linton. GAP: groups, algorithms, programming. ACM Communications in Computer Algebra, 41(3): 108–109, September 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Liu:2015:RGB

[Liu15]  Yue Liu.  Research on

Gröbner basis of the defining ideal of polynomial ring in quaternionic variables. ACM Communications in Computer Algebra, 49(3): 108, September 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Liu:2019:BDI

[Liu19]  Jiang Liu.  Block dis-

tance invariant method for monoterm canonicalization of Riemann tensor polynomials. ACM Communications in Computer Algebra, 53(3):134–137, September 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Li:2008:GBM

[LLG08] Dongmei Li, Jinwang Liu, and Shuhong Gao. The w-Gröbner bases and monomial ideal under polynomial composition (abstract only). ACM Communications in Computer Algebra, 42(1–2):81–82, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

LlopisdeTrias:1984:NGS

[Llo84] R. Llopis de Trias. A new generation of symbolic and algebraic computation systems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18 (2):12–14, May 1984. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lauster:2018:SCM

[LLT18] Florian Lauster, D. Russell Luke, and Matthew K. Tam. Symbolic computation with monotone operators. ACM Communications in Computer Algebra, 52(4):139–141, December 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lavasani:2008:IFA

[LM08] Ahmad Lavasani and Reza Mohammadi. Implementing a feasible attack against ECC2K-130 Certicom Challenge (abstract only). ACM Communications in Computer Algebra, 42(1–2):61– 62, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Labahn:2012:GWC

[LM12] George Labahn and Johannes Middeke. Gröbner walk for computing matrix normal forms over Ore polynomials. ACM Communications in Computer Algebra, 46(3–4):106–107, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Law:2016:CCP

[LM16] Marshall Law and Michael Monagan. Computing characteristic polynomials of matrices of structured polynomials. ACM Communications in Computer Algebra, 50(4):182–184, December 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Llorente:2012:NSM

[LMF12] Alberto Llorente and Jorge Mozo-Fernández. Numericsymbolic methods for computing the Liouvillian solutions of differential equations and systems. ACM Communications in Computer Algebra, 46(3–4):112– 113, September 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Li:2008:MLB

[LMRS08] Xin Li, Marc Moreno Maza, Raqeeb Rasheed, and Éric Schost. The modpn library: bringing fast polynomial arithmetic into MAPLE. ACM Communications in Computer Algebra, 42(3): 172–174, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lebreton:2012:CCC

[LMS12] Romain Lebreton, Esmaeil Mehrabi, and Éric Schost. On the complexity of computing certain resultants. ACM Communications in Computer Algebra, 46(3–4): 108–109, September 2012.

CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Laskari:2005:TTC

[LMTV05] E. C. Laskari, G. C. Meletiou, D. K. Tasoulis, and M. N. Vrahatis. Transformations of two cryptographic problems in terms of matrices. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(4):127–130, December 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lemaire:2005:RLM

[LMX05] F. Lemaire, M. Moreno Maza, and Y. Xie. The RegularChains library in MAPLE. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39 (3):96–97, 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). ISSAC 2005 poster abstract.

Loos:1976:GTC

[LN76] Rüdiger Loos and J. Neubüser.

Group theoretical computations in mathematical crystallography. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 10 (4):11–18, November 1976.

CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lobo:1996:FRA

[Lob96] Austin Lobo. Formal review of articles in the Bulletin. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (3):2–3, September 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lobo:1997:CLA

[Lob97] Austin Lobo. Computer linear algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):3, September 1997. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Loos:1972:AAD

[Loo72a] Rüdiger Loos. Algebraic algorithm descriptions as programs. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (23):16–24, July 1972. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Loos:1972:ATT

[Loo72b] Rüdiger Loos. Analytic treatment of three similar Fredholm integral equations of the second kind with Reduce 2. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (21):32–40, January 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Loos:1974:TFI

[Loo74] Ruediger G. K. Loos. Toward a formal implementation of computer algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(3):9–16, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Loos:1975:PCP

[Loo75] Rudiger Loos. Problem #10: the cattle problem of Archimedes. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9 (3):11–12, August 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Loos:1977:MUC

[Loo77a] Rüdiger Loos. 1977 MAC-SYMA users’ conference. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11(2):7– 13, May 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Loos:1977:A

[Loo77b] Rüdiger Loos. Abstracts. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11(1):3–7, February 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Loos:1977:AA

[Loo77c]  Rüdiger Loos.  AB-

STRACTS (Amsterdam 78). SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11– 12(4–1):3–6, November and February 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Loos:1977:ARN

[Loo77d] Rüdiger Loos. Abstracts (REDUCE newsletter 1.1978). SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11–12(4–1):86– 88, November and February 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Loos:1977:ASA

[Loo77e] Rüdiger Loos. Author’s solution of Archimedes’ cattle problem. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11(1):4–7, February 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lounesto:1984:CAV

[Lou84] Pertti Lounesto. A computeraided vector algebra system for use in high schools. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18–19(4–1):25, November/February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Li:2008:SSF

[LOW08] Ziming Li, Martin Ondera, and Huaifu Wang. Simplifying skew fractions modulo differential and difference relations. ACM Communications in Computer Algebra, 42(3):164–165, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lewis:2015:ADR

[LPA15] Robert H. Lewis, Bela Paláncz, and Joseph Awange. Application of Dixon resultant to maximization of the likelihood function of Gaussian mixture distribution. ACM Communications in Computer Algebra, 49(2):57, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lamban:1999:SSE

[LPR99] L. Lambán, V. Pascual, and J. Rubío. Simplicial sets in the EAT system. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):17, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Labahn:1998:PSA

[LR98] George Labahn and Mohamed Omar Rayes. Poster session abstracts of the 1998 International Symposium on Symbolic and Algebraic Computation, IS-SAC ’98, Rostock, German, August 13–15, 1998. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(2):49–51, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lemaire:2022:IEM

[LR22] François Lemaire and Louis Roussel. Integral equation modelling and deep learning. ACM Communications in Computer Algebra, 56 (2):51–55, June 2022. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572867.3572874.

Leler:1985:IGI

[LS85] Wm Leler and Neil Soiffer. An interactive graphical interface for Reduce. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19(3):17–23, August 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Levandovskyy:2003:UCA

[LS03] Viktor Levandovskyy and Hans Schönemann. Using computer algebra system SINGULAR:PLURAL for computations in noncommutative polynomial algebras. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37 (3):77–78, September 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lecerf:2016:SSK

[LS16] Grégoire Lecerf and Joelle Saadé. A short survey on Kantorovich: like theorems for Newton’s method. ACM Communications in Computer Algebra, 50(1): 1–11, March 2016. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

LaScala:2018:CNH

[LT18] Roberto La Scala and Sharwan K. Tiwari. Computing noncommutative Hilbert series. ACM Communications in Computer Algebra, 52(4):136–138, December 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lefevre:2010:LTL

[LTdD^{+}10] Vincent Lefévre, Philippe Théveny, Florent de Dinechin, Claude-Pierre Jeannerod, Christophe Mouilleron, David Pfannholzer, and Nathalie Revol. LEMA: towards a language for reliable arithmetic. ACM Communications in Computer Algebra, 44(2):41–52, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lubbes:2011:MFC

[Lub11] Nils Lubbes. Minimal families of curves on surfaces. ACM Communications in Computer Algebra, 45(3–4): 238–239, September 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lubbes:2015:CS

[Lub15] Niels Lubbes. Circles on surfaces. ACM Communications in Computer Algebra, 49(2):53–54, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ludeman:1969:PUP

[Lud69] M. M. Ludeman. PL/I utility programs for Formac applications. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(12):53–58, July 1969. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Luneburg:1984:WB

[Lün84] Heinz Lüneburg. Write a book! SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18 (2):34, May 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Liu:1994:HCBa

[LW94a] Zhuojun Liu and Paul S.

Wang. Height as a coefficient bound for univariate polynomial factors (Part I). SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 28(2):20–27, August 1994. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Liu:1994:HCBb

[LW94b] Zhuojun Liu and Paul S.

Wang. Height as a coefficient bound for univariate polynomial factors (Part II). SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 28 (3–4):1–9, December 1994. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Li:1995:CPP

[LW95] Liwei Li and Paul S. Wang. The CL–PVM package. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(3–4):2–8, December 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lewis:1999:CPO

[LW99] Robert H. Lewis and Michael Wester. Comparison of polynomial-oriented computer algebra systems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(4):5–13, December 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lai:2008:GWB

[LW08] Xun Lai and Paul S. Wang. GeometryEditor: a Webbased interactive geometry manipulative authoring system (abstract only). ACM Communications in Computer Algebra, 42(1–2):78– 79, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Labahn:2016:TPC

[LW16] George Labahn and Stephen Watt. Thematic program in computer algebra: July 1–Dec 31, 2015. ACM Communications in Computer Algebra, 50(1):25–26, March 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Li:2007:TLD

[LWZ07] Ziming Li, Min Wu, and Dabin Zheng. Testing linear dependence of hyperexponential elements. ACM Communications in Computer Algebra, 41(1–2):3– 11, March–June 2007. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Li:2012:SDR

[LYG12] Wei Li, Chun-Ming Yuan, and Xiao-Shan Gao. Sparse differential resultant for Laurent differential polynomials. ACM Communications in Computer Algebra, 46(3–4):110–111, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Li:2005:CHF

[LZ05] Ziming Li and Dabin Zheng. Computation with hyperexponential functions. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(3):84–85, 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). ISSAC 2005 poster abstract.

Moritsugu:2008:GTP

[MA08] Shuichi Moritsugu and Chisato Arai. Geometry theorem proving by Gröbner bases: using ideal decompositions. ACM Communications in Computer Algebra, 42(3):158– 159, September 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

MacCallum:1989:CPA

[Mac89] Malcolm A. H. MacCallum. Comments on the performance of algebra systems in general relativity and a recent paper by Nielsen and Pedersen. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23 (2):22–25, April 1989. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Madore:1992:RAM

[Mad92]  Blair F. Madore.  Re-

port on the 10th Annual Maple Retreat, June 13– 17, 1992. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26 (3):7–8, August 1992. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Maddah:2015:FSC

[Mad15a] Suzy S. Maddah. Formal solutions of completely integrable Pfaffian systems with normal crossings. ACM Communications in Computer Algebra, 49(1):26–27, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Maddah:2015:FRS

[Mad15b]  Suzy S. Maddah.  On

the formal reduction of singularly-perturbed linear differential systems. ACM Communications in Computer Algebra, 49(1):31, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Maeder:1987:SBV

[Mae87a] R. E. Maeder. Solving boundary-value problems with perturbations. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21(3):16–18, August 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Maeder:1987:CPM

[Mae87b] Roman E. Maeder. A collection of projects for the mathematical laboratory. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21(3):12–94, August 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Maignan:1999:RSI

[Mai99] A. Maignan. Real solving of ill-conditioned sinepolynomials equations. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):17, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Maier:2015:PDE

[Mai15] Annette Maier. Parameterized differential equations and patching. ACM Communications in Computer Algebra, 49(1):27, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Manders:1977:ICT

[Man77] Kenneth L. Manders. The impact of complexity theory on algorithms for sparse polynomials. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11 (1):8–15, February 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Manubens:2010:AOP

[Man10] Montserrat Manubens. Abstract only: Parametric polynomial system discussion: canonical comprehensive Gröbner systems. ACM Communications in Computer Algebra, 44(1): 11–12, March 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Mantzaflaris:2011:RAM

[Man11] Angelos Mantzaflaris. Robust algebraic methods for geometric computing. ACM Communications in Computer Algebra, 45(3–4):237, September 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Manganiello:2015:NCS

[Man15a] Felice Manganiello. Network coding via skew polynomials. ACM Communications in Computer Algebra, 49(1):15, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Manon:2015:HCH

[Man15b] Chris Manon. Horospherical contraction and Hamiltonian functions on an algebraic group. ACM Communications in Computer Algebra, 49(2):59, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Mathews:2010:AOE

[Mat10] Deborah Mathews. Abstract only: An empirical study of parallel big number arithmetic. ACM Communications in Computer Algebra, 44(2):25, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Matveiakin:2021:DMP

[Mat21] Andrei Matveiakin. Discovering multiple polylogarithm equations via symbolic computations. ACM Communications in Computer Algebra, 55(3):112– 116, September 2021. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3511528.3511539.

Maurischat:2016:NFI

[Mau16] Andreas Maurischat. Nonfree iterative differential modules. ACM Communications in Computer Algebra, 50(4):126–135, December 2016. CODEN ????

ISSN 1932-2232 (print), 1932-2240 (electronic).

May:2011:ECC

[May11] John May. East Coast Computer Algebra Day 2011 abstracts. ACM Communications in Computer Algebra, 45(2):100–106, June 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Mazzarella:1985:ISO

[Maz85] Giuseppe Mazzarella. Improved simplification of odd and even functions in RE-DUCE. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19 (2):29–30, May 1985. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Maza:2015:DAG

[Maz15] Marc Moreno Maza. Doing algebraic geometry with the RegularChains library. ACM Communications in Computer Algebra, 49(2): 54, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Mordukhai-Boltovskoi:1981:GII

[MBKP81] D. Mordukhai-Boltovskoi, Boris Korenblum, and Myra Prelle. A general investigation of integration in finite form of differential equations of the first order article 1. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15 (2):20–32, May 1981. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Marco-Buzunariz:2023:EHU

[MBRD23] Miguel A. Marco-Buzunáriz, Ana Romero, and Jose Divasón. Effective homology of universal covers. ACM Communications in Computer Algebra, 57(2):47– 50, June 2023. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3614408.3614414.

McCleeary:2015:LAA

[MBS15] Ryan McCleeary, Martin Brain, and Aaron Stump. A lazy approach to adaptive exact real arithmetic using floating-point operations. ACM Communications in Computer Algebra, 49(3):83–86, September 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Montes:1995:SLF

[MC95] Antonio Montes and Jordi Castro. Solving the load flow problem using the Gröbner basis. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(1):1–13, January 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

McCarthy:1973:PKP

[McC73] D. P. McCarthy. Problem #6 — the Kiang problem. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(28):12, December 1973. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

McCallum:1997:FIR

[McC97] Scott McCallum. Factors of iterated resultants and discriminants. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):33–34, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

McGeer:1984:DIB

[McG84] Patrick C. McGeer. A discussion and implementation of Brown’s REX simplification algorithm. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(1):10–19, February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

McGettrick:2004:OGB

[McG04] Michael McGettrick. Online Gröbner basis [OGB]. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(1):19–21, March 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

McIsaac:1985:PMA

[McI85] Kevin McIsaac. Pattern matching algebraic identities. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19 (2):4–13, May 1985. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Moir:2015:UPR

[MCJ15] Robert H. C. Moir, Robert M. Corless, and David J. Jeffrey. Unwinding paths on the Riemann sphere for continuous integrals of rational functions. ACM Communications in Computer Algebra, 49(1):35, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

McNamee:2004:CMA

[McN04] John Michael McNamee. A comparison of methods for accurate summation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(1):1–7, March 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (elec-

tronic). URL http://www.
acm.org/sigs/sigsam/bulletin/
repository/issue147.

Marquez-Corbella:2015:ECP

[MCP15] Irene Márquez-Corbella and Ruud Pellikaan. Errorcorrecting pairs: a new approach to code-based cryptography. ACM Communications in Computer Algebra, 49(1):21, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Magron:2018:RMP

[ME18] Victor Magron and Mohab Safey El Din. RealCertify: a Maple package for certifying non-negativity. ACM Communications in Computer Algebra, 52(2):34–37, June 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Medvedev:2015:DDA

[Med15] Alice Medvedev. Dimensions of difference-algebraic groups. ACM Communications in Computer Algebra, 49(1):27, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Meng:2016:ARD

[Men16] Lingchuan Meng. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 50(1): 40–42, March 2016. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Metzlaff:2023:SAB

[Met23] Tobias Metzlaff. Symmetry adapted bases for trigonometric optimization. ACM Communications in Computer Algebra, 57(3):137– 140, September 2023. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3637529.3637535.

Meunier:2003:MFS

[Meu03] Ludovic Meunier. MAD: a flexible system for authoring mathematical documents. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37 (3):79–81, September 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Marti:1983:RCM

[MF83] Jed Marti and John Fitch. REDUCE 2 for CP/M. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(1):26–27, 37, February 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Moritsugu:1989:NPF

[MG89] Shuichi Moritsugu and Eiichi Goto. A note on the preconditioning for factorization of homogeneous polynomials. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23 (1):9–12, January 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Moufek:2015:MCB

[MG15] Hamza Moufek and Kenza Guenda. McEliece cryptosystem based on punctured convolutional codes and the pseudo-random generators. ACM Communications in Computer Algebra, 49(1):21, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

May:2008:AVO

[MGR^{+}08] John P. May, Mark Giesbrecht, Daniel Roche, Marc Moreno Maza, and Yuzhen Xie. Automatic variable order selection for polynomial system solving (abstract only). ACM Communications in Computer Algebra, 42(1–2):83, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Miguel:2021:CNS

[MGRR21] Daniel Miguel, Andrea Guidolin, Ana Romero, and Julio Rubio. Constructing new spectral systems from simplicial fibrations. ACM Communications in Computer Algebra, 55(3):87–91, September 2021. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3511528.3511534.

Marti:1985:RLB

[MH85] Jed B. Marti and Anthony C. Hearn. REDUCE as a Lisp benchmark. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 19(3):8–16, August 1985. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Marti:1980:SLR

[MHGG80] J. Marti, A. C. Hearn, M. L. Griss, and C. Griss. Standard Lisp report. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14(1):23–43, February 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Middeke:2011:CVN

[Mid11] Johannes Middeke. A computational view on normal forms of matrices of Ore polynomials. ACM Communications in Computer Algebra, 45(3–4):190–191, September 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Middeke:2013:AMI

[Mid13] Johannes Middeke. Annihilating monomials with the integro-differential Weyl algebra. ACM Communications in Computer Algebra, 47(3–4):116–117, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Mignotte:1977:HFA

[Mig77] Maurice Mignotte. How to find algebraic integers whose conjugates lie near the unit circle? SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11–12(4–1):9– 10, November and February 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Mignotte:1980:FUP

[Mig80a] Maurice Mignotte. Factorization of univariate polynomials: a statistical study. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14(4):41–44, November 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Mignotte:1980:TP

[Mig80b] Maurice Mignotte. Tests for polynomials. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14 (1):21–22, February 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Mignotte:1987:IAF

[Mig87] Maurice Mignotte. Inequalities about factors of integer polynomials. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21(4):24, November 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Minimair:2004:MMR

[Min04] Manfred Minimair. MR: Macaulay Resultant package for Maple. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(1):26–27, March 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Minimair:2015:CCA

[Min15] Manfred Minimair. Collaborative computer algebra systems. ACM Communications in Computer Algebra, 49(2):56–57, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Miola:1974:USC

[Mio74] Alfonso Miola. The use of symbolic computation in solving free boundary problems by the Ritz–Galerkin method. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8 (3):103–104, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Miola:1981:BSW

[Mio81] A. M. Miola. A bibliography of Soviet works in algebraic manipulations. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15(1):5–7, February 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Miola:1982:CHC

[Mio82] Alfonso M. Miola. The conversion of Hensel codes to their rational equivalents (or how to solve the Gregory’s open problem). SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16(4):24–26, November 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Miola:1984:CIA

[Mio84] Alfonso Miola. Complexity issues in algebraic algorithms: relevant topics for mathematical education. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18– 19(4–1):26–30, November/ February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Miola:1988:RPA

[Mio88] A. Miola. Report on previous activities at AAECC. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22(1):32–43, January 1988. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Meng:2013:TPG

[MJ13] Lingchuan Meng and Jeremy Johnson. Towards parallel general-size library generation for polynomial multiplication. ACM Communications in Computer Algebra, 47(3–4):106–107, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Middeke:2014:FFF

[MJ14] Johannes Middeke and David J. Jeffrey. Fractionfree factoring revisited. ACM Communications in Computer Algebra, 48(3/4): 130–132, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Meng:2010:AOS

[MJF^{+}10] Lingchuan Meng, Jeremy Johnson, Franz Franchetti, Yevgen Voronenko, Marc Moreno Maza, and Yuzhen Xie. Abstract only: SPIRALgenerated modular FFTs. ACM Communications in Computer Algebra, 44(2): 25–26, June 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Manocha:1996:SAS

[MK96] Dinesh Manocha and Shankar Krishnan. Solving algebraic systems using matrix computations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (4):4–21, December 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Moritsugu:1989:NNE

[MM89] Shuichi Moritsugu and Makoto Matsumoto. A note on the numerical evaluation of arctangent function. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23 (3):8–12, July 1989. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Martinez-Moro:2015:MCP

[MMNR15] Edgar Martínez-Moro, Alejandro P. Nicolás, and Ignacio F. Rúa. Multivariable codes in principal ideal polynomial quotient rings. ACM Communications in Computer Algebra, 49(1): 15, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Marcelo:1999:RPS

[MMR99] A. Marcelo, F. Marcelo, and C. Rodríguez. Radicals of primary submodules. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):17, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Martinez-Moro:2015:LCZ

[MMSY15] Edgar Martínez-Moro, Steve Szabo, and Bahattin Yildiz.

Linear codes over Z_{4}[x]/⟨x^{2}− 2x⟩: dual preserving maps and images as codes over Z_{4}. ACM Communications in Computer Algebra, 49 (1):16, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Moenck:1978:LLB

[Moe78] Robert Moenck. Is a linked list the best storage structure for an algebra system? SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 12(3):20–24, August 1978. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Moeller:1999:GNZ

[Moe99] Michael Moeller. Good nonzeros of polynomials. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):10–11, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Monagan:1993:GFC

[Mon93a] Michael B. Monagan. von zur Gathen’s factorization challenge. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27 (2):13–18, April 1993. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Montes:1993:NCS

[Mon93b] Antonio Montes. Numerical conditioning of a system of algebraic equations with a finite number of solutions using Gröbner bases. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27(1):12–19, January 1993. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Monagan:1996:MSI

[Mon96] M. Monagan. MapleTech special issue. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(1):63, March 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Montes:1999:BAS

[Mon99] A. Montes. Basic algorithms for specialization in Gröbner bases. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):18, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Montes:2001:ASS

[Mon01] Antonio Montes. Abstracts of the Sixth Spanish Meeting on Computer Algebra and Applications: EACA-2000 [Barcelona, Spain, September 6–8, 2001]. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35(4):10–23, December 2001. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (elec-

tronic). URL http://www.
unirioja.es/dptos/dmc/
eaca2000.

Montes:2011:SCG

[Mon11] Antonio Montes. Software for computing the Gröbner cover of a parametric ideal. ACM Communications in Computer Algebra, 45(3–4): 180–182, September 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Monagan:2020:RDJ

[Mon20] Michael Monagan. The 2019 Richard D. Jenks memorial prize. ACM Communications in Computer Algebra, 54(2):61, September 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3427218.3427226.

Monagan:2021:CBP

[Mon21] Michael Monagan. A call to build a publicly accessible library of lecture recordings in computer algebra. ACM Communications in Computer Algebra, 55(3):65–67, September 2021. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3511528.3511529.

Moritsugu:1987:PSS

[Mor87] S. Moritsugu. On the power series solution of a system of algebraic equations. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21(4):14–23, November 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Mora:1989:AAA

[Mor89]  T. Mora, editor.  Ap-

plied Algebra, Algebraic Algorithms and Error-Correcting Codes. 6th International Conference, AAECC-6, Rome, Italy, July 4– 8, 1988. Proceedings, volume 357 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51083-4. LCCN QA268 .A35 1988.

Conference held jointly with ISSAC ’88.

Morales:1999:ACE

[Mor99] G. Fleitas Morales. Algorithms to compute the eigenvalues of a p-adic matrix. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):15, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Mora:2002:IPI

[Mor02] Teo Mora, editor. ISSAC 2002: Proceedings of the 2002 International Symposium on Symbolic and Algebraic Computation, July 07–10, 2002, Université de Lille, Lille, France. ACM Press, New York, NY, USA, 2002. ISBN 1-58113-484-3. LCCN QA76.95.

URL http://www.lifl.
fr/ISSAC2002/.

Moritsugu:2010:RCI

[Mor10] Shuichi Moritsugu. Radius computation for an inscribed pentagon in Sanpou-Hakki (1690). ACM Communications in Computer Algebra, 44(3):127– 128, September 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Moses:1969:ICS

[Mos69] Joel Moses. The integration of a class of special functions with the Risch algorithm. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (13):14–27, December 1969. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Moses:1970:FFL

[Mos70a] Joel Moses. The function of FUNCTION in LISP or why the FUNARG problem should be called the environment problem. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(15):13–27, July 1970. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Moses:1970:SFS

[Mos70b] Joel Moses. Summary of the 1969 FJCC SIGSAM– SIGART technical session. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(14):4–10, January 1970. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Moses:1974:MFY

[Mos74] Joel Moses. MACSYMA — the fifth year. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8 (3):105–110, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Mourrain:2001:IJU

[Mou01] Bernard Mourrain, editor.

ISSAC 2001: July 22–25, 2001, University of Western Ontario, London, Ontario, Canada: proceedings of the 2001 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY, USA, 2001. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Monagan:2010:SPM

[MP10] Michael Monagan and Roman Pearce. Sparse polynomial multiplication and division in Maple 14. ACM Communications in Computer Algebra, 44(4):205– 209, December 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Maza:2011:SBP

[MP11] Marc Moreno Maza and Wei Pan. Solving bivariate polynomial systems on a GPU. ACM Communications in Computer Algebra, 45(2): 127–128, June 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Monagan:2012:PNP

[MP12] Michael Monagan and Roman Pearce. POLY: a new polynomial data structure for Maple 17. ACM Communications in Computer Algebra, 46(3–4):164– 167, September 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Monagan:2014:DMS

[MP14] Michael Monagan and Roman Pearce. The design of Maple’s sum-of-products and POLY data structures for representing mathematical objects. ACM Communications in Computer Algebra, 48(3/4):166–186, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Monagan:2016:FBR

[MP16] Michael Monagan and Roman Pearce. Fermat benchmarks for rational expressionals in Maple. ACM Communications in Computer Algebra, 50(4):188– 190, December 2016. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Maurer:1997:LLC

[MPW97] Markus Maurer, Thomas Pfahler, and Susanne Wetzel. LiDIA — a library for computational number theory. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):54–57, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Muller-Quade:1998:CIF

[MQB98] Jörn Müller-Quade and Thomas Beth. Computing the intersection of finitely generated fields. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(2):62, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Abstract only.

Mora:1984:IBC

[MR84] Ferdinando Mora and Lorenzo Robbiano. The interplay between commutative algebra and computer algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(2):15, May 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

McGrail:2008:TCM

[MS08] Robert W. McGrail and Mary Sharac. Tricoloring as a corrective measure (abstract only). ACM Communications in Computer Algebra, 42(1–2):83– 85, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Maignan:2016:FGL

[MS16a] Aude Maignan and Tony C. Scott. Fleshing out the generalized Lambert W function. ACM Communications in Computer Algebra, 50(2):45–60, June 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Middeke:2016:DBH

[MS16b] Johannes Middeke and Carsten Schneider. Denominator bounds for higher order systems of linear recurrence equations. ACM Communications in Computer Algebra, 50(4):185– 187, December 2016. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Middeke:2018:TDM

[MS18] Johannes Middeke and Carsten Schneider. Towards a direct method for finding hypergeometric solutions of linear first order recurrence systems. ACM Communications in Computer Algebra, 52(3):89–91, September 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

McPhedran:2023:KLC

[MSM23] Ross McPhedran, Tony C.

Scott, and Aude Maignan. The Keiper–Li criterion for the Riemann Hypothesis and generalized Lambert functions. ACM Communications in Computer Algebra, 57(3):85– 110, September 2023. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3637529.3637530.

Myllari:2005:VDS

[MSP05] Aleksandr Mylläri, Tapio Salakoski, and Alexey Pasechnik. On the visualization of the DNA sequence and its nucleotide content. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(4):131–135, December 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Mehta:2018:UBN

[MSS18] Ranjana Mehta, Joydip Saha, and Indranath Sengupta. Unboundedness of Betti numbers of curves. ACM Communications in Computer Algebra, 52(3): 104–107, September 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Morel:2008:LRB

[MSV08] Ivan Morel, Damien Stehlé, and Gilles Villard. From an LLL-reduced basis to another. ACM Communications in Computer Algebra, 42(3):142–143, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Mayr:2021:VGD

[MSV21] Ernst W. Mayr, Werner M. Seiler, and Evgenii V. Vorozhtsov. Vladimir Gerdt dies at 73. ACM Communications in Computer Algebra, 55(1):1–2, March 2021. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3466895.3467684.

Monagan:2015:USI

[MT15] Michael Monagan and Baris Tuncer. Using sparse interpolation to solve multivariate diophantine equations. ACM Communications in Computer Algebra, 49(3):94–97, September 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Murao:2000:AAE

[Mur00]  Hirokazu Murao.  2W -

ary algorithm for extended problem of integer GCD. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):25, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Murakami:2009:CFT

[Mur09] Hiroshi Murakami. A continued fraction type method to find a rational number in a given closed interval whose denominator is minimal. ACM Communications in Computer Algebra, 43(3):88–90, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Murakami:2010:SRC

[Mur10] Hiroshi Murakami. Sturm root counting using Chebyshev expansion. ACM Communications in Computer Algebra, 44(3):129– 131, September 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Murakami:2014:CRN

[Mur14] Hiroshi Murakami. Calculation of rational numbers in an interval whose denominator is the smallest by using FP interval arithmetic. ACM Communications in Computer Algebra, 48(3/4):134–136, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Maza:2011:IMM

[MV11] Marc Moreno Maza and Paul Vrbik. Inverting matrices modulo regular chains. ACM Communications in Computer Algebra, 45(2): 129–130, June 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Monagan:2017:FPM

[MW17] Michael Monagan and Alan Wong. Fast parallel multi-point evaluation of sparse polynomials. ACM Communications in Computer Algebra, 51(1):12–14, March 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Mou:2018:CSL

[MW18] Chenqi Mou and Dongming Wang. On W - characteristic sets of lexicographic Gröbner bases. ACM Communications in Computer Algebra, 52(4): 142–144, December 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Maza:2009:BDP

[MX09] Marc Moreno Maza and Yuzhen Xie. Balanced dense polynomial multiplication on multi-cores. ACM Communications in Computer Algebra, 43(3):85–87, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Miola:1974:CAH

[MY74] Alfonso Miola and David Y. Y. Yun. Computational aspects of Henseltype univariate polynomial greatest common divisor algorithms. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8 (3):46–54, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Myllari:2005:SEH

[Myl05] Aleksandr Mylläri. Stability of expanding homographic configurations: 3D case. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(4):136–138, December 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Mitrofanov:1997:SMB

[MZ97] V. A. Mitrofanov and E. V.

Zima. Standalone MCRbased numerical engine. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):59–60, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Nabeshima:2007:PPC

[Nab07] Katsusuke Nabeshima. PGB: a package for computing parametric Gröbner and related objects. ACM Communications in Computer Algebra, 41(3):104– 105, September 2007. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Nagasaka:2004:TMA

[Nag04]  Kosaku Nagasaka.  To-

wards more accurate separation bounds of empirical polynomials. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(4):119–129, December 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Nagasaka:2007:SPM

[Nag07] Kosaku Nagasaka. SNAP package for Mathematica. ACM Communications in Computer Algebra, 41(3): 105–106, September 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Nagasaka:2008:APG

[Nag08] Kosaku Nagasaka. Approximate polynomial GCD over integers. ACM Communications in Computer Algebra, 42(3):124–126, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Nagasaka:2012:BEA

[Nag12] Kosaku Nagasaka. Backward error analysis of approximate Gröbner basis. ACM Communications in Computer Algebra, 46(3–4): 116–117, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Nagasaka:2017:SBA

[Nag17] Kosaku Nagasaka. Seeking better algorithms for approximate GCD. ACM Communications in Computer Algebra, 51(1):15–17, March 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Nagasaka:2021:AGR

[Nag21] Kosaku Nagasaka. Approximate GCD by relaxed NewtonSLRA algorithm. ACM Communications in Computer Algebra, 55(3):97– 101, September 2021. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3511528.3511536.

Naiman:2022:AVS

[Nai22] Aaron E. Naiman. Automated vector space proofs using Mathematica. ACM Communications in Computer Algebra, 56(1):1–13, March 2022. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572865.3572866.

Namba:1986:SIU

[Nam86]  K. Namba.  Some im-

provements on Utah standard LISP. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 20 (1–2):29–36, February/May 1986. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Nataf:1992:ASN

[Nat92] Jean-Michel Nataf. Algorithm of simplification of nonlinear equations systems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26 (3):9–16, August 1992. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Nair:2019:LCE

[NDSM19] Akshar Nair, James Davenport, Gregory Sankaran, and Scott McCallum. Lazard’s CAD exploiting equality constraints. ACM Communications in Computer Algebra, 53(3):138– 141, September 2019. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Neidleman:1971:UEF

[Nei71] L. D. Neidleman. A user’s examination of the formula manipulation language — SYMBAL. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (20):8–24, October 1971. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Neubuser:1969:SSC

[Neu69] Joachim Neubüser. From the SIGSAM Subcommittee on Computation of Algebraic Structures. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(12):17, July 1969. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Neubuser:1980:SRP

[Neu80] J. Neubüser. Some remarks on a proposed taxonomy for algebraic computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14 (1):19–20, February 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Nievergelt:1969:NML

[NFIS69] J. Nievergelt, F. Fischer, M. I. Irland, and J. R. Sidlo. NUCLEOL — a minimal list processor. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(12):40–52, July 1969. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ng:1974:SIC

[Ng74] Edward W. Ng. Symbolic integration of a class of algebraic functions. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(3):99–102, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ng:1979:SAC

[Ng79] Edward W. Ng, editor. Symbolic and algebraic computation: EU-ROSAM ’79, an International Symposium on Symbolic and Algebraic Manipulation, Marseille, France, June 1979, volume 72 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNCSD9. ISBN 0-387-09519-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E88 1979.

Ngo:2011:RGS

[Ngô11] L. X. Châu Ngô. Rational general solutions of first-order algebraic ODEs. ACM Communications in Computer Algebra, 45(3–4): 238, September 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Nievergelt:1970:SGP

[Nie70] J. Nievergelt. Software for graph processing. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(14):11–12, January 1970. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Nievergelt:1984:CGO

[Nie84] J. Nievergelt. Computing with geometric objects. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(2):19, 18, May 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Nagasaka:2013:RQC

[NM13] Kosaku Nagasaka and Takaaki Masui. Revisiting QRGCD and comparison with ExQRGCD. ACM Communications in Computer Algebra, 47(3–4):88– 89, September 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Noda:1997:AMC

[NMS97] Matu-Tarow Noda, Isao Makino, and Toshiaki Saito. Algebraic methods for computing a generalized inverse. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):51–52, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Newman:1991:WLS

[NO91] M. F. Newman and E. A. O’Brien. The Wielandt length of some 3-groups. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(1):50–51, January 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Nagasaka:2023:CGB

[NO23] Kosaku Nagasaka and Ryo Oshimatani. Conditional Gröbner basis: Gröbner basis detection with parameters. ACM Communications in Computer Algebra, 57(3): 160–164, September 2023.

CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3637529.3637540.

Noda:1996:R

[Nod96a] M-T Noda. RIMS ’94. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(1):50–56, March 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Noda:1996:AJC

[Nod96b] Matu-Tarow Noda. Abstracts of Japanese computer algebra meeting in Kyoto. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (2):36–43, June 1996. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Norman:1975:SP

[Nor75a] A. C. Norman. Solutions to problem 8. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(1): 7–9, February 1975. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Norman:1975:SPS

[Nor75b] Arthur C. Norman. The SCRATCHPAD power series package. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9 (1):13–20, February 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Norman:1978:TRS

[Nor78] Arthur Norman. Towards a REDUCE solution to SIG-SAM problem 7. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 12(4):14–18, November 1978. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Norman:1979:TWS

[Nor79]  A. C. Norman.  Test-

ing word-sized numbers for primality. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 13 (4):19–20, November 1979. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Norton:1980:NAL

[Nor80] Lewis M. Norton. A note about Laplace transform tables for computer use. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14(2):30–31, May 1980. CODEN SIGSBZ.

ISSN 0163-5824 (print), 1557-9492 (electronic).

Norfolk:1982:SCR

[Nor82] Timothy S. Norfolk. Symbolic computation of residues at poles and essential singularities. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16 (1):17–23, February 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Nagloo:2018:CPP

[NOT18] Joel Nagloo, Alexey Ovchinnikov, and Peter Thompson. Commuting planar polynomial vector fields for conservative Newton systems. ACM Communications in Computer Algebra, 52(3):59–62, September 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Novocin:2008:FUPa

[Nov08]  Andrew Novocin.  Fac-

toring univariate polynomials over the rationals (abstract only). ACM Communications in Computer Algebra, 42(1–2):85, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Nielsen:1988:NEM

[NP88] Birger Nielsen and Hendrik Pedersen. A note on Einstein metrics — a simple application of a symbolic algebra system. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22(1):7–11, January 1988. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Nakayama:2009:DDB

[NS09] Hiroki Nakayama and Hiroshi Sekigawa. Determining divisibility between polynomials with inexact coefficients. ACM Communications in Computer Algebra, 43(3):91–94, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Numahata:2017:ASN

[NS17] Dai Numahata and Hiroshi Sekigawa. An algorithm for symbolic–numeric sparse interpolation of multivariate polynomials whose degree bounds are unknown. ACM Communications in Computer Algebra, 51(1):18–20, March 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Nagashima:2018:EIS

[NS18a] Hiroki Nagashima and Kiyoshi Shirayanagi. Effect of the interval-symbol method with correct zero rewriting on the δ-LLL algorithm. ACM Communications in Computer Algebra, 52(2):24–31, June 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Numahata:2018:RAS

[NS18b] Dai Numahata and Hiroshi Sekigawa. Robust algorithms for sparse interpolation of multivariate polynomials. ACM Communications in Computer Algebra, 52(4):145–147, December 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Nakos:1997:FFA

[NTW97] George C. Nakos, Peter R.

Turner, and Robert M. Williams. Fraction-free algorithms for linear and polynomial equations. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):11–19, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Novocin:2008:FUPb

[NvH08] Andy Novocin and Mark van Hoeij. Factoring univariate polynomials over the rationals. ACM Communications in Computer Algebra, 42(3):157, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Norman:1983:CVR

[NW83] Arthur C. Norman and Paul S. Wang. A comparison of the Vaxima and Reduce factorization packages. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(1):28–30, February 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Noro:2003:IPD

[NY03] Masayuki Noro and Kazuhiro Yokoyama. Implementation of prime decomposition of polynomial ideals over small finite fields. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37 (3):81–82, September 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

OBrien:1991:PEA

[O’B91] E. A. O’Brien. Providing electronic access to group descriptions. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25 (1):52–56, January 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Okumura:1997:AGB

[Oku97] Kohshi Okumura. An application of Gröbner bases to the classification of nonlinear circuits. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):61, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Ollongren:1984:CEI

[Oll84] Alexander Ollongren. Classroom experience with interactive formula manipulation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18– 19(4–1):31–37, November/ February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Oren:1984:FSA

[Öre84] Tuncer I. Ören. A forum for the synergy of ACM special interest groups. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(3):31–32, August 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Orecchia:1999:IGU

[Ore99] F. Orecchia. Implicitization of a general union of parametric varieties. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):18, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Osoekawa:2006:HF

[OS06] Takeshi Osoekawa and Tateaki Sasaki. Hensel fan. ACM Communications in Computer Algebra, 40(2): 48–49, June 2006. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ocansey:2016:RHP

[OS16] Evans Doe Ocansey and Carsten Schneider. Representation of hypergeometric products in difference rings. ACM Communications in Computer Algebra, 50(4):161–163, December 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Okuda:2023:UIS

[OS23] Kazuki Okuda and Kiyoshi Shirayanagi. Using the interval-symbol method with zero rewriting to factor polynomials over algebraic number fields. ACM Communications in Computer Algebra, 57(2):21– 30, June 2023. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3614408.3614409.

Ostrowski:1999:STC

[Ost99] A. Ostrowski. On the significance of the theory of convex polyhedra for formal algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (1):5, March 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Orozco:2015:OSC

[OVB15] Ana Lucila Sandoval Orozco, Luis Javier García Villalba, and Mario Blaum. Optimum shortened cyclic codes for multiple burst-error correction. ACM Communications in Computer Algebra, 49(1):22–23, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ovchinnikov:2014:KSD

[Ovc14] Alexey Ovchinnikov. Kolchin seminar in differential algebra. ACM Communications in Computer Algebra, 48(2): 67–69, March 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Padovani:2002:MIC

[Pad02]  L. Padovani.  MathML

International Conference 2002: MathML and Technologies for Math on the Web June 28–30, 2002 Hickory Ridge Conference Center, Chicago, Illinois. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36(1):29–30, March 2002. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Page:2007:AOS

[Pag07] William S. Page. Axiom: open source computer algebra system. ACM Communications in Computer Algebra, 41(3):114, September 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Pages:2022:FDO

[Pag22] Raphaël Pagès. Factoring differential operators over algebraic curves in positive characteristic. ACM Communications in Computer Algebra, 56(2):60– 63, June 2022. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572867.3572876.

Pan:2005:PRF

[Pan05] Victor Pan. Polynomial root-finding with matrix eigen-solving. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(3):87, 2005. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). ISSAC 2005 poster abstract.

Paramonov:2015:UUT

[Par15] Sergey Paramonov. Undecidability of the uniqueness testing problem for analytic solutions of PLDE with boundary conditions. ACM Communications in Computer Algebra, 49(1): 27, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Patton:1996:RBC

[Pat96] Charles M. Patton. A representation of branch-cut information. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (2):21–24, June 1996. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Paulhus:2015:DJV

[Pau15] Jennifer Paulhus. Decomposing Jacobian varieties. ACM Communications in Computer Algebra, 49(2): 59, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Phisanbut:2010:GBC

[PBD10] Nalina Phisanbut, Russell J. Bradford, and James H. Davenport. Geometry of branch cuts. ACM Communications in Computer Algebra, 44(3): 132–135, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Prieto:2000:MOA

[PDP00] Hélène Prieto, Stéphane Dalmas, and Yves Papegay. Mathematica as an Open-Math application. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(2):22–26, June 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Pearce:1984:ICB

[Pea84] P. D. Pearce. The introduction of a computerbased tool for solving mathematical problems analytically to an undergraduate course. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18– 19(4–1):38–39, November/ February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Pech:2007:STG

[Pec07] Pavel Pech. Selected topics in geometry with classical vs. computer proving. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2007. ISBN 981-270-942-8 (hardcover). xi + 239 pp. LCCN QA445 .P42 2007. URL http://

www.loc.gov/catdir/toc/
ecip0722/2007028204.html.

Peikert:1987:ATP

[Pei87a] R. Peikert. Automated theorem proving: the resolution method. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21 (3):61–68, August 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Peikert:1987:FPS

[Pei87b] R. Peikert. Formal power series. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21 (3):24–32, August 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Perez:1999:NPA

[Pér99] P. D. González Pérez. A Newton–Puiseux algorithm for the ring C{X_{1}, …, X_{d}}[Y]. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):16, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Perry:2015:SYF

[Per15] John Perry. The skeletons you find when you order your ideal’s closet. ACM Communications in Computer Algebra, 49(2): 62–63, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Petkovsek:2008:SML

[Pet08] Marko Petkovsek. Solving multivariate linear recurrences in wedges (abstract only). ACM Communications in Computer Algebra, 42(1–2):18–19, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Petkovsek:2023:MP

[Pet23] Eva U. Petkovsek. Marko Petkovsek 1955–2023. ACM Communications in Computer Algebra, 57(1):10– 18, March 2023. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3610377.3610380.

Pfeifle:2021:LFP

[Pfe21] Julian Pfeifle. Large final polynomials from integer programming. ACM Communications in Computer Algebra, 55(3):82–86, September 2021. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3511528.3511533.

Pearce:1981:AAO

[PH81] P. D. Pearce and R. J.

Hicks. The application of algebraic optimisation techniques to algebraic mode programs for REDUCE. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15(4):15–22, November 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Pearce:1983:DSE

[PH83] P. D. Pearce and R. J.

Hicks. Data structures and execution times of Algebraic mode programs for REDUCE. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17 (1):31–37, February 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Pigquette:1990:TSF

[Pig90] Jean C. Pigquette. Table of special function integrals. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 24 (4):8–21, October 1990. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Piquette:1989:SFI

[Piq89] J. C. Piquette. Special function integration. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23(2):11–21, April 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Piquette:1990:TSF

[Piq90] Jean C. Piquette. Table of special function integrals. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 24 (4):8–21, October 1990. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Plemmons:1969:SCA

[Ple69] Robert J. Plemmons. A survey of computer applications to semigroups and related structures. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(12):28–39, July 1969. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Pearce:2007:MLH

[PM07] Roman Pearce and Michael Monagan. A Maple library for high performance sparse polynomial arithmetic. ACM Communications in Computer Algebra, 41(3):110–111, September 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Paluck:2019:NBH

[PM19] Garrett Paluck and Michael Monagan. New bivariate Hensel lifting algorithm for n factors. ACM Communications in Computer Algebra, 53(3):142–145, September 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Pascual-Ortigosa:2019:AAR

[POdCW19] Patricia Pascual-Ortigosa, Eduardo Sáenz de Cabezón, and Henry P. Wynn. Algebraic algorithms for the reliability analysis of multistate k-out-of-n systems. ACM Communications in Computer Algebra, 53(3): 146–149, September 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Pohst:1981:CLV

[Poh81] Michael Pohst. On the computation of lattice vectors of minimal length, successive minima and reduced bases with applications. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15(1):37–44, February 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Poinsot:2015:DLA

[Poi15] Laurent Poinsot. Differential (lie) algebras from a functorial point of view. ACM Communications in Computer Algebra, 49(1): 28–29, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ponder:1987:AHA

[Pon87] Carl D. Ponder. Applications of hashing in algebraic manipulation (an annotated bibliography). SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21(4):10–13, November 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ponder:1988:PPS

[Pon88a] Carl G. Ponder. Parallel processors and systems for algebraic manipulation: current work. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22 (3):15–21, July 1988. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ponder:1988:PAA

[Pon88b] Carl G. Ponder. Parallelism and algorithms for algebraic manipulation: current work. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22 (3):7–14, July 1988. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Popov:1997:SFA

[Pop97] Bogdan A. Popov. A system for function approximation properties investigation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):29–30, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Popov:1998:SCP

[Pop98] Atanas Popov. Symbolic computation of potential energy functions. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(2):12–18, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Popov:2000:OSA

[Pop00] Bogdan A. Popov. Optimal starting approximation and iterative algorithm for inverse error function. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):25–26, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Porat:2009:VQA

[Por09] Moshe Porat. A vectorquantization approach to coding systems. ACM Communications in Computer Algebra, 43(3):94–96, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Postma:2013:NGT

[Pos13] E. J. Postma. The new group theory package in Maple 17. ACM Communications in Computer Algebra, 47(3–4):174–178, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Phillips:2015:SNA

[PR15] Bo Phillips and Jace Robinson. Some new almost difference sets via finite fields. ACM Communications in Computer Algebra, 49(1): 21–22, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Prather:1990:RLC

[Pra90] Ronald E. Prather. The relation of Lorenzen calculus to formal language theory. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 24 (2):9–18, April 1990. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Previato:2015:DAA

[Pre15a] Emma Previato. Differential algebraic aspect of orthogonal polynomials and modular forms. ACM Communications in Computer Algebra, 49(1):29–30, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Previato:2015:ECW

[Pre15b] Emma Previato. Evaluation codes and Weierstrass semigroups. ACM Communications in Computer Algebra, 49(1):22, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Prince:2008:CDU

[Pri08] Matthew Prince. The concept of derivatives using abstract algebra (abstract only). ACM Communications in Computer Algebra, 42(1–2):85–86, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

project:2012:PPS

[pro12] The HPC-GAP project. Parallel programming support in GAP. ACM Communications in Computer Algebra, 46(3–4):161–163, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Pritchard:1998:IVP

[PS98] F. Leon Pritchard and William Y. Sit. On initial value problems for ordinary differential-algebraic equations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (2):61, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Abstract only.

Pillwein:2015:EPD

[PS15] Veronika Pillwein and Miriam Schussler. An efficient procedure deciding positivity for a class of holonomic functions. ACM Communications in Computer Algebra, 49(3):90–93, September 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Piehl:1997:GAE

[PST97] Valerie Piehl, Jonathan Sorenson, and Neil Tiedeman. Genetic algorithms for the extended GCD problem (work in progress). SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):34–35, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Peralta:1999:GCG

[PT99] J. Peralta and B. Torrecillas. Graded codes by G-sets. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):18, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Pudwell:2008:HCC

[Pud08] Lara K. Pudwell. How to cleverly count patternavoiding words (abstract only). ACM Communications in Computer Algebra, 42(1–2):18, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Purtilo:1980:IAP

[Pur80] Jim Purtilo. On implementing arbitrary precision arithmetic in NIL: an exercise in data abstraction. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14(1):14–18, February 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Pernet:2003:LBA

[PW03] Clément Pernet and Zhendong Wan. LU based algorithms for characteristic polynomial over a finite field. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37 (3):83–84, September 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Quinn:2008:CSS

[QH08] Robert Quinn and Hoon Hong. Connectivity in semialgebraic sets (abstract only). ACM Communications in Computer Algebra, 42(1–2):86, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Qi:2020:CMC

[Qi20] Jiayue Qi. A calculus for monomials in Chow group of zero cycles in the moduli space of stable curves. ACM Communications in Computer Algebra, 54(3):91–94, September 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3457341.3457344.

Qi:2022:IMC

[Qi22] Jiayue Qi. An identity on multinomial coefficients. ACM Communications in Computer Algebra, 56(2): 68–71, June 2022. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572867.3572878.

Quinn:1997:BRT

[Qui97] D. Dane Quinn. Book review: Topics in Nonlinear Dynamics with Computer Algebra, by Richard H. Rand, Gordon and Breach Scientific Publishers, 1994, ISBN 2-8849-113-9 (hardcover), 2-8849-114-7 (softcover). SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (1):27–28, March 1997. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Quintin:2012:DLL

[Qui12] Guillaume Quintin. The decoding Library for list decoding. ACM Communications in Computer Algebra, 46(3–4):168–170, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Raab:2011:IFT

[Raa11] Clemens G. Raab. Integration in finite terms of non-Liouvillian functions. ACM Communications in Computer Algebra, 45(2): 133–134, June 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Raab:2015:RRN

[Raa15] Clemens G. Raab. Rewrite rules for nested integrals. ACM Communications in Computer Algebra, 49(1): 34, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Radu:2010:AOA

[Rad10] Silviu Radu. Abstract only: An algorithmic approach to Ramanujan’s congruences and related problems. ACM Communications in Computer Algebra, 44(1):11–12, March 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ragulskis:2015:GMO

[Rag15] Minvydas Ragulskis. The generalized multiplicative operator of differentiation for the construction of analytic solitary solutions to nonlinear differential equations. ACM Communications in Computer Algebra, 49(1):32, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Raichev:2011:NSC

[Rai11] Alexander Raichev. New software for computing asymptotics of multivariate generating functions. ACM Communications in Computer Algebra, 45(3–4):183– 185, September 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Rajaram:1980:ROA

[Raj80] Navaratna S. Rajaram.

Random oscillations arising in the MACSYMA implementation of Khachian-type algorithms in linear programming. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14 (2):32–35, May 1980. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Rothstein:1976:STE

[RC76] Michael Rothstein and B. F. Caviness. A structure theorem for exponential and primitive functions a preliminary report. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 10(4):25–28, November 1976. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Rezvani:2005:NPG

[RC05] Nargol Rezvani and Robert M. Corless. The nearest polynomial with a given zero, revisited. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39 (3):73–79, September 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Recio:1997:CCA

[Rec97] Tomas Recio. A course on computer algebra applications. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (2):32, June 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Recio:1999:CWD

[Rec99] T. Recio. Computing Weil’s descente variety. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):12, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Reeves:1991:WOA

[Ree91] Alyson A. Reeves. The worst order is not always the lexicographic order. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(4):18–19, October 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Reeves:1992:EWO

[Ree92] A. A. Reeves. Erratum:

“The Worst Order Is Not Always The Lexicographic Order”. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26 (1):13, January 1992. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Reeves:1997:PW

[Ree97] A. Reeves. PetaFLOPS workshop. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (2):1, June 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Reid:2000:IIS

[Rei00] Greg Reid. ISSAC’2001: International symposium on symbolic and algebraic computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34 (1):32, March 2000. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Reid:2002:SCA

[Rei02]  G. J. Reid.  Symbolic

Computational Algebra 2002: Fields Institute special meeting on Symbolic and Numeric Computation in Geometry, Algebra and Analysis, July 15–19, 2002, The University of Western Ontario, London, Ontario, Canada

http://www.orcca.on.ca/sca2002.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36(1):31, March 2002. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Renbao:1982:AAC

[Ren82] Zhong Renbao. An algorithm for avoiding complex numbers in rational function integration. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16(3):30–32, August 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ren:2015:CTV

[Ren15] Yue Ren. Computing tropical varieties over fields with valuation using classical standard basis techniques. ACM Communications in Computer Algebra, 49(4):127–129, December 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ren:2017:CTV

[Ren17] Yue Ren. Computing tropical varieties using Newton’s polygon method. ACM Communications in Computer Algebra, 51(3):83–85, September 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Roche:2008:DPP

[RG08] Daniel Roche and Mark Giesbrecht. Detecting polynomial perfect powers (abstract only). ACM Communications in Computer Algebra, 42(1–2):87, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Richardson:2009:EPT

[Ric09] David Richardson. Efficient programming techniques for the SACLIB computer algebra library. ACM Communications in Computer Algebra, 43(4):120, December 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Riley:2008:OMA

[Ril08]  Sean Riley.  Origami

and mathematics (abstract only). ACM Communications in Computer Algebra, 42(1–2):86, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Rimey:1984:PSR

[Rim84] Ken Rimey. Problem section (rotating fluids): a system of polynomial equations and a solution by an unusual method. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(1):30–32, February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Rich:1996:FEB

[RJ96] Albert D. Rich and David J. Jeffrey. Function evaluation on branch cuts. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(2):25–27, June 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Roanes-Lozano:1998:AIA

[RL98] E. Roanes-Lozano. Announcements: 1999 IMACS Applications of Computer Algebra Conference (IMACS-ACA ’99), Euroforum, El Escorial, Madrid, Spain, 24–27 June 1999 (Scientific Program 25–27 June 1999). SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (3):31, September 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Roach:2008:SIQ

[Roa08] Kelly Roach. Solving integrals with the quantum computer algebra system. ACM Communications in Computer Algebra, 42(3): 167–168, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Robino:2001:ASS

[Rob01] Julio Robino. Abstracts of the Seventh Spanish Meeting on Computer Algebra and Applications: EACA-2001 [Ezcaray, La Rioja, Spain, September 12–14, 2001]. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35 (4):24–39, December 2001. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). URL

                                        http://www.unirioja.es/
                                        dptos/dmc/eaca2001.

Roche:2013:AEC

[Roc13] Dan Roche. Abstracts of the 2013 East Coast Computer Algebra Day. ACM Communications in Computer Algebra, 47(1–2):64– 70, March 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Rodionov:1984:WNC

[Rod84] A. Ya. Rodionov. Work with non-commutative variables in the Reduce-2 system for analytical calculations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18 (3):16–19, August 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Roques:2015:GGD

[Roq15]  Julien Roques.  Galois

groups of difference equations on elliptic curves. ACM Communications in Computer Algebra, 49(1): 27–28, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Roune:2009:SAK

[Rou09] Bjarke Hammersholt Roune.

A slice algorithm for Koszul simplicial complexes on the LCM lattice of monomial ideals. ACM Communications in Computer Algebra, 43(3):96–98, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Rowan:1981:EPS

[Row81] William H. Rowan. Efficient polynomial substitutions of a sparse argument. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15(3):17–23, August 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Royle:1991:CVT

[Roy91] Gordon F. Royle. Constructing the vertex-transitive graphs on 24 vertices. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(1):56–59, January 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Renschuch:2003:CCP

[RRRA03] Bodo Renschuch, Hartmut Roloff, Georgij G. Rasputin, and Michael Abramson.

Contributions to constructive polynomial ideal theory XXIII: forgotten works of Leningrad mathematician N. M. Gjunter on polynomial ideal theory. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(2):35– 48, June 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Translated by Michael Abramson. Originally appeared in Wiss. Z. d. Pädagogische Hochschule Potsdam 31 (1987), 111— 126.

Romero:2020:NKM

[RRSS20] Ana Romero, Julio Rubio, Francis Sergeraert, and Markus Szymik. A new Kenzo module for computing the Eilenberg–Moore spectral sequence. ACM Communications in Computer Algebra, 54(2):57–60, September 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3427218.3427225.

Rochon:1975:SSP

[RS75] M. Rochon and H. Strubbe. Solution to SIGSAM problem no. 5 using SCHOONSCHIP. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(4):30–38, November 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Rowney:1989:FFM

[RS89] K. T. Rowney and R. D. Silverman. Finite field manipulations in Macsyma. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23(1):39–48, January 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Rubio:1991:PCH

[RS91] Julio Rubio and Francis Sergeraert. A program computing the homology groups of loop spaces. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(4):20–24, October 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Rueda:2015:AAP

[RSS15] S. L. Rueda, J. Sendra, and J. R. Sendra. An approximate algorithm to parametrize algebraic curves. ACM Communications in Computer Algebra, 49(2): 54–55, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Roch:1989:CAM

[RSSRV89] Jean-Louis Roch, Pascale Senechaud, Françoise Siebert-Roch, and Gilles Villard. Computer algebra on MIMD machine. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23(1):16–32, January 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Raulefs:1979:SSS

[RSSU79] P. Raulefs, J. Siekmann, P. Szabó, and E. Unvericht. A short survey on the state of the art in matching and unification problems. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 13(2):14–20, May 1979. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Reid:2002:GCD

[RSV02] Greg Reid, Chris Smith, and Jan Verschelde. Geometric completion of differential systems using numeric-symbolic continuation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36 (2):1–17, June 2002. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ronveaux:1988:PSR

[RT88] A. Ronveaux and G. Thiry.

Polynomial solution of recurrence relation and differential equation. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22(4):9–19, October 1988. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Rueda:2005:FFA

[Rue05] Sonia L. Rueda. Finite fans, actions of tori and D-modules. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39 (3):90–92, 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). ISSAC 2005 poster abstract.

Rueda:2010:LDI

[Rue10] Sonia L. Rueda. Linear differential implicitization and differential resultants. ACM Communications in Computer Algebra, 44(3):136– 137, September 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Rueda:2015:DED

[Rue15] Sonia L. Rueda. Differential elimination by differential specialization of Sylvester style matrices. ACM Communications in Computer Algebra, 49(1):29, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Rump:1977:RRI

[Rum77] Siegfried M. Rump. Real root isolation for algebraic polynomials. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11 (2):2–3, May 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Robertson:2011:ACH

[RVMH11] Aaron Robertson, Vince Vatter, Gizel Maimon, and Qing-Hu Hou. Abstracts of conferences in honor of Doron Zeilberger’s 60th birthday. ACM Communications in Computer Algebra, 45(1):49–61, March 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Rahkooy:2011:URI

[RZ11] Hamid Rahkooy and Zafeirakis Zafeirakopoulos. Using resultants for inductive Gröbner bases computation. ACM Communications in Computer Algebra, 45(2):135–136, June 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Rojas:2020:CCS

[RZ20] J. Maurice Rojas and Yuyu Zhu. A complexity chasm for solving sparse polynomial equations over p-adic fields: extended abstract. ACM Communications in Computer Algebra, 54(3): 86–90, September 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3457341.3457343.

Sakakihara:1997:IBT

[Sak97] Michio Sakakihara. Iterative blow-up time approximation for initial value problems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):36, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Sakata:2019:ERS

[Sak19] Kosuke Sakata. An efficient reduction strategy for signature-based algorithms to compute Gröbner basis. ACM Communications in Computer Algebra, 53(3):81–92, September 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Salvy:1991:EAA

[Sal91] Bruno Salvy. Examples of automatic asymptotic expansions. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25 (2):4–17, April 1991. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Salvy:1993:EPM

[Sal93] Bruno Salvy. Efficient programming in Maple: a case study. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27 (2):1–12, April 1993. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Salem:2011:PCE

[Sal11] Fatima K. Abu Salem. Parallel and cache-efficient Hensel lifting. ACM Communications in Computer Algebra, 45(2):107– 108, June 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Salem:2012:COE

[Sal12] Fatima K. Abu Salem. A cache-oblivious engineering of the G2V algorithm for computing Gröbner bases. ACM Communications in Computer Algebra, 46(3– 4):72–73, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sandewall:1971:PSF

[San71] Erik Sandewall. A proposed solution to the FUNARG problem. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (17):29–42, January 1971. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sandling:1991:GRP

[San91] Robert Sandling. A group ring package for Cayley. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(1):60–64, January 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sanchez:2015:PLE

[Sán15] Omar León Sánchez. Parametrized logarithmic equations and their Galois theory. ACM Communications in Computer Algebra, 49(1):26, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sanchez:2017:UBO

[Sán17] Omar León Sánchez. An upper bound for the order of a differential algebraic variety. ACM Communications in Computer Algebra, 51(1):

23–25, March 2017. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sanuki:2018:IEG

[San18] Masaru Sanuki. Improvement of EZ-GCD algorithm based on extended Hensel construction. ACM Communications in Computer Algebra, 52(4):148–150, December 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Saouter:2010:HCA

[Sao10] Yannick Saouter. Heuristics for Cannon’s algorithm with an application to Lyons sporadic group. ACM Communications in Computer Algebra, 44(3):138– 139, September 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sarabia:2015:OAI

[Sar15] Manuel González Sarabia. An overview on algebraic invariants and the main parameters of some parameterized codes. ACM Communications in Computer Algebra, 49(1):19–20, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sasaki:2004:TSC

[Sas04] Tateaki Sasaki. A theorem for separating close roots of a polynomial and its derivatives. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38 (3):85–92, September 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sasaki:2022:BBE

[Sas22] Tateaki Sasaki. A bridge between Euclid and Buchberger: (an attempt to enhance Gröbner basis algorithm by PRSs and GCDs). ACM Communications in Computer Algebra, 56(3): 97–124, September 2022. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3594252.3594253.

SanJuan:1999:SPS

[SASG99] J. F. San Juan, A. Abad, S. Serrano, and A. Gavín. Solving problems symbolically by using Poisson Series Processors. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):18–19, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sato:1997:SCS

[Sat97] Yosuke Sato. Set constraint solver — Gröbner bases for non-numerical domains.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):41–42, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Sato:2000:PCB

[Sat00]  Yosuke Sato.  Parallel

computation of Boolean Gröbner bases. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):27–28, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Saunders:1980:SAS

[Sau80] B. David Saunders. A survey of available systems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14(4):12–28, November 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Saunders:1981:A

[Sau81a] B. David Saunders. Abstracts. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15 (2):6–7, May 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Saunders:1981:AMM

[Sau81b] B. David Saunders. Algebra made mechanical. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15(3):23–25, August 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Saunders:2010:AOE

[Sau10] David Saunders. Abstract only: Exact linear algebra. ACM Communications in Computer Algebra, 44(2):24, June 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Schou:1989:RAM

[SB89] W. C. Schou and K. A.

Broughan. The Risch algorithms of MACSYMA and SENAC. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23 (3):19–22, July 1989. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sit:1997:JSC

[SB97] W. Sit and M. Bronstein.

Journal of Symbolic Computation special issue on differential equations and differential algebra. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(1):32, March 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sevyeri:2017:LSF

[SC17] Leili Rafiee Sevyeri and Robert M. Corless. Linearization of a specific family of Bézout matrices. ACM Communications in Computer Algebra, 51(1): 21–22, March 2017. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Schnupp:1971:PFM

[Sch71] Peter Schnupp. Practical formula manipulation for theoretical chemistry. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(17):27–29, January 1971. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Schwartz:1975:WPS

[Sch75] J. T. Schwartz. What programmers should know. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(3):26–27, August 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Schrufer:1982:IEC

[Sch82] E. Schrüfer. An implementation of the exterior calculus in REDUCE: a status report. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16 (4):27–31, November 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Schwarz:1983:RPS

[Sch83] Fritz Schwarz. A REDUCE package for series analysis by Hadamard’s theorem and QD schemes. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(1):38–44, February 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Schrufer:1988:CNE

[Sch88] Eberhard Schrüfer. A comment on “A note on Einstein metrics”. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 22(3):22–26, July 1988. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Schneider:1991:RTC

[Sch91a] Gerhard J. A. Schneider. Representation theory in CAYLEY: tools and algorithms. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25 (1):64–70, January 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Schwarz:1991:MOG

[Sch91b]  Fritz Schwarz.  Mono-

mial orderings and Gröbner bases. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25 (1):10–23, January 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Schwarz:1994:EFL

[Sch94] Fritz Schwarz. Efficient factorization of linear ODE’s. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 28(1):9–17, March 1994. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Schicho:1999:PPA

[Sch99] J. Schicho. The parameterization problem for algebraic surfaces. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):13, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Schenck:2002:CLS

[Sch02] H. Schenck. CBMS lecture series at Texas A&M University: Solving systems of polynomial equations, May 20–24, 2002, by Bernd Sturmfels (UC Berkeley). SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36 (1):23, March 2002. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Schwarz:2007:AW

[Sch07]  Fritz Schwarz.  ALL-

TYPES in the web. ACM Communications in Computer Algebra, 41(3):109– 110, September 2007. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Schneider:2008:MSD

[Sch08a] Carsten Schneider. Multisummation in difference fields (abstract only). ACM Communications in Computer Algebra, 42(1–2):19, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Schwarz:2008:AW

[Sch08b]  Fritz Schwarz.  ALL-

TYPES in the Web. ACM Communications in Computer Algebra, 42(3):185– 187, September 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Schwarz:2016:DSQ

[Sch16] Fritz Schwarz. Decomposing and solving quasilinear second-order differential equations. ACM Communications in Computer Algebra, 50(3):73–82, September 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Su:2008:SOE

[SCWL08] Wei Su, Chuan Cai, Paul S. Wang, and Lian Li. A solution for online entering and editing mathematical formulas (abstract only). ACM Communications in Computer Algebra, 42(1–2):87– 88, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sadeghimanesh:2022:SSN

[SE22] AmirHosein Sadeghimanesh and Matthew England. An SMT solver for nonlinear real arithmetic inside Maple. ACM Communications in Computer Algebra, 56(2):76–79, June 2022.

CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572867.3572880.

Sedjelmaci:2008:SLP

[Sed08] Sidi Mohamed Sedjelmaci.

A straight line program computing the integer greatest common divisor (abstract only). ACM Communications in Computer Algebra, 42(1–2):62– 64, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sedjelmaci:2013:FPG

[Sed13] Sidi M. Sedjelmaci. Fast parallel GCD algorithm of many integers. ACM Communications in Computer Algebra, 47(3–4):92– 93, September 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Seiler:2020:JCD

[Sei20] Werner M. Seiler. Jacques Calmet dies at 77. ACM Communications in Computer Algebra, 54(1):14–15, August 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3419048.3419050.

Sekigawa:2023:AAN

[Sek23] Hiroshi Sekigawa. An approximation algorithm for the nearest decomposable polynomial in the Hamming distance. ACM Communications in Computer Algebra, 57(3):119– 125, September 2023. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3637529.3637532.

Sendra:1990:HMC

[Sen90] Juan R. Sendra. Hankel matrices and computer algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 24 (3):17–26, July 1990. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Senda:2003:IPI

[Sen03] J. Rafael Senda, editor. IS-SAC 2003: Proceedings of the 2003 International Symposium on Symbolic and Algebraic Computation, August 3–6, 2003, Drexel University, Philadelphia, PA, USA. ACM Press, New York, NY, USA, 2003. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Sergeraert:1999:CAT

[Ser99] F. Sergeraert. Constructive algebraic topology. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):13, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Scott:2013:ASG

[SFG13] Tony C. Scott, Greg Fee, and Johannes Grotendorst. Asymptotic series of generalized Lambert W function. ACM Communications in Computer Algebra, 47(3–4):75–83, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Scott:2014:NGL

[SFGZ14] Tony C. Scott, Greg Fee, Johannes Grotendorst, and W. Z. Zhang. Numerics of the generalized Lambert W function. ACM Communications in Computer Algebra, 48(2):42–56, March 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Schulzky:2000:RSR

[SFH00] Christian Schulzky, Astrid Franz, and Karl Heinz Hoffmann. Resistance scaling and random walk dimensions for finitely ramified Sierpinski carpets. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(3):1–8, September 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Gao:2006:WTW

[sG06]  Xiao shan Gao.  Wen-

Tsun Wu’s academic career. ACM Communications in Computer Algebra, 40(2): 60–64, June 2006. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Shayanfar:2009:CAE

[SH09] Nikta Shayanfar and Mahmoud Hadizadeh. Computer algebraic efficiency of matrix polynomials for a system of integral equations. ACM Communications in Computer Algebra, 43(3):101–102, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Shaska:2003:CAA

[Sha03] Tanush Shaska. Computational algebra and algebraic curves. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(4):117–124, December 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Shapeev:2005:ACC

[Sha05] Vasily Shapeev. Application of CAS to constructing schemes of superhigh order of accuracy for heat conduction equation. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(4):139–142, December 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Shaska:2015:MMS

[Sha15a] T. Shaska. Minimal models for superelliptic curves over their minimal field of definition. ACM Communications in Computer Algebra, 49(2):58, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Shaska:2015:GHC

[Sha15b] Tony Shaska. Genus 3 hyperelliptic curves with (2, 4, 4)-split Jacobians. ACM Communications in Computer Algebra, 49(2):55, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Shemyakova:2007:SAM

[She07] Ekaterina Shemyakova. Symbolicalgebraic methods for linear partial differential operators. ACM Communications in Computer Algebra, 41(3):100, September 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Shemyakova:2010:IDT

[She10] Ekaterina Shemyakova. Xand Y -invariants for Darboux transformations. ACM Communications in Computer Algebra, 44(3):149– 150, September 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Shirshov:1999:CAP

[Shi99] A. I. Shirshov. Certain algorithmic problems for Lie algebras. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (2):3–6, June 1999. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Shochat:1982:EMM

[Sho82] David D. Shochat. Experience with the muSIMP/ muMATH-80 symbolic mathematics system. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16(3):16–23, August 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Shor:2015:WWB

[Sho15]  C. Shor.  On the q-

Weierstrass weights of branch points of superelliptic curves. ACM Communications in Computer Algebra, 49(2):58, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sado:1996:PPE

[SHS96] Takahiro Sado, Tomohiro Halaikawa, and Tateaki Sasaki. Pseudo-parallel execution of modular computation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (1):4–8, March 1996. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Shtokhamer:1977:ALS

[Sht77a] R. Shtokhamer. Attempts in local simplification of non nested radicals. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11(1):20–40, February 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Shtokhamer:1977:ULS

[Sht77b] R. Shtokhamer. The use of “LET” statements in producing short comprehensible outputs. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11– 12(4–1):20–22, November and February 1977. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sasaki:2000:HCS

[SI00] Tateaki Sasaki and Daiju Inaba. Hensel construction of F (x, u_{1}, …, u_{ℓ}), ℓ ≥ 2, at a singular point and its applications. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34 (1):9–17, March 2000. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sasaki:2008:CDS

[SI08] Tateaki Sasaki and Daiju Inaba. Convergence domain of series expansions of multivariate algebraic functions. ACM Communications in Computer Algebra, 42(3):127–128, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sasaki:2017:EEH

[SI17] Tateaki Sasaki and Daiju Inaba. Enhancing the extended Hensel construction by using Gröbner basis. ACM Communications in Computer Algebra, 51(1): 26–28, March 2017. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sasaki:2005:ASE

[SIK05] Tateaki Sasaki, Daiju Inaba, and Kentaro Katamachi. An approach to singularity from the extended Hensel construction. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(3):94–95, 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). ISSAC 2005 poster abstract.

Sasaki:2014:SPS

[SIK14] Tateaki Sasaki, Daiju Inaba, and Fujio Kako. Solving parametric sparse linear systems by local blocking. ACM Communications in Computer Algebra, 48 (3/4):137–139, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sit:1989:SCT

[Sit89] W. Y. Sit. Some comments on term-ordering in Gröbner basis computations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23 (2):34–38, April 1989. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sit:2002:ECCa

[Sit02a] William Sit. East Coast Computer Algebra Day.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36(1):19–21, March 2002. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sit:2002:ECCb

[Sit02b]  William Y. Sit.  East

Coast Computer Algebra day: ECCAD 2002 poster and demonstration abstracts. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36 (3):1, September 2002. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sit:2006:ASS

[Sit06]  William Sit.  Abstracts

of special session on differential algebra: American Mathematical Society, Eastern Section Spring Meeting: Stevens Institute of Technology, April 14-15, 2007. ACM Communications in Computer Algebra, 40(3–4):88–107, September–December 2006. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sit:2014:KSD

[Sit14] William Sit. The Kolchin seminar in differential algebra at the Graduate Center of the City University of New York: Summer — 2014. ACM Communications in Computer Algebra, 48(2):70–73, March 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL http:/

/www.sci.ccny.cuny.edu/
~ksda/gradcenter.html.

Stein:2005:SSA

[SJ05] William Stein and David Joyner. SAGE: System for Algebra and Geometry Experimentation. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(2):61–64, June 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Strotmann:2000:OCA

[SK00] Andreas Strotmann and Ladislav Kohout. Open-Math: compositionality achieved at last. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(2):66–72, June 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sasaki:2010:CFP

[SK10] Tateaki Sasaki and Fujio Kako. Computing floatingpoint Gröbner bases accurately. ACM Communications in Computer Algebra, 44(3):142–143, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sasaki:2016:MPM

[SK16] Tateaki Sasaki and Hiroshi Kai. Memories on Professor Matu-tarow Noda. ACM Communications in Computer Algebra, 50(2): 69, June 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Saouter:2010:FIC

[SL10] Yannick Saouter and Gérald Le Mestre. A FPGA implementation of Chen’s algorithm. ACM Communications in Computer Algebra, 44(3):140–141, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Lee:2017:MCA

[sL17] Wen shin Lee. Milestones in Computer Algebra (MICA 2016): a workshop celebrating the research of Erick Kaltofen. ACM Communications in Computer Algebra, 51(2):57–65, June 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Slattery:1991:CDG

[Sla91] Michael C. Slattery. Character degrees of p-groups: a case study. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25 (1):71–74, January 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sun:2015:ISB

[SLW15] Yao Sun, Dongdai Lin, and Dingkang Wang. On implementing signature-based Gröbner basis algorithms using linear algebraic routines from M4RI. ACM Communications in Computer Algebra, 49(2):63–64, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Smit:1974:IN

[Smi74] Jaap Smit. Introduction to NETFORM. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(2): 31–36, May 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Smith:1984:IPT

[Smi84a] C. J. Smith. Implementation of a package of tools for manipulation of sums. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(1):20–24, February 1984. CODEN SIGSBZ.

ISSN 0163-5824 (print), 1557-9492 (electronic).

Smith:1984:PPR

[Smi84b] Carolyn J. Smith. Procedures for polynomial and rational function recognition. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18 (3):2–4, August 1984. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sakakihara:1997:SNM

[SN97] Michio Sakakihara and Shigekazu Nakagawa. Symbolic Newton method for two-parameter eigenvalue problem. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (3):36, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Sasaki:2009:AFP

[SO09] Tateaki Sasaki and Yasutaka Ookura. Approximate factorization of polynomials in Z[x]. ACM Communications in Computer Algebra, 43(3):99–101, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Soiffer:1997:MPR

[Soi97]  Neil Soiffer.  MathML:

a proposal for representing mathematics in HTML. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):44–45, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract

only. See http://www.w3.
org/pub/WWW/TR/WD-math/

for the proposal details.

Sorgatz:1997:DMS

[Sor97] Andreas Sorgatz. Dynamic modules: Software integration in MuPAD. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):45–47, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Slavyanov:1999:DSF

[SPA^{+}99] S. Yu. Slavyanov, A. B.

Pirozhnikov, A. M. Akopyan, W. Lay, and Z. B. Yazik. Database of the special functions of mathematical physics. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (1):21–27, March 1999. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sperner:2008:CTM

[Spe08] Emanuel Sperner. On a combinatorial theorem of Macaulay and its applications to the theory of polynomial ideals. ACM Communications in Computer Algebra, 42(4):192– 201, December 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). Translation by Michael Abramson of Über einen kombinatorischen Satz von Macaulay und seine Anwendungen auf die Theorie der Polynomideale, Abhandlungen aus dem Mathematische Seminaren der Universität Hamburg 7 (1930): 149– 163.

Sage:1968:IPT

[SS68] F. H. Sage and D. V. Smith. The introduction of a paging technique into the symmetric list processor, SLIP. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (9):34–46, April 1968. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Smith:1983:IMM

[SS83] Paul Smith and Leon Sterling. Of integration by man and machine. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(3–4):21–24, August/November 1983. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sasaki:1992:IRC

[SS92] T. Sasaki and M. Sasaki.

On integer-to-rational conversion algorithm. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26(2):19–21, April 1992. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Shallit:1993:BAJ

[SS93] Jeffrey Shallit and Jonathan Sorenson. A binary algorithm for the Jacobi symbol. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27 (1):4–11, January 1993. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sasaki:1997:PRS

[SS97] Tateaki Sasaki and Mutsuko Sasaki. Polynomial remainder sequence and approximate GCD. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):4–10, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sasaki:1998:CEM

[SS98] Tateaki Sasaki and Tomoyuki Sato. Cancellation errors in multivariate resultant computation with floating-point numbers. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (4):13–20, December 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Suzuki:2007:ICC

[SS07] Akira Suzuki and Yosuke Sato. Implementation of CGS and CGB on Risa/Asir and other computer algebra systems using Suzuki–Sato algorithm. ACM Communications in Computer Algebra, 41(3):107–108, September 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Shirayanagi:2009:NMR

[SS09] Kiyoshi Shirayanagi and Hiroshi Sekigawa. A new method of reducing exact computations to obtain exact results. ACM Communications in Computer Algebra, 43(3):102– 104, September 2009. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sekigawa:2010:SBP

[SS10] Hiroshi Sekigawa and Kiyoshi Shirayanagi. Solvability of bivariate polynomial systems under perturbation. ACM Communications in Computer Algebra, 44(3): 147–148, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sekigawa:2011:IRZ

[SS11] Hiroshi Sekigawa and Kiyoshi Shirayanagi. Isolated real zero of a real polynomial system under perturbation. ACM Communications in Computer Algebra, 45(2): 131–132, June 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sekigawa:2012:MPP

[SS12a] Hiroshi Sekigawa and Kiyoshi Shirayanagi. Maximal perturbation for preserving the number of solutions of a polynomial system. ACM Communications in Computer Algebra, 46(3–4):120– 121, September 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Shaska:2012:ECC

[SS12b] Tony Shaska and Daniel E.

Steffy. East Coast Computer Algebra Day 2012 abstracts. ACM Communications in Computer Algebra, 46(1–2):55–59, March/June 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Schost:2013:NLI

[SS13] Éric Schost and Pierre-Jean Spaenlehauer. Newton-like iteration for determinantal systems and structured low rank approximation. ACM Communications in Computer Algebra, 47(3–4):96– 97, September 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Slavyanov:2017:ADH

[SS17] S. Yu. Slavyanov and O. L. Stesik. Antiquantization of deformed Heun class equations as a tool for symbolic generation of Painlevé equations. ACM Communications in Computer Algebra, 51(1):29–31, March 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

SanSegundo:2005:OPC

[SSS05] F. San Segundo, J. R. Sendra, and Juana Sendra. Offsets from the perspective of computational algebraic geometry. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39 (3):87–90, 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). ISSAC 2005 poster abstract.

Salvy:2011:PFF

[SSS^{+}11] Bruno Salvy, Bob Sedgewick, Michele Soria, Wojciech Szpankowski, and Brigitte Vallee. Philippe Flajolet: the father of analytic combinatorics. ACM Communications in Computer Algebra, 45(2):90–91, June 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sasaki:1997:AAC

[SST^{+}97] Tateaki Sasaki, Kousuke Shiihara, Akira Terui, Yuichi Ozaki, and Fujio Kako. Approximate algebraic computation: Practice and problems. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):32, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only.

Sasaki:2002:FSS

[ST02] Tateaki Sasaki and Akira Terui. A formula for separating small roots of a polynomial. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36 (3):19–23, September 2002. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Starkey:1974:SKP

[Sta74] J. Denbigh Starkey. A solution to Kiang’s problem #6. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(1): 2–3, February 1974. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Stan:2010:ASF

[Sta10] Flavia Stan. Algorithms for special functions: computer algebra and analytical aspects. ACM Communications in Computer Algebra, 44(3):156–157, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Staff:2020:RUE

[Sta20]  Staff.  Recent and up-

coming events. ACM Communications in Computer Algebra, 54(2):62–64, September 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3427218.3427227.

Suchen:2022:UST

[STC22] Jazz G. Suchen and Josué Tonelli-Cueto. Ultrametric Smale’s α-theory. ACM Communications in Computer Algebra, 56(2):56– 59, June 2022. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572867.3572875.

Steen:1981:CCL

[Ste81] Lynn Arthur Steen. Computer calculus: long perceived as merely ‘number crunchers,’ computers are now moving into the realm of elegant mathematics. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15 (3):26–27, August 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Steinberg:1982:MSM

[Ste82] Stanly Steinberg. Mathematics and symbol manipulation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16 (3):11–15, August 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Stetter:1996:MEH

[Ste96] Hans J. Stetter. Matrix eigenproblems are at the heart of polynomial system solving. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (4):22–25, December 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Stetter:1999:NPG

[Ste99] Hans J. Stetter. The nearest polynomial with a given zero, and similar problems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(4):2–4, December 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Steinwandt:2000:CST

[Ste00] Rainer Steinwandt. On computing a separating transcendence basis. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(4):3–6, December 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Stetter:2006:ACA

[Ste06] Hans J. Stetter. “Approximate Commutative Algebra”: an ill-chosen name for an important discipline. ACM Communications in Computer Algebra, 40(3–4):79–81, September– December 2006. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Stefanus:2008:BFD

[Ste08] Lim Yohanes Stefanus. De Boor–Fix dual functionals for transformation from polynomial basis to convolution basis. ACM Communications in Computer Algebra, 42(3):146– 148, September 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Steffy:2010:ESL

[Ste10] Daniel E. Steffy. Exact solutions to linear systems of equations using output sensitive lifting. ACM Communications in Computer Algebra, 44(4):160–182, December 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Stickel:1983:NLI

[Sti83] Mark E. Stickel. A note on leftmost innermost term reduction. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(3–4):19–20, August/November 1983. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Stoutemyer:1976:ASA

[Sto76] David R. Stoutemyer. Automatic simplification for the absolute-value function and its relatives. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 10(4):48–49, November 1976. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Stoutemyer:1979:CSM

[Sto79] David R. Stoutemyer. Computer symbolic math & education: a radical proposal. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 13(3):8–24, August 1979. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Stoutemyer:1980:PES

[Sto80] David R. Stoutemyer. PICOMATH-80: an even smaller computer algebra package. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14(3):5–7, August 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Stoutemyer:1984:RPC

[Sto84] David R. Stoutemyer. A radical proposal for computer algebra in education. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18–19(4–1):

40–53, November/February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Stoutemyer:1998:TSI

[Sto98] David R. Stoutemyer. A TI-92 solution to the ISSAC ’97 Challenge Problems. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(1):38–55, March 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Stoutemyer:2007:UCN

[Sto07] David R. Stoutemyer. Useful computations need useful numbers. ACM Communications in Computer Algebra, 41(3):75–99, September 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Stoutemyer:2008:MPF

[Sto08] David R. Stoutemyer. Multivariate partial fraction expansion. ACM Communications in Computer Algebra, 42(4):206–210, December 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Stoutemyer:2009:UNM

[Sto09] David R. Stoutemyer. Unit normalization of multinomials over Gaussian integers. ACM Communications in Computer Algebra, 43(3):73–76, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Stoutemyer:2011:WIC

[Sto11] David R. Stoutemyer. Ways to implement computer algebra compactly. ACM Communications in Computer Algebra, 45(3–4):199– 224, September 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Stoutemyer:2012:SC

[Sto12a] David R. Stoutemyer. Series crimes. ACM Communications in Computer Algebra, 46(1–2):19–52, March/June 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Stoutemyer:2012:SM

[Sto12b] David R. Stoutemyer. Series misdemeanors. ACM Communications in Computer Algebra, 46(3–4):134– 153, September 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Stoutemyer:2013:CAU

[Sto13a] David R. Stoutemyer. A computer algebra user interface manifesto. ACM Communications in Computer Algebra, 47(3–4):130– 165, September 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Stoutemyer:2013:SPF

[Sto13b] David R. Stoutemyer. Simplifying products of fractional powers of powers. ACM Communications in Computer Algebra, 47(1–2): 26–58, March 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Stoutemyer:2017:APC

[Sto17] David R. Stoutemyer. AskConstants proposes concise nonfloats close to floats. ACM Communications in Computer Algebra, 51(1):32–34, March 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Strubbe:1974:PSS

[Str74]  H. Strubbe.  Presenta-

tion of the SCHOONSCHIP system. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8 (3):55–60, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Studzinski:2010:CDN

[Stu10] Grischa Studzinski. Computation of dimensions of non-commutative algebras. ACM Communications in Computer Algebra, 44(3): 151–152, September 2010.

CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Studenicova:2023:AGC

[Stu23] Katarína Studenicová. Arithmetics in generalised Cantor base systems. ACM Communications in Computer Algebra, 57(3):156– 159, September 2023. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3637529.3637539.

Sundblad:1972:S

[Sun72] Y. Sundblad. SYMBAL. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(24):18–19, October 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sundblad:1973:OUO

[Sun73] Yngve Sundblad. One user’s one-algorithm comparison of six algebraic systems on the Y_{2n}-problem. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(28):14–20, December 1973. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sundblad:1974:SMS

[Sun74] Yngve Sundblad. Symbolic mathematical systems now and in the future. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(3):1–8, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sutherland:2017:DCG

[Sut17]  Nicole Sutherland.  A

demonstration of computing Galois groups of polynomials in Magma. ACM Communications in Computer Algebra, 51(3):81–82, September 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sutherland:2018:CGG

[Sut18] Nicole Sutherland. Computation of geometric Galois groups and absolute factorizations. ACM Communications in Computer Algebra, 52(4):151–153, December 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Suzuki:2009:CGB

[Suz09] Akira Suzuki. Computing Gröbner bases within linear algebra and its implementation. ACM Communications in Computer Algebra, 43(3):104–105, September 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Shevchenko:1993:AND

[SV93] Ivan I. Shevchenko and Nikolay N. Vasiliev. Algorithms of numeric deduction of analytical expressions. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27 (1):1–3, January 1993. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sendra:1999:ORP

[SV99] J. R. Sendra and C. Villarino. Optimal reparameterization of polynomial algebraic curves. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):19, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Schlomiuk:2010:ISC

[SV10] Dana Schlomiuk and Nicolae Vulpe. Invariants and symbolic calculations in the theory of quadratic differential systems. ACM Communications in Computer Algebra, 44(3):144– 146, September 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Smit:1981:NCO

[SvH81] J. Smit, J. A. van Hulzen, and B. J. A. Hulshof. NET-FORM and code optimizer manual. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15 (4):23–32, November 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sokolsky:1995:GNR

[SVV95] Andrej G. Sokolsky, Akmal A. Vakhidov, and Nickolay N. Vasiliev. Generating a new recurrence relations system for elliptic Hansen coefficients by means of computer algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(2S (special issue)): 16–18, June 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Szanto:2006:SNC

[SVZ06] Agnes Szanto, Jan Verschelde, and Zhonggang Zeng. Symbolic-numeric computation and applications: list of abstracts. ACM Communications in Computer Algebra, 40(1): 15–17, March 2006. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Shearer:1985:ASS

[SW85] J. M. Shearer and M. A. Wolfe. ALGLIB, a simple symbol-manipulation package. Communications of the ACM, 28(8):820–825, August 1985. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stetter:1996:JSI

[SW96] Hans J. Stetter and Stephen M. Watt. JSC special issue CFP. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (3):32, September 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Stetter:1997:JSC

[SW97] H. J. Stetter and S. M. Watt. Journal of Symbolic Computation special issue on symbolic-numeric algebra for polynomials. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(1):30, March 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Shemyakova:2006:AFL

[SW06] Ekaterina Shemyakova and Franz Winkler. Approximate factorization of linear partial differential operators: full system of invariants for order three. ACM Communications in Computer Algebra, 40(2): 34–35, June 2006. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Storjohann:2014:RAO

[SY14] Arne Storjohann and Shiyun Yang. A relaxed algorithm for online matrix inversion. ACM Communications in Computer Algebra, 48(3/4):140–142, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Steininger:2022:EAS

[SY22] Jakob Steininger and Sergey Yurkevich. Extended abstract for: Solving Rupert’s problem algorithmically. ACM Communications in Computer Algebra, 56(2):32–35, June 2022. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572867.3572870.

Tabuguia:2023:OAF

[Tab23] Bertrand Teguia Tabuguia.

Operations for D-algebraic functions. ACM Communications in Computer Algebra, 57(2):51–56, June 2023. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3614408.3614415.

Tamayo:2015:AFH

[Tam15]  Matthew Tamayo.  Al-

gebraic full homomorphic encryption and resisting Gröbner basis cryptanalysis. ACM Communications in Computer Algebra, 49 (2):63, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Tzoumas:2005:ACC

[TE05] George M. Tzoumas and Ioannis Z. Emiris. Apollonius circle conflict. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(4):143–146, December 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Macsyma:1998:MSI

[Tec98] Technical Staff at Macsyma, Inc. Macsyma solutions to the ISSAC ’97 Challenge Problems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (1):18–37, March 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Tec:2011:SFG

[Tec11] Loredana Tec. A symbolic framework for general polynomial domains in Theorema: applications to boundary problems. ACM Communications in Computer Algebra, 45(3–4):189, September 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Teodorescu:2015:MIP

[Teo15] Iuliana Ciocănea Teodorescu. The module isomorphism problem for finite rings and related results. ACM Communications in Computer Algebra, 49(1):14, March 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Terui:2010:GIM

[Ter10] Akira Terui. GPGCD, an iterative method for calculating approximate GCD, for multiple univariate polynomials. ACM Communications in Computer Algebra, 44(3):153–155, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Terui:2011:ARW

[Ter11] Akira Terui. Abstracts from RIMS workshop on developments in computer algebra research. ACM Communications in Computer Algebra, 45(3–4):162–165, September 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Terui:2013:CAG

[Ter13] Akira Terui. Calculating approximate GCD of multiple univariate polynomials using approximate syzygies. ACM Communications in Computer Algebra, 47 (3–4):110–111, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Thanatipanonda:2008:CGC

[Tha08] Thotsaporn Aek Thanatipanonda. Computergenerated conjectures(!) and proofs(!!) in combinatorial game theory (II) (abstract only). ACM Communications in Computer Algebra, 42(1–2):20, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

SCIE:2010:SCS

[The10a] The SCIEnce project. Symbolic computation software composability protocol and its implementations. ACM Communications in Computer Algebra, 44(4):210– 212, December 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

SCIE:2010:SPP

[The10b] The SCIEnce project. SymGrid-Par: parallel orchestration of symbolic computation systems. ACM Communications in Computer Algebra, 44(4):213–216, December 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Thiery:2000:AIGa

[Thi00a] Nicolas M. Thiéry. Algebraic invariants of graphs: a computer aided study. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):29, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Thiery:2000:AIGb

[Thi00b] Nicolas M. Thiéry. Algebraic invariants of graphs; a study based on computer exploration. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34 (3):9–20, September 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Thiery:2008:IAC

[Thi08] Nicolas M. Thiery. Implementing algebraic combinatorics: Some feedback from the development of MuPAD–Combinat (abstract only). ACM Communications in Computer Algebra, 42(1–2):20, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Thome:2009:ACW

[Tho09] Emmanuel Thomé. Abstracts of the CADO workshop on integer factorization. ACM Communications in Computer Algebra, 43(1):11–14, March 2009. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Tiffany:2015:EMC

[Tif15] Patrice Geary Tiffany. Effectively melding computer algebra systems into the calculus curriculum. ACM Communications in Computer Algebra, 49(2):49–50, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Tian-Jiao:2003:OP

[TJ03] Wu Tian-Jiao. On optimization problems. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(3):87–88, September 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Tabuguia:2022:FAE

[TK22] Bertrand Teguia Tabuguia and Wolfram Koepf. FPS in action: an easy way to find explicit formulas for interlaced hypergeometric sequences. ACM Communications in Computer Algebra, 56(2):46–50, June 2022. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572867.3572873.

Trochez:2022:AML

[TMCP22] Juan Pablo González Trochez, Marc Moreno Maza, Matt Calder, and Erik Postma. Algorithms for multivariate Laurent series. ACM Communications in Computer Algebra, 56(2):64– 67, June 2022. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572867.3572877.

Telen:2018:TNF

[TMV18] Simon Telen, Bernard Mourrain, and Marc Van Barel. Truncated normal forms for solving polynomial systems. ACM Communications in Computer Algebra, 52(3):78–81, September 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Tertychniy:1997:GCA

[TO97] S. I. Tertychniy and I. G. Obukhova. GRGEC: Computer algebra system for applications to gravity theory. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(1):6–14, March 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Tobey:1970:HNG

[Tob70]  R. G. Tobey.  Histori-

cal note on gradient evaluation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (14):35, January 1970. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Toon:2015:ICD

[Too15] Andrew Toon. Investigating complex data and dynamics via computer algebra. ACM Communications in Computer Algebra, 49(2):50, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Tot:2015:SDP

[Tot15] Jonathan Tot. Spinning double pendulum: equilibria and bifurcations. ACM Communications in Computer Algebra, 49(2):57, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Traverso:2000:IAU

[Tra00] Carlo Traverso, editor. IS-SAC 2000: 7–9 August 2000, University of St. Andrews, Scotland: proceedings of the 2000 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY, USA, 2000. ISBN 1-58113-218-2. LCCN QA76.95.I59

2000. URL http://www.
acm.org/pubs/contents/
proceedings/issac/345542/

. ACM order number 505000.

Trager:2006:PIS

[Tra06] Barry Trager, editor. Proceedings of the 2006 International Symposium on Symbolic and Algebraic Computation, Genoa, Italy July 09–12, 2006. ACM Press, New York, NY, USA, 2006. ISBN 1-59593-276-3. LCCN ???? ACM order number 505060.

Trager:2015:GRP

[Tra15] Barry Trager. Good reduction of plane curves. ACM Communications in Computer Algebra, 49(2): 58, June 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Trevisan:1990:RHP

[Tre90] Vilmar Trevisan. Recognition of Hurwitz polynomials. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 24 (4):26–32, October 1990. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Trinks:1984:CNT

[Tri84a] W. Trinks. Comments from number theory on computer algebra. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18 (2):22–23, May 1984. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Trinks:1984:IAR

[Tri84b] W. Trinks. On improving approximate results of Buchberger’s algorithm by Newton’s method. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(3):7–11, August 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Triulzi:2000:OSU

[Tri00] Arrigo Triulzi. OpenMath support under CSL-hosted Reduce. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34 (2):27–30, June 2000. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Trinks:2011:BMS

[Tri11] Wolfgang Trinks. On Buchberger’s method of solving systems of algebraic equations. ACM Communications in Computer Algebra, 45(3–4):150–161, September 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Trotter:1982:SFP

[Tro82] Hale F. Trotter. Statistics on factoring polynomials modp and p-adically. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16(3):24–29, August 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Trott:1997:MSI

[Tro97] M. Trott. Mathematica solutions to the ISSAC system challenge 1997. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(4):2–35, December 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Tsarev:1998:FLP

[Tsa98] S. P. Tsarev. Factorization of linear partial differential operators and Darboux integrability of nonlinear PDEs (ISSAC’98 poster). SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(4):21–28, December 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Tsai:2000:WCI

[Tsa00] Harrison Tsai. Weyl closure of a D-ideal. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):29–30, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Tsigaridas:2008:ACA

[Tsi08] Elias P. Tsigaridas. Algebraic computations and applications to geometry (abstract only). ACM Communications in Computer Algebra, 42(1–2):22, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Tsigaridas:2016:SSR

[Tsi16] Elias Tsigaridas. SLV: a software for real root isolation. ACM Communications in Computer Algebra, 50(3):117–120, September 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Tuncer:2017:ARD

[Tun17] Yusuf Baris Tuncer. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 51(3):98, September 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Tuttens:1973:SPS

[Tut73] W. Tuttens. Solution to problem #2: with “a subroutine package for polynomial manipulation” (Polman). SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (28):13, December 1973. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wu:1999:ATC

[tW99] Wen tsün Wu. Asian Technology Conference in Mathematics: Applications of Technology in Mathematics Research and Teaching for the 21st Century, December 17–21, 1999, Guangzhou Normal University (GNU), Guangzhou, P. R. China. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(1):32, March 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ugalde:2012:PRF

[Uga12] Esteban Segura Ugalde. Polynomial root-finding via structured matrices. ACM Communications in Computer Algebra, 46(3–4):118– 119, September 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ucoluk:1982:PER

[ÜHK82] G. Üçoluk, A. Hacınlıyan, and E. Karabudak. A proposal for extensions to REDUCE. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16 (2):4–14, May 1982. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ullrich:2006:CFF

[Ull06] Peter Ullrich. Closed-form formulas for projecting constructible sets in the theory of algebraically closed fields. ACM Communications in Computer Algebra, 40(2): 45–48, June 2006. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Usevich:2012:VPM

[UM12] Konstantin Usevich and Ivan Markovsky. Variable projection methods for approximate GCD computations. ACM Communications in Computer Algebra, 46(3–4):122–124, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Umeda:2006:CDR

[US06] Yasushi Umeda and Tateaki Sasaki. Computing determinants of rational functions. ACM Communications in Computer Algebra, 40(1):2–8, March 2006. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Ulmer:1995:NKA

[UW95] Felix Ulmer and Jacques-Arthur Weil. Note on Kovacic’s algorithm. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(2):10–11, April 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Valluri:2020:CPT

[Val20] Maheswara Rao Valluri.

Combinatorial primality test. ACM Communications in Computer Algebra, 54(4): 129–133, December 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3465002.3465004.

vanHulzen:1980:CPP

[van80] J. A. van Hulzen. Computational problems in producing Taylor coefficients for the rotating disk problem. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14(2):36– 49, May 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

vandenHeuvel:1986:ASR

[van86] Pim van den Heuvel. Adding statements to RE-DUCE. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 20 (1–2):8–14, February/May 1986. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Varbanova:2015:CCM

[Var15] Elena Varbanova. CAS in the context of methodology of mathematics education. ACM Communications in Computer Algebra, 49(2):48, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Vasconcelos:1999:BCI

[Vas99] W. V. Vasconcelos. Bounds in the computation of the integral closure. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):13, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

vanderHoeven:2004:GT

[vdH04] Joris van der Hoeven. GNU TeXmacs. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38 (1):24–25, March 2004. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

vanderHoeven:2013:GTS

[vdHGG^{+}13] Joris van der Hoeven, Andrey Grozin, Massimiliano Gubinelli, Grégoire Lecerf, François Poulain, and Denis Raux. GNU TEXmacs: a scientific editing platform. ACM Communications in Computer Algebra, 47(1–2): 59–61, March 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

vanderHoeven:2014:SPI

[vdHL14] Joris van der Hoeven and Grégoire Lecerf. Sparse polynomial interpolation in practice. ACM Communications in Computer Algebra, 48(3/4):187–191, September 2014. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

vanderHoeven:2018:FGB

[vdHL18] Joris van der Hoeven and Robin Larrieu. Fast Gröbner basis computation and polynomial reduction for generic bivariate ideals. ACM Communications in Computer Algebra, 52(3):55–58, September 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

vanderHoeven:2020:UCN

[vdHL20] Joris van der Hoeven and Grégoire Lecerf. Ultimate complexity for numerical algorithms. ACM Communications in Computer Algebra, 54(1):1–13, August 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3419048.3419049.

vanderHoeven:2021:SIR

[vdHL21] Joris van der Hoeven and Grégoire Lecerf. On sparse interpolation of rational functions and gcds. ACM Communications in Computer Algebra, 55(1):1–12, March 2021. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3466895.3466896.

vanderHoeven:2011:MQM

[vdHLM^{+}11] Joris van der Hoeven, Grégoire Lecerf, Bernard Mourrain, Philippe Trébuchet, Jérémy Berthomieu, Daouda Niang Diatta, and Angelos Mantzaflaris. Mathemagix: the quest of modularity and efficiency for symbolic and certified numeric computation? ACM Communications in Computer Algebra, 45(3–4):186–188, September 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

vanderHoeven:2020:COB

[vdHM20] Joris van der Hoeven and Michael Monagan. Computing one billion roots using the tangent Graeffe method. ACM Communications in Computer Algebra, 54(3):65–85, September 2020. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3457341.3457342.

vandeRiet:1973:ASP

[vdR73] R. P. van de Riet. The automatic solution of partial differential equations by means of Taylor series using formula-manipulation methods. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (28):33–36, December 1973.

CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Verbaeten:1974:ACP

[Ver74] P. Verbaeten. The automatic construction of pure recurrence relations. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(3):96–98, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Verbaeten:1975:CRZ

[Ver75] P. Verbaeten. Computing real zeros of polynomials with SAC-1. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(2): 8–10, May 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Verschelde:2010:PHC

[Ver10] Jan Verschelde. Polynomial homotopy continuation with PHCpack. ACM Communications in Computer Algebra, 44(4):217– 220, December 2010. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

vanHulzen:1974:FTW

[vH74] J. A. van Hulzen. FOR-MAC today, or what can happen to an orphan. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(1):5–7, February 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

vanHulzen:1975:SPD

[vH75] J. A. van Hulzen. Special purpose differentiation algorithms. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(2): 14–18, May 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

vanHulzen:1982:EAP

[vH82] J. A. van Hulzen and B. J. A. Hulshof. An expression analysis package for REDUCE. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16 (4):32–44, November 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

vanHulzen:1984:SNI

[vH84] J. A. van Hulzen. The symbolic-numeric interface. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(2):27– 28, May 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

vanHoeij:2010:AOC

[vH10] Mark van Hoeij. Abstract only: The complexity of factoring univariate polynomials over the rationals. ACM Communications in Computer Algebra, 44(2): 24, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Velez:2017:LQP

[VHC17] Juan D. Velez, Juan P. Hernandez, and Carlos A. Cadavid. Limits of quotients of polynomial functions in three variables. ACM Communications in Computer Algebra, 51(2):42–56, June 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

vanHoeij:2014:TSS

[vHI14] Mark van Hoeij and Erdal Imamoglu. _{2}F_{1}-type solutions of second order differential equations. ACM Communications in Computer Algebra, 48(3/4):143– 144, September 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

vanHoeij:2012:FFT

[vHK12] Mark van Hoeij and Vijay Jung Kunwar. Finding 2F_{1} type solutions of differential equations with 5 singularities. ACM Communications in Computer Algebra, 46(3–4):96–97, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

vandenHeuvel:1987:SSP

[vHvH87] P. van den Heuvel, B. J. A.

Hulshof, and J. A. van Hulzen. Some simple pretty-print facilities for REDUCE. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21 (1):14–17, February 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Vialla:2013:BWA

[Via13]  Bastien Vialla.  Block

Wiedemann algorithm on multicore architectures. ACM Communications in Computer Algebra, 47(3–4): 102–103, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Villard:1998:BSS

[Vil98] Gilles Villard. Block solution of sparse linear systems over GF(q): the singular case. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (4):10–12, December 1998.

CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Vejdemo-Johansson:2008:ESU

[VJ08] Mikael Vejdemo-Johansson. Enumerating the Saneblidze– Umble diagonal in Haskell (abstract only). ACM Communications in Computer Algebra, 42(1–2):20, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Volpi:1995:RIM

[Vol95] Aldo Volpi. p-th roots for integers modulo p^{n} and for p-Adic integers. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(2S (special issue)): 9–15, June 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Volcheck:1996:TRC

[Vol96] Emil Volcheck. Technical report column: INRIA and RISC-Linz. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (3):26–28, September 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Volcheck:1997:TRCc

[Vol97a] Emil Volcheck. Technical report column. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(4):50–53, December 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Volcheck:1997:TRCa

[Vol97b] Emil Volcheck. Technical report column: INRIA and RISC-Linz. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31 (1):24–25, March 1997. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Volcheck:1997:TRCb

[Vol97c] Emil Volcheck. Technical report column: INRIA-Rocquencourt, Research Institute for Symbolic Computation, University of Passau, Germany. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(2):22–31, June 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Volcheck:1998:PSC

[Vol98] E. Volcheck. Parallel Symbolic Computing Workshop, October 1–3, 1998, Mathematical Sciences Research Institute, Berkeley, CA, USA. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (4):29–30, December 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Voros:1980:SCS

[Vor80] A. Voros. Symbol calculus by symbolic computation and semi-classical expansions. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14 (3):7, August 1980. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

vonTschirnhaus:2003:MRA

[vTG03] Ehrenfried Walter von Tschirnhaus and R. F. Green. A method for removing all intermediate terms from a given equation. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37(1):1–3, March 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Verschelde:2015:PHC

[VY15] Jan Verschelde and Xiangcheng Yu. Polynomial homotopy continuation on GPUs. ACM Communications in Computer Algebra, 49(4):130–133, December 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

vonzurGathen:1992:PFC

[vzG92] Joachim von zur Gathen.

A polynomial factorization challenge. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26 (2):22–24, April 1992. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wall:1989:CS

[Wal89] B. Wall. On the computation of syzygies. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23(4):5–14, October 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1975:FMP

[Wan75] Paul S. Wang. Factoring multivariate polynomials over algebraic number fields in MACSYMA. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(3):21–23, August 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1976:FLM

[Wan76a] Paul S. Wang. Factoring larger multivariate polynomials. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 10 (4):42, November 1976. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1976:ISC

[Wan76b] Paul S. Wang. Implications of symbolic computation for the teaching of mathematics. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 10 (3):15–18, August 1976. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1977:ESD

[Wan77] Paul S. Wang. An efficient squarefree decomposition algorithm. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 11(2):4–6, May 1977. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1978:Aa

[Wan78a] Paul S. Wang. Abstracts. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 12(2):8–9, May 1978. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1978:Ab

[Wan78b] Paul S. Wang. Abstracts.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 12(2):13– 17, May 1978. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1978:SIS

[Wan78c] Paul S. Wang. Seminar on Introduction to Symbolic and Algebraic Manipulation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 12 (2):10–12, May 1978. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1978:SNM

[Wan78d] Paul S. Wang. SIAM 1978 national meeting: invited lectures. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 12 (3):6–7, August 1978. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1979:A

[Wan79a] Paul S. Wang. Abstracts.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 13(1):12–13, February 1979. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1979:ECP

[Wan79b]  Paul S. Wang.  EU-

ROSAM’79 conference proceedings. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 13 (4):8–15, November 1979. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1980:EGA

[Wan80a] Paul S. Wang. The EEZ-GCD algorithm. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14(2):50–60, May 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1980:TAMb

[Wan80b]  Paul S. Wang.  Titles

and abstracts: mathematical symbolic manipulation on the computer. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14(2):20–22, May 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1980:TAMa

[Wan80c]  Paul S. Wang.  Titles

and abstracts: mathematical symbolic manipulation on the computer and applications. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 14(1):6–7, February 1980. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1981:SPA

[Wan81] Paul S. Wang, editor. SYM-SAC ’81: proceedings of the 1981 ACM Symposium on Symbolic and Algebraic Computation, Snowbird, Utah, August 5–7, 1981. ACM Press, New York, NY, USA, 1981. ISBN 0-89791-047-8. LCCN QA155.7.E4 A28 1981. US$23.00. ACM order no. 505810.

Wang:1989:PCL

[Wan89]  D. Wang.  A program

for computing the Liapunov functions and Liapunov constants in Scratchpad II. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23 (4):25–31, October 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1991:TMI

[Wan91] Dongming Wang. A toolkit for manipulating indefinite summations with application to neural networks. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25 (3):18–27, July 1991. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1992:PII

[Wan92] Paul S. Wang, editor. Proceedings of ISSAC ’92. International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY, USA, 1992. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992. ACM order number: 505920.

Wang:1998:UPD

[Wan98]  Dongming Wang.  Un-

mixed and prime decomposition of radicals of polynomial ideals. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (4):2–9, December 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wan:2006:IPAa

[Wan06a] Zhendong Wan. ISSAC 2005 poster abstracts. ACM Communications in Computer Algebra, 40(1):9–14, March 2006. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Wan:2006:IPAb

[Wan06b] Zhendong Wan. ISSAC 2006 poster abstracts. ACM Communications in Computer Algebra, 40(3–4):86– 87, September–December 2006. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Wan:2013:HGF

[Wan13] James G. Wan. Hypergeometric generating functions and series for 1/π. ACM Communications in Computer Algebra, 47(3–4):114– 115, September 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Wang:2016:PSS

[Wan16] Fei Wang. Proof of a series solution for Euler’s trinomial equation. ACM Communications in Computer Algebra, 50(4):136– 144, December 2016. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Watt:1991:IPI

[Wat91] Stephen M. Watt, editor.

ISSAC ’91: proceedings of the 1991 International Symposium on Symbolic and Algebraic Computation, July 15–17, 1991, Bonn, Germany. ACM Press, New York, NY, USA, 1991. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Watt:1996:MC

[Wat96] S. M. Watt. Message from the chair. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (1):1, March 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Watt:1997:SEE

[Wat97a] S. M. Watt. SIGSAM executive elections. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(1):28, March 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Watt:1997:MPC

[Wat97b] Stephen M. Watt. Message from the Past Chair. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):2, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Watt:1998:RSN

[Wat98a] Stephen Watt. Report of the SIGSAM Nominating Committee. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (4):1, December 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Watt:1998:PSM

[Wat98b] Stephen M. Watt. Positions in symbolic mathematical computation at the University of Western Ontario. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (3):1, September 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Watt:2005:AG

[Wat05] Stephen M. Watt. Algebraic generalization. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(3):93–94, 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). ISSAC 2005 poster abstract.

Watt:2008:FDM

[Wat08] Stephen M. Watt. On the functional decomposition of multivariate Laurent polynomials (abstract only). ACM Communications in Computer Algebra, 42(1–2):89–90, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Winkler:1983:CEU

[WB83] F. Winkler and B. Buchberger. A criterion for eliminating unnecessary reductions in the Knuth-Bendix algorithm (abstract). SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(3–4):20, August/November 1983. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wolf:1995:ICR

[WB95] Thomas Wolf and Andreas Brand. Investigating DEs with CRACK and related programs. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(2S (special issue)): 1–8, June 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wilson:2012:RCE

[WBD12] David J. Wilson, Russell J. Bradford, and James H. Davenport. A repository for CAD examples. ACM Communications in Computer Algebra, 46(3–4):67– 69, September 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Wu:2012:EAF

[WCF12] Wenyuan Wu, Jingwei Chen, and Yong Feng. An efficient algorithm to factorize sparse bivariate polynomials over the rationals. ACM Communications in Computer Algebra, 46(3–4): 125–126, September 2012. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Weber:1990:EHP

[Web90] Ken Weber. An experiment in high-precision arithmetic on shared memory multiprocessors. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 24 (2):22–40, April 1990. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Weber:1996:CRE

[Web96] Andreas Weber. Computing radical expressions for roots of unity. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30 (3):11–20, September 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Weeks:1974:FRA

[Wee74] Dennis Weeks. Formal representations for algebraic numbers. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8 (3):91–95, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Weispfenning:1987:AOL

[Wei87] Volker Weispfenning. Admissible orders and linear forms. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21 (2):16–18, February 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Weispfenning:1997:IR

[Wei97] Volker Weispfenning. IS-SAC ’98 Rostock. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):62–63, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Weispfenning:2000:DLE

[Wei00] Volker Weispfenning. Deciding linear-exponential problems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34 (1):30–31, March 2000. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Weimann:2015:CGM

[Wei15] Martin Weimann. Computing gonal maps of algebraic curves. ACM Communications in Computer Algebra, 49(2):55, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Wester:1996:A

[Wes96] Michael Wester. ACA ’97.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(3):29, September 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1982:ARR

[WGD82] Paul S. Wang, M. J. T. Guy, and J. H. Davenport. P - adic reconstruction of rational numbers. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16 (2):2–3, May 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wheeler:1987:BP

[Whe87] F. S. Wheeler. Bell polynomials. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 21 (3):44–53, August 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

White:1978:LST

[Whi78] Jon L. White. LISP/370: a short technical description of the implementation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 12(4):23–27, August 1978. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wibmer:2015:JHT

[Wib15]  Michael Wibmer.  A

Jordan–Hölder theorem for difference algebraic groups. ACM Communications in Computer Algebra, 49(1): 28, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Williamson:1984:TSS

[Wil84] Clifton Williamson, Jr. Taylor series solutions of explicit ODE’s in a strongly typed algebra system. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(1):25–29, February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Willcock:1992:TCL

[Wil92] Diane M. Koorey Willcock.

Tlisp: A Concurrent Lisp for the Transputers. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26(4):15–23, November 1992. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Willis:2001:EDE

[Wil01] Barton L. Willis. An extensible differential equation solver. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 35 (1):3–7, March 2001. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Winkler:1984:CRP

[Win84]  Franz Winkler.  The

Church–Rosser property in computer algebra and special theorem proving: an investigation of criticalpair/completion algorithms (Ph.D. thesis). SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(3):22, August 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Winkler:1997:YGB

[Win97] Franz Winkler. 33 years of Gröbner bases. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3):64, September 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Winkler:1998:YGB

[Win98] F. Winkler. 33 years of Gröbner bases. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(2):38–39, June 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wise:1984:RMQ

[Wis84] David S. Wise. Representing matrices as quadtrees for parallel processors: extended abstract. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(3):24–25, August 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Watt:1999:EM

[WL99] Stephen M. Watt and Xuehong Li. Examples of MathML. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (1):1–4, March 1999. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:2008:FPD

[WL08] Mingsheng Wang and Jinwang Liu. Factorization problems on n-D polynomial matrices (abstract only). ACM Communications in Computer Algebra, 42(1–2):88–89, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Watanabe:1990:IPI

[WN90] Shunro Watanabe and Morio Nagata, editors. IS-SAC ’90: proceedings of the International Symposium on Symbolic and Algebraic Computation: August 20–24, 1990, Tokyo, Japan. ACM Press and Addison-Wesley, New York, NY, USA and Reading, MA, USA, 1990. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

Wolf:2005:PCS

[Wol05] Thomas Wolf. The package CRACK for solving large overdetermined systems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(3):95–96, 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). IS-SAC 2005 poster abstract.

Wolf:2008:SLS

[Wol08a] Thomas Wolf. On solving large systems of polynomial equations appearing in discrete differential geometry. ACM Communications in Computer Algebra, 42(3):131–134, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Wolfe:2008:DTA

[Wol08b] Randall Wolfe. Divisibility tests (abstract only). ACM Communications in Computer Algebra, 42(1–2): 90, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Wolfram:2013:RRC

[Wol13]  Stephen Wolfram.  Re-

membering Richard Crandall (1947–2012). ACM Communications in Computer Algebra, 47(1–2):14– 15, March 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Wang:1973:FMP

[WR73] Paul S. Wang and L. Preiss Rothschild. Factoring multivariate polynomials over the integers. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (28):21–29, December 1973. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Weber:1979:PSR

[WR79] Lawrence A. Weber and Gerhard Rayna. Problem #11 solved in REDUCE: a case study in program translation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 13 (4):21–24, November 1979. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wright:2000:IMW

[Wri00] Francis J. Wright. Interactive mathematics via the Web using MathML. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(2):49–57, June 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wester:1999:IAC

[WRL99] Michael Wester and Eugenio Roanes-Lozano. IMACS Applications of Computer Algebra Conference 1999: Summary. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (2):53–57, June 1999. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:2017:FMR

[WRW17] Fei Wang, Greg Reid, and Henry Wolkowicz. Finding maximum rank moment matrices by facial reduction on primal form and Douglas–Rachford iteration. ACM Communications in Computer Algebra, 51(1):35–37, March 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Wester:1983:EMC

[WS83] Michael Wester and Stanly Steinberg. An extension to MACSYMA’s concept of functional differentiation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 17(3–4):25–30, August/November 1983. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wolf:2013:LSS

[WSW13] Thomas Wolf, Eberhard Schrüfer, and Kenneth Webster. A linear sparse systems solver (LSSS) applied to the classification of integrable non-abelian Laurent ODEs. ACM Communications in Computer Algebra, 47(3–4):122–123, September 2013. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Wiesinger-Widi:2011:GBG

[WW11] Manuela Wiesinger-Widi.

Gröbner bases and generalized Sylvester matrices. ACM Communications in Computer Algebra, 45(2): 137–138, June 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Xambo:1999:UOE

[Xam99] S. Xambó. Using OMEGA for the effective construction, coding and decoding of block error-correcting codes. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):13, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Xia:2007:DTS

[Xia07] Bican Xia. DISCOVERER:

a tool for solving semialgebraic systems. ACM Communications in Computer Algebra, 41(3):102– 103, September 2007. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Xie:2007:FAM

[Xie07]  Yuzhen Xie.  Fast al-

gorithms, modular methods, parallel approaches and software engineering for solving polynomial systems symbolically. ACM Communications in Computer Algebra, 41(3):101, September 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Xue:2015:PIS

[Xue15] Michael Xue. Prove inequalities by solving maximum/minimum problems using a computer algebra system. ACM Communications in Computer Algebra, 49(2):50, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Yanami:2007:SMT

[YA07] Hitoshi Yanami and Hirokazu Anai. SyNRAC: a Maple toolbox for solving real algebraic constraints. ACM Communications in Computer Algebra, 41(3): 112–113, September 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Yan:1991:RFA

[Yan91]  Tak W. Yan.  A ra-

tional function arithmetic and simplification system in Common Lisp. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(4):4–6, October 1991. CODEN SIGSBZ.

ISSN 0163-5824 (print), 1557-9492 (electronic).

Yasuda:2008:DLP

[Yas08] Masaya Yasuda. The discrete logarithm problem on elliptic curves defined over Q (abstract only). ACM Communications in Computer Algebra, 42(1–2):64– 66, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Yasuda:2015:MQC

[YDH^{+}15] Takanori Yasuda, Xavier Dahan, Yun-Ju Huang, Tsuyoshi Takagi, and Kouichi Sakurai. A multivariate quadratic challenge toward post-quantum generation cryptography. ACM Communications in Computer Algebra, 49(3):105– 107, September 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Yao:2019:ISP

[YFJS19] Shanshan Yao, Yifei Feng, Xiaohong Jia, and Li-Yong Shen. ISSAC 2019 software presentations communicated by Yue Ren: a package to compute implicit equations for rational curves and surfaces. ACM Communications in Computer Algebra, 53(2): 33–36, June 2019. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Yuhasz:2008:CMG

[YK08] George Yuhasz and Erich Kaltofen. Computing minimal generators of integer matrix sequences (abstract only). ACM Communications in Computer Algebra, 42(1–2):91, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Youse:2010:AOB

[YS10] Bryan Youse and B. David Saunders. Abstract only: Bitslicing with matrix algorithms oblivious to the data compression. ACM Communications in Computer Algebra, 44(2):26, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Yun:1973:ASS

[Yun73] David Y. Y. Yun. On algorithms for solving systems of polynomial equations. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(27):19–25, September 1973. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Yun:1974:ADR

[Yun74] David Y. Y. Yun. A p-adic division with remainder algorithm. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8 (4):27–32, November 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zassenhaus:1984:PSE

[Zas84] H. Zassenhaus. Position statement for EUROCAL 85: the efficiency of mathematical models. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(2):8–9, May 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zeilberger:1991:MPP

[Zei91] Doron Zeilberger. A Maple program for proving hypergeometric identities. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(3):4–13, July 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zeilberger:2012:FWH

[Zei12] Doron Zeilberger. Farewell to “W” (Herbert Saul Wilf), a true VISIONARY for whom EVERYTHING was INTERTWINED. ACM Communications in Computer Algebra, 46(1–2):12– 13, March/June 2012. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zeller:2008:IPC

[Zel08] Dean Zeller. An introduction to programming curriculum including elements of mathematics, art, and creative writing (abstract only). ACM Communications in Computer Algebra, 42(1–2):91, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zeng:2004:MPC

[Zen04] Zhonggang Zeng. A Matlab package computing polynomial roots and multiplicities. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38 (1):28–29, March 2004. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zeng:2008:AST

[Zen08] Zhonggang Zeng. ApaTools: a software toolbox for approximate polynomial algebra. ACM Communications in Computer Algebra, 42(3):177–179, September 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zidaric:2023:FFF

[ZGA^{+}23] Nusa Zidaric, Guang Gong, Mark Aagaard, Aleksandar Jurisić, and Olexandr Konovalov. FFCSA — finite field constructions, search, and algorithms. ACM Communications in Computer Algebra, 57(2):57– 64, June 2023. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3614408.3614416.

Zhao:2013:GKD

[Zha13] Xiangui Zhao. Gelfand– Kirillov dimensions of differential difference modules via Gröbner bases. ACM Communications in Computer Algebra, 47(3–4):94– 95, September 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zhong:1982:AAC

[Zho82] R. Zhong. An algorithm for avoiding complex numbers in rational function integration. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 16 (3):30–32, August 1982. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zhou:2007:SCT

[Zho07] Wenqin Zhou. Symbolic computation techniques for solving large expression problems from mathematics and engineering. ACM Communications in Computer Algebra, 41(4):125– 126, December 2007. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zhou:2008:EOB

[Zho08] Wei Zhou. Efficient order basis computation (abstract only). ACM Communications in Computer Algebra, 42(1–2):91, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zhou:2015:CAB

[Zho15] Jie Zhou. The computations and applications based on comprehensive Groebner system. ACM Communications in Computer Algebra, 49(3):109–110, September 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zhu:2015:CAB

[Zhu15] Wei Zhu. The computations and applications based on comprehensive Groebner system. ACM Communications in Computer Algebra, 49(3):110, September 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zidaric:2023:HUC

[Zid23] Nusa Zidaric. How to use a CAS for hardware design automation. ACM Communications in Computer Algebra, 57(3):141– 147, September 2023. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic). URL https://

dl.acm.org/doi/10.1145/
3637529.3637536.

Ziegler:2015:ARD

[Zie15] Konstantin Ziegler. Abstracts of recent doctoral dissertations in computer algebra. ACM Communications in Computer Algebra, 49(2):67, June 2015. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zimmer:1976:FPA

[Zim76] Horst G. Zimmer. Factorization of polynomials according to a method of Zassenhaus’. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 10 (4):43–45, November 1976. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zimmer:1984:CNE

[Zim84a] H. G. Zimmer. Congruent numbers — from elementary to algebraic number theory. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18– 19(4–1):54–62, November/ February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zimmer:1984:AAN

[Zim84b] Horst G. Zimmer. Algorithms in algebraic number theory. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18 (2):25–26, May 1984. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zimmer:1997:TEI

[Zim97] Horst G. Zimmer. Three examples of S-integral points on elliptic curves over Q. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(2):10– 21, June 1997. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zima:1998:ACPa

[Zim98a] Eugene Zima. Announcements: Call for posters: IS-SAC ’99 Vancouver. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(3):32, September 1998. CODEN SIGSBZ.

ISSN 0163-5824 (print), 1557-9492 (electronic).

Zima:1998:ACPb

[Zim98b] Eugene Zima. Announcements: Call for posters: IS-SAC ’99 Vancouver. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(4):32, December 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zima:1999:IPA

[Zim99] Eugene Zima. ISSAC’99 poster abstracts: Session I. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33(3):20–27, September 1999. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zima:2000:IPA

[Zim00] Eugene Zima. ISSAC’99 poster abstracts: Session II. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(1):18–31, March 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zimmermann:2003:SCT

[Zim03] Burkhard Zimmermann. A Sister Celine type algorithm for definite summation and integration. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 37 (3):89, September 2003. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zima:2014:DIS

[Zim14] Eugene V. Zima. Direct indefinite summation. ACM Communications in Computer Algebra, 48(3/4):145– 147, September 2014. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zima:2018:AIH

[Zim18] Eugene V. Zima. Accelerating indefinite hypergeometric summation algorithms. ACM Communications in Computer Algebra, 52(3):96–99, September 2018. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zippel:1975:MSP

[Zip75] Richard Zippel. A MAC-SYMA solution to problem 8. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9 (1):10–12, February 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zippel:1984:FCA

[Zip84] Richard Zippel. The future of computer algebra. SIG-SAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(2):6–7, May 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zengler:2011:BGB

[ZK11] Christoph Zengler and Wolfgang Küchlin. Boolean Gröbner bases in SAT solving. ACM Communications in Computer Algebra, 45(2): 141–142, June 2011. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zengler:2011:NAB

[ZKK11] Christoph Zengler, Andreas Kübler, and Wolfgang Küchlin. New approaches to Boolean quantifier elimination. ACM Communications in Computer Algebra, 45(2):139–140, June 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zeng:2013:NMT

[ZL13] Zhonggang Zeng and Tien-Yien Li. NAClab: a Matlab toolbox for numerical algebraic computation. ACM Communications in Computer Algebra, 47(3–4):170– 173, September 2013. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zeilberger:2008:CGC

[ZT08] Doron Zeilberger and Thotsaporn Aek Thanatipanonda. Computer-generated conjectures(!) and proofs(!!) in combinatorial game theory (abstract only). ACM Communications in Computer Algebra, 42(1–2):19– 20, March/June 2008. CO-DEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zhou:2019:FAF

[ZvH19] Yi Zhou and Mark van Hoeij. Fast algorithm for factoring difference operators. ACM Communications in Computer Algebra, 53(3):150–152, September 2019. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Zwillinger:1984:CP

[Zwi84] Daniel Zwillinger. Completing the L-th power in Z[x]. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18(3):20–22, August 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zhang:2015:GKD

[ZZ15] Yang Zhang and Xiangui Zhao. Gelfand–Kirillov dimension of differential difference algebras. ACM Communications in Computer Algebra, 49(1):32, March 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).